A Complete Bibliography of IEEE Transactions on Computers (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2025 Version 1.22

Title word cross-reference

#14 [Fre74b].

(d, k) [Sto70d]. (d, r) [Spi73]. −2 [KO73, Yue75b, Zoh73d]. 1 [Dav78]. 16 [Ath75]. 2 [Agr77, BR71, Chi77, GBMC72, Liu75, SB77a, SB79, TL73b]. 2^{n} × 2^{n} [Ari79]. 3 [Ano76t, Ano77u, Ano78-42, Dav78, SK78a]. 4 [Ano78-43, CSY75]. 40 [GSB^{+}76]. ^{2} [McC79]. A = (r − 1)p^{2} [Rao78]. B [Hwa72]. C [Ben71]. d [PHG77]. ẏ = Ay^{n} [Max77]. H [MS73b, MA71, RRNA74]. j [PHG77]. k [AM79c, FN75, Hwa72, TOS71, Wyn71]. L^{2} [Pav75b]. log_{2} x [KP75]. LU [SK74]. M [AH76, AM73, AM79c, MF78]. GF(2^{m}) [LR71b, Sar78]. N [Boy74, HN71, LSB77, RH75b, Sin72, TM71, TI72a, WM71, AM73, Goo70, MF78, Pea77, RW74, RH74, Wyn71]. n ≥ 1 [Dru71a]. O(n) [CW79b]. P [RW74, Kar74, Kri77, Swa74a]. Q [SB74b, HC78]. QR [SK77a, SK77b]. R [Kun75, SV74, Ulm70, PHG77]. r = 2^{n} [Dru71a]. T [HK77, HK78, RH75b, PHG77]. X [Ath75]. x^{P} + x^{k} + 1 [AM79c]. Y [Ath75]. Z [Cor77].

-2 [Zoh76b]. -adic [Kri77]. -Algorithm [SK74]. -ary [HC78]. -Asummability [SB77a, SB79]. -Class [Ben71]. -Cube [Pea77, Wyn71]. -Dimensional [Goo70]. -Fail-Safe [TM71, TI72a]. -Gate [HK78, HK77]. -Grams [RH74]. -Graphs [HN71]. -Input [SK78a]. -level [Liu75]. -Multiple [Kar74]. -Nearest [FN75]. -Ordered [Swa74a]. -Out-of- [AM73, Dav78, MF78, RW74]. -out-of-n [TOS71]. -Output [SK78a]. -Sequences [AM79c]. -Space [LSB77]. -Stable [AH76]. -Summability [GBMC72, Hwa72]. -Transform [Ulm70]. -Universal [SB74b]. -Valued [SV74, Sin72]. -Vector [Ath75].

1 [Har70a, Har70b, HA70, Ric74a, Sil73, Su75, Wex71]. 10 [Arb70, Gel73a, Rem72, Ros74a]. 11 [Fai73, Kin72, Poe74, Yam70]. 12 [Bib74, Rot70, Su73]. 13 [Fre74b, Win70a]. 14 [Cut70, Jos74]. 15 [Bar74c, She70b]. 16 [Lud74, Skl70]. 17 [Fri70, Lew74, Rip72]. 17-Bit [Pez71a, Pez71b]. 18 [Bar74b, Goo70, Jos72a]. 19 [Ano70a, Cox72, Gil74a, New70]. 1969 [Shi70]. 1972 [Ano72-40]. 1973 [Ano73f]. 1976 [Ano76v, Ano76z, Ano76y, Ano76w]. 1977 [Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78t].

2 [And72, And70, Bas71a, Bon74, DLL^{+}73, Fos73, Gor75, Hop70, Tab70b]. 20 [Ano71a, Den74, Luc72, She70c]. 21 [Ano72a, Arn70, Pav74, Wan72]. 22 [Ano73a, Lom74, Swa72b, Tur70]. 220 [SB71]. 23 [Ano74a, Har74a, Kam70a, McL72]. 24 [Ano75a, Dee72, Rob70b, Vic74]. 25 [Ano76b, Ben70, Coo72, Mil74a, SB71]. 254 [Ano79b]. 26 [Ano77a, Duk74, Flo72, Yoe70a]. 27 [Ano78b, Ber70a, Gag72, Wul74]. 28 [Ano79a, Ham74, Pre70]. 29 [Jum70, Lei74]. 2nd [Jos72a, Luc72].

3 [Aar70, Cha74c, Fly72a, Pag70a, Pav75a, Rux71, Uhr73]. 30 [Fu74, Joh71, Su70]. 31 [Kuo74]. 32 [Voi74, Wag70]. 33 [Bar74a, Paz70a]. 34 [Bar70, Fre74a]. 35 [Kra74, Tsi70]. 36 [Hed70, Mah74]. 360 [Cam72, SB71, Wal74]. 360/25 [SB71]. 37 [Mui74, She70a]. 370 [Flo72]. 38 [BS74, Lev70]. 39 [Kam70b, Wal74].

4 [Bre71c, Cam72, Geo70, Jos72b, Pla73, Sch74b, Sto70a, Voi75]. 40 [Ake70, Sie74]. 40-ns [Pez71a]. 40ns [Pez71b]. 41 [Cha70, Rei74b]. 42 [Dav70a, Rei74a]. 43 [Kim74, Yeh70]. 44 [Brz70, Gri74]. 45 [Paz70b, Ris74]. 46 [Mic74].

5 [Bre73a, Gil74b, Gra70, de 70]. 5th

[Ano79c, Ano79d].

6 [Bas71b, Hob73, Jor74b, Kar70, Tab70b].

7 [Gil70b, Jor74a, Roo73]. 70 [Wak70c]. 74 [Bar74a, Duk74, Fre74a, Fu74, Ham74, Kuo74, Lei74, Voi74, Wal74, Wul74]. 74-26 [Duk74]. 74-27 [Wul74]. 74-28 [Ham74]. 74-29 [Lei74]. 74-30 [Fu74]. 74-31 [Kuo74]. 74-32 [Voi74]. 74-33 [Bar74a]. 74-34 [Fre74a]. 74-39 [Wal74]. 79 [Ano79m, Ano79n, Ano79o, Ano79p, Ano79q].

8 [Els74, Ker73, You70].

9 [CDO78, Feu72, Law70, Ram73a, Swa74b]. 9-V [CDO78]. 900 [Guf71].

A/D [Mor75, Zoh73d]. Abbreviated [CM79c]. ABE [AT70]. Abelian [Cai71, Kar77b]. Abort [FD74a]. Abstract [BT73, Kim79, Poe74, de 70]. Abstracts [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano73e]. Academic [Tab70b]. Accelerated [Cha71]. Acceptors [Har74b]. Access [AU74, Bat77, DI71, Hay75, Kau73, KH77b, KH77c, Law75, Len77, NTA78, Nai79, Nel74, SD79, Sri78, Sti78a, Sto72, Sto75, Wil78]. Accessing [MVW79]. Accuracy [Boh77, Kul78, Mar73c]. Accurate [AS74]. ACDL [BT74]. Achieve [BCH^{+}73]. Achieving [AN74]. Acting [Man71]. Activity [KR75]. Adaptation [Dee72, Mak71]. Adaptive [BI72, BR76b, Bur74, GW77, Ios75, Koh74, Mil77, Mor70a, TI72b]. Addendum [LE79]. Adder [Gad70, She77, Svo70]. Adders [BK71a, LM79, LHSM74, Pra74, Rao72, WL74, WM71]. Addition [AR78, AO79, Ban74b, Bre70a, DS72b, Gar78, HC73, KO73, SW71a, SW73]. Additional [CF70b, Den76, Gou77, WA73]. Additive [EGG72]. Address [HPP77, Sch78c, Sie77]. Addressable [LC79]. adic [Kri77]. Adjacencies [Myk75, Van73a]. Adjacency [Cur77]. Adjunct [Har74b]. Admissibility [GW77]. Adoption [PB71]. Advance [Ano78c, Ano78f, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j]. Advanced [BCA^{+}75]. Advantage [CS70]. Advantages [Feu73]. Aerospace [Mal72]. AFA [GG73]. Affecting [GM76]. Affine [DN77]. AGC [Mor75]. AI [EFLR76]. Aid [Ber70b, KR71a, Kar77a, Lan78]. Aided [EZS78, Lar71, Mar78, SN71]. Air [PS78]. Aircraft [DBM77]. Algebra [DM75, DM76, Eps73b, Lee76, SAPR70, SS72, Svo73, Tis71, VLS70]. Algebraic [BF75a, CSJ74, Den76, Gou77, GM76, HC78, Pra71a, RS71a, Rud74, SG74, UB74, Whi71c]. Algebras [Eps73a, NSR74, Yeh70]. Algol [Fre74a]. Algorithm [AS74, AA77, All75a, All75b, Ami77, Ath75, BC71a, BW73, BI72, Ber72, Bla74, BE70a, BG79, Bro77, Bur74, But71, Car71a, CM74, CDO78, Cha74a, CL77, Cor77, Das71c, Den76, DH79, Eks73, EO78, FT76, FD77a, Fre79a, FC77, FBS75, FV79, FK70b, FO71a, FN75, Ful72, GM72a, GL70, God71, Gou77, Gri78a, HPP79, HO79, Hoe76, HW75, IR76, Jac73, JB76a, JLH73, Kau74a, Ken74, KH77b, KH77c, KS73, KF72a, KNF75, Kro74, LNL72, LH76, Lee78b, Len78, LK75, Lig77, Lin70, Lit74, MT74a, MM73b, MM78b, MF76d, MC70, NK74, Nai79, NF77, Nec79, NW79, OS76, Pap78, Par79, PM77b, Pic74, Pre78b, PR71, Ram75, Ram77a, RT79, Rub74, Rub75a, Rub76a, Rub75c, SG74, SK77a, SK77b, SCK73b, SB77a, SV75, SK74, SM70a]. Algorithm [SMW71, Sil79, Sil77, SCL70, Smi76, SO79b, Sto74b, SKK79, Sur74, SH79, Tab70a, Tan78, TI72b, Tom70, Tom73, Tru71, TE76, Van77, Wak70b, Wan75a, WG73, Weg72, Wet78, Wil70, Wil77b, WL77, YT76, YT71c, Yue77d, Bel76, FT74]. Algorithmic [Bre71d]. Algorithms [Agr75a, AOR74, ABK76, Avi73, Bak75b, BB72, Bar77, BS72, BS78, BF78, BO79, BHP^{+}71, BG78, BV79, Che73, CF78b, Des79, DTI78, Dow73, DW73b, Far73, Fer77, FL75, GHJ77, GC77, Gel73b, Geo70, GL74, GF78, Han75, KKK74, Kri71a, LT73, Liu75, Mah72, Mar72, Mor70b, MM70, Muk79, NTA78, PBC74, RT73b, Ros74a, SCK73a, SR72a, Smi79e, Sta74, Tab74, Tom74, TD70, Tri76, TE77, Tri77a, Yan74a, Yan75a, Bak75a]. Aligned [SP72a]. Alignment [Law75]. Allocation [BE70a, BE70b, GSB^{+}76, RC74, Ung77a]. Allocator [Von75]. Almost [Hal74, Kau74a]. Almost-Optimal [Kau74a]. Alpha [Gri76]. Alphanumeric [Ern72]. Alternate [She78a]. Alternate-Data [She78a]. Alternating [RM78]. Alternative [BCB74, But71, O’C75]. Alternatives [Ano77b, Ano77c]. Ambiguity [BEGO71, MS77a]. Amenable [Buz73]. Among [New70]. Amplitude [SP72b, SP72c]. Analog [Aar70, All75b, BS77, Gil70b, Gra70, KES72, KR71a, Kon71, KV76, Lew77, LC72, Mar78, New75b, RS72a, RC71c, TP72, WK73, Wil77b, You70, Yue78]. Analog-Computer [KES72]. Analog-to-Digital [Aar70]. Analog-to-Gray [Yue78]. Analog/Digital [Gil70b]. Analog/Hybrid [KV76]. Analysis [AD75, AA75, AKT75, AC70, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, ABK76, BW76a, Bha75a, BO70, BCL75, BH71a, Bre71a, BG78, CR72d, CM73b, CA77, CS79, Cle73, CDL78, Cov74, Dav72b, DB74, FB74, FT79, GS79, Geo70, GGMM71, Gil70b, HND75, HO74, HE73, HF71, HI72, KES72, KF79, Kim70, KKA71, KF72b, KF72c, KNF76, Kor79, KS79, Kru71, KM75b, KPS77, LB74, Led72, LF72, Low73, LF77, Mal72, MAB78, Mat71c, Mey76, MMB77, Mor70a, NA74, OH79, PM75a, Pav74, PM78b, RH75a, Rat76, Rip72, RT71, RL72, Sam70a, SF71, Sch79b, Shr71, Sie77, SR71, SR77, Tho74, Tho77, UM77, WKP75, Wal76, WN76, Wel78, Whi72, Whi71c, FT74, NA76b]. Analytic [FS74, Smi79a]. Analytical [Hoe74, Var79]. Analytically [RS71b]. Analyzed [Met73]. Analyzer [DEA73, MN70, MB75]. and-Exclusive-or [Fis74]. Angle [RJ73, RW75]. Angles [Dav77b]. Announcement [Ano76-28, Ano78g, Ano78c, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j]. Announcements [Ano78f]. Announces [Ano78w, Ano78v, Ano78x]. Announcing [Ano76c, Ano76d, Ano77b, Ano77c, Ano78h, Ano78i, Ano78j, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79k, Ano79l]. Anomalies [YRT75]. Anomalous [CM73a, FD79, Pec76]. Answering [BCSW76]. Anticipatory [WY73]. Any [Bar71, CS70, Dru71a, Len77, MF78]. Aperture [Wei75]. Apes [WBRG76]. APL [Har71, TM70, Wal74]. APL/360 [Wal74]. Apparent [Tun70a]. Application [ARS71a, Arm73, Avi71, BM77b, BB74a, CH72a, CH73b, CC73, CM79d, Dav72b, DW72a, DVS74, DV78b, Edw75, Ern72, Fel79, FK70a, GL70, HHJ^{+}76, HK77, HK78, HA70, Hou72, Hun73, KT72, LF72, Mag70, Mam77, MM73b, RS71b, Sar71, Sch73, SF72, Sho74, SB72, Sil79, Su70, TM76, TD73, The75, Tis71, TE76, WG73, Wei75, Yan75b, Yue76, ZHR77a, ZHR77b]. Application-Oriented [Fel79]. Applications [Ale71, Ano72n, BR76a, Bha78, Brz70, CS75, Duk74, HLD70, Har74a, HSS78, HP72, Lem70, LJ79, RS71a, Rem72, RH73, SNOS79, Tri76, Wal74, YT70]. Applicative [FW78]. Applied [Moh71]. Applying [BT73]. Approach [AM79a, Agr79c, Agr75b, AKT75, BK76, BO70, CM77, CCB70, Coo72, Das71b, DK72, Das73, Eks74, Gel73b, Hu75, JB76b, Kam75, KS78a, KNF76, KSW74, KV73, LD75, Lec70b, Man77, MA71, Mil74a, MS75b, MF76c, Mon74, Mur70, PW79b, RM71, Ram76b, Reu75, RV73, RH71, Rud74, SB79, SD74, SR72b, SB77b, Sto77, SC72b, Swa72b, TBY78]. Approaches [AO79, Bre73a, Gil74b, SM71b]. Approximate [BM70, Fos73, HLD70, Mam77, Mar72]. Approximating [CY76]. Approximation [ADMR78, Kar70, KP75, LT70, Pav73, Pav75b, PR74a, Tom74, Van75, Vil74, Yue75e]. Approximations [BB78c, Bre71b, Chu74, Hun71, Pav77]. Arbiter [CM73a, PFL75]. Arbiters [Plu72]. Arbitrarily [Sut74]. Arbitrary [DEA73, Han72, Hen71, KN70, LH76, MS77b, OW72, OY70, Pap77b, PB76a, Sik70, Smi71a, SP72b, SP72c, YT71b]. Architectural [FSLC78, HS75, SNEL79]. Architecture [AAK74, BV78, BS75, Chu79, Col77, Feu73, KK79a, RA78, Sri76]. Architectures [EFLR76, MM78b]. Arcs [CY76, SKK79]. Arctan [SK71]. Area [LH79, Stu75]. Argument [Hoe74]. Arising [CB72, GHJ77]. Arithmetic [Agr75a, Agr77, Agr79a, Ano77g, Atk70, AG73, AT70, Avi71, Avi73, Ban74c, Bar77, BM78, BM77b, BKM73, CR72c, CH72b, Cod73, Dev75, Gar76, Gri78b, HF72, HL86, KKK74, KS77c, Kri77, KPS77, Kul77, LW78, Maj76, Man72b, Meo75, NR75, OW73a, PA78, RPL73, Rao70, RT70, Rao78, Red79, SCK73a, Shi70, Spi73, SM71c, Sut74, SCNS83, Swa90, Tri77b, Tri78, Tun70b, Wak75b, Wil77a, YC76, Yoh73, Yue75b, NA76b]. Arithmetical [TG72a]. Arithmetics [MM73a]. Array [Ake72, BBC72a, BBC72b, BW73, Bla74, CH73a, DS70, Dia76, FL77b, Gar74, HC73, Hua71, HK74, Hwa79, KSA74, KS78c, Kro74, Law75, LR71b, LC79, Lip77, MK71a, MK71b, Maj72a, Maj73, Nut77, Ogu74, PM77a, Pea77, Pez71a, Red79, Rot78, SC74, Tom75, Tri76, Tri77a, Woo79, Pez71b]. Arrays [Agr79c, Agr79a, AB78, BM77a, Ber77, Cur79, Deu70, Dev75, Dir72, Fle79, FM72b, Goo70, GT78a, Ham71, Hou72, HI72, IH72, IC78, Jul78, JF72, Jum74, Kam79, KAN79, LY71, Lar71, Maj76, Man77, MF71, Nic71, OH79, Pap72, Pra78d, PG75, SR74a, Smi79c, Stu75, Sun76, YO70]. Arrivals [Chu70, MN78]. Art [Fre74b, Hob73, Kra74, Lom74, Mui74, Pla73]. Artificial [Arb76, BW76b, Coo72, FL77a, Hol76, Rei74b]. ary [HC78]. Asbestos [PS78]. Aspect [SC74]. Aspects [BR76b, Fre79b, FW78, TL73a, Yip76, Yoe70a]. Assembler [Hus73]. Assembles [EUY^{+}72]. Assembly [Kau74a, WG75]. Assessing [Mal73]. Assessment [Kla71]. Assignment [FG71c, GLL79, HA74, HB70, KR78a, MT70, MT71, MSJ72, Noe73, NR74, NR76b, RSH79, Sau72b, SHR72, TB74]. Assignments [Cur73a, FK78a, Kin71c, MN78, NT78, NT79, Par72a, Pra78a, RN77, Smi74, Tan71]. Associated [KR73]. Associative [CCF79, DB74, FL77b, FS71, Kin71b, Koh74, Lan77, LC79, RTW78, RS71a, SNN70, SH77, SNOS79]. Associative-Search [LC79]. Asummability [SB77a, SB79, Sur74]. Asymmetric [Hu75]. Asymptotes [Win70b]. Asymptotic [BB78c, CF70b, KF72b]. Asymptotically [CW79c]. Asynchronous [Bat70, BH71b, BH73, Bre75, BA71, BW70, CD73, Chu76, Dav77a, Fri70, FG71b, FG71c, FG76, HD71, HC71b, HA76, Jum74, KM73, Kin70, Kin71c, KR78a, LF74, Mag71, MT70, MT71, MS73a, MS74, Mar77, MT74b, NT78, NT79, NM77, PM74, PFL75, Plu72, PR73, Pra78a, Pra78b, PR71, RV74, Sau72a, Sau72b, SM74, Saw74, SM75a, SM75b, Ser73, She70a, SY75, SS71, Sin71, SR71, Smi74, Tan71, Ung71, Wak75a, WM74, Yen79, BT74]. Asynchronously [FM71]. Atomic [PC78]. Atrubin [Goy76]. ATS [SP72a]. Attainable [Bre73b]. Auditing [Ano76u]. Augmented [CH73a, Gar74, KM73, SK74]. August [Ano76v, Ano77w]. Author [Agr79b, All77, Ano70n, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Bab74, BH74a, CM79a, Che72, CBD70, Chu72, CD75, Coh70, Das72, DM76, Edw76, Ekl73, En79, FCC72, FO71b, GRR74, GR73a, GM72b, HC74, HF70, KAA77, Koh72a, Koo78, LM71a, Mar75, Mar76, MB79, Men70a, PR75, Pra71b, Rao78, RS73a, RS73b, SM79, SS75b, To76, Yue77a, Yue77b, Zoh71]. Authors [Ano70-38, Ano70-39, Ano70-40, Ano70-41, Ano70-42, Ano70-43, Ano70-44, Ano70-45, Ano70-46, Ano70-47, Ano70-48, Ano71-37, Ano71-38, Ano71-39, Ano71-40, Ano71-41, Ano71-42, Ano71-43, Ano71-44, Ano71-45, Ano71-46, Ano71-47, Ano71-48, Ano72-65, Ano72-66, Ano72-67, Ano72-68, Ano72-69, Ano72-70, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano74-27, Ano75-27, Ano78-44, Ano79-51, HW78]. Autocorrelation [Mar73b, NM76b, NM76a]. Automata [Arb70, Ban74a, Brz70, Cho74, Cut70, EGG71, Gaj75, Hed70, Hop70, Kam70a, Kam70b, KW70, LF78, Moo71, Pag70b, Par72b, Paz70a, PR74a, Poe74, Pu72, RS71a, SIF76, Tsi70, Wag70, Yan74a, Yan75b, YM78, de 70, de 71]. Automated [CRB72, HND75, KJ79, MOW76, Mor76b, SP72a]. Automatic [Abe72, AA72a, Bec71, Dee72, OS76, PS78, RS74, RJ72, Rei76, Ros75, SW71b, SH72, WN76]. Automation [Cha74c, NR76a]. Automaton [Arn70, Bar71, HN78, MF76d, RS72a]. Automorphism [TV79]. Autonomous [IH72, Kri73, MR72, Wan71]. Autoregressive [Tjo77]. Autotesting [Cio78]. Auxiliary [Dru71a]. Availability [CLL78]. Available [Ano72-50, Ano73h, Ano79-53]. Average [Bri73, Yue74a]. Averaged [Dun75]. Avoidance [HD71]. Axiomatization [Eps73a]. Axioms [San76].

B [Bar74a, Duk74, Fre74a, Fu74, Ham74, Kuo74, Lei74, Voi74, Wal74, Wul74]. B70 [Aar70, And70, Har70a, Sto70a, Tab70b, de 70]. B70-1 [Har70a]. B70-2 [And70]. B70-3 [Aar70]. B70-4 [Sto70a]. B70-5 [de 70]. B70-6 [Tab70b]. B71 [Bas71a, Bas71b, Bre71c, Rux71, Wex71]. B71-1 [Wex71]. B71-2 [Bas71a]. B71-3 [Rux71]. B71-4 [Bre71c]. B71-6 [Bas71b]. B715 [Har71]. B72 [And72, Cam72, Coo72, Cox72, Dee72, Feu72, Flo72, Fly72a, Gag72, Jos72a, Jos72b, Kin72, Luc72, McL72, Rem72, Rip72, Swa72b, Wan72]. B72-10 [Rem72]. B72-11 [Kin72]. B72-17 [Rip72]. B72-18 [Jos72a]. B72-19 [Cox72]. B72-2 [And72]. B72-20 [Luc72]. B72-21 [Wan72]. B72-22 [Swa72b]. B72-23 [McL72]. B72-24 [Dee72]. B72-25 [Coo72]. B72-26 [Flo72]. B72-27 [Gag72]. B72-3 [Fly72a]. B72-4 [Cam72, Jos72b]. B72-9 [Feu72]. B73 [Bre73a, Fai73, Fos73, Gel73a, Hob73, Ker73, Pla73, Ram73a, Roo73, Sil73, Su73, Uhr73]. B73-1 [Sil73]. B73-10 [Gel73a]. B73-11 [Fai73]. B73-12 [Su73]. B73-2 [Fos73]. B73-3 [Uhr73]. B73-4 [Pla73]. B73-5 [Bre73a]. B73-6 [Hob73]. B73-7 [Roo73]. B73-8 [Ker73]. B73-9 [Ram73a]. B74 [Bar74b, Bar74c, BS74, Bib74, Bon74, Cha74c, Den74, Els74, Fre74b, Gil74b, Gil74a, Gri74, Har74a, Jor74a, Jor74b, Jos74, Kim74, Kra74, Lew74, Lom74, Lud74, Mah74, Mic74, Mil74a, Mui74, Pav74, Poe74, Rei74a, Rei74b, Ric74a, Ris74, Ros74a, Sch74b, Sie74, Swa74b, Vic74]. B74-1 [Ric74a]. B74-10 [Ros74a]. B74-11 [Poe74]. B74-12 [Bib74]. B74-13 [Fre74b]. B74-14 [Jos74]. B74-15 [Bar74c]. B74-16 [Lud74]. B74-17 [Lew74]. B74-18 [Bar74b]. B74-19 [Gil74a]. B74-2 [Bon74]. B74-20 [Den74]. B74-21 [Pav74]. B74-22 [Lom74]. B74-23 [Har74a]. B74-24 [Vic74]. B74-25 [Mil74a]. B74-3 [Cha74c]. B74-35 [Kra74]. B74-36 [Mah74]. B74-37 [Mui74]. B74-38 [BS74]. B74-4 [Sch74b]. B74-40 [Sie74]. B74-41 [Rei74b]. B74-42 [Rei74a]. B74-43 [Kim74]. B74-44 [Gri74]. B74-45 [Ris74]. B74-46 [Mic74]. B74-5 [Gil74b]. B74-6 [Jor74b]. B74-7 [Jor74a]. B74-8 [Els74]. B74-9 [Swa74b]. B75 [Gor75, Pav75a, Su75, Voi75]. B75-1 [Su75]. B75-2 [Gor75]. B75-3 [Pav75a]. B75-4 [Voi75]. Balakrishnam [Tab70b]. Balanced [FL75, KV73, LP76, LP78]. Balancing [Cha74d, CK79]. Band [GSJ^{+}76, Sok72]. Bandpass [Wai70]. Bandwidth [CKL77, Rau79a, Rav72]. Bandwidths [CBS71]. Base [Agr75a, Agr77, Cha74a, CR72d, Chi77, Jos74, KO73, LJ79, Mat70, Mui74, OW73a, O’K75, OW72, SCK73a, SCK73b, WM71, Yue75b]. Base- [Agr77, Chi77]. Based [AKT^{+}77, Ber71, Che71b, Che73, DM78b, Gel76, JP73, KH77a, Mey76, MS76, OTK76, PW79a, Per78, Rub79, Sah74, SR72a, TG79, YC76, RT73a]. Bases [GS79]. Basic [Jor74a]. Basis [HA77]. Batch [Ano72-39, MN78, Par71]. Batch-Fabricated [Ano72-39, Par71]. Baye [Tou75]. Bayes [Dev74, GY77]. Bayesian [Hun77]. BCD [Ben77, BR71, Maj73]. Be [SO73, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Kin71a, LF74]. Beam [Wan77]. Behavior [BF72, CM73a, Chu70, CW75, FD79, Hei79, HI72, Rau79b, She77]. Benes [Len78]. Berkeley [DS70]. Bernoulli [EGG72]. Beta [Gri76]. Better [Ano75c, Ano75d, Ano75e, Ano75f]. Between [DM75, DM76, Eps73b, Goo71, Kun79, Lan76, MR72, Moo71, RHR76a, RHR76b, Rob70a, Rus72, Stu75, SS72, Tak76, Tou72e, Ull72, Van73a]. Bias [DF78]. Bidirectional [NT70]. BIFORE [ARS71b]. Big [SO73]. Bilateral [GT78a, Maz71]. Bilevel [JR76]. Bilinear [Lam79a]. Binary [Agr76, AR78, Ake78, AO79, BBC72b, Ban78, Bar77, Ben77, BF75a, Boy74, Bre70a, BP72, BL75, Byr71, CH73a, CS70, CMS79, Cha74d, Che71b, Dav71, Dor72, EI77, FL75, Gad70, Gar72b, GD72, Gra71, GP74, HLD70, Har73, HW79, Hor71a, IC78, JB76a, Klu77, Kri73, LM79, LF74, LW78, Lyn72, MK71b, Maj72a, Maj73, Mar72, Met73, MBM77, NR77b, Obe74, Pak71, PW79a, PM77b, Pea77, PC72, RC71b, RT70, Rat76, Rhy70, RE71, RH74, Rom78, Sag77, SV74, SB72, Sit74, Sok72, Ste72, Sur70a, Tom75, Toy71b, WL74, Wan75c, WC74, Bak76]. Binary-Coded-Decimal [LW78]. Binary-to-Decimal [Hor71a, Rhy70]. Biresidue [Rao70]. Bit [CW79b, EF79, Lin70, LHSM74, Pez71a, PBC74, Rob79a, Pez71b]. Bit-Sequential [CW79b]. Bitonic [NS79]. Bits [JW73]. Black [Mor74]. Black-White [Mor74]. Blake [Bha78]. Blobs [AL71, SCL76]. Block [LR71a, Sto70c]. Blocking [RM71]. Blocks [BT72, MC78, O’K73, Sch78b]. Blurred [Jai77b]. BN [CR72c]. Board [SO73]. Boards [AA72a]. Bodies [SF77, SF78]. Book [Ano72u, Ano72v, Ano72x, Ano72w, Ano72y, Ano72z, Ano73b, Ano73c, Ano73d, Shi75]. Books [Ano77d, Ano78k]. Boolean [Bis73, CM79d, DDSM76, DM75, DM76, Eps73b, MST^{+}75, Yue76, AB78, Bar70, Bau72, Bis70, Bis71, BG78, Bro70b, CRB72, Chi77, Cur76, DS75, Edw75, GGMM71, GBMC72, GC73, GC71b, HG78a, Hay71b, Hig73, HO72, HC71b, Hwa74, Kli74, KM75b, Lee76, Mar71, MPS78, Pap77a, Pic74, PD79, PC78, PM71, Pre71a, PM76b, PMB77, RS71a, Rud74, Rud76, SB77a, She70c, Spi73, SS72, Sur74, Svo73, TD73, Ulr71, Win70a]. Booth [Rub75c]. Bound [Du72, FN75, Han72, Hol72a, KNF75, NF77, TB75, Van72, Vil70, Wet78]. Boundaries [Nev76]. Boundary [ACJ75, FD73, FC77, Hu73, NJ77, PCM70, SB72, Tho77, Vic71, WK73]. Boundary-Value [WK73]. Bounded [NA72, SF78]. Bounds [BCKT79, Bau71, CK75, CK77, Chu71, Chu74, Der73, Dev74, FS70, FB73a, Fre79b, Moo71, Red73b, Rub75b, Tou71, Tou72b, Tou75, Wei72, Yue72]. Bragg [WKP75]. Bragg-Diffraction [WKP75]. Brain [Rei74b]. Branch [FN75, KNF75, NF77, Wet78]. Bridge [Ios78, WS76]. Bridging [Mei74]. Bruijn [Lem70]. Bubble [BD71, Bha75b, BCH^{+}73, CW79a, Che78, Gar72a, KSM76, LC79, LJ79, MBSS75, Min75, WC76]. Bubble-Memory [BCH^{+}73]. Bubbles [Whi71a, Wil77a]. Buddy [Von75]. Buffer [Chu70, Hei79, Hil79, MH70a]. Building [AT70, O’K73, ZT75]. Buried [YRT75]. Burroughs [SB71]. Bus [FT79, Nut77]. Business [Luc72]. Bypass [Kau70a]. Byte [Gri78b, NR75, Red78a]. Byte-Organized [Gri78b, NR75]. Byte-per-Card [Red78a].

C

[Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. C-19 [Ano70a]. C-20 [Ano71a]. C-21 [Ano72a]. C-22 [Ano73a]. C-23 [Ano74a]. C-24 [Ano75a]. C-25 [Ano76b]. C-26 [Ano77a]. C-27 [Ano78b]. C-28 [Ano79a].

Cache [BCB74, Cho76]. Calculation [Fos73, Van75]. Calculus [Lee76, San76, TD73]. Call [Ano77e, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Ano79w, Ano79x, Ano79z, Ano79-27, Ano79-28, Ano79y]. can [LF74]. Canonic [PP75, SR75, SO79a]. Canonical [KS75, KS77b]. Capabilities [PRD71, RM78]. Capability [HPP77]. Capable [Gra73, MKO70]. Capacity [Bit77, Cho76]. Card [Red78a]. Cardiac [CC73]. Cards [AB79b]. CARE’S [Vin78]. Carlo [AA76, SF74]. Carry [Bri73, Chi77, Pra74, She77, Yue74a]. Carry-Save [Pra74]. Cascade [BPC75, Cut70, DSBC74, Mat71b, MBSS75, Pap78, Pap76, SK78a, SK79b]. Cascaded [GC71a]. Cascades [But75, HE79, Kau70a, Kol72, Pap77b, Yoe70b]. Case [Bri73, Cha73b, CDL78, FL77a, MH70b, Rub75b, Yue74a]. CASSM [SNEL79]. Catastrophic [Sri71b]. Categorization [TG72b]. Cause [BH71a, LB74]. Cause-Effect [BH71a, LB74]. Cayley [Vil73a]. CC [Cur71a, HS70, VVA78]. CC-Tables [Cur71a, VVA78]. CCD [FM78]. Cell [BM77a, BB73, Nel74, Ogu74, SM73a, Sri71a]. Cells [But78b, CK74, CW79c, Coy79, Gec77, GT78a, Led72, Liu77a, MH73, PG75]. Cellular [Agr79c, BBC72a, BM77a, But76, CL72, Gar72a, HN78, Hu72, Hua71, HK74, Hwa79, Kau70a, KS78c, Kol72, KC73, LR71b, MK71a, Maj72a, Maj73, Man77, NK75, Pap72, RTW78, RW74, SR74a, Skl70, SCL76, SC74, Sun76, TM70, Tom75, Yam70, YO70, Yoe70b]. Cellular-Array [LR71b]. Certain [BB74a, Erc73, Hen78, Hol72b, Lam79a, Pag72, SK75, Sik70]. Certifying [Lee75]. Chain [FD77a, God73]. Chain-Coded [FD77a]. Chaining [TL73b]. Chains [Wyn71]. Chamber [BD71]. Chaney [Wor79]. Change [Ano73e, BH73, Bre75, CD73, Lil72]. Changes [Chu76, Ung71]. Channel [Gil70a, GLL79, Ric72, Smi79a]. Channel-Assignment [GLL79]. Character [ARS71a, Dos77, Dut74, Gra72, HTD72, Koz71, Nay71, Wil70, Win75]. Character-Recognition [Dos77]. Characteristic [Has74]. Characteristics [BV78, BB73, Cod73, LS71, Pea76]. Characterization [BPC75, DM78b, HA74, Kin71c, Wan75c]. Characterizing [Dir72, TB75]. Characters [DT70, TG72b, TD70]. Charge [PM76a]. Charge-Coupled [PM76a]. Chart [Lan78]. Charts [BM70]. Chebyshev [VD76b]. Check [AM73, BT70b, KM73]. Checked [OW73b]. Checker [Dav78, MF78, Usa75]. Checkers [AR77, Red74]. Checking [AM73, AR77, BCH^{+}73, Bou74a, Bou74b, BG79, CDJ71, CDS73, CHSS73, CSST73, Dav78, DAA79, Dor72, Fon71, FM73, Hol72a, Hsi71b, MF78, Red74, She78b, TB75, Usa75, Wak74, WY75]. Chest [BS76]. Chief [Ano71-49]. Chip [Woo79]. Choice [Dos77, Dui76, Sok72]. Choosing [MG71b, Tou72a]. Chromosome [KKA71]. Chronology [HH76]. Cineangiography [CC73]. CINTEL [DWT70]. Cipher [Rub79]. Circle [MB75]. Circuit [BW76a, BT74, CDJ71, CDJ73, CK77, DTG79, DW73a, FS70, Fri74, GC71b, HH77, Hou72, Kau70b, Mut76, NT70, Obe72, Ogu75, PM78a, Rus72, Saw74, SR73, SM76, SO73, TW77, WR77]. Circuit-to-Pin [Rus72]. Circuitry [BT70b]. Circuits [AB79a, AM73, Ano75-28, Arm72, ATN70, BC71b, Bet71, BF75a, BHP^{+}71, BH73, Bre70b, Bre71d, Bre73c, BW70, CM74, CM77, Cer78, CDO78, CM73a, Cha70, CE75, Cio78, Cur71b, Dan73, Dav77a, Dia75, DAA79, Du71, DW73b, EGG72, EI77, Fai73, FM74, Fis74, FD79, Fri70, Fri74, FG71b, FG71c, FG76, GR72a, GR73b, HD71, Hay76c, Hu73, HA76, Iba74, JM70, Jum70, Kai70, Kau70b, Kin70, KSM76, Kje72, KS71, LR72, Mag71, Mag73, MT70, MT71, Mar79, Mat71b, McC79, MJ77, MJ78, MT74b, MS70, NA72, NA76a, Ols70, PM75a, PR71, RHR76a, RHR76b, Red72a, Red77b, Rip72, Riv77, Rot70, Roy74, SK78a, SK78b, Saw75, SM72b, Ser73, Ser75, She70a, Sin71, Smi78, SHR72, SC72b, Su73, TM76, Tis71, To73, Toh74, TV76, Ung71, Ung77d, Ung77b, Ung77c]. Circuits [Voi77a, Voi77b, Wak74, WM74, WA73, Wyn71, YT71c, YY75, Yua74, ZKK79, Spi73]. Circuits-General [GR72a]. Circular [Red73b, SKK79]. Circulating [Myk75, Van73a]. Class [AOR74, BS72, Ben71, BV79, BR74, Che71a, Cor71a, Dev74, Ehr72, Gel73b, GT74, Har74b, Has73, Has74, HP72, IH72, KS73, Mar73a, Mat71a, PCM70, Red78a, RS71b, Tou72c, Tou75]. Classes [Bau72, Bro70a, DH79, Har73, HI72, JM71, Sag77, Tsi70, Ull72, YT76]. Classification [BC71a, BT70a, BB78b, CH72a, CH73b, DTI78, Edw75, Fri77, FK71, FO71c, FF75, Glu71, Hus72, Hus74, Kit78, LT70, LF72, MM73b, MS75b, NFS77, Oza78, PR78, PF72, Pav74, Sam70a, SH72, Yue76]. Classifier [Sto74a]. Classifiers [Cha74b, Kul78, Tou72d]. Classify [Uhr72]. Classifying [GD72]. Clause [DK72, SD74, Sic76]. Clause-Column [DK72]. Clipped [Sok72]. Clisp [Tei76]. Clock [BH71b, HA76]. Clocked [Sin77]. Closed [DBC71, Kau70b, VN78, Yan74b, ZR72]. Closed-Loop [VN78]. Closure [DW72a, RB76, RK75a, Tsi70, Yan73, YB75, YT76]. Cloud [HEWB72, SP72a]. Cluster [BCL75, KF72c, KNF76]. Clustering [Ake70, FS78, GL70, JP73, KF72b, KNF75, PS71, RL72, Wet78, WL77]. Clusters [SV75, Zah71]. Coaxial [Ano72-39, Par71]. Cobol [Sil73]. Code [Agr76, Ano70o, Dev79, En77, FGR71, FR72, Gre75, IC78, Kau70c, MF78, Obe74, Paa79, PB76b, TOS71, Yue78]. Coded [Avi73, FD77a, FL75, LW78, Wei75]. Codes [AM73, AR77, Avi71, BM74, BM78, Bha78, BCC78, CR72c, CD78, FG71a, Gri78b, HP72, KM73, LW78, NR75, PA78, Rao70, Rao78, Red78a, Sri71a, Sri71b, Sun79, TL73b, Wak75b]. Coding [LR72, RC70, Sti78a]. Coherence [CF78a]. Coherent [FHTS78, LH75, Sri76, Win75]. Cold [Eas78]. Cold-Start [Eas78]. Collection [Tru76]. Collinearity [OC76]. Color [SM72a]. Colored [Jai77a]. Coloring [Ake74]. Column [DK72]. Columns [CW79a]. Combinational [AM77, AA75, BK76, BC71b, BK73, Bos79, BH71a, BCS76, CM77, Cer78, CDO78, CK77, Coy79, DB76, Dia75, DAA79, Dor72, DW73b, ES77, EZS78, FM74, FD74b, FD77b, Fri74, GRR72, GS74, GM71, GT78a, GT78b, Hay71a, Hay76c, Kau70b, KS78b, KS71, KK72, KB75, KK77a, KK77b, KC73, LB74, LD74, Mar79, MC71, NA76a, Ogu75, Pap77a, PM75b, PG75, Red77a, Riv77, Rot70, Roy74, RK71, SM72b, SD74, SM75c, Smi71b, Smi78, Smi79b, SS77, Sto77, SC72b, SC74, To73, Tun70b, TV76, TBY78, Ung77d, Wan75a, Wan75b, Whi71c, YT71c, YY75, ZKK79]. Combinatorial [Abr75, Gol76, Par79, Tom73]. Combinatoric [Jac73]. Combined [IC78]. Comment [Ano70o, Bas73, Bel76, Bro70a, BW70, Car71a, Hen70, Hor71a, Lec70a, Paw71, Ree70, SW70a, Sch79a, She77, Swa79a, Wad71, Whi72, Yue74a]. Comments [Ada78, Agr77, Agr79c, Ale71, Bar78, Bau77, BT70a, Bis73, Bla79, Bla74, Bos79, Bre75, Car71b, CH73b, Cer76, Cha79, Chi72, Col77, CM79c, CM79d, Das71c, DDSM76, Das78, DTG79, Dev79, Du71, DM75, DM76, Eps73b, FHTS78, Flo75, FM72b, Gar74, GS74, Gou77, Gri78b, HO79, Hla78, HL86, Iba74, Kan73a, KS72, Kro74, Kru71, LB74, LD72a, MH78, Maj72b, Maj76, MS73a, Mas78, MMB78, MST^{+}75, Nai79, New75a, NA76a, NRS74, Paa79, RB76, Rao78, Red77a, Ric72, RS73b, Sag77, SM75a, Sch73, Sha75, Shi75, Smi79b, Sur70a, Sur73, Sur78, Tab74, Tou72a, Tou72b, TV72, Tou72c, Tou74, Tou75, Tru71, Van73b, Van77, Wak75a, WY75, WK71, Wel74, Wet78, WZ72, Wil77b, Wor79, YB75, YM78, Yue76, Zoh74]. Committee [Osb77]. Common [Gri78a, JM73, Sch79b]. Communicating [SN79]. Communication [Hei79, LTE77, Mah74, MF76d, Ris74, Sah77]. Communications [Kle76]. Commutative [Kar77a]. Compact [But72, Los78, SKG77]. Compaction [AN74]. Comparative [WM76]. Compared [CS70]. Comparison [Bar75a, CCES74, Che71a, FD75, GG75, GF78, Mil73, MG71b, Sie79, Tou72a, Tou72c]. Comparisons [Dig73]. Compatibility [Bro70a]. Compatible [Rub75b, Tom72, YT76]. Compatibles [Han75, RS72b, RS73b, Sto74b, Sur73, Yan75a]. Compcon [Ano79b, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano77f]. Compensating [Sie75]. Compensation [ASS72, KES72]. Compilation [RPL73]. Compiler [And72, Feu72]. Compiling [Lew74]. Complement [BW73, Bla74, DM75, DM76, Eps73b, GG75, Hwa79, JW76, Kro74, MG75, SS72]. Complementary [CCB70, Kri71a]. Complementation [HO72]. Complemented [Vin78]. Complete [Ano76t, Ano77u, Ano78-42, Ano79-47, Ano79-48, Ano79-49, Ano79-50, DDSM76, Fon71, IS75, Mar71, MST^{+}75, Pu72, Red73a, Rob79a, Tom70, YT71c, Ano78-43]. Completely [LH76, Set70a]. Completeness [KF78]. Completion [Ung77d]. Complex [ARS71b, Bau77, CHSS73, Hig73, YT77, Yue75c]. Complexities [Liu77a]. Complexity [Arn70, AO79, BV79, CJ74, FK78b, Ham71, Han72, Hay78b, Lam79a, Mor70b, Ric74a, SM73b, Smi71b, VD76a, Wak70a]. Component [AO79, CK77]. Components [BF75a, FP75]. Composed [Ish77a]. Composer [BL75]. Compound [Hus74]. Computable [Ehr76]. Computation [Ano73w, Arb76, Ban74c, Boh77, Buz73, Cha73b, Dru71a, Eas78, God71, God73, Goo70, Hen70, Hun72, KV76, Kun75, Lec70a, Lit73, Mey76, Mil73, Muk79, Pra72, SF71, SR73, Spi73, Ulm70, VD76b, WS76]. Computation-Based [Mey76]. Computational [BV79, CY76, CDL78, DW73b, ES77, FK78b, Hel72, Hel74, Mas78, Moo76, Mor70b, PBC74, PD77, SG74, TL73a, VD76a, Wel74]. Computations [CA74, Der73, Erc77, Kri77, Ric74a, Roo73, Vee78]. Compute [Gri78a, Spi71]. Computer [AAK74, AD75, AB76, Agr79c, All75a, AK75, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72-50, Ano72-57, Ano72-58, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n]. Computer [Ano76c, Ano76d, Ano76h, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano76t, Ano76u, Ano77b, Ano77c, Ano77d, Ano77g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77u, Ano78k, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34, Ano78w, Ano78-35, Ano78x, Ano78-36, Ano78-42, Ano78-43, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Ano79-53, Ath75, AG73, AGM^{+}71, BHK79, Bar74b, Bas71a, BS75, Bha75b, Bha78, Bib74, BS77, BO70, Bon74, BE70b, BB74a]. Computer [BR77, BW77, Bre73a, CG74, CS73, Cas73, CM74, CR72a, Cha73b, Chi78, CA77, Col77, CSY75, DRS75, Ekl72, EZS78, Els74, Erc77, EFLR76, FD74a, Flo72, Fly72a, Fly72b, Fre77, Gar76, Gar78, GGMM71, GM72a, GSB^{+}76, Gil74b, GD72, Guf71, Hal74, Ham74, HC70, Hei79, Hob73, HP72, Hun73, HB75b, JP76, KES72, Kar70, KR71a, Kar77a, KC71, Kim74, Kje72, KKA71, Koc71, Kon71, KP71, Kor76, KK70, Kra74, KF73, Kul77, Kuo74, Lam79b, Lar71, Lew77, Lin70, Lom74, Lou70, Lud74, Mai71, Maj72b, Man77, Mar78, MB75, MA77, MJTW75, MR74, MG71a, Mil75, MM70, Mui74, NR77a, NS79, New75a, NS75, NK75, PM78b, Pla73, Ple77, PGS70, RM71, Ram73a, Ram76a, Rao70, Rau79a, RA78, RF78, Ric74a, RC71c, Rip72, Ris74, Ros74a, Ros71]. Computer [RH73, ROW71, SF74, SW70a, SW70b, SK75, Sch73, SF78, Shi70, Sho74, Sta75b, Sto70b, SN71, SC72c, SSR72, Swa74b, Swa90, TI72b, TM70, TK74, Tri77b, Tri78, Ung77a, Van77, Vic74, Voi75, WBRG76, Whi70, WG75, RT73a, DWT70]. Computer-Aided [EZS78, Lar71, Mar78, SN71]. Computer-Communication [Ris74]. Computer-Controlled [Ros71, WBRG76]. Computer-Drawn [Lou70]. Computer-Maintained [Agr79c, Man77]. Computer-Oriented [Bre73a, Gil74b]. Computers [Ano75b, Ano76e, Bar75a, Bas71b, BS78, BH76, Bla79, BCJ^{+}71, Cor71a, DBC73, Des74, ET75, Feu72, FM72a, Gra70, GM70, HG78b, Hec73, HHJ^{+}76, Hou72, Jos72a, Jos72b, Luc72, Mon74, Mor79, Orc76, RC71a, RS71b, Sar71, SGR78, Wex71, Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73a, Ano73m, Ano73n, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. Computerss [Ano70-37]. Computing [Agr78a, Ano73h, Bak79, Bea78, Ber71, Car73, CH75, CM79c, FN75, Har70b, Hay76a, Hob73, Hun71, HH76, Ken74, Kim75, Lei74, LR75, MN70, Met73, MR76, Ram71, Red78b, RT79, Sch74a, SM71b, SK74, Stu75, Tan78, TM78a, Yue75a, Ano72-50, Tab70b]. Concave [VD76b]. Concavities [SCL76]. Concavity [Skl72]. Concentrators [TWB78]. Concept [Arn73, Hop71]. Concepts [UM77, Vai71]. Concerning [Maj76, YM78]. Concurrency [TF73]. Concurrent [RH75a]. Conditional [FG71a, KO73, RF72]. Conditional-Sum [KO73]. Conditions [Hu73, Smi79d]. Cone [Uhr72]. Conference [Ano79c, Ano79d, Ano79-54, Shi70]. Confidentiality [Ple77]. Configuration [BW76a, HB75a, Sch79b, Sti78b, Wel78]. Conflict [Kur74, Nut77]. Conflicts [DI71, FT79]. Congruences [Reu72]. Congruential [Gal78, NM76b]. Connected [ADL72, NS79, RS72a]. Connection [FK78a, HA74, Rub74, Rub76a, Tho78, Tom70]. Connections [Abe72, HO74, KJ79]. Connectives [Set70b]. Consecutive [WG74]. Consensus [CM79d, Mag70]. Consequences [RSL76]. Conservative [SK78a, SK78b, SK79a]. Considerations [DR74, HB75b, Kle70]. Constant [Chu70, MH70a, TL73b]. Constant-Weight [TL73b]. Constrained [BI72, Bro75a, BW70, CK74, Hun73, Ram77a, Whi71c]. Constrained-Input [Bro75a]. Constraint [PMB77]. Constraints [DT70, SN71]. Construct [PR76]. Constructed [MF76b]. Constructing [BF78, PM77b]. Construction [Ban78, Ben70, Feu72, HC78, Oja79, Sto70d]. Constructions [JW76]. Content [LC79, Red73b]. Content-Addressable [LC79]. Contentions [FT79]. Context [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46, Gre75, Har70b, Sel74, Tho75]. Context-Free [Gre75, Har70b, Tho75]. Contextual [Chu71, DT70, Dos77, EK75, RE71, RH74, Tou72b]. Continuation [WK73]. Continued [BB74a, BB74b, Tri77b, Tri78, VV74, RT73a]. Continuity [Tun70a]. Continuous [AOR74, Gou71, JR76, Lew77, SKM78, Tom74]. Continuous-Parameter [SKM78]. Contour [DT70, TD70]. Contour-Traced [DT70, TD70]. Contradiction [Hwa72]. Contributions [Gar76]. Contributor [Ano73f]. Contributors [Ano70p, Ano70q, Ano70r, Ano70s, Ano70t, Ano70u, Ano70v, Ano70w, Ano70x, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r, Ano71s, Ano71t, Ano71u, Ano71v, Ano71w, Ano72-40, Ano72-27, Ano72-28, Ano72-29, Ano72-30, Ano72-31, Ano72-32, Ano72-33, Ano72-34, Ano72-35, Ano72-36, Ano72-37, Ano72-38]. Control [Ast78, BK79, BT70b, BW76b, BA71, Cam72, CHSS73, CSST73, Cov74, DBC73, HG78b, Har74a, JB76b, JM73, Jum74, KP71, Kor76, LS76a, LNL72, LT73, Len78, Les71, LM78, Man72a, MG71a, NM77, PR72, Red78a, RS75, RS78, Svo70, Swa79b, Tab70a, Tab74, Toy71a, WY73]. Controllability [Cer78, Gal78]. Controlled [Ath75, Bre75, CD73, Ros71, WBRG76, WG75]. Controllers [Smi79a]. Controls [Gil70b]. Conventional [Ish77a]. Convergence [GW77, GH72, RF75]. Convergent [MC70]. Converging [NW79]. Conversational [Tei76]. Conversion [Aar70, Ben77, Hor71a, IC78, Maj73, Mat70, Mor75, Nic71, Paw71, Rhy70, San76, Wad71, Yue78, Zoh70, Zoh73d]. Conversions [BR71, Kri71a]. Converter [Mar74, TS75, TP72]. Convertor [WS76]. Convex [Sch78a]. Convolution [DV78b, PH73, Rem74, Zal71]. Convolutions [Rad72, Rad73]. Cooperation [Dos77]. Cooperative [Arb76]. Coordinate [BF78, EO78]. Copy [DLL^{+}73, GS79]. Copyright [Ano79-45, Ano79-46]. CORDIC [Des74]. Core [Mcd72, RB78]. Coreless [Pin71]. Corner [FD77a]. Corner-Finding [FD77a]. Corners [Dun75]. Correct [Cha74a, Kin71a, Max77, Ogu75, Pat76]. Correcting [BM73, BCC78, Dev79, Gri78b, KR71a, Kau70c, LW78, LF77, LF78, NR75, Rao70, TF78, WL74]. Correction [Ano70y, Ano72-39, Ano72-40, Ano73f, Ano78t, Ano78s, BM74, Bei71, BK71a, Cat77, CH72b, En77, HK78, Hla70, LL75, MF76a, Man72b, MJ78, OTK76, Paa79, PB76b, Rad73, Ram77b, Rao72, RH74, Rub76a, SK77a, Saw77, Ser73, Ser75, She78a, Sri71a, Sri71b, Tho75, Tho76, WSB79, YM78, YL73, ZHR77b]. Corrections [MH78, RHR76a, RL78, Tri78]. Correctly [Lam79b]. Correctness [Hus72, RS74]. Correlated [MM73a]. Correlation [Koh72b, RF75]. Correlations [Ken74, Yue75d]. Correlator [Win75]. Correspondence [Hwa72, Rob70a]. Corresponding [San76]. COS [MJ78, MJ77]. COS/MOS [MJ78, MJ77]. Cosine [ANR74, Har76, Sha75, SP77, TM78a, Bak76]. Cost [Avi71, CF73, CY76, CDL78, KAN79, LL79, Mas78, ML76, PA78, SA74, She71, Wak75b, WM71]. Cotes [Hun71]. Count [Hay76c, Hay78a]. Counter [Ano70o, Hol72a, Toy71b]. Counters [Byr71, CM79b, Dor72, HK77, HK78, KO78, MF76b, O’K73, RC70, Swa73a, VV74]. Counting [BO79, Byr71, Dev79, FS71, Kau70c, Myk75, PS78, Red77b, Zoh73a, Zoh73b, Zoh74]. Coupled [Den76, Gou77, PM76a, SG74]. Coupling [MMO74]. Course [Fre74a, Wal74]. Cover [Ano78t, Das71b]. Coverability [Sil79]. Coverage [AM79a, AM79b, Arn73, Sti78b]. Covering [DW72a, Du72, Reu75, Sch75, Tom73]. Covering-Closure [DW72a]. Creases [Huf76]. Creating [Mil75]. Criteria [Che71a, MH70c, Tou72c]. Criterion [Deu71, FE78, FK70b, VV71]. Critical [BH71b, Koh75]. Criticalities [Wan75b]. Cross [MMO74, Wan77]. Cross-Coupling [MMO74]. Cross-Section [Wan77]. Crossing [NS75]. Cube [Pea77, Wyn71]. Cubic [All75b, Wil77b]. Current [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano77v, Cat77]. Currents [Lei74]. Cursive [EK75, Far79]. Curvature [BD75a, Huf76]. Curve [But71, Can71, Dan70, DR77, RT71, RTL72, Sha78b, Skl78, ZHR77a, ZHR77b]. Curved [AB76, Gou71, MA77]. Curves [Dub79, FHTS78, FD77a, JLH73, LH75, MF75, PH74, Ram76c, RJ73, RW75, Rub76b, ZR72, Bel76]. Curvilinear [CA77]. Customized [Lar71]. Cut [Cur79, Das73]. Cut-Sets [Das73]. Cutpoint [Pap72, YO70]. Cybernetics [Rei74b]. Cycle [HO79, Sil77, Smi71a, WR77]. Cycles [Hol72a, Per70, Van73a]. Cyclic [RRNA74, SAPR70, SF77, TL73b, YC76, Yue75d, Zal71]. Cyclic-Order [SF77]. Cyclically [WY73].

D [Hil79, Mar74, Moo76, Mor75, NR76b, Zoh73d]. D-Script [Moo76]. D/A [Mar74]. Damage [Smi75]. Damage- [Smi75]. Data [Anz71, Bas71b, BH76, BT70b, BCL75, CR72d, Chu79, DR75, DB74, Dow73, Dru71a, DB73, Far75, Fen74, FK70b, FO71a, FO71c, GS79, GSB^{+}76, Gol72, HND75, HEWB72, Hig76, How75, HTD72, JM71, Jos74, KR73, Kru71, Law75, Len78, LJ79, LG74, McL72, Mui74, NW79, OF73, Pat72, Ram73a, Ram77b, RNR75, Rum77, Sch78b, Sch74b, Sha78b, She78a, Sto70c, Sto72, SSR72, Tru71, Wan72, Whi72, WL77, FT74]. Data-Storage [Anz71]. Database [BHK79, BH76, De 79, FSLC78, Lan79]. Databases [BHK79, SNOS79]. DBC [BHK79]. DDA [MMB78, HO74, HW76, Max77, MB75]. Deadlock [LL79]. December [Ano76w, Ano78t]. Decentralized [GR72b]. Decimal [Hor71a, LW78, Rhy70, SW71a, Yue77c]. Decimal-to-Binary [Hor71a, Rhy70]. Decimations [AM79c]. Decision [Ake78, BS76, Bit77, CMS79, Chi71, Ehr76, Fri77, KS78a, KKA71, MM73b, PCM70, PS71, PM77b, WL77]. Decision-Directed [PCM70, WL77]. Decisions [EF79]. Declustering [FE78]. Decoder [CDJ71, Wym73]. Decoders [GT78b, Voi77a]. Decoding [FF75, Lin70, SM72a, Sun78]. Decomposability [Kol72, New70]. Decomposable [DF75]. Decomposition [Cur76, Cut70, DF70, FP75, GB71a, HSP78, Hig73, Kan76, Kin70, Kje72, Lan78, Mar70, Paz71, Per70, Ram73b, RS72b, RS73b, Sch78a, SM70a, SMW71, Sur73, Vic71, Wil75, Yoe70a]. Decompositions [Paz70b, Rud76, Tur70]. Deconvolution [Dep70]. Decrypting [Rub79]. Deductive [Arm72, BCSW76, CCES74, MC78, Rei76]. Defects [Sri71a]. Defined [Dav77a, Smi71b]. Definite [Brz70, Pre70]. Definition [But72, Car71b, Dav72a, Ern76, KS72, LM71b, O’C75, Pag70b, Van73b]. Definition-Driven [Ern76]. Definitions [DM75, DM76, Eps73b, SS72]. Degraded [Jai77a]. Degrading [BF75b]. Degree [AN74]. Degrees [AC71]. Delay [Ake70, Bre74, BW70, Dav70b, Hla70, PM71, Sch79b, Ser73, She70a, Vil70]. Delay-Free [BW70]. Delayed [LV75, MBM77, Sin71, Sur78]. Delayed-Input [Sin71]. Delays [Bat70, Bre74, BW70, SS71]. Delegates [Ano72n]. Deletion [Gel76]. Deletion-Directed [Gel76]. Delta [Lee78a]. Demand [Yan74a]. Denominator [Gri78a]. Density [AB79b, Dui76, FS74, SM71b, Woo79]. Dependence [BCL75, Das71a, HG78a, TV72, Vil72, Vil73b]. Dependent [CB72, Hus74, RC71b]. Depth [PMB77]. Derivation [Ben71, Bet71, Dan73, DDSM76, Hil79, Mar71, MST^{+}75]. Derivatives [Tha78]. Deriving [BG79, Cur79, Cur77, YT76]. Describe [Uhr72]. Describing [Bar75a, SN72a, Zah71]. Description [AB76, BEGW72, Ber77, GL74, JK77, SF78, Sta70b]. Descriptions [Moo76]. Descriptor [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46]. Descriptor-in-Context [Ano70z, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46]. Descriptors [Tan78, ZR72]. Design [AM73, And74, Arm73, AM79c, Ast78, Atk70, Aus70, AT70, Avi71, AGM^{+}71, Bat70, BC71b, BT74, Ber70b, Bha77, BT76, CJW^{+}71, CR72b, CDS73, Cha74c, CD78, CS79, CF70a, CSST73, Cor71a, CLL78, Coy79, DR74, Dav77a, DAA79, EZS78, Elm76, Els74, FM72b, FK74, GHJ77, GD72, Gon70, Gra70, GT78b, HSS78, HP75, Hol79, KH74, Kam79, Kin71b, Kin72, LY71, Lar71, Lec70b, LD72b, Lem70, Lud74, Maj76, MA71, MH70c, MF78, MPS78, McC79, MJ77, MJ78, MI72, Nay71, New76, NM77, O’K73, PR74b, Pra78a, Pre71a, RTW78, Red72a, RSL76, Saw75, SM71a, SF72, Sho74, SY75, SG71, SM73a, Sie74, Sto74a, Sto77, Su75, ST76, TG78, TM70, Toh74, TW77, TK74, TS79, Toy71a, Toy71b, TG79, Usa75, VV74, Voi74, YC76, KD79]. design [RT73a]. Designed [SM74, Wak75a]. Designer [FS70]. Designing [BF75a, Far73, Hu75, KS71]. Designs [Abd71, IS76]. Detect [Dun75]. Detecting [BDR77, DS71, Gad70, GL70, MF75, SR75, Zah71]. Detection [AB79a, BM74, BK73, Bos79, BH71a, BHP^{+}71, Bre71d, DF75, DB76, DR75, DV78a, DW73b, Ehr77, FM74, Far73, FD75, FC77, GRR72, GS74, Gon70, GT78a, Gri73, HF74, Hay75, HC71b, HPP77, IS75, Ios78, KP74, Kar77a, KS71, KK72, KR78b, LL75, LB74, Lil72, MF71, NR77a, OS76, PA78, Pra78d, PG75, RH75a, Ram76b, RM78, RT71, RTL72, RJ73, RW75, Ros75, SM74, Sha78b, SV74, SCL76, Skl78, Smi79c, SS77, Su73, SKM78, TF70, Toy71a, Toy71b, TM78b, TBY78, Usa75, Wak75a, Wak75b, WR77, YY75, Zis76, ZKK79]. Detector [DF78]. Detectors [RJ72, Sur70a]. Determination [Bro70a, Bub72, Cho76, DBC71, IR76, Maj72c, MT74a, MBM77, Mor70c, NRS74, Sar71, Sik70, Sto74b]. Determine [Set70b]. Determining [BK72, Bor72, HEWB72, Han75, PC72, RS72b, RS73b, SV75, SO79b, Sur73, SB74a, Tom70, Tom72]. Deterministic [Dav72b, Kri73, Moo71, San76, SCK73b]. Developing [Ben77]. Development [Pat76]. Developments [Cox72]. Devices [Bak79, Bar74a, DS72a, Ern72, Fai73, Flo75, HH76, Kar77a, LC79, PM76a, Stu75]. Diagnosability [AKT75, BGM76, Hay74, RK75a, RK75b]. Diagnosable [CS79, FM72b, FK74, FK78a, FK78c, HA74, Jos78, LY71, MH76, MH78, MM78a, Pai73, SM71a]. Diagnosing [Mar79]. Diagnosis [Ake74, Cha70, CRB72, Fri74, FK78b, GST79, Hay71a, Kam75, KF79, Kim70, Kim79, KB75, KK77a, KK77b, MH76, MH78, MS75a, MM78b, MKO70, RM71, RC71a, Rot70, Roy74, RK71, RK75a, RK75b, SK77c, Smi79e, ST76, Toi76, YO70]. Diagnostic [BK72, MK79, Rob70b, Sel74]. Diagnostics [HF71]. Diagram [MS73b, MA71, RRNA74, WR77]. Diagrams [Ake78]. Dialogues [Els74]. Diameter [TS79]. Dielectric [YRT75]. Difference [CRB72, HC71b, KM75b, Wes78]. Differences [DDSM76, Mar71, MST^{+}75]. Different [DM75, DM76, Eps73b, Ges70, LHSM74, SS72]. Differential [All75b, Bak79, BW76a, BM77b, DEA73, Fra78, Kar70, MB75, MN70, New75b, RS71b, SF74, TD73, Uhr73, Wil77b, Wit70, Wor76]. Diffraction [WKP75]. Digit [Tun70b]. Digital [Aar70, Agr75b, Ake70, AP76b, AP76a, ACJ75, Ano78v, Ano78w, Ano78x, Ath75, Avi71, Bar75a, BS72, BGM76, BV78, BCA^{+}75, BB74a, BdB70, Bre70b, Bre73c, But79, Cas73, CS75, CR78, CDO78, Cha74c, DBC73, DTG79, Dav71, DEA73, DRS75, EH78, Erc77, Fai73, Feu72, Fre79a, GSB^{+}76, GSJ^{+}76, Gil70b, GM70, HND75, HLD70, HO79, HC70, Hun72, Hun73, HB75b, Hun77, Jai77a, JA78, KR71a, Kar77a, KF79, Kim74, Kim70, Kim79, KV76, LL75, LC72, Lit74, Maj72b, Mar73c, Mat71c, MMO74, MN70, MK79, Mon74, Mor76a, ND73, NA74, PGS70, RHR76a, RHR76b, Red78a, RC71c, RS71b, Rob70a, RJ73, Ros74b, RW75, RH71, SF74, SAPR70, Sar71, SM72b, SM71b, SM76, SC77, SG71, Sie74, Sil77, SP72a, Sta75b, SP77, SM70b, Stu75, Su73, Swa74b]. Digital [ST76, TL73a, TP72, TW77, Tri77b, Tri78, TE76, Var79, Voi74, Wai70, Wex71, Whi70, Yue77d, Zoh73a, Zoh73c, Zoh76a, NA76b, MB75]. Digital-to-Analog [Aar70, TP72]. Digital/Optical [Cas73]. Digitally [CB72]. Digitized [Ami77, AN74, SCH72]. Digits [Sok72, Sut74]. Digrams [RE71]. Digraphs [DS73, PC78]. Dimensional [Ano72-39, BDR77, CC73, DR75, DM78b, FO71c, Goo70, HND75, IH72, Jai77c, Kun79, MS78, NA74, NW79, Par71, Roe78, SR74a, Sar71, SM72a, UM77, Vil73a, Wei75, NA76b]. Dimensionality [CA74, Ehr76, FO71a, JA74, SV75, Tru71, Tru76]. Dimensions [Gra71]. Direct [AK75, BH71c, FHTS78, LH75, Les71, LR75, Sho74, SB77b, Whi71b, De 79]. Directed [CR72d, Gel76, PCM70, Tho75, WL77]. Directly [Bae74, Hoe74]. Discrete [AP76b, BLZ75, Boo70, CF79, DF70, Dot70, DV78b, Duk74, Kun79, Law70, Rad72, Rad73, RT79, Rei74a, Sha75, Sto74a, TL73a, TD77, Tou72d, Van75, VR79, WL77, ANR74, Har76, Hun71, TM78a]. Discrete-Time [DF70, Dot70]. Discrete-Valued [Tou72d]. Discretization [HO74]. Discriminant [Cha73a, FS75, Kit77, Kit78, Sam70b, VD76b]. Discrimination [Vil72, ZRD75]. Discussion [Han75, Hel74]. Disease [SH72]. Disjoint [Eps73a]. Disjunctive [But75, Hig73, SM70a, SMW71]. Disk [FS74]. Dispatching [FR72]. Displacement [Ano77v, Cat77]. Display [FO71c, HSS78, HB75b, JR76, KF73, LH79, LG74, New76]. Displaying [JR76]. Dissent [HP73]. Distance [Dev79, Kau70c, KF72a, LR76, TL73b, Tou72e, VV71, ZT75]. Distance- [TL73b]. Distinct [Ban78, RN77]. Distinguishing [Bou74a]. Distributed [Coo77, RSH79, Svo70]. Distribution [ABK76, BT72, JW73, Obe72, Rom78, SP72b, SP72c, Wei75]. Distribution-Free [ABK76, BT72]. Distributions [BD75b, Hal74]. Diversified [RS75]. Diversity [WL75]. Divide [Cha74a, She71]. Divide-and-Correct [Cha74a]. Divided [Wes78]. Divider [Ste72]. Dividers [Boy74]. Division [AKT^{+}77, BJ75, BKM73, CH73a, Cha74a, Chi72, Fly70, Gar74, Jac73, KKK73, Kri70a, Kri70b, Kri71b, Lig77, Rhy71, Rob70a, SCK73b, SR72a, Sit74, TE77, Tun70b]. Divisors [Jac73]. Does [Ano75c, Ano75d, Ano75e, Ano75f]. Doesn’t [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50]. Domain [KH74]. Domains [Gar72a]. DON’T [Vin78]. Dotted [FM72a]. Double [Cur73b, GH78, HC70, MF76a, Myk75, PB76b, Van73a]. Double-Precision [HC70]. Double-Rank [Cur73b, GH78]. Down [Har74b]. Drawing [RL72]. Drawings [EUY^{+}72]. Drawn [Lou70]. Driven [Ern76]. Drivers [Wym73]. Drum [Ful72]. Dual [FS74]. Dual-Density [FS74]. Dualize [HPP79]. Due [CB72]. Dyadic [GC77, Pea75, RRNA74]. Dynamic [AU74, BBLT73, BE70a, BH74b, CJW^{+}71, CK79, Cod73, CF70a, DI71, GSB^{+}76, Hla78, KK79a, LT73, Len77, MVW79, RA78, RS78, RH73, Si78, Sto72, Sto75, Tab74, Uhr73, Vic74]. Dynamical [HN78]. Dynamically [BR76b, OW73b]. Dynostat [GM72a].

E-Nets [NN73]. E2 [Sil79]. Easier [Ano75c, Ano75d, Ano75e, Ano75f]. Easily [Coy79, Fri73, FNSK75, Jos78, Kod74, Red72b, RW74, RH75b, SR74a]. Easy [Lio76]. Economic [SK71]. Economical [Kri71b]. ed [Jos72a, Luc72]. Edge [DF78, FD75, Gri73, Nie75, PR78, RT71, RTL72, SR71]. Edge-Sensitive [SR71]. Edges [ADL72, DR75, Dun75, Ehr77, OC76, Sto77]. Editing [Wil70]. Editor [Ano70-27, Ano71x, Ano71y, Ano71-49, Ano72-47, Ano72-48, Ano72-49, Ano73g, Ano74g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76f, Ano77i, Ano77j, Ano77h, Ano78u, Ano79-29, Ano79-30, Ano79-31, Ano79-32, Shi75]. Editor-in-Chief [Ano71-49]. Editorial [Ano73e]. Editors [Ano76h]. Eds [Tab70b]. Education [Ram76a]. Effect [Arn73, BH71a, BCHL73, Gra70, LB74, Mar77]. Effective [CKL77, JA78]. Effectiveness [Avi71, Fly72b]. Effects [Bre74, FLBK76, Hun71, Sie77]. Effects-of [FLBK76]. Efficiency [Bit77, FLR78, Kan75, KR71a, Los78]. Efficient [DH79, EO78, God73, Gri78a, Har76, Hur74, KS73, Los76, MF78, MM78b, NTA78, NS75, PGS70, PB76a, PM76b, RC71b, Sha78a, STG71, Sil79, SO79b, Weg72, YT71c, Bak75a]. EFL [Elm76]. Eigenvalue [Has74, The75]. Eigenvectors [BC71a]. Eight [MTB70]. Ek [Hil79]. Ek/D/1 [Hil79]. Eklundh [Ram75, TE76]. Electrocardiogram [Ges71]. Electrochemical [Nel74]. Electronic [CDS73, CHSS73, KES72]. Element [AT70, HK77, HK78, KT72, Nel74, Tak76]. Elementary [Bak75b, Erc73, Mar72, Meo75, Oja79, Bak75a]. Elements [Ama72, Dav70b, DVS74, GC73, Gra73, Gri74, Hen71, Hla70, HS75, Ish77a, JK77, Lyo70, OY70, SK78b, SK79a, Ser73, She70b, SR72b, Win70a]. Eliminate [BH71b]. Eliminating [BH74b, Hla78]. Elimination [HD71, Jol76, Kan73b]. Ellipses [TM78b]. Emerging [FLBK76]. Emitter [ET75]. Emitter-Function [ET75]. Empirical [Gri76]. Employing [HA76, NA76b]. Emulation [FLBK76, Hoe74, Les71]. Emulator [SB71]. Encoding [KM73, KKA71, SM72a, STG71, SB72]. Encryption [Ple77, KD79]. End [She77]. End-Around-Carry [She77]. Engineering [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Boe76, Jor74b, Lom74, Ram76a, Ano78-39, Ano78-40, Ano78-41]. English [Hub76]. Enhance [FR72]. Enhanced [Sto72]. Enhancement [Hal74, HB75a, KCLB71, KS77a, ZHR77a, ZHR77b]. Enhancing [WA73]. Ensemble [WL75]. Entirely [MF76b]. Entropy [CF73, Mas78, WD77]. Entry [Ano73f]. Enumeration [AA72b, MTB70]. Enviromnent [Sch79b]. Environment [AI78, BD75a, BCHL73, KM78, Tun70a]. Environments [LD75, Li77, NC78, Nev76]. Equal [KP75, Per70]. Equation [BM77b, BLZ75, Kar70, MR75, Sch73]. Equational [Bro74, Bro75b, Cer76, Eps73a]. Equations [AH76, Bak79, BB74b, Bro70b, Bro71, CW77, Den76, Fra78, Gou77, Has73, Hwa72, Kar77a, KS73, MS73a, Max77, MN70, MC70, New75b, Pic74, RS71b, Rud74, SF74, SG74, Sau72a, She70c, Sur70b, Uhr73, Ulr71, Weh74, Wit70, Wor76]. Equivalence [Bau72, Kun79, Mat71a, MC71, Moo71, MS70, NSR74, RS74, Roy74, Yan75a]. Equivalent [Bar71, BT72]. Equivalents [Gaj75, GB71b]. Erasure [Sun78, WSB79]. Erlang [MH70a]. Errico [Ano73z]. Error [Avi71, Avi73, BM73, BM74, BCC78, CB72, CH72b, Chu71, Chu74, DTG79, Dev79, DTI78, En77, FK71, Gad70, GY77, Gri78b, HO74, KR71a, Kar77a, Kau70c, LW78, LF77, LF78, MF76a, Man72b, Man72a, MM73a, Nad71, NR75, NA74, NA76b, Paa79, PA78, Pea76, PR72, PB76b, RH75a, Rao70, Rao72, Red78a, RH74, SM76, She78a, Sri71a, Sri71b, SR77, Sun78, TF78, Tou71, Tou72b, Toy71a, Toy71b, WSB79, WL74, YL73, Yue75a]. Error-Coded [Avi73]. Error-Control [PR72]. Error-Correcting [Gri78b, KR71a, LW78, LF77, LF78, NR75, Rao70, TF78]. Error-Correction [MF76a, PB76b]. Error-Detecting [Gad70]. Errors [HC71a, KES72, KP75, Lit73, PA78, Sri71b, Tho75, Usa75, Wak75b]. Essential [Hla70, Maj72c]. Estimates [MR75]. Estimating [BD75b]. Estimation [AOR74, Boo70, BR74, Dav72b, DM78b, FK71, GY77, Hun73, KS77a, Maj72b, MS76, NJ77, PCM70, TL73a, Tru76, Whi70]. Estimators [Dui76]. Evaluating [BW77, FV79]. Evaluation [AS74, BB74a, BG78, BKM73, Erc73, Erc77, FD75, Gay78, Gel73b, Kim74, Koh75, Kor79, Mar72, Mar73d, Moh71, Nec79, PD79, PM76b, SB77b, Tou72e, Vil73b, Win76, Yue74b]. Events [Paz70b, Paz70c, Paz70a, Rey70, Whi71a, Yam70]. Every [LF74]. Exact [Kri77]. Exactly [RW74]. Examination [CSJ74, KAN79]. Example [MR74, REFN76]. Examples [Fly72a]. Exchange [Ano72n, Che73, Lan76, LS76a]. Excitation [Bre75, CD73]. Exclusive [Fis74, MS70, RR74]. Exclusive-or [RR74]. Executable [Bae74, KMC72]. Executed [Hoe74]. Executes [Lam79b]. Execution [Cha72, FR72, GR72b, LM78, Tho74, TF70]. Exhibiting [Smi78]. Existence [FK78c, RS72a, Smi78]. EXOR [Pra78d]. Expandable [Kau73]. Expanded [Ano72-50, Ano73h]. Expanding [Jos72b, RB72]. Expansion [FK70a, Ges71, God71, Gri78a, KT72, NFS77, SO79a, TM76, VV74]. Expansions [AP76a, Swa72a]. Experience [MM70]. Experimental [CM76, CW75, Dut74, Smi79a, WG75]. Experiments [AB79b, Bou74a, Bou74b, BG79, DF75, EK75, Far73, FT76, Gon70, Hol72a, Hsi71b, MOW76, RTL72, TB75, WKP75, WY75]. Explanation [Obe74]. Exploration [Gri76, SF72]. exploratory [FT74]. Exposure [RK75b]. Expression [Hwa74]. Expressions [BEGO71, BKM73, GLR71, PM78a, Pat71, PC78, PM76b, RPL73, SB77b]. Expressive [Set70b]. Extended [MN70, Mcd72, Roc70]. Extending [Ano79b]. Extension [LG74, OW73a, O’K75, SD74, Sli70, TE76]. Extensions [She70a]. External [Dru71a, Pat73]. Externally [DS74]. Extra [FK74, FNSK75]. Extract [Ges71].

Extraction [ADL72, Bab72, Cal70, DW72a, FE78, GP74, HKD^{+}71, Hey71, KF72a, Koz71, MK71a, Maj72a, Mar78, Pre71b, The75, Tjo77, TG72b, Tou74]. Extractors [GY77, You71].

Fabricated [Ano72-39, Par71]. Face [SF77]. Facility [MCRD72]. Factor [FLR78]. Factoring [HW75]. Factorization [Dru71b, Jai77b]. Factors [PB76a, RK71, Yue73a]. Fail [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TM71, TI72a, TOS71, Toh74]. Fail-Safe [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TOS71, Toh74]. Failsafe [TKH71]. Failure [Koc71, LR72, RC70, TA71]. Failure-Tolerant [LR72, RC70, TA71]. Failures [BCC78, CLL78, Rot70, Sie75, SS77, Wak75c]. Fall [Ano79b, Ano77f]. Fan [Ast78, CK74, FM74, KS78b, NA72, PMB77, SN71, Wan77]. Fan-Beam [Wan77]. Fan-In [CK74, PMB77, SN71]. Fan-Out [Ast78, FM74, KS78b, NA72, SN71]. Fanout [BB78c, BK73, Mar79, SK79b]. Fanout-Free [BB78c, BK73, Mar79, SK79b]. Fast [ARS71b, BS72, BF78, BM70, BB74a, Bro77, Buz73, Car71a, Cor77, Des79, DV78b, Ekl72, FA76, FC77, GM72a, GLMR71, GD73, HW70, Hab77, Hen70, Jai77a, Lec70a, Len77, Lig77, Maj72b, Nec79, NW79, Nus78, O’K75, PBC74, PD77, RTW78, RT79, RNMP77, Sar78, Sch73, SMW71, Sto75, Sur75, Whi70, Zal71, Zoh73b, Zoh74, Bak76, Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, Gla70, Goo71, GW70, Kar77b, MS73b, PB76a, SR77, WK71]. Fast-Converging [NW79]. Fast-Division [Lig77]. Fault [AB79a, AM77, Ake74, Ake76, AGM^{+}71, Avi76, BC71b, BCA^{+}75, BK73, Bos79, BH71a, BCJ^{+}71, Bre71d, Bre72a, Car73, CM76, Cer78, CCES74, CRB72, DF75, DB76, DTG79, Dia75, DV78a, DW73b, FM74, Far73, FM72a, FD74b, FD77b, GRR72, GS74, Gon70, Gra73, GM76, GT78a, GCP^{+}77, GF78, Hay71a, HF74, Hay76a, Hec73, Hop71, HS75, HC71b, HPP77, IS75, Ios75, Kim75, Kim79, KS71, KB75, KK77b, KM75b, LB74, MH76, MH78, MS74, Mat71c, MC71, MF71, MC78, Mey71, MR76, MM78b, MKO70, NA76a, Ogu74, OW73b, OH79, PR73, PR74b, Pra74, Pra78a, Pra78b, Pra78d, PG75, Ram71, Red72a, Red77a, Red78b, RM78, Roy74, RK71, RK75a, RK75b, SR75, SM74, SM75a, SM71a, Sch74a, SM76, SV74, SG71, Smi75, SM78, Smi79d, Sri78, Sti78b, SC72b, Su73]. Fault [To73, TBY78, Wak75a, Wei72, Whi71c, YO70, YY75]. Fault-Detection [DF75, Far73, HPP77, KS71]. Fault-Locatable [Red72a]. Fault-Secure [Pra78a]. Fault-Tolerance [Ogu74]. Fault-Tolerant [AGM^{+}71, Avi76, BCJ^{+}71, Car73, CM76, FM72a, GM76, GCP^{+}77, Hay76a, Hec73, Hop71, HS75, Ios75, Kim75, MS74, Mat71c, MR76, OW73b, PR73, PR74b, Pra74, Pra78b, Ram71, Red78b, SM75a, Sch74a, SG71, Smi75, Sti78b]. Faults [Arm72, BHP^{+}71, Bou74b, Bre73c, Bre74, BCS76, CSJ74, CLL78, Fri74, GM71, GST79, Hay75, Hu75, Ios78, KP74, Kam75, Kau74b, KH77c, KK72, KK77a, KS79, KR78b, MM78a, Mar79, Mei74, MS75a, Nai79, PM75a, SM72b, SK77c, Smi79c, SKM78, Sun79, Var79, Wan75b]. Faulty [Cha70, Cle73, Dir72]. FDP [GLMR71]. Feasibility [Far75]. Feature [ARS71a, And71, Cal70, Che71a, Chi71, Das71a, Deu71, FE78, FP75, FK70a, GY77, GP74, HKD^{+}71, Hey71, KCLB71, Kit77, KKA71, KF72a, Koz71, LD75, Moh71, NF77, OC76, Pre71b, RJ72, The75, Tjo77, TG72b, TV72, Tou72c, Tou72e, Vil73b, WL75, You71]. Features [Bab72, Bec71, BB78a, Chi77, Ges71, Ram76b, STG71, SNEL79, Tou72d, Tou74].

Featuring [CLL78]. Feed [Liu77a, Liu77b]. Feed-Forward [Liu77a, Liu77b]. Feedback [Brz71, Fri70, HC72, LNL72, Lem70, LF74, Mar73a, Mar73b, MAB78, Mor75, Per70, Riv77, Zei71]. Feedback-Control [LNL72]. Feedforward [Ols70]. Ferroelectric [Kau73]. FET [Pai73]. Fewer [Bar71]. FFT [HL86, But79, Dru71b, SCNS83, TJM79]. Fibers [PS78]. Field [Red79, RTKH77, Sar71]. Figural [VV71]. Figure [Hec73]. File [Ste77, WG74]. Files [Anz71, CW79a, RC71b]. Filling [But71, Fre77]. Filter [MS78, Zoh73a, Zoh76a]. Filtered [CH72a, CH73b]. Filtering [BV78, But79, Dav71, Far75, Gol72, Hab77, HLD70, Hun72, Jai77c, NA74, Pra72, SM72a, Yue77d]. Filters [HC70, KH74, LL75, Lit74, SF72, Zoh73c, NA76b]. Final [de 71]. Find [Du72]. Finding [Cha74b, Das73, FD77a, FBS75, FO71a, OC76, PHG77, RB78, Tru71, Whi71a, Yan75a]. Fingerprint [MJTW75, MF76c, RB78]. Finite [BV78, BF72, Cai71, Cha72, CL75, DV78b, Fre79b, Gaj75, Gil70a, Goo70, HN78, KYO70, KW70, Kar77a, Kar77b, Kri73, LF74, MM73a, MB77, Moo71, Paz70a, Pfl73, Pu72, Red79, RTKH77, RS72a, Ric72, Vai72, Vil70, Yan74a, Yan75b, YM78, Yeh70]. Finite-Memory [KYO70]. Finite-Precision [MM73a]. Finite-State [BF72, Cha72, CL75, Gil70a, Goo70, Kri73, LF74, MB77, Pfl73, Ric72, Vai72, Yeh70]. FIR [MS78]. Firing [Wak70c]. Firmware [Pag72, Pat76]. First [She70a, Weg76, Shi70]. Fit [Lio76]. Fitting [Can71]. Fixed [Jac73, LK75, SR77]. Fixed-Increment [LK75]. Fixed-Integer [Jac73]. Fixed-Point [SR77]. Flexibility [MH75]. Flexible [CK74, KKA71, MH73, Obe72]. Flight [FD74a]. Flip [BH73, Bre75, Bro71, Brz71, CD73, Cur70, Cur71b, EW77, FD79, ISN76, MF76b, NR76b, Rub79, Sin77, SR71, Sur70b, Toh74, TB74]. Flip-Flop [Bro71, Cur70, Cur71b, EW77, NR76b, Sin77, Sur70b]. Flip-Flops [BH73, Bre75, Brz71, CD73, FD79, ISN76, MF76b, Rub79, SR71, Toh74, TB74]. Flipping [Chi77]. Floating [Bau77, Boh77, Bre73b, Cod73, HW76, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, Pak71, Yoh73, Yue75c, NA76b]. Floating-Point [Boh77, Bre73b, Cod73, HW76, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, Pak71, Yoh73, NA76b]. Flop [Bro71, Cur70, Cur71b, EW77, NR76b, Sin77, Sur70b]. Flops [BH73, Bre75, Brz71, CD73, FD79, ISN76, MF76b, Rub79, SR71, Toh74, TB74]. Flow [CCF79, Rej72, Rum77]. Fluctuating [BG75]. Focus [LE79, Swa79a, LE77]. Folding [Lee78a, NA76a, To73]. Following [SB72]. Forests [KLCC77]. Foreword [Ano76g, Ano77g]. Forgeries [NR77a]. Form [AM79a, Ano79-45, Ano79-46, Kun75, MT74a, Max77, Mor70c, Ulm70]. Formal [TF78]. Formalism [GL74]. Formalization [Mat70]. Format [Anz71, Hig76]. Formed [RF75]. Forms [CM79c, HG78a, Hen78, KS75, KS77b, Lam79a, PP75, Sri75]. Formula [SA74]. Formulas [Ben70, Pin71]. Formulating [Nie75]. FORTRAN [BCKT79, KMC72, Roc70, Swa72b, Fre74a, LG74, Rej72]. FORTRAN-Like [BCKT79, KMC72]. Forward [Liu77a, Liu77b]. Foundation [Fu74, Kul77]. Four [CYNM79, CM79b, Dir72, NK75]. Four-Phase [Dir72]. Four-State [NK75]. Four-Valued [CM79b]. Four-Variable [CYNM79]. Fourier [Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, FHTS78, Gla70, Goo71, GW75, GW70, Hen70, Hun71, Kar77b, LH75, Lec70a, PB76a, SR77, Zoh74, CF79, Des74, Des79, DV78b, Gra72, GW75, Kun79, RT79, Rej72, Sar78, Tan78, Yue75a, ZR72, Zoh73b].

Fourth [Gag72]. Fraction [VV74]. Fractions [BB74a, BB74b, God71, Tri77b, Tri78, RT73a]. Frame [SH77]. Free [ABK76, BT72, BB78c, BK73, BW70, FM74, Gre75, Har70b, HK77, HK78, Mar79, SK79b, Tho75, WN76]. Freedom [AC71]. Freehand [NR77a]. Frequency [FV79, GF78, KH74, New70, RC71b, Sok72, TS75]. Frequency-Dependent [RC71b]. Frequently [Len78]. Frills [JW76]. Front [Ano78t]. Full [LHSM74]. Fully [Mor79]. Function [AB78, ASS72, Bar70, Bor72, BT76, CCB70, DW73a, Edw75, EH78, ET75, GC71b, HD71, HG78a, HPP79, Han72, Hen71, HA76, KAN79, LH76, Lee78a, Lyn72, Mar73b, ML76, OY70, Pap77b, RB72, Van75, YT71b, Yue75e, Yue76, ZT75]. Function-Unit [BT76]. Functional [AM79a, AK75, Ber70b, Bre70b, Fly70, Gar71, Kar70, KK70, LS71, MC78, Pav73, Rud76, SF71, VW72]. Functions [AA72b, AA71, And70, Bak75b, Bau71, Bau72, BCSM72, BB78c, BR76a, BR78, Bis70, Bis71, Bis73, Bit77, Boh77, BB74a, BG78, BV79, Bro75b, Bro75c, But72, But78a, But75, Can71, Car71b, CR78, Cha73a, CL77, CM73b, CYNM79, Cur76, CM79c, CM79d, Dah73, Das71b, DS71, DK72, Dav72a, DS75, Dui76, Eks73, Erc73, Erc77, Fen74, FD73, FB74, GBMC72, GC73, God71, GM76, HF72, Har70a, Hay71b, Hig73, HE79, HC73, Hol72b, HO72, HA77, Iba74, Kan73a, Kan73c, Kan76, KT72, KS72, Kit78, Kod74, KS77b, KS78c, Kol72, KN70, LM71b, LL72, Lew77, Mag70, Mag73, Mar72, Mat71b, MS77a, MTB70, Nec79, Orc76, OW72, Pav75b, Pin71, PD79, PH73, PP75, Pra78c, PM71, Pre71a, Red72b, Red73a, RW74, Red73b, Red71b, RS71a, RNMP77, RH75b, Rud76]. Functions [SA70, SB77a, SK79b, Sch75, Set70a, SM70a, SMW71, Sin72, SC72a, Smi71b, Spi71, SP77, SC72c, Sur74, Sur75, TG72a, TD77, Tom74, VD76b, Van73b, Vin78, VW72, Wes78, Win70a, Wri73, YT71a, Yeh70, Yue75d, Bak75a, Spi73]. Fundamental [BR78]. Fundamentals [Swa74b]. Further [Cur71a, RB76, RTL72]. Future [Cox72, TP73]. Fuzzy [BD75b, Cha72, GL70, Kan73a, Kan73c, Kan76, Mam77, Paz70a, SC72a, Sri75, Su70].

G101 [PRD71]. Galois

[BR76a, BR78, Pra78c, Wes78]. Gap [Nel74]. Garbled [OTK76]. Gate [BY79, CC72, HK77, HK78, Iba71, Ken77b, NTK72, Nak79, Obe76, RM71, RH75b, Rot70]. Gate-Interconnection [Iba71]. Gate-Minimum [Nak79]. Gates [BCSM72, CYNM79, Ham73, Hu75, Iba71, IM71, LM79, LHSM74, NT70, TG79]. Gaussian [Mar73c]. General [Bar70, Brz70, CH72a, CH73b, Dru71a, Erc77, GR72a, HP72, Hoo77, HA76, KKK73, KS73, Mur70, PM75b, Rey70, Smi71a, Sri76, Win70a, Zis76]. Generalization [CM79d, FC77, Gla70, Ram75]. Generalized [AA77, AC70, CL79, CF79, FS78, Hsi71a, KSA74, ML76, Pra72, SK74, Swa72a, Tho78, VD76a, Wak70c]. Generate [Pap77b]. Generated [CL71, Rao78]. Generating [Agr75b, AM79c, Ber70a, CL72, Das71c, DK72, Fre79a, Hwa74, MAB78, MS77b, Mur70, Myk75, SCL70, Sok72, YT71c]. Generation [AA72a, AA75, Ake76, Bau72, BCC78, BH71b, BH71c, Bre71d, Bre72a, Bre74, BH74b, CSJ74, CDO78, Che78, Dan70, Dav72a, FP75, HLD70, HO79, Hay78a, Hla78, How75, Hub76, Hur74, JLH73, KN70, Lee78b, LE71, Lew77, MB75, Mut76, OH79, Pap77a, Pat71, Qua74, RJ72, Reu75, Ros71, Rub76b, Sil77, Smi74, SKK79, Ung77d, Van73b, Wan75a, WC74, WC76, Yan74b, Bel76, Gag72]. Generator [And72, Arm73, HPP77, MH70b, MH70c, Mar73c, PW79a, SLC75, Bak76]. Generators [NM76b, NM76a, Yue77e]. Geometric [Ban74a, BO79, Mor75, Ram77b]. Geometry [Fer77, Wan77]. Gestalt [Zah71, ZT75]. Getting [Ano74-28, Ano75-29, Ano75-30]. Global [BD71, Ehr77, Hwa79, MN78]. GNAF [LR75]. Goal [LS76b]. Goes [Ano74b, Ano74c, Ano74e, Ano74f, Ano74d]. Golay [Pre71b, PC72]. Golomb [Rob79b]. Good [Hur74, She70a]. Gracefully [BF75b]. Gradients [Ros75]. Grammars [MB77, Pat71, Tho76]. Grammatical [Gai79]. Grams [RH74]. Graph [Ake74, Das73, Hay76a, KNF76, Law70, Lem70, MR74, Rej72, Rub75a, Sau72b, Yan75b, YM78, Zah71]. Graph-Theoretic [KNF76, Law70]. Graph-Theoretical [Zah71]. Graphic [Arm73, New76, PRD71, SA70]. Graphical [DLL^{+}73, Fon71, MA71]. Graphics [Bib74, KC71, LH79, MB75, New75a, SAFV71]. Graphs [BEGO71, Dav77a, Har70a, HN71, LR71a, Reu72, ROW71, Sic76, Sto70d, Tur70, Vil70]. Gray [Ano70o, Mor74, PR78, Yue78]. Gray-Scale [Mor74]. Great [WBRG76]. Greatest [Gri78a]. Grey [Ami77]. Grey-Level [Ami77]. Grid [Fre79a, LR76]. Group [Dun75, Kol72, Lev70]. Grouping [FK70b, ZT75]. Groups [Cai71, Kar77a, Kar77b, YC76]. Guidance [LS76b]. Guide [Har71]. Guided [Rei76]. Guidelines [Ano76h].

Haar [RNR75, Rej72]. Hadamard [Car71a, MS73b, WK71, AK72, CL77, FA76, GC77, Hen78, Kun75, Kun79, RNR75, RRNA74, SO79b, Ulm70]. Halftone [KC71]. Halftones [Ros71]. Halting [Nie70b]. Hamilton [Vil73a]. Hand [Gra72]. Handling [Ram77b]. Handprinted [DT70, HTD72, TD70]. Handwritten [Dut74, TG72b]. Hard [DLL^{+}73]. Hard-Copy [DLL^{+}73]. Hardware [BW77, CS75, Chu79, CLL78, Des79, Erc77, Gad70, GSJ^{+}76, GL74, Ham74, Mar72, Muk79, Pag72, Rot77, Sie74, SW76, SP72c, SP77, Von75, Zoh73b, Zoh73c, Zoh74, RT73a]. Hardware-Oriented [Erc77]. Hardware-Shared [Gad70]. Harmonics [RB72]. Harrison [Nam70]. Hartley [Hil79]. Having [Ges70, God71, Sti78b]. Hazard [HK77, HK78, Hla70, Kan73b, Ser73, Ser75, Zis76]. Hazardless [Fra72]. Hazards [BH74b, HD71, Hla78, ZKK79]. Head [Kam70a]. Hearsay [FL77a, REFN76]. Hearsay-I [REFN76]. Height [LP76, LP78]. Height-Balanced [LP76]. Heterogeneous [CK79]. Heuristic [AAK74, BD75b, Col77, Coo72, Gri76, PR71, RL72, ROW71]. Heuristics [AI78]. Hexagonal [Deu70, LR76, Pre71b]. Hidden [Lou70, Wri73]. Hidden-Line [Lou70]. Hierarchical [Kul78, Nad71, WH78]. Hierarchy [Bae74, Cho76, Wel78]. High [Agr79a, AN74, BCH^{+}73, CH73a, CL75, Chi78, CH72b, Cor71b, Gar74, Glu71, Hou72, JM70, KES72, KG72, Ken77b, Kri70a, Kuc70, LE71, Lin70, Lit74, LM70, Loo70, Mar73c, NW79, Pla73, Qua74, RGK72, RT74, SW71a, SW76, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75, Woo79]. High-Dimensional [NW79]. High-Level [RT74]. High-Order [Glu71]. High-Speed [Agr79a, CH73a, Chi78, CH72b, Cor71b, Hou72, JM70, KES72, KG72, Ken77b, Kri70a, Kuc70, Lin70, Lit74, LM70, Loo70, Qua74, RGK72, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75]. Higher [Atk70, DV78a, EH76]. Highlights [Ken77a]. Highly [Agr79c, GLW74, Los76, Mai71, Man77]. Hilbert [But71]. Hole [LS76b]. Holographic [BL75, GW75]. Holography [SNN70]. Homomorphic [Zei71]. Homomorphism [Lem70]. Homomorphisms [Kam70b]. Horizontal [RT74, TG76]. Horn [Hen76b]. Hough [Skl78, TM78b]. Huffman [Par79]. Human [FD75]. Hybrid [AK75, CS73, Cas73, CM70, Deu71, GM72a, Gil70b, IS76, KK70, KV76, Lit73, MJTW75, MG71a, Ogu74, RT79, RH73, SF71, SM73a, SM73b, TL73a, Wit70]. Hybrid-Computer [KK70]. Hyperresolution [Win76].

I/O [VN78]. IBM [Joh71, SB71]. IC [Gra70]. Ideal [Hoe74]. Identification [Ada78, Bar78, Bis70, Bis73, Bou74a, BCS76, Cor77, DT76, Das78, Dep70, DBM77, Gar72b, GH72, Kel71, MJTW75, Moh71, WZ72, YT71a, You70]. Identified [Bab72, Tou74]. Identifier [Ano70-28, Ano72-51, Ano72-52, Ano72-53, Ano72-54, Ano72-55, Ano72-56]. Identifying [Eks73]. IEEE [AG73, Shi70, Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-37, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72n, Ano72-50, Ano72-57, Ano72-58, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano73a, Ano73m, Ano73n, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano74a, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano75a, Ano76c, Ano76d, Ano76h, Ano76k, Ano76l]. IEEE [Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano76b, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77a, Ano78h, Ano78i, Ano78j, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34, Ano78w, Ano78-35, Ano78x, Ano78-36, Ano78b, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79k, Ano79l, Ano79-45, Ano79-46, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79a, Ano79-53]. IEEE/Popov [Ano72n]. If [Ano76t, Ano77u, Ano78-42, Ano79-47, Ano79-48, Ano79-49, Ano79-50, Ano78-43]. II [Cur70, FL77a, Lew74, Loo70, SP72c, Sri71b, WN76]. III [Atk70]. ILLIAC [Atk70, AA72a, Dow73, Orc76]. Image [AP76a, ACJ75, Ano78v, Ano78w, Ano78x, BS72, CDL78, DM78b, FR76, Hal74, HA77, Hun73, Hun77, JA74, KF73, MMB77, NJ77, PR78, PD77, PC72, RNR75, RTKH77, Roe78, Sha78b, SC77, Sta75a, TE76, VM78, WD77]. Image-Processing [SC77]. Imagery [ASS72]. Images [AD75, Ami77, AP76b, CB72, CA77, Ehr77, GSJ^{+}76, Gra71, HKD^{+}71, HE73, HB75b, Jai77a, Jai77b, Jai77c, JR76, KC71, KS77a, Ram77b, ST79, Sta75b]. Imaging [Eks73, Saw77, WKP75, Wei75, YRT75]. Impact [Chi78, FD74a]. Imperfect [SF78, She78b]. Imperfectly [Bab72, TG75, Tou74]. Implement [Har76]. Implementation [Ban74b, Ber72, CM76, Des79, EI77, Gad70, GM72a, GL74, GC71b, HL86, Kar74, MS78, Orc76, Red71b, SP72c, SNEL79, SCNS83, TJM79, Win76]. Implementations [Alm78, Fen74]. Implemented [BB75, Dev79, Kau70c]. Implementing [CM79b]. Implicant [BM70, HW75, IR76, Mor70c]. Implicants [BH71c, Bub72, Das71c, DK72, Hwa74, Maj72c, NRS74, Reu75, SCL70, Voi77b]. Implication [VVA78]. Implications [EFLR76, Gar71, PM76a]. Implicit [SM70b]. Importance [Mal73]. Important [Ano74-28, Ano75-29, Ano75-30]. Improve [Hay74, IR76]. Improved [BK72, Ben71, BG79, Hol72a, JLH73, MSJ72, RHR76a, RHR76b, RW75, Rub75a, Sto70d, TH79, TB75, Van72, Vil70, Bel76]. Improvement [CF70b, Eks74]. Impulse [BV78]. Incompatibility [Das73]. Incompatible [SB74a]. Incomplete [Ehr72, Kit78, RB76, Yan73, YB75]. Incompletely [Bis74, DH79, GD73, Hig73, Kel70, Pfl73, RB75, Wil75]. Incorporating [Boy74]. Increasing [KR71a, SM71b]. Increment [LK75]. Incremental [BR77, Dan70]. Independent [CJ74, Cio78, Kel74, TF70, Tou72d]. Indeterminate [She77]. Index [Ano70n, Ano70z, Ano70-28, Ano70a, Ano71a, Ano71-51, Ano72o, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46, Ano72a, Ano72-51, Ano72-52, Ano72-53, Ano72-54, Ano72-55, Ano72-56, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a, BM78, IA76, Man71, RT70]. Indirect [Pea77]. Industrial [NR76a, Per78, YT77]. Inequalities [Cha71, Men70b, NK74, Tou72e, WG73]. Inertial [Ser73]. Inessential [HC71a]. Inexact [Pea76]. Inference [MB77, Vee78]. Inferencer [Gai79]. Inferred [SF71]. Information [And70, Ano70-38, Ano70-39, Ano70-40, Ano70-41, Ano70-42, Ano70-43, Ano70-44, Ano70-45, Ano70-46, Ano70-47, Ano70-48, Ano71-37, Ano71-38, Ano71-39, Ano71-40, Ano71-41, Ano71-42, Ano71-43, Ano71-44, Ano71-45, Ano71-46, Ano71-47, Ano71-48, Ano72-65, Ano72-66, Ano72-67, Ano72-68, Ano72-69, Ano72-70, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano74-27, Ano75-27, Ano78-44, Ano79-51, Anz71, BR74, Dot70, Gri73, Hop71, Key77, Low73, Mat71a, SM72a, Sri76, Ste77, TA71, TW76, Vil74, Wei75]. Information-Lossless [Dot70]. Informational [Mat71b]. Infotech [Fre74b]. Inhibition [RF72]. Initial [Kar70, MR75].

Initialization [Kri73]. Initializing [Gra73]. Injection [DMR77]. Ink [Ern72]. Inner [Bla79, SGR78, Wak70b]. Input [AH76, Ano75-28, BH73, Bre75, Bro75a, CW79b, CD73, Chu76, DDSM76, Fre79b, FG71b, FG71c, GSB^{+}76, Hig71, Hu75, KN72, Mar71, MH70a, MST^{+}75, PM75a, Pec76, SK78a, SK78b, Sin71, Smi75, Sur70b, TB75, Ung71, Vai71, Vai72, Vin78]. Input-Memory [Fre79b, Vai71, Vai72]. Input-Output [TB75]. Input/Output [GSB^{+}76, Smi75]. Inputs [FNSK75, Ges70, Mar77]. Instruction [KV73]. Instructions [Ken77a, TF70]. Instrumentation [WBRG76]. Integer [Jac73, LR75, MI72, Sik70, Wil77a]. Integral [Lee78a]. Integrals [Fos73]. Integrated [Agr79c, BT70b, DMR77, EI77, Man77, Sri71a, WA73]. Integration [Ben70, Max77, RH71]. Intelligence [Arb76, BW76b, Coo72, FL77a, Hol76, Rei74b, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l]. Intelligent [EUY^{+}72, LJ79, SAFV71]. Intensity [Ern72, Saw77]. Intensity-Modulated [Ern72]. Interactive [AKM^{+}71, Bib74, BO70, CJW^{+}71, Nay71, New75a, PS71, Sam70a, WN76]. Intercell [MF76d]. Intercommunication [Tho78]. Interconnected [Hur74, Pat73]. Interconnecting [Mor79]. Interconnection [CW79c, Gec77, GC71b, Ham71, Iba71, JM70, Sie77, Sie79, TW77]. Interconnections [Ano72-39, DEA73, Lan76, Par71, Swa74a]. Interconnectivity [Sic76]. Interdisciplinary [Ham74]. Interface [DWT70, NS75, Stu75]. Interfaces [Tro75]. Interference [Bha75a, Hoo77, Rav72, SD79]. Interleaved [CBS71, KR75, Rau79a, Rau79b, Rav72]. Interlock [Cha79, Wor77, Wor79]. Intermittent [Bre73c, KP74, Kam75, KK77a, KS79, MM78a, SKM78, Var79].

Internal [HB70, Smi74, Toh74]. International [Ano79c, Ano79d, Hob73, Kra74, Lom74, Mui74, Pla73]. Interpolation [AP76b]. Interpolator [Kor76]. Interpretation [FG71a, SW71b]. Interpreter [Chu79]. Interrelationships [AM77]. Interruption [Mcd72]. Intersections [BO79]. Interval [NS75, She78b]. Intervals [KP75]. Interwoven [Abr75]. Intrinsic [CA74, FO71a, Tru71, Tru76]. Introduction [Ano73w, Ano76g, AG73, Bas71a, Car73, CH75, Den74, Fen77, Fle79, FR71, Ham74, Hen76a, Kim75, Kin72, Lan79, Les71, Luc72, MR76, Ram71, Ram73a, Red78b, Rei74a, Rei74b, Sch74a, Sil73, Vra77]. Invariant [Eks73, Pea75]. Invariants [DBM77, Mat71a, WGS78]. Inverse [AA77]. Inverses [Ols70, SK74, Yua74]. Inversion [Che73, NC78]. Inversions [Wak70a]. Inverted [Ste77]. Inverters [BS77]. Invertible [SSK72]. Investigation [AGM^{+}71, BCB74, Hen71]. investigations [RT73a]. Invited [Ano72n]. Irredundant [AA75, Bub72, CM79c, DBC71, HG78a, Hay76b, Mor70c, NA76a, NRS74, To73]. Irrelevant [BB78a]. Isobaricity [GBMC72]. Isolation [BCA^{+}75]. Isomorphism [Yan75b, YM78]. Issue [Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78t, CH75]. Issues [Bha77, TW76]. Iteration [Fly70]. Iterations [Des74]. Iterative [ARS71a, BBC72b, Bar74c, Bor72, CH73a, CM70, CW77, Coy79, Dev75, Dia76, Eks74, FM72b, Fri73, Gar74, GR72a, GR73b, Goo70, Goy76, Ham71, HI72, IH72, Kon71, Kri70a, Kri71b, LY71, Mag70, MK71b, MF71, Nic71, Ogu74, RGK72, SM73a, Sri79, Ung77d, Ung77c]. Iterative-Improvement [Eks74]. Iteratively [ATN70, Hu73, RF75]. IV [AA72a, Dow73, Orc76, Swa72b]. IV-Type [Orc76].

J [Wor79, EW77, MF76b, Obe76, Rub79, TB74]. James [Ano73z]. January [Ano73f, Ano76x]. Jets [Ern72]. Job [Cam72, Mal72, SW76]. Job-Mix [Mal72]. Join [Tha78]. Joining [Sta74]. July [Ano77x]. Jumps [RF72]. June [Shi70, Ano72-40].

Karhunen [FK70a, Jai77a]. Keyhole [Oza78]. Keys [Sel74]. Kind [She70a]. Knee [ADL72]. Knowledge [PS71]. Kronecker [Dru71b].

L [Tab70b, McC79]. Labeling

[DR77, ZHR77a, ZHR77b]. Ladder [BS76]. Ladder-Structured [BS76]. Landmarks [HEWB72]. Language [BT74, Fel79, HP75, KS77c, RT74, RC71c, Tho76, Cam72]. Languages [Arn70, Bar75a, BW76b, BT73, Gre75, Har74b, Har70b, Hoe74, Pag70a, Pla73, Ros74a, Sta70b, TF78, Tho75, Weg76]. Laplace [MR75]. Large [Ari79, BHK79, Ben77, BM70, CM70, GC71b, KO78, New70, RC71b, Ram75, Ram77b, Smi74, WA73]. Large-Scale [CM70, GC71b, WA73]. Latency [DTG79, FM78, SM76]. Latin [HB75a]. Lattice [CW79a, Yeh70]. Layer [Tak76]. Layered [Uhr72]. Layout [Lar71]. Learning [Ama72, CF70b, Dee72, DTI78, Fu74, LD75, Mic74, MS75b, MS76, Ull72, UC75]. Least [All75b, Gol72, HA77, Hun73, Wil77b]. Lee [Rub76a, Hoe76, Rub74]. Length [Bak79, BC71b, But79, Hil79, Hol72a, LE71, MH70a, MBM77, PW79a, Per70, Qua74, Smi71a, Wei72]. Level [Ami77, BK76, BS75, CC72, CCB70, EZS78, ET75, Far79, HP75, Ios78, KV73, MK79, Nak79, OY70, PR78, Pla73, PR74b, RT74, RR74, RH75b, RV73, Liu75]. Levenshtein [OTK76]. Library [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50].

Like [BCKT79, Cur73a, KMC72, PP75]. Limitations [Sha78a]. Limited [Mak71, RSH79]. Limits [Ano77b, Ano77c]. Line [Ada78, Agr75b, All75b, Ano70y, Bar78, BLZ75, BW70, BL75, DT76, Das78, Fre79a, HPP77, Kau74a, Lou70, Lud74, Mac70, Meo75, MS75a, RL72, Ros74b, SW70a, SW70b, Sta74, SSR72, TE77, Wei75, Wil70, Wil77b, Wri73, You70, ZHR77a, ZHR77b]. Line-Drawing [RL72]. Linear [Bre72a, Can71, Cha71, Cha73a, Che74, CK75, CW77, Das71a, DS72a, Dav70a, DM78b, Dep70, Dun75, EGG72, FL77b, Gal78, GR71a, GR72a, GR73b, HO74, Has73, HI72, IH72, Kar77a, KCLB71, KP75, Kri77, Mar70, Mar73a, MAB78, Men70b, MBM77, NK74, Ols70, Pat71, PR74a, Pea76, Per70, Red78a, SR73, SK77c, SIF76, Tom74, TV72, VD76b, Van75, WG73, You71, Yua74, RS71b]. Linearity [GCC74, KD73]. Linearly [Hur74]. Lines [CY76, Hor71b, Hor72, MMO74, SKK79]. Linguistic [Mam77, WBRG76]. Linkages [CR72b]. Lisp [Tei76]. List [Ano76a, Ano78s, Ano78a, Ano79-52, Hol79, Mor70b, MM70, Weg72]. Literature [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano70m, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m]. Little [Yue77d]. Load [CK79]. Loaded [SW70a, SW70b]. Loading [Hou72, Wym73]. Local [Gra71, RS75, TS79]. Localization [FS78]. Localized [SD79]. Locatability [GM71]. Locatable [Red72a]. Locating [Nev76]. Location [FM74, Sri78, SC72b]. Loeve [Jai77a, FK70a]. Logarithm [HL86, SK71, SA75, SCNS83]. Logarithmic [DG72]. Logarithms [BB75, HLD70, Mar72]. Logic [Abr75, AM77, AA72a, AA75, AA76, Agr77, Ake72, Ake73, Ake76, Ale71, Alm78, AB78, Arm72, Bar77, Bar70, BW76a, BCA^{+}75, Ber70b, BF75a, BHP^{+}71, Bre72a, Bre72b, Bro74, CJW^{+}71, CS70, CM74, Cer76, CE75, Chi77, Cle73, CM79b, Dan73, DR74, DMR77, Dia75, Dia76, Dir72, Dor72, DVS74, Edw75, EZS78, ET75, Elm76, EH76, FS70, Fis74, Fle79, FM72a, GW77, Gar72a, GC73, GLR71, GT74, GT78b, Hay71a, Hay74, Hay76b, Hay76c, Hay78b, HC73, Hu75, HK74, Ios75, Ios78, ISN76, IC78, JM70, Kam79, KH77a, KSM76, Kli71, KAN79, Kod74, KS75, KS78c, KS71, KK72, LR71a, LV75, Lec70b, LC79, LM70, Loo70, Lyo70, Mam77, MA71, MC71, McC79, MR74, MF71, MBSS75, Min75, Mor70a, ML76, Ops72, Osb77, OW73b, OH79, Pap78]. Logic [PM75a, PW79b, Pat73, PR72, Pre72, RT70, Red72b, Red73a, RW74, Red77b, Ree70, RM78, Rot78, ROW71, Rus72, SR74a, SR74b, SK78a, SK78b, SK79a, Si78, Sin72, Skl70, Sli70, Smi78, SM78, Smi79c, Sri79, Sto77, Sto70b, Su70, Su75, SB74b, Sur78, TM76, TD77, Tis71, TKH71, Tom75, TK74, Toy71a, TG79, Voi77a, Voi77b, Vra77, Wan75a, WR77, Wea75, Weh74, WA73, Win70a, Win70b, Woo79, YI74, YT70, YT71c, YY75, Yue75b, Yue76]. Logic-in-Memory [ET75, Sto70b]. Logical [Aus70, Bet71, CF73, GL74, KF78, KT72, MH73, MH75, Mas78, MS70, TM71, Wak74]. Logics [Mor76b]. Long [Sut74, Wil70]. Lookaside [CDJ73]. Lookup [MS77b]. Loop [Gal78, PB76a, VN78]. Loopfree [RS74]. Loops [BCKT79, Gil70b, Kau70b, Lit73, Pag72]. Lossless [Dot70]. Low [PA78, She71, Wak75b]. Low-Cost [PA78, She71, Wak75b]. Lower [Du72, FS70, Han72, Van72]. LR [Jol76]. LSI [Chi78, HB75a, Rus72, Toy71a]. LX [HA70]. LX-1 [HA70].

M. [Ano73z]. Machine [Abd71, Alm78, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, Arb76, Ber71, Cur73b, Fel79, God71, Gra73, Ham71, HC72, Har70b, JK77, Kel71, Koz71, Kru73, LV75, LF74, Osb77, PM74, PHG77, RN77, Sie77, SM71c, Sur78, Tom72, TK74, WZ72, YT77, Yan74b, Yoe70a]. Machines [BF72, Bis74, Bou74b, BH71b, Bre75, Brz70, Cha72, CL75, Che74, CD73, Chu76, CD78, Cur70, Cur73a, DF75, Dav70a, Dav70b, DSBC74, DS73, DGC74, DAA79, DW72b, DH79, ES77, Ehr72, Far73, Fre79b, FM73, FK74, FNSK75, GH78, Gel71, GR71a, Ges70, Gil70a, Goo70, GB71a, GB71b, GD73, GCC74, GCP^{+}77, HSP78, HG79, HN73, HC71a, Hol72a, HC71b, Hsi71b, HI72, Hu72, Hwa73, IH72, Jos78, KYO70, Kan75, Kar74, KM73, Kel70, KN72, Kri73, KR78a, Lan79, MS73a, MS74, Man72a, Mar70, Mey71, MR72, MKO70, NT79, New70, Nie70a, PM77a, Par72a, Par72b, Paz70c, Pfl73, Pre70, RB75, RB76, RV74, Rey70, Ric72, Roo73, RT73b, Sau72a, Sau72b, SM74, SM75b, She70b, SV74, SG71, Sie79, Sil79, Smi74, SB74a, TI72a, Tan71, TA71, TOS71]. Machines [TB74, Vai71, Vai72, Wak75a, WY75, Wan71, Wil78, Wil75, Yan73, YB75, Yeh70, Yen79, Vee78, Lyn72]. Macro [Aus70, NN73]. Made [Lio76]. Magnetic [Bha75b, CW79a, Che78, Gar72a, Hoa76, KSM76, LJ79, MBSS75, Min75, PB76b, Wil77a, WC76]. Magnitude [BM78, MMO74, RT70]. Main [JM73, Tho74]. Maintained [Agr79c, CLL78, Man77]. Maintenance [CHSS73, Gay78]. Majority [HP73, RR74, Sie75]. Make [Lam79b]. Making [KKA71, PS71]. Maloperation [Cha79, Wor77, Wor79]. Man [Arb76, Bon74, Els74]. Man-Computer [Els74]. Management [Ano76u, DB74, De 79, Dow73, LJ79, Mui74, Wan72]. Management-International [Mui74].

Manipulating [Fen74, Wil70]. Manned [WKD70]. Many [Dig73, DVS74, DV78a, KT72, VLS70]. Many-to-Many [Dig73]. Many-Valued [DVS74, DV78a, KT72, VLS70]. Mapping [EH78, Kru71, LSB77, NW79, ROW71, Sik70, Whi72]. Mappings [MR74]. Maps [ACJ75]. Maritsas [Hil79]. Markets [Jos72b]. Marking [PC72]. Markov [Hab77, SKM78]. Maryanski [Gai79]. Maryland [AG73]. Masking [Dia75, RK75b]. Masks [Sie77]. Mass [PA78]. Master [FG76]. Master-Slave [FG76]. Match [And74]. Matching [Cho76, DW78, DN77, FE73, Oza78, VR77, WH78]. Mathematical [Bre71c, Dir72, Duk74, FS70, Kul77]. Mathematics [Har74a]. Matings [And76]. Matrices [All75a, Ari79, Ban78, DS74, Har73, Hen78, Oja79, Ram75, Sag77, SK77a, SK77b, SSK72, TF73, Van77]. Matrix [ARS71b, AC71, Ast78, Cap76, Car71a, Che73, CL79, Den76, Dru71b, Ekl72, FA76, God73, Gou77, Hwa72, Koh72b, KR73, Kri77, NC78, Ram77a, Rat76, RF75, SG74, SR73, Sch73, SK74, Tom70, Tom72, TE76, VR79]. Max [MMB77]. Max-Min [MMB77]. Maximal [Ada78, ADMR78, Bar78, DT76, Das78, HP72, LE71, MBM77, PW79a, Qua74, Rub75b, Yan75a]. Maximal-Length [PW79a]. Maximin [Hed70]. Maximization [Ram77a]. Maximum [Boh77, Bro70a, Han75, KP75, RS72b, RS73b, Sto74b, Sur73, TH79, WD77]. Maxwell [Cat77, Ano77v]. May [AG73, Ano77y]. Mean [All75b, AOR74, Ano75c, Ano75d, Ano75e, Ano75f, Kul78, Maj72b, Whi70, Wil77b]. Means [AN74, Pag72]. Measure [Das71a, Hel72, Hel74, JP73, KF79, Kor79, Mas78, MMB77, TV72, Wel74]. Measured [Ung77a]. Measurement [BD75a, BCC78, JP76, Kim74, NS75, Whi71b]. Measurements [CJ74, Glu71, You70].

Measures [Bea78, BT73, FGR71, SH72, Tou72e, Vil73b]. Measuring [FLR78, Skl72]. Mechanical [HF71, WG75]. Mechanism [MF75, SN72a]. MECRA [Mai71]. Media [BG75]. Medium [Red71a]. Meet [Tha78]. Megasamples [GSB^{+}76]. Megasamples/s [GSB^{+}76]. Members [Ano72-50, Ano73h]. Membership [Ano72-57, Ano72-58, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74k, Ano74l, Ano74m, Ano74n, Ano74o, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano76i, Ano76j]. Memories [AU74, Bae74, Bha75b, BD77, BK71b, CKL77, Che78, CBS71, DB74, FM78, GSB^{+}76, GM70, Hay75, JM73, Klu77, KH77b, KH77c, Koh72b, Len77, LJ79, Mon74, MVW79, NTA78, Nai79, PA78, Pra78b, Rau79b, RF75, Sha78a, SY75, Sri71a, Sti78a, Sto72, Sto75, Sun78, Sun79, Swa74a, Wil71, WC76]. Memory [BK79, Bat77, Bha75a, BCH^{+}73, BE70a, BE70b, BB75, CDJ73, CM76, Cov74, Cur70, Cur71b, DI71, DBC73, Dig73, Dru71a, ET75, FS71, Fre79b, FT79, Gar71, GSB^{+}76, GLW74, Gra73, HK77, HK78, Hoo77, Hou72, HB75a, KYO70, Kar76, Kau73, Kin71b, KR75, Koh74, Kuc70, Kur74, Lan76, LC79, LM78, MJ77, MJ78, Nel74, Nut77, Pea76, RTW78, RSH79, Rau79a, Rav72, SNN70, SD79, Sho74, SW76, Sri78, Sri71a, Sri71b, Sto70b, Sto70c, Tho74, TI72b, Toh74, Vai71, Vai72, VM78, Von75, WSB79, Wel78, Wil78, WY73, Yua74]. Memoryless [CM74]. Merging [Gol76, Hol79, Lam79a, Par79]. Merit [Hec73]. Mersenne [Rad73, Rad72]. MESFET [TG79]. Mesh [NS79]. Mesh-Connected [NS79]. Message [Sah74]. Message-Based [Sah74]. Metaphorical [Rei74b]. Metastable [CW75]. Method [AA75, And74, ACJ75, Ano75-28, AB78, Arm72, Ban74b, Ben71, BF75a, BCL75, BM70, BB74b, Bub72, Cha71, Che73, Cur79, Cur77, DS74, DB73, Ekl72, Erc77, Fon71, FG71c, GLR71, Gon70, HB70, Hwa72, Hwa74, Jai77b, Kit77, Kla71, KO78, Koh75, KP71, KK70, LSB77, Lee78a, LR75, Maj72b, Mar78, MR75, MT74b, Nie75, NRS74, OTK76, Oza78, Pap72, Pap77b, Pav77, RB76, RW75, RS72b, RS73b, SA70, Sch73, Set70b, Sha78b, SS75a, SW71b, SN72b, Sur73, SSR72, The75, Tom72, Vic71, WK73, Whi70, Whi71b, Yan73, YB75, Zis76]. Methodology [CM77, How75]. Methods [AK75, AKT^{+}77, Atk70, AT70, Bar70, Bar74c, CSJ74, CCES74, EH78, GBMC72, HEWB72, Hun77, Kar70, KM78, KPS77, Mag71, Mil75, MS77b, Mor76b, NC78, Pav75a, RGK72, RS75, Rob70a, Tab70b, TH79, UB74, Wes78, WS76, Wor76, Wul74, You70, YRT75, Zah71]. Metric [OTK76]. Microcellular [Ber77]. Microcode [Rob79a]. Microcomputer [Kar76, Les71]. Microdiagnostics [Guf71, Joh71, RC72]. Microelectronic [Ano72-39, Par71]. Microfilm [Ros71]. Microinstructions [JB76b]. Micromultiprocessing [RV73]. Microprocessor [BBLT73, CF70a, FT79, HA70, Pea77, Roe78, TW77, VN78]. Microprocessors [Ano79b, JM73, Lip77]. Microprogram [Age76, CSST73, Pat76, Sta70a]. Microprogrammed [AAK74, BK79, BT70b, CDS73, CHSS73, Col77, DBC73, GM70, HG78b, JF72, KSW74, Mon74, Red71a, SAFV71, Swa79b]. Microprogramming [AR74, Ber70b, FR71, FLBK76, Gar71, Hig78, JM74, RT74, RA78, Rux71, Shr71]. Microprograms [Ada78, Bar78, DT76, Das78, KR71b, Lee75, RS74, Tho74, TG76]. Microwave [YRT75]. millions [Ano76e]. MIMD [PM78b]. Min [MMB77]. Minicomputers [Kra74, Sch74b, Tro75]. Minicomputers-International [Kra74]. Minimal [AB78, BF78, Bor72, Bre71b, CMS79, Cur77, Dan73, DDSM76, Du72, Fra72, FD74b, FD77b, HG78a, Hay71b, Hig71, KLCC77, Kau70b, KS77b, Lee78b, MT74a, Mar71, MST^{+}75, NTK72, Nie70a, Red77a, Reu72, Riv77, TM76, Vai72, Vin78, Yua74]. Minimal-State [Nie70a]. Minimalization [Arb70]. Minimally [SR74b]. Minimization [BK79, BI72, BS77, Bis71, Bis74, BV79, Che74, CE75, CM79d, DBC73, DH79, Ehr72, GR73b, GM70, GD73, HG78b, Iba71, KW70, Kan73a, Kan73c, Kel70, Koh71, LD72a, Meo75, MS70, ML76, Pap79, RB75, RNMP77, SO79a, SC72a, SC72c, Sur75, Voi77b, YI74, Yue75a]. Minimize [Ake70]. Minimizes [JB76b]. Minimizing [FM78, Hun72, RB76, Yan73, YB75]. Minimum [BC71b, Bet71, CC72, CDJ73, DTI78, IM71, LM79, LL79, Nak79, Nie75, Pat72, SK78b, SCH72, Voi77a]. Minimum-Perimeter [SCH72]. Minneapolis [Shi70]. Minnesota [Shi70]. Minor [Gri78a]. Miss [Eas78]. Mission [FD74a]. Mix [Mal72]. Mixed [Yen79]. Mixed-Operation [Yen79]. Mixture [BD75b]. MLP [Guf71]. MLP-900 [Guf71]. Modal [JW76]. Mode [Byr71, MS76, Ung77b, Yen79]. Model [Arn73, Ban74a, BW76b, BF75b, BP72, BY79, Chi71, FS74, Gel71, Hay71a, Hay76a, Hoo77, IS76, Jai77c, Joh71, KP74, Kim70, Kim79, Law70, Low73, Mat71b, MK79, Mut76, Per78, Rau79a, Sie79, SN79, SKM78, Ung77a, VN78, Yam70]. Model-Based [Per78]. Modeling [BCSW76, BCJ^{+}71, BCH^{+}73, BT76, Mal72, Mat71c, RC72, Sie75, ST76, Var79, Yan74a]. Modelling [JA74]. Models [CK79, CLL78, Dir72, Fre77, MH76, MH78, Mil73, Pea75, Reg73]. Modem [McL72, Swa72b]. Modern [SM71b]. Modification [NR74]. Modified [Agr76, DS72a, Das73, Gra73, Hil79, Kit78, LK75, Rub75c, Smi76, Tom70, TM78b]. Modifying [Hay74]. Modular [AS74, AC71, Cer78, Dav77a, DM78a, DW72b, GR71a, GST79, Hwa79, KK78, KS79, NM77, OW73a, SM75c, Sut74, TG78, TK74, Toy71a, Wak75c]. Modularity [O’K73]. Modulated [Ern72]. Modulation [Kon71]. Module [Ake70, BK76, KH77a, PFL75, Sie75, Voi77a, Voi77b]. Module-Level [BK76]. Modules [Ale71, Alm78, BEGW72, FM71, GT74, HK74, Kel74, Lan76, LV75, NA72, Ops72, OW72, Pat73, Pre72, SN79, SB74b, Sur78, TW77, Wak75c, YT70]. Moduli [O’K75]. Modulo [Boy74, CR72c, Pap79, YC76]. Modulo- [Boy74]. Modulo-2 [Pap79]. Molecular [Fre77]. Moment [DBM77]. Moments [SW71b]. Monitored [de 71]. Monotone [Iba74, Lam79a, Mag73, Riv77]. Monotonic [BG78, Dan73, Mat71b]. Monotonicity [Fon71]. Monte [AA76, SF74]. Morphisms [Yan75b]. MOS [MJ78, Dir72, EZS78, Lar71, Liu75, Liu77a, Liu77b, MJ77]. Mosaic [Skl72]. Mosaics [DS75]. Motions [HEWB72]. Movement [CW79a]. Moving [AD75, CA77]. MSI [PGS70]. MU5 [LTE77]. Muller [KS75, KS77b, PP75, SR75, SO79a, Swa72a]. Multi [Kam70a, Nak79, SD74]. Multi-Head [Kam70a]. Multi-Output [Nak79, SD74]. Multi-Tape [Kam70a]. Multicache [CF78a]. Multicategory [Glu71]. Multicell [SNEL79]. Multichannel [Gil70a, Ric72, Win75]. Multiclass [Tou71]. Multicode [KR78a, NT79]. Multicomputer [KK79a, LTE77, She70d]. Multiconnected [TWB78]. Multidimensional [And71, Bat77, CF79, GR72a, GR73b, Kun79]. Multifont [Sch78c]. Multiform [KD73]. MultiFunction [Ham73]. Multigate [Win70a]. Multihypothesis [Dev74, Tou75]. Multilevel [Glu71, Liu77b, Ram76b, SN71, TG72b]. Multilingual [LS71]. Multioperand [AO79]. Multiple [AR78, And74, Bos79, BH71a, BT76, BH73, Bre75, BCS76, CSJ74, CK79, Chu70, CD73, Chu76, CBS71, DS75, Dos77, DW73b, FM74, Fos73, FT76, FT79, GRR72, GS79, Ges70, GLR71, God71, GS74, HC73, ISN76, JM73, KH77a, Kar74, KK72, KM75b, Lan77, LB74, Lin70, Mag70, Maj72c, MM78b, Mil74b, Mor79, Pra78d, PG75, RT73b, SM71a, SW73, Smi78, Smi79a, Smi79d, SN71, TG79, UC75, VVA78, Vra77, Wak75b, YY75]. Multiple-Bit [Lin70]. Multiple-Choice [Dos77]. Multiple-Fault [FM74]. Multiple-Input [BH73, Bre75, CD73, Chu76]. Multiple-Microprocessor [FT79]. Multiple-Output [DS75, Mag70, SN71]. Multiple-Response [Lan77]. Multiple-Stream [Mil74b]. Multiple-Valued [GLR71, ISN76, KH77a, TG79, Vra77]. Multiple-Word [CBS71]. Multiplexed [HW76, MMB78]. Multiplexer [TM76]. Multiplexers [Lan78]. Multiplexing [LC72]. Multiplication [AC71, BBC72b, BW73, Bla74, BH71c, BB75, Che71b, CF78b, DS72b, JB76a, KG72, Kro74, Lev70, Lin70, MK71b, RGK72, Rub75c, SR72a, SW73, SKG77, Tom75, TE77]. Multiplicative [Bak75b, Bar77]. Multiplier [CW79b, Goy76, Ken77b, LR71b, MG75, MS78, Obe72, Pez71a, Rob70a, Swa73b, Pez71b]. Multipliers [CL72, GG75, HW70, Hwa79, NR77b, Nus78, Sik70]. Multiplies [HO79, Sil77]. Multiply [She71]. Multiplying [Mar74, TP72]. Multipoint [CR72b]. Multipole [Ges71]. Multiprocess [Lam79b]. Multiprocessing [Bae76, BCHL73, KV73, Reg73, RV73]. Multiprocessor [Bha77, Cov74, De 79, EFLR76, FB73a, HS75, Koh75, Kur74, Lam79b, Mal72, RCG72, Rau79a, Rum77, SK75, Sch79b, SD79]. Multiprocessors [Bha75a, Hoo77]. Multiprogrammed [RC74]. Multiprogramming [AI78, BCHL73].

Multipurpose [GHJ77]. Multiserver [Reg73]. Multispectral [Deu71]. Multistep [Ben70]. Multitalker [KM78]. Multithreshold [DVS74, DV78a, GC71a, GC73, Ish77a, Ish77b, Set70a, SR72b]. Multivalued [BF75a, DMR77, DM75, DM76, Eps73b, HE79, Ish77a, KS75, KS78c, Kol72, LL72, Mar78, McC79, NSR74, PP75, Sin77, SS77, Sri79, SC72c, SS72, VW72, Yoe70b]. Multivariate [BCL75, FO71c]. Multiwire [MMO74]. Munson [HTD72]. Mutual [BR74, Vil74]. Myopic [BB78b].

n [TOS71, KS79, Rat76]. N-Modular [KS79]. N-Tuple [Rat76]. N.Y [Tab70b]. NAND [CCB70, Cur79, Hay71a, LM79, LD72b, Pap78, SN71]. Natural [BDR77, GC77, Mur70]. Near [AB78, Bou74b, Hay71b, JP73, Pat73]. Near-Minimal [Hay71b]. Near-Optimal [Bou74b, Pat73]. Nearest [Cha74b, FBS75, FN75, GY77]. Nearly [Cur79, Yue74b]. Necessary [Smi79d, Spi71]. Necessity [Kau70b, Riv77]. Needs [Jos72b]. Negabinary [Agr78a]. Negative [Agr75a, Cha74a, Iba71, IM71, Kri71a, NTK72, Nak79, Paw71, SCK73a, SCK73b, Wad71, Zoh70]. Neighbor [Cha74b, GY77]. Neighborhood [O’C75]. Neighborhoods [ADMR78]. Neighbors [FBS75, FN75, JP73]. Nested [Hop70]. Nets [Ama72, BB73, Mor79, NN73, TKH71, Tun70b]. Network [Bos79, CG74, CCB70, CR72b, CCF79, CF73, EFLR76, Hen71, Hol79, KO78, Lan76, LS76a, LD72b, Lee78b, Len78, Mar77, Mas78, MPS78, OY70, SA74, Smi75, Tom70]. Networks [AS74, Abr75, AM77, AM79a, AA75, AA76, Ake70, Ake73, BK76, BCSM72, Bei70, Bei71, Ben71, BK73, BH71a, Bre72a, BCS76, BA71, BY79, But75, But76, But78b, CK74, CW79c, Cle73, CYNM79, DR74, DB76, Dav74, DW72b, DVS74, DV78a, EZS78, Fra72, FD74b, FD77b, GRR72, Gec77, GC71a, GS74, GM71, Har72, Hay71a, Hay74, Hay76b, Hay78b, HW79, Iba71, IM71, Ios75, IA76, Ish77a, Ish77b, KH77a, KK78, Kau74b, Kin71c, Kle76, KS78b, Koh71, KK72, KB75, KK77a, KK77b, Lam79a, LB74, LD72a, LD74, Liu75, Liu77a, Liu77b, Mah74, MH73, MH75, Maz71, MC71, Meo75, MI72, ML76, NTK72, Nak79, Pai73, Pap77a, PM75b, PR73, PR74b, Pra78b, PMB77, Ram73b, Red77a, Ree70, Ris74, RR74, RH75b, RK71, Sah74, Sah77, SR74b, SR75]. Networks [SM75a, SM71a, SD74, SK77c, She78b, SY75, Si78, Sie77, Sie79, SR71, SM78, Smi79b, SN79, SS77, SN71, SH79, Tak76, TWB78, Tho78, TS79, TBY78, Uhr72, Van72, Vin78, Wan75a, Wan75b, Wea75, Whi71c, YI74, KD79]. Networls [SM75c]. Neumann [Yam70]. Neustadt [Tab70b]. Newton [Hun71, Pav77]. Next [AH76, Ano75b, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, MS73a, Sau72a]. Next-State [AH76, MS73a, Sau72a]. Nine [Mut76]. Nine-Valued [Mut76]. Noise [Boy74, CY76, EGG72, Jai77a, Mar78, MMO74, Sok72]. Noise-Perturbed [CY76]. Noisy [DR75, KS77a, LF77, Tru76, You70]. Non [Kar77b, MH70b, MH70c, RS71b]. Non-Abelian [Kar77b]. Non-linear [RS71b]. Non-Poisson [MH70b, MH70c]. Nonclassical [Mor76b]. Nondeterministic [DW72b, KW70, Moo71]. Nonfundamental [Ung77b]. Nonlinear [Bar74c, CA74, Du71, Far75, FB74, GH72, Hor72, Hun77, Kai70, KF72a, Kru71, LT73, Max77, ND73, NW79, OF73, Tab70a, Tab74, TL73a, WK73, Whi72]. Nonlinearities [Saw77]. Nonlinearly [AA71]. Nonnumeric [Muk79]. Nonorthogonal [Cal70]. Nonparametric [BT70a, DTI78, Fri77, GY77, JLH73, KF72b, KF72c, KNF76, MS76, Ram76c, Rub76b, Whi71b, Bel76].

Nonrecursive [Dav71, Zoh73c].

Nonredundant [Wei75]. Nonresetting [JM71]. Nonrestoring [BJ75, Chi72, MK71a, Rhy71]. Nonsquare [All75a, Van77]. Nonsupervised [CF70b]. NOR [BCSM72, LM79, Wan71]. Normal [BD75b, CM79c, HG78a, Mor70c]. Normalization [KCLB71]. Notation [Ber77]. Note [Agr77, Ano77h, Bre72b, Cha79, Ehr72, Gon79, Goy76, Gra73, Har70b, KO73, Kan73b, Kod74, KS77b, Mcd72, O’K75, Ols70, Pre70, Qua74, Red74, Red77b, Reu72, Ros75, SR72a, SIF76, SB74b, Ulr71, Wor77, Wor79, Wyn71, Yue75b, Zal71]. Notes [CR72c, Pag72]. Notice [Ano70-27, Ano71x, Ano71y, Ano72-47, Ano72-48, Ano72-49, Ano73g, Ano74g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76f, Ano77i, Ano77j, Ano78u, Ano79-29, Ano79-30, Ano79-31, Ano79-32]. Novel [Agr78a, Ban74b, NT70, Toy71b]. November [Ano76y, Ano77z]. NP [Rob79a]. NP-Complete [Rob79a]. NPN [Bau72]. ns [Pez71a]. Number [Ano75-28, BB72, Ban74b, BM73, BM74, BM78, Bau71, BB78c, Bor72, Bre73b, But75, DS72b, FB73a, Har73, Hay71b, HE79, IM71, Jul78, KKK73, KKK74, KMC72, LE77, LE79, NM76b, NM76a, New70, Nus78, O’K75, PW79a, Par72a, PB71, RN77, Sag77, SK79b, SB77b, SA75, Swa79a, Tom75, TJM79, YL73, Yue77e]. Numbers [AR78, BBC72b, Bau77, Bre70a, MK71b, Maj72a, Mar73c, MS77b, Yue75c, Yue77c]. Numerals [Yue73b]. Numeric [KS77c, Mat70]. Numerical [Cod73, Eks74, Eks73, Fer77, Har74a, Kor76, MR75, Pav75a, Pic74, PH73, RH73, Shr71, TM76, Wor76].

O [VN78]. Object [Nev76]. Objective [HEWB72]. Objects [AB76, BG75, CC73, EUY^{+}72, MA77]. Observability [Cer78, MR72]. Obtained [Dav71, HTD72, SC77]. Obtaining [SK71, SS75a]. Occurrence [MR72]. Octal [PB71, Yue73b]. October [Ano76z, Ano77-27]. Office [HSS78]. On-Line [All75b, HPP77, Lud74, MS75a, SSR72, TE77, Wei75, Wil70, Wil77b, You70]. On-Set [DGC74]. One [Bar71, BDR77, DR75, Geo70, IH72, Kun79, LHSM74, Pav75b, PB76a, Tom74]. One-Bit [LHSM74]. One-Dimensional [BDR77, DR75, IH72, Kun79]. One-Way [Bar71]. Ones [KO78]. Only [BB75, Dav70b, Mon74, Pra78b, SY75, GM70]. Op [FGR71]. Op-Code [FGR71]. Open [Gon79]. Operand [AR78, SW73]. Operands [Avi73]. Operating [Bas73, BS74, BR76b, CW75, Fre74b, FM71, Mil74a, WKD70]. Operation [FG71a, Hig78, Ung77b, Yen79]. Operations [Ano76u, Cha73b, Deu70, FL75, Jul78, KMC72, SC77, Wak74]. Operator [Jai77b]. Operators [KR73, Mor70c, The75]. Optical [Cas73, CH75, CS75, CR78, Far75, FHTS78, HND75, Koc75, Koz71, MJTW75, Sta75b, Stu75, Win75, GW75, LH75]. Optical-Digital [HND75, Sta75b]. Optical/Digital [CS75, CR78]. Optimal [BS78, BCSM72, BD75b, Bou74a, Bou74b, Can71, CG74, CW79c, CS79, CYNM79, Cur79, FB73a, FS75, Ful72, GLL79, Hay78a, Ios78, Kau74a, KH77c, Kri70a, LT73, LHSM74, MS77a, MG71a, Mon74, MI72, Nai79, NFS77, Noe73, NR74, Pap78, Pap77a, Pat73, Pav75b, PM77b, Pea75, RCG72, Sam70b, SH77, SF72, Tab74, Tan78, TW77, WG73]. Optimization [Age76, AK75, Bar74c, Bur74, DS75, JM71, KR71b, KK70, LNL72, MM73b, Rob79a, Rot78, Saw74, Tab70b, TWB78, TG76]. Optimum [GM76, NR76b, Ops72, PC72, Rob79b, Sok72, SHR72]. Order [Car71a, EH76, GC77, Glu71, RS71b, SF77, WK71]. Ordered [Bro77, GY77, Kun75, RTW78, Swa74a, Ulm70]. Ordering [Abe72, But78a, CL77, FK70a, KJ79, TF73].

Orders [Pea76]. Ordinary

[Fra78, FL75, New75b, RS71b, Wor76]. Organization [Bar74b, Bas71a, BCH^{+}73, BD77, BK71b, De 79, Flo72, GLW74, HC70, Hsi71a, PM78b, RC71b, Ram73a, Sie74, Sto70c, Tho74, Voi75, WG74]. Organizations [BCB74, Cov74, EFLR76, Fly72b, SK75, VN78]. Organized [Cha73b, Gri78b, NR75, Red78a]. Organizing [Ama72, JB76b]. Orientation [DF78]. Oriented [AM77, BBLT73, Bre73a, Erc77, Fel79, Fis76, Gil74b, RA78]. Orthogonal [And70, CC73, CL79, HB75a, KT72, NFS77]. Other [Jul78]. our [Ano79b, DM76]. Outcome [CF70b]. Outer [AP76a]. Outlook [BG75]. Output [Ano70y, Bou74b, CW79b, DS75, Fre79b, GSB^{+}76, Kri73, Mac70, Mag70, Maj72c, MR72, Nak79, Ogu75, PRD71, PM78a, SK78a, SD74, Smi75, SN71, TB75, Vai71, Vai72]. Output-Memory [Fre79b, Vai71, Vai72]. Outputs [Chu70, Dav70a, FK74, Obe72, SM70b]. Overflow [BK71a, CF78b, LL75]. Overflows [Gar78]. Overlapped [Hig78]. Overlapping [AL71]. Overview [AG73, Ram71, Vra77].

Package [BCC78]. Packaging

[Chi78, Hou72]. Packet [CG74, MN78]. Packet-Switched [CG74]. Page [GF78]. Paged [TI72b]. Paged-Memory [TI72b]. Pages [Pat72]. Paging [ABK76, GF78, Smi76, Tri77a, Tun70a, Yan74a]. Pair [Pag70b, Yeh70]. Pairs [PHG77, SB74a, Tom72, YT76]. Paper [Ano71-52, Huf76]. Papers [Ano73e, Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77e, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano78v, Ano78w, Ano78x, Ano79r, Ano79s, Ano79t, Ano79u, Ano79v, Ano79w, Ano79x, Ano79z, Ano79-27, Ano79-28, Ano79y]. Parallel [Agr75b, Ano73w, AM79c, Bak75b, BCKT79, BS78, BW73, Ber72, Bla74, Bre71a, BKM73, BD77, BK71b, Buz73, Cha73b, CCES74, Cha74d, CKL77, CK75, CW79b, CW77, CSY75, CM79b, DF75, DEA73, DS72b, Deu70, Fen74, Fen77, FR72, Fra78, FW78, GG75, GR71b, GR72b, IC78, KO78, Koc75, KS73, Kro74, Kru73, LM79, LT73, Len78, Les71, Li77, Man71, Mar73d, MF75, Mil73, NS79, NN73, PM76b, Pre78a, RC71a, RPL73, RFL76, RL78, SK77a, SK77b, Sha78a, SN72a, SCL76, SB77b, Sta74, Sta75a, Sta75b, Ste72, SKG77, Sto70c, Sto71, Svo73, Swa74a, Swa73a, Tab74, Tho78, TP73, TF70, Toy71b, WC74, Wor76]. Parallel-Acting [Man71]. Parallel-Decomposable [DF75]. Parallel-Pipelined [BD77]. Parallel-Sorting [Pre78a]. Parallel-to-Serial [IC78]. Parallel/Pipelined [Li77]. Parallelism [Ada78, Bar78, DT76, Das78, FL77a, RF72]. Parameter [Boo70, Bro71, LNL72, SKM78, SSR72]. Parameters [BPC75, BD75b, Dui76]. Parametric [New75b]. Parity [Hen71, KM73, OY70, Toy71b]. Parsers [Jol76, TF78]. Parsing [Lew74]. Part [Cur70, LM70, Loo70, SP72b, SP72c, Sri71a, Sri71b]. Partial [BM77b, CRB72, Cov74, DS71, God71, Jol76, Kar70, KD73, MA77, SF74, Uhr73, Wit70]. Partial-Fractions [God71]. Partially [Sch75, Smi71b, Wak74]. Partition [GC73, Pag70b, PHG77, RB76, Yan73, YB75]. Partitionable [GB71b]. Partitioned [Mor70b, MM70, Pat72]. Partitioning [ARS71b, BT70a, Bre70b, DM78b, Fri77, GHJ77, MM73b, PGS70, RH73, ROW71, Sto77]. Partitions [BD75b, LR71a, MR74, Yan74b]. Parts [Per78, YT77]. Parzen [Dui76]. PAS-II [WN76]. Past [TA71]. Patch [RC71c]. Patches [Mah72]. Patching [RC71c]. Patern [Pav74]. Path [CRB72, Hay78b, Koh75, Rub74, Rub76a]. Paths [KK79b]. Pattern [AA77, Ama72, Ano78h, Ano78i, Ano78j, Ano78-37, Ano78-38, Ano79k, Ano79l, Ano79c, Ano79d, BC71a, Bab72, BD71, BT70a, BT72, BB78a, Bre73a, BP72, BR74, Cal70, CDO78, Cha73a, Che71a, Chi71, CR72d, Chu74, Dev74, DN77, FE78, FP75, FB73b, FR76, FF75, Gil74b, Gil74a, Hay75, HF71, Hus72, Kit78, LT70, LNL72, LS72, LF72, LF78, MM73b, Mic74, MF76c, MG71b, NC78, Nad71, PM76a, Pre71b, Rem74, Sam70a, SLC75, Sel74, STG71, SN72b, Sto74a, Tak76, Tou71, Tou72a, Tou72c, Tou74, Tou75, Vil72, VV71, WG73, Wea75, WL75]. Pattern-Recognition [Tak76]. Pattern-Sensitive [Hay75]. Patterns [Ama72, BDR77, BL75, Glu71, GD72, Hus74, KKA71, LF77, NFS77, Rat76, TG75, Wan75c]. Paull [Pra71a]. PC [AB79b]. PCN [Mat71a]. Peaks [NM76b]. Pen [BdB70]. Penalties [Low73]. People [Ano74-28, Ano75-29, Ano75-30]. Percolation [FR72]. Perfect [Sto71]. Performance [Bea78, Bha75b, Bha77, CM73b, CBS71, FD75, Gri76, GF78, Kla71, Rau79b, Rus72, SK75, Tri77a]. Performance-Related [Bea78]. Performing [Wak74]. Perimeter [SCH72]. Periodic [Gaj75, GB71a, GB71b, Hwa73, Law70, RS72a, SW70a, SW70b, Set70a, Usa75]. Peripheral [Bar74a, Flo75, MF76a]. Peripherals [Ano73x, Ano73y]. Permanent [KS79]. Permutable [WY73]. Permutation [CW79a, Har72, Orc76, Ram73b, KD79]. Permutations [Che78, Len78, WC76]. Permuter [BBC72a]. Perspective [AR74, Kim75]. Perturbed [CY76]. Phase [Dir72, MAB78]. Phase-Shift [MAB78].

Photo [SW71b]. Photographs [BD71]. Photomosaicking [Mil77]. Photomosaics [Mil75]. Phrases [KM78]. Physical [Key77]. Pi [IR76]. Pi-Algorithm [IR76]. Pictorial [Ano70y, FE73, Mac70, Ram76b]. Picture [AL71, KS78a, KM75a, Kru73, Mor74, ND73, OC76, SB72, ZRD75]. Pictures [ADMR78, Hun72, MF75, SC77, SP72a]. Piecewise [ADMR78, Can71, Cha73a, KP75, Lee78a, Pav75b, PH73, Tom74, VD76b, Van75]. Piecewise-Linear [Tom74]. Pin [LR71a, Rus72]. Pipeline [Coo77, Dev75, GW70, KSA74, Maj76]. Pipelined [BD77, Li77, Mor79]. Pipelines [Roe78]. Pipelining [HF72, JA78, KV73]. Pixel [PR78]. PL [Jor74b, Rem72, Wex71]. PL/I [Jor74b, Rem72, Wex71]. PLA [OH79]. Placement [Bae74, HF74]. Plan [Ano76c, Ano76d, EUY^{+}72]. Planar [Ano72-39, CA77, Par71]. Planarity [Rub75a]. Plane [CY76, NW79, PH74, Sam70b, ZR72]. Plateaus [BDR77]. Plotter [Ath75]. Plotting [Wri73]. PLZT [BL75]. Pneumoconiosis [JP76]. Point [Bau77, Boh77, Bre73b, Cod73, HW76, HF74, Hig76, Ken77a, KKK74, KPS77, Mat70, MMB78, Mor71, O’C75, Pak71, RB78, SV75, SR77, Vic71, Wei75, Yoh73, Yue75c, NA76b]. Points [FD73, LSB77, OC76, Sta74, WA73]. Poisson [Buz73, Chu70, MH70b, MH70c, MN78, Sch73]. Polarized [MT74a]. Poles [God71]. Policies [BB78b]. Policy [Ano73e]. Polish [Hus73]. POLLY [Bec71]. Polygonal [AD75, Pav77]. Polygons [FP75, Sch78a, SCH72]. Polyhedra [Lou70]. Polylinear [Cur71b, HSP78]. Polynomial [BB74b, Gri78a, Lee78a, MT74a, MC70, Pav75b, PH73, Rem74, Tou72d]. Polynomially [IS75]. Polynomials [Has74, Mar73d, VD76b]. Popov [Ano72n]. Positive [HG79]. Post [Eps73a, NSR74]. Postal [Sch78c]. Postcorrection [BJ75, Rhy71]. Postcorrections [Chi72]. Posteriori [TH79]. Postfix [Ber77]. Postprocessing [Dos77, RH74]. Potential [RF72]. Power [Set70b, WGS78, YI74, Yue74b]. pp [Tab70b]. Practical [TBY78]. Practice [AGM^{+}71, Bre71c]. Practices [Rux71]. Pragmatic [Mil74a]. Precisely [SP72a]. Precision [Bre73b, HC70, MM73a]. Predicate [JW76, San76]. Predicate-Calculus [San76]. Prediction [Boo70, Toy71b]. Preemptive [Gon79, VN78]. Preface [Ano71-50, Ano72-71, Ano76-27, AG73, Shi70]. Preliminary [Dub79, Koh75, Lar71, Pin71]. Prepaging [Tri76]. Preprocess [Uhr72]. Preprocessing [Gra72, HKD^{+}71, Kuc70]. Presence [Ast78, CM79c, Mar78]. Preserved [WL74]. Preserving [Yan75b]. Preset [FT76]. Press [Tab70b]. Previous [Hen71]. Prime [Ben71, BM70, BH71c, Bub72, Das71c, DK72, DBC71, HW75, Hwa74, IR76, Maj72c, Mor70c, NRS74, Pap77b, Reu75, SCL70]. Primer [Huf76]. Primitives [Kli71, MBSS75]. Principally [BF75a]. Principle [Koh74]. Principles [Bib74, New75a, Rux71]. Print [Gra72]. Printed [AKM^{+}71, AN74, SO73]. Printer [Ano70y, Mac70]. Printing [Ern72, Yue73b]. Priori [Gri73]. Privacy [TW76]. Probabilistic [AA75, Hu75, Kri73, MH76, MH78, MB77, PM75b, Paz70a, Tho76, Vil73b, Yoe70a, de 71]. Probabilistically [FK78a, FK78c]. Probabilities [MPS78, PM75a, PM78a, SD79]. Probability [BT73, Dui76, Hus74, Ogu75, Tou71]. Problem [Ake74, Bro75a, CH72a, CH73b, DTI78, Du72, FL77a, Fis76, FS78, GHJ77, Geo70, GLL79, Har70a, Kau74a, Kur74, Lou70, Nam70, Nie70b, PS71, PR74a, PHG77, RA78, Reu75, RC74, Sch75, She70a, Sri76, SIF76, VD76a, Wak70c, Wex71, Wri73]. Problem-Oriented [Fis76, RA78]. Problems [Bar74c, CF78a, CCF79, Gar72b, Has74, IS75, Kar70, Lam79a, Lee75, MOW76, MG71a, Par72b, Sto74a, Tab70b, Tom73, Vic71, WK73, YI74]. Procedure [ARS71a, BT70a, Chu71, DS72a, Dut74, EH78, Fis76, Gel76, HTD72, KP74, Kor79, LF72, MT71, Noe73, NR74, Red72a, Rob70b, ROW71, TG78, Tou72b]. Procedures [BK72, BH74b, Cur70, FG76, Gol72, Hla78, RC72, Rob70a, SKM78]. Proceedings [Ano79b, Ano79-54]. Process [Mai71, REFN76, Tab70a]. Processable [RPL73, RFL76, RL78]. Processed [CB72]. Processes [MVW79, SK71]. Processing [AP76a, Ano78v, Ano78w, Ano78x, AL71, BG75, CCF79, CDL78, Cor71b, DRS75, Far75, Fen77, FW78, FR76, GSJ^{+}76, GR71b, HF71, Hop71, HA70, HB75b, Koc75, KSW74, KF73, Kru73, LT73, Li77, MJTW75, ND73, PC72, RNR75, RTKH77, Roe78, Sch78b, Sch74b, SC77, Sta75a, Sta75b, Sto71, Svo73, SH79, Tab74, TP73, Tun70a, TE76, VM78]. Processor [Aus70, BCKT79, Cap76, CS75, CR78, CDS73, CK75, CK79, CW77, CL72, DS70, FL77b, GLMR71, Guf71, HND75, HSS78, Law75, LS71, MM78b, Mil74b, Nut77, RSH79, Red71a, Red79, RSL76, SH77, SNOS79, Sie77, SN79, Ste77, Tho78]. Processors [AT70, BK79, BT76, BD77, Fen74, Fen77, FB73a, GC77, Gri78b, Kri77, Lan76, McL72, NR75, PR72, Red71a, Ros74a, SB77b, Swa79b]. Product [AP76a, BM74, Bla79, Dru71b, Mar73a, SGR78, Yue75a]. Productions [Jol76]. Products [AB78, God73, HLD70, Oja79, Pap79, Wak70b]. Professionals [Ano77d, Ano78k]. Program [Ano78c, Ano78f, Ano78d, Ano78e, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Bae74, Cov74, EH76, FLR78, GR71b, How75, Kje72, LM78, Rau79b, Rej72, Sch78b, Wul74]. Programmable [AB78, Fle79, GSB^{+}76, GLMR71, HND75, Kam79, KAN79, NR76a, OH79, PM77a, PW79b, Smi79c, Woo79]. Programmed [Fre74a, Rot78]. Programming [BS77, BW76b, Bre71c, Coo72, Den74, Fel79, FW78, Geo70, Gri74, Jor74a, Jor74b, Ker73, MI72, Pap77a, Rem72, RH73, Sil73, Uhr73, Voi75, Weg76, Wex71]. Programs [Bas71b, BBP75, CR72a, Cha72, Kar77a, Kin71a, KMC72, Lam79b, MOW76, PB76a, San76]. Progress [SG71]. Progressive [ST79]. Projection [AM79b, SC77, The75, FT74]. Projections [Cal70, CC73, KM75a, SF78, Wan75c]. Projectors [Oja79]. Proof [Gel76, Pra71a, Rub75c, Vil73a]. Proofs [Moo71]. Propagation [MM73a, RHR76a, RHR76b, SW70a, SW70b]. Properties [BM73, Bre71b, CL71, CF79, Dav71, Fis74, Gal78, GR72a, Gra71, GM76, HN78, Hay76b, KK79b, Kri73, Mar73a, MG71b, Pra78a, RGK72, SC77, Sun79, TD77, The75, Tou72a, Wan75b, Wea75]. Property [DF70, HN71, Hwa72, SF77, She78b, SB72, WG74]. Proposed [KK70]. Protection [FSLC78]. Protocol [WN76]. Prototype [EUY^{+}72]. Prototypes [Cha74b]. Prover [Ern76]. Provided [CF70b]. Proving [Abd71, Fis76, Kin71a, MOW76, Mor76b, Rei76]. Pseudo [Kli74, PW79a, Pap77a]. Pseudo-Boolean [Kli74, Pap77a]. Pseudo-Random [PW79a]. Pseudodivision [SK71]. Pseudoinverse [PD77]. Pseudonoise [Hur74, SP72b, SP72c]. Pseudorandom [Dav71, MH70b, MH70c, Mar73b, Mar73c, MAB78, NM76b, NM76a]. Pseudothreshold [Bau71]. Publications [Ano72-57, Ano72-58, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78-34, Ano78-35, Ano78-36, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42]. Publications [Ano79-43, Ano79-44]. Published [Ano76v, Ano76z, Ano76y, Ano76w, Ano76x, Ano77y, Ano77x, Ano77w, Ano77-28, Ano77-27, Ano77z]. Pulmonary [SH72]. Pulse [BH71b, FG71b, FG71c, GD72, Hor71b, Kon71, Obe72, RHR76a, RHR76b, SW70a, SW70b]. Pulse-Ratio [Kon71]. Pulses [MH70b, MH70c]. Pure [Myk75]. Purging [Los76]. Purpose [DRS75, GGMM71, HC70, SF74]. Purposes [Ben70]. pursuit [FT74]. Push [Har74b]. Push-Down [Har74b]. Pushdown [Arn70, Kam70a].

Quadratic [CY76]. Quadric [Mah72, SF78]. Qualified [Win76]. Quality [FLR78]. Quantitative [DB74, DF78, FD75]. Quantities [Mat71a]. Quantization [CB72, CJ74]. Quantized [BD75a]. Quarter [JB76a]. Quarterly [Ano78h, Ano78i, Ano78j, Ano79k, Ano79l]. Quasi [MG75, Swa73b]. Quasi-Serial [MG75, Swa73b]. Question [BCSW76]. Question-Answering [BCSW76]. Queue [Sch79b]. Queueing [Reg73]. Quinary [PW79a]. Quotients [HLD70].

R [AH76, Wak70c]. R-70-31 [Wak70c]. R70 [Ake70, Arb70, Arn70, Bar70, Ben70, Ber70a, Brz70, Cha70, Cut70, Dav70a, Fri70, Geo70, Gil70b, Goo70, Gra70, Har70b, Hed70, Hop70, Jum70, Kam70a, Kam70b, Kar70, Law70, Lev70, New70, Pag70a, Paz70b, Paz70a, Pre70, Rob70b, Rot70, She70a, She70b, She70c, Skl70, Su70, Tsi70, Tur70, Wag70, Win70a, Yam70, Yeh70, Yoe70a, You70]. R70-1 [Har70b]. R70-10 [Arb70]. R70-11 [Yam70]. R70-12 [Rot70]. R70-13 [Win70a]. R70-14 [Cut70]. R70-15 [She70b]. R70-16 [Skl70]. R70-17 [Fri70]. R70-18 [Goo70]. R70-19 [New70]. R70-2 [Hop70]. R70-20 [She70c]. R70-21 [Arn70]. R70-22 [Tur70]. R70-23 [Kam70a]. R70-24 [Rob70b]. R70-25 [Ben70]. R70-26 [Yoe70a]. R70-27 [Ber70a]. R70-28 [Pre70]. R70-29 [Jum70]. R70-3 [Pag70a]. R70-30 [Su70]. R70-32 [Wag70]. R70-33 [Paz70a]. R70-34 [Bar70]. R70-35 [Tsi70]. R70-36 [Hed70]. R70-37 [She70a]. R70-38 [Lev70]. R70-39 [Kam70b]. R70-4 [Geo70]. R70-40 [Ake70]. R70-41 [Cha70]. R70-42 [Dav70a]. R70-43 [Yeh70]. R70-44 [Brz70]. R70-45 [Paz70b]. R70-5 [Gra70]. R70-6 [Kar70]. R70-7 [Gil70b]. R70-8 [You70]. R70-9 [Law70]. Races [BH71b, Bre74]. Rademacher [Yue76, Edw75]. Radiation [BG75]. Radiographic [HKD^{+}71, HE73]. Radiographs [ADL72, BS76]. Radix [Atk70, BR71, CSY75, Dru71a, DV78a, Erc73, Kri71a, Nic71, Paw71, Wad71, Zoh70, Zoh73d, Zoh76b, Bak75a]. Radix- [BR71, CSY75]. Radix-16 [Erc73]. radix-2 [Bak75a]. Rado [Lyn72]. Rahman [Bha78]. RAM [Zoh76a]. Random [AA75, Agr78b, AU74, BI72, Bre71d, DB76, EF76, En77, GD72, Hay75, Kau73, KH77b, KH77c, Len77, Los78, MS77b, MVW79, Mur70, NTA78, Nai79, Paa79, PW79a, Sch79a, SLC75, Sok72, Sri78, Sti78a, Sto75, Wil78, Yue77e]. Random-Access [Hay75, NTA78, Sri78, Sti78a]. Random-Error-Correction [En77, Paa79]. Random-Pulse [GD72]. Randomized [LS72]. Randomly [BG75]. Range [Kri70b, Kri71b, MH75, Sik70]. Range-Transformation [Kri70b, Kri71b].

Rank [Cur73b, GH78, SK74]. Rank-Augmented [SK74]. RAP.2 [SNOS79]. Rapid [AU74, MC70, Roc70]. Raster [HSS78, ST79]. Rate [CL75, MH70a, MR72, NR77b, Obe72, Sok72]. Rates [Nad71]. Ratio [Hus74, Kon71, Rus72]. Rational [SB77b, VV74]. Ratios [Eas78]. Ray [Wei75]. Reach [Ano79b]. Read [BB75, Mon74, Pra78b, SY75, GM70]. Read-Only [BB75, Mon74, Pra78b, SY75, GM70]. Readily [Dev79, Ehr76, Kau70c]. Reading [AKM^{+}71, JM74, Sch78c]. Readings [Fly72a]. Real [AK72, Dep70, Dow73, Gay78, GSJ^{+}76, Goo70, Goy76, HO79, Hen78, HA70, Koc75, Mar78, Pol71, Saw77, Sil77, WKD70]. Real-Time [AK72, Dep70, Dow73, Gay78, GSJ^{+}76, Goo70, Goy76, HA70, Koc75, Mar78, Pol71, Saw77, WKD70]. Realizability [BPC75]. Realizable [Dav70b, Gel71, HE79, Mat71b]. Realization [AA71, CL75, Cur71b, DSBC74, DGC74, DVS74, EGG71, FT76, GH78, GC73, GCC74, GCP^{+}77, HG79, Har70a, HC70, Hen71, Hwa73, KT72, Mag71, MT74b, OY70, PC78, SK78a, SK78b, She70b, Sin72, Sri75, Sri79, TOS71, Zoh76a]. Realizations [Bro75b, Bro75c, CC72, GR71a, GM76, HC72, Hay71b, Kod74, Lan78, MBSS75, OW73b, PM77a, Red72b, RW74, RT73b, Saw74, Ung77c, Wil78, Zei71, Zoh73c]. Realize [PM71]. Realized [ATN70, But75, Hu73, LF74, Paz70a]. Realizers [Han72]. Realizing [Cur70, OW72]. Reasoning [Mam77]. Rebalancing [LP78]. Rebuttal [DM76]. Received [BG75]. Reciprocal [Ric74b]. Recoding [Rob70a]. Recognition [ARS71a, AA77, Ano79c, Ano79d, BD71, BB78a, Bre73a, BP72, BR74, Cal70, Cha73a, Che71a, Chi71, CR72d, Chu71, Chu74, Dev74, DT70, Dos77, Dut74, EK75, Far79, FE78, FP75, FB73b, FR76, Gil74b, Gil74a, Gra72, Har70b, HF71, Hus72, HTD72, JK77, KM78, Koz71, LF77, LF78, Mak71, Mat71b, MA77, Mic74, MF76c, MG71b, NC78, Nay71, PM76a, Pol71, RPL73, REFN76, Rem74, RE71, Sch78c, Sel74, STG71, SN72b, Sto74a, Tak76, TG75, Tjo77, TG72b, Tou71, Tou72a, Tou72b, Tou72c, Tou75, Uhr72, Ull72, Vil72, VV71, WG73, Wea75, Win75]. Recognizer [Nad71]. Recognizing [BS76, CY76, TD70]. Reconfigurable [Mai71]. Reconfiguration [GT74]. Reconstruction [CC73, DB73, KM75a, Wan77, WD77, YRT75]. Reconstructors [JM71]. Record [Hsi71a]. Recorder [Ros71]. Recording [BL75, Hoa76]. Recovery [CJW^{+}71, CR72a, LL79]. Rectangular [Ake72, Skl72]. Recurrence [CK75, KS73, RB72]. Recurrent [Kar77a]. Recursion [TV76]. Recursive [AM79b, AOR74, Bro77, Fri77, Jai77c, Maj72b, Mor70c, NC78, Whi70, Zoh73a, NA76b]. Redefinition [RA78]. Reduced [Bro70b]. Reducing [Ano75-28, HS70]. Reduction [Ben71, Cur71a, JA74, KLCC77, Pfl73, Pin71, PMB77, VVA78]. Redundancy [AS74, AB79a, Atk70, DR74, DM78a, FM72a, Hey71, IS76, Kla71, KS79, LR72, LD74, Los76, Ogu74, SM73a, SM73b, Sie75, Smi78, Smi79b, Wak75c, YT71a]. Redundant [Abr75, BM73, Dah73, DW78, FD74a, FD77b, Lyo70, NA76a, Si78, To73, Toi76, YL73]. Reed [KS75, KS77b, PP75, SR75, SO79a, Swa72a]. Referees [Ano71-52, Ano78s, Ano78a, Ano79-52]. Refinement [ST79]. Refinements [WM76]. Reflected [Agr76, Obe74]. Reflection [Dav72b, Hor71b]. Refutations [And76]. Region [CW75]. Register [Bar75a, BEGW72, HC78, HP75, JW73, Kar76, KK78, Klu77, LF74, Mar73a, Mar73b, MAB78, Max77, Myk75, NR77b, Per70, RT73b, Smi71a, Van73a]. Register-Transfer [BEGW72]. Registering [HEWB72]. Registerless [Mil74b]. Registers [AM79c, Hur74, Lem70, Man71, O’K73]. Registration [ACJ75, BS72]. Regular [Ber70a, HPP79, PM78a, Pat71, Paz70b, Paz70c]. Reject [Nad71]. Related [Bea78]. Relational [De 79, EF76, Kam70b]. Relations [Kli74]. Relationship [DM75, DM76, Eps73b, Goo71, LR71a, MR72, SS72, Tak76]. Relaxation [Cha71, DR77, ZHR77a, ZHR77b]. Reliability [AS74, Abr75, Arn73, Bea78, BF75b, BCJ^{+}71, BCH^{+}73, CLL78, FD74a, HB75a, IS76, IA76, Kor79, KS79, Mal73, Mar77, Mat71c, Mey76, Par72b, RH75a, Sie75, Sti78b, You71]. Reliable [Ano79-54, Hu75, Mai71, Saw75, WSB79]. Remark [Gre75, Vai71]. Remarks [Noe73, OW73a]. Remote [Arm73, PRD71]. Removal [MH70a]. Repair [CS79, GT74, RK75a, RK75b, YO70]. Repairable [Arn73]. Repairing [AGM^{+}71]. Repeatable [MH70b, MH70c]. Repeated [Bor72]. Repetitive [Brz70, Rey70]. Rephrased [Nie70b]. Replacement [EF79, Gel73b, SM75c, TI72b]. Replication [Cov74, LR72]. Reply [Agr79b, All77, Ano73z, Bab74, Bar75b, Bib75, BH74a, CM79a, Che72, CBD70, Chu72, CD75, Coh70, Das72, DM76, Edw76, Ekl73, En79, FCC72, FO71b, GRR74, GR73a, GM72b, HC74, HW78, HF70, KAA77, Koh72a, Koo78, LM71a, Mar75, Mar76, MB79, Men70a, PR75, Pra71b, Rao78, RS73a, RS73b, SM79, SS75b, To76, Yue77a, Yue77b, Zoh71]. Report [Fre74b, Hob73, Kra74, Lom74, Mui74, Pla73]. Reporting [BO79]. Representation [Bau77, Das71b, EF76, FE73, HC73, KR73, KN70, KS77c, Lee78a, Mor74, Mor71, NN73, OF73, SM72b, Sel74, SB74a, TL73a, TF73, Tou72d, Yam70, Yue75c, Yue77c]. Representations [GB71a, GB71b, Pea76]. Representative [Bau72]. Represented [Glu71]. Required [Lev70, Nie70b, PM71]. Requirements [FLBK76]. Requiring [Hay71b]. Research [BBLT73]. Research-Oriented [BBLT73]. Resident [Gar72a]. Resident-Bubble [Gar72a]. Residue [BB72, Ban74b, Ban74c, BM73, BM74, BM78, HC73, Jul78, KKK73, KKK74, Man72b, O’K75, RT70, TJM79, YL73]. Resistor [Pak71]. Resolution [AM77, BK72, DI71, Gel76, Hen76a, Hen76b, Lan77, LS76b, MN70, Pic74, WM76, YK76]. Resolution-Based [Gel76]. Resolution-Oriented [AM77]. Resolver [MS77a]. Resolvers [And74]. Resource [Mil74b, PM78b, SN79, Ung77a]. Responders [FS71]. Response [BV78, FS74, Lan77, Pec76, Sah74, Sah77]. Restoration [HA77, Hun73, Hun77, JA74, Jai77a, Jai77b, PD77, TH79]. Restricted [But76, FM73, KS78b]. Restrictions [KN72]. Restructurable [RF78]. Result [DW72a]. Resulting [KMC72]. Results [ABK76, Bri73, CYNM79, Cut70, DW73b, EZS78, HTD72, Lyn72, SR77, Ung77a, YK76, Yue74a]. Retrieval [CM73b, RTW78, Ste77, WG74]. Retrieving [Wei75]. Retry [She78a]. Reversal [PBC74]. Review [Ano72u, Ano72v, Ano72w, Bas73, Flo75, New75a]. Reviewer [Bar75b, Bib75]. Reviewers [Ano76a]. Reviews [Ano72x, Ano72y, Ano72z, Ano73b, Ano73c, Ano73d, Ano73h, Shi75, Ano72-50]. Revision [Hil79]. Riccati [Gou77, Den76, SG74]. Right [Pat71, She71]. Right-Linear [Pat71]. Rings [DV78b]. Risk [Dev74, Tou75]. Robot [EUY^{+}72]. Robust [Yue75a]. Rollback [CR72a, She78b]. ROM [Jul78]. Root [SK71]. Rooting [RGK72]. Roots [MK71a, Maj72a, Rud76]. Rotary [JK77]. Rotate [NT70]. Rotating [FV79]. Rotation [EO78]. Rotations [And71]. Rounding [KPS77, Zoh76b]. Roundings [Yoh73]. Router [AB79b]. Routing [Abe72, AKT^{+}77, CG74, HW79, KJ79, Lam79a, Sah74, Sah77, SH79]. RST [Sur70b]. RTM [BEGW72]. Rule [Ehr76, Fri77, HS70, KK78]. Rulers [Rob79b]. Runs [Rom78].

s [GSB^{+}76, AH76]. Safe [Chu76, CD78, DGC74, DAA79, GCP^{+}77, HG79, MT74b, PM74, SM75b, TM71, TI72a, TOS71, Toh74, WM74]. sample [Vee78]. Sampled [All75b, ASS72, DB73, SSR72, Wil77b]. Sampled-Data [SSR72]. Samples [Bab72, BF72, PS78, Tou74]. Sampling [Sok72]. Satellite [HEWB72, MN78]. Satisfying [KK78]. Save [Pra74]. Scale [CM70, GC71b, Mor74, VV74, WA73]. Scaling [Jul78, SM71c]. Scan [AN74, HB75b]. Scan-Digitized [AN74]. Scanner [HND75]. Scanner-Processor [HND75]. Scanning [Bec71, EF79, Stu75, Wan77]. Scene [Pav74, RT71, WH78]. Scenes [Gri73]. Schedules [FB73a, Gon79]. Scheduling [AI78, Ful72, Kau74a, Koh75, Li77, RCG72, RFL76, RL78, Sto70a, VD76a]. Schemata [Tho75]. Scheme [AKT^{+}77, Che71b, Kla71, LM70, Loo70, Los76, MF76a, SKG77, Toy71b, WSB79, WL74]. Schemes [BT70b, DF78, FD75, Hus72, Kri70a, Kri71b, Ple77, Pre78a]. Science [Fre74a, Jor74b, Ram76a]. SCIRTSS [HH77]. SCR [WS76]. Script [EK75, Far79, Moo76]. Scroll [Wil70]. Sea [Whi71a]. Search [BI72, Bur74, Cha74d, Dig73, Fis76, Gel76, HH77, LNL72, LS72, LC79, Pag70b, Pat72, Sic76, TS79, WM76, WC74]. Searching [Gri76, RC71b]. Second [JM74, RS71b, AG73]. Second-Order [RS71b]. Section [Ano72-40, Fle79, Wan77]. Sectioning [Hun72]. Sections [Ano76-28]. Secure [Pra78a, SM78]. Security [FSLC78, TW76]. Seeking [KF72c]. Segment [Sta74]. Segmentation [DR77, PR78, PF72, Pav73, PH74, RC71a, Wil71, ZRD75]. Segmented [Sah77]. Segments [Ros74b]. Seismic [SF72]. Seismology [Dav72b]. Selcuk [Sah74, Sah77]. Selected [Ano78v, Ano78w, Ano78x]. Selecting [Chi71, Rob70b]. Selection [ARS71a, And71, Che71a, Das71a, Deu71, FK70a, Kit77, LD75, MT70, MSJ72, NF77, SW76, TV72, Tou72c, WNR74, Whi71b]. Selector [Hol72b]. Self [Ama72, AM73, AR77, BCH^{+}73, CDJ71, CDS73, CHSS73, CS79, CSST73, Dav78, DAA79, Dor72, HA76, Los76, Mai71, MF78, OW73b, Red74, RV74, She78b, Ung77b, Usa75, Wak74, AGM^{+}71]. Self-Checked [OW73b]. Self-Checking [AM73, AR77, BCH^{+}73, CDS73, CHSS73, CSST73, Dav78, DAA79, Dor72, MF78, Red74, She78b, Usa75, Wak74]. Self-Diagnosable [CS79]. Self-Organizing [Ama72]. Self-Purging [Los76]. Self-Reconfigurable [Mai71]. Self-Synchronization [HA76]. Self-Synchronized [RV74]. Self-Synchronizing [Ung77b]. Self-Testing [CDJ71, AGM^{+}71]. Semantic [BCSW76, Hen76b, Wal76]. Semantically [Rei76]. Semantics [JW76, Ros74a]. Semi [Sar78]. Semi-Fast [Sar78]. Semicausal [Jai77c]. Semiconductor [BD77, NTA78, Sri78, Sun78]. Seniority [Osb77]. Sensitive [BH73, Hay75, SR71]. Sensitization [CRB72]. Sensitized [KK79b]. Sentences [Hub76]. Separable [AA71, AR77, CD78]. Separating [Sto77]. September [Ano77-28]. Sequence [Kri73, Sur70a]. Sequences [Ama72, AM79c, CL71, DDSM76, Dav71, FM73, Gal78, Hab77, HC78, HH77, JW73, LE71, Mar71, Mar73a, MH70b, MH70c, MAB78, MST^{+}75, MBM77, PW79a, Qua74, Smi71a, Sut74, TB75]. Sequency [Bro77, CL77, GC77, KH74, Red73b, SO79b]. Sequency-Ordered [Bro77]. Sequential [Abd71, AB79a, Agr75b, AU74, Alm78, Ano75-28, ATN70, Bei70, Bei71, BB78b, Ben71, Bis74, Bou74b, BH71b, Bre71d, Che74, CW79b, Chi71, CD78, Cio78, Cur70, Cur71b, Cur73b, Cur73a, DF75, DTG79, Dav70a, Dav70b, DSBC74, DGC74, DAA79, DW72b, Du71, DH79, Ehr72, EGG72, Far73, Fel79, Fre79b, Fri70, FM73, FG71b, FG71c, FK74, FNSK75, GH78, Gel71, Ges70, GB71a, GB71b, GD73, GCC74, GCP^{+}77, HD71, HSP78, HG79, Har72, HC71a, HC72, HH77, Hol72a, HC71b, Hsi71b, Hu72, Hu73, HA76, Hus72, Hus74, Hwa73, Iba74, Jos78, Jum70, Kai70, KYO70, Kar74, KM73, Kel70, Kel71, KS77a, KN72, Kin70, Kin71c, KK77b, KR78a, KC73, LR72, LV75, LSB77, Len77, Mag71, Mag73, MT70, MT71, MS73a, MS74, Man72a, Mar70, Mar77, Mey71, MVW79]. Sequential [MJ77, MJ78, MT74b, MKO70, NT79, NA72, Nie70a, Ols70, OS76, PM77a, Par72a, Par72b, PM78a, PM74, Paz70c, PHG77, PD79, Pre70, Ram76b, RB75, RB76, RV74, RN77, RT73b, Sau72a, Sau72b, SM74, Saw74, Saw75, SM75b, SR73, Ser73, Ser75, SM76, She77, She70b, She70c, SV74, SY75, Sil79, Sin71, SR71, Smi74, SN72b, Sto74b, Sto75, SHR72, Sun76, SB74a, Sur78, TI72a, Tan71, TA71, TOS71, Toh74, Tom72, TB74, Ulr71, Ung71, Vai71, Voi77a, Wak75a, Wak75c, WY75, Wan71, WM74, Wil78, WZ72, Wil75, WH78, Yan73, Yan74b, YB75, Yua74, Vee78]. Serial [BK71a, Hor71a, IC78, Lit73, MG75, Rhy70, Sit74, Swa73b, Tro75, Wil77a]. Series [Ric74b, Yue75e]. Set [AM79a, Bub72, DGC74, DW72a, FS75, GHJ77, GF78, Lig77, MBSS75, Moo71, Nie75, NRS74, RS71a, Set70b, SF78, Smi76, WGS78, Yan74b]. Sets [AM79b, Ake73, Bet71, Bou74a, Dan73, Das73, DDSM76, DBC71, GY77, GL70, Gre75, HG78a, Hen76b, KS71, Mar71, MST^{+}75, Pra78d, Red73a, SR75, Sch78a, SC77, Wan75a, YT71c, de 71]. Setting [OS76]. Seven [MG71b, Tou72a]. Shading [Gou71, LH79]. Shape [Dav77b, UM77]. Shaped [Oza78]. Share [PRD71]. Shared [Gad70, JP73, Kle70, Mil74b, OW73b, PM78b]. Shared-Resource [Mil74b]. Sharing [Gel73a]. Shift [CF78b, Dav74, Eks73, HC78, Hur74, JW73, Kar76, KK78, Klu77, Lem70, LF74, Mar73a, Mar73b, MAB78, Myk75, NR77b, NT70, Per70, RT73b, She71, Smi71a, TG72a, Van73a]. Shift-Invariant [Eks73]. Shift-Register [JW73, Kar76, KK78, Klu77, LF74, Mar73a, Mar73b, Smi71a]. Shifts [RRNA74]. Ship [SW71b]. Shop [Gon79]. Short [Bak79, Cur79, Fri74]. Short-Circuit [Fri74]. Short-Cut [Cur79]. Short-Word-Length [Bak79]. Shortened [Sun79]. Should [SO73]. SHR [Noe73, NR74]. SHR-Optimal [Noe73, NR74]. Shuffle [Lan76, LS76a, Sto71]. Shuffle-Exchange [Lan76, LS76a]. Sides [Dav77b]. Sieve [Mar78]. Sift [DM78a]. Sift-Out [DM78a]. Sigma [Lyn72]. Sign [Chi77, HL86, SA75, SCNS83]. Sign-Flipping [Chi77]. Sign/Logarithm [HL86, SA75, SCNS83]. Signal [CB72, Cor71b, DRS75, GLMR71, HF71, HA70, Kor79, KSW74, NA72, PM75a, Rej72, SH79, Tru76]. Signal-Dependent [CB72]. Signals [All75b, Kuo74, LC72, Mar78, Ung77d, Usa75, Wil77b, Yue74b]. Signature [HF71]. Signed [Agr76, AR78, BBC72b, Chi77, DS72b, MK71b, Tom75, Tun70b]. Signed-Carry [Chi77]. Signed-Digit [Tun70b]. Signed-Number [DS72b]. Significance [BM77b]. Significant [Ges71].

Silhouettes [SCH72, SN72a]. SIMD [Sie77, Sie79]. Similar [Liu77a]. Similarity [JP73, OS76]. Simple [BJ75, BV78, CDJ71, Gar72b, GBMC72, Gri73, HTD72, IR76, Jol76, Maj72c, MBM77, MC70, PC72, Rhy71, SN79, TB75, Von75]. Simpler [FP75]. Simplex [Sil79]. Simplification [KN72, VW72]. Simplified [Car71b, Cur76, HK74, KS72, LM71b, LS76a, Zis76]. Simplify [AB78, HF74]. Simplifying [Das71b, Pin71]. Simply [Has73, SSK72]. Simulated [LF74]. Simulating [Arm72]. Simulation [Ben70, Bre70b, Bre72b, CCES74, Fre77, Gor75, JM71, KP71, KV76, MCRD72, MC78, Mor76a, RHR76a, RHR76b, RC71c, SW70a, SW70b, Sch74b, She70d, SM70b, Stu75, ST76, Ung77a, Vic74, Wai70]. Simulations [RH73]. Simulative [MM73a]. Simulator [BW77, Gra73]. Simultaneous [Has73]. Simultaneously [KMC72]. Sine [SP77, Bak76]. sine/cosine [Bak76]. Single [Bro71, Dev79, FS74, Gil70a, Har70b, Kau70c, KT72, KR78a, Mar79, MS78, NT78, Ric72, Rot70, SS77, WSB79, WL74]. Single-and [FS74]. Single-Channel [Gil70a, Ric72]. Single-Error-Correcting [Dev79, Kau70c, WL74]. Single-Gate [Rot70]. Single-Parameter [Bro71]. Single-Tape [Har70b]. Single-Threshold [KT72]. Sinusoids [HO79, Sil77]. Size [Len77, Moo71]. Sized [Red71a]. Skip [Lig77]. Skip-and-Set [Lig77]. Skit [O’K73]. Slave [FG76, Wil71]. Small [MBSS75, RV73, SB77b, TS79]. Small-Diameter [TS79]. Smaller [KO78]. Smooth [CL71]. Smoothing [Boo70, Dui76, Gol72, KS78a]. Society [Ano72n, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano72-50, Ano72-57, Ano72-58, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74j, Ano74p, Ano74k, Ano74q, Ano74l, Ano74r, Ano74m, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74n, Ano74y, Ano74o, Ano74z, Ano75o, Ano75p, Ano75q, Ano75i, Ano75r, Ano75j, Ano75s, Ano75k, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75l, Ano75y, Ano75m, Ano75z, Ano75n, Ano76c, Ano76d, Ano76h, Ano76k, Ano76l, Ano76i, Ano76m, Ano76j, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano78y, Ano78z, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78-32, Ano78-33, Ano78v, Ano78-34]. Society [Ano78w, Ano78-35, Ano78x, Ano78-36, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79-44, Ano79-53]. Software [Ano78-39, Ano78-40, Ano78-41, Boe76, Gay78, Ham74, Lom74, Pag72, Ano79-54]. Solid [Stu75]. Solid-State [Stu75]. Solution [Abr75, Bak79, BM70, CF78a, CM70, CW77, DTI78, Du72, Eks74, Fra78, GLR71, GLL79, Kar70, Kli74, KS73, Lou70, Maz71, Men70b, MG71a, MR75, MC70, NK74, Nam70, New75b, SF74, Sch73, Ulr71, WG73, Wit70, Wor76, Wri73]. Solutions [Bro70b, Bro71, Hig71, WK73]. Solvable [Has73]. Solve [RS71b]. Solver [Buz73]. Solving [BB74b, FL77a, Max77, MN70, SG74, Sri76, Tom73, Wex71]. Some [ABK76, Bak75b, Bha77, Bha78, Bre71b, Bri73, BB73, Chi72, Chu74, Cut70, DF78, DH79, Fly72b, FK78c, Gar76, HN78, Hoe76, HI72, Hus72, JW76, KF78, Kin71c, Kri73, Lee75, LS76b, Maj76, Mil73, Nay71, New75b, Pag72, RGK72, RS71a, SC77, SR77, Tou71, Tou72e, Ung77a, Vin78, Wea75, WK71, YM78, Yip76, Yue74a, Bak75a]. Sonic [BdB70]. Sort [NS79]. Sorting [BS78, Lam79a, Mor79, Pre78a, Roc70, SM70b, Van72]. Source [Mar73b]. Sources [Boy74]. Soviet [SG71]. Space [But71, Der73, Fre77, Hec73, Hop71, LSB77, PR78, SIF76, Weg72, Wri73].

Space-Efficient [Weg72]. Space-Filling [But71, Fre77]. Spaceborne [WSB79]. Spaceflight [WKD70]. Spaces [BF78]. Spanning [BF78, KLCC77]. Sparing [BF75b]. Sparse [Gri78a]. Spatial [SM72a]. Speaker [Mak71, Moh71]. Special [Ano76-28, Bec71, CH75, DRS75, DH79, Ehr72, Fle79, GGMM71, HC70, SF74]. Special-Purpose [DRS75, GGMM71, HC70]. Specialized [Swa79b]. Specification [Dub79, Mon74]. Specifications [Ano79-54]. Specified [Bis74, DH79, GD73, Hig73, Kel70, Pfl73, RB75, Sch75, TG75, Wil75]. Spectral [AC70, CL71, Hun71]. Spectroscopy [GW75]. Spectrum [WGS78, Yue74b]. Speech [EFLR76, Mak71, REFN76, Wal76]. Speed [Agr79a, CH73a, Chi78, CH72b, Cio78, Cor71b, Gar74, Hou72, JM70, KES72, KG72, Kel74, Ken77b, Kri70a, Kuc70, LE71, Lin70, Lit74, LM70, Loo70, Mar73c, Qua74, RGK72, SW71a, SW76, Sri71a, Sri71b, Ste72, SKG77, Sto70c, SKK79, Tom75]. Speed-Independent [Cio78, Kel74]. Speeding [BBP75, Pag72]. Speedup [KMC72]. Spherical [RB72]. Spline [Dub79, HA77, Lio76]. Spoken [KM78, Pol71]. SPOOF [Cle73, SS75a]. Spread [Eks73, Wyn71]. Squad [Wak70c]. Square [HB75a, JB76a, MK71a, Maj72a, RGK72, Rud76, SK71]. Square-Rooting [RGK72]. Squares [Agr78a, All75b, Gol72, HA77, Hun73, Maj72a, Wil77b]. Squaring [Che71b]. SR [Brz71]. Stability [GW77, Gil70b, SIF76]. Stabilization [Tab70a]. Stable [AH76, HO79, Sil77]. Stack [Hop70]. Stage [Ast78, VR77]. Standard [GSB^{+}76, Hen78, Guf71]. Standby [BF75b]. Standby-Sparing [BF75b]. Standing [Ano76c, Ano76d]. STAR [AGM^{+}71]. STARAN [Bat77]. Stare [DLL^{+}73]. Start [Eas78]. Stastical [Tru76]. State [AH76, Ben71, BF72, Bis74, Bou74a, Cha72, CL75, CW79c, Cur73a, DH79, Ehr72, Fre74b, FG71c, Gec77, Gil70a, Goo70, GD73, Hob73, HB70, KW70, Kel70, Kin71c, Kra74, Kri73, KR78a, LF74, Lom74, MT70, MT71, MSJ72, MS73a, MB77, Moo71, Mui74, NT78, NT79, Nie70a, Noe73, NR74, NR76b, NK75, Par72a, Pfl73, Pla73, Pra78a, RN77, Ric72, Sau72a, Sau72b, Smi74, SHR72, Stu75, Tan71, TB74, Vai72, Wai70, WR77, Yan74b, Yeh70, de 71]. State-Set [Moo71]. State-Variable [Wai70]. States [Bar71, New70, Per70, Tom72, YT76]. Static [BE70b, BH74b, Cod73, HK77, HK78, Hla78, Zis76]. Static-Hazard-Free [HK77, HK78]. Statistical [BW76a, BCL75, EZS78, Mar73a, Moh71]. Statistically [BT72, Hur74, Hus74]. Status [AR74, Ano77b, Ano77c, RT73a]. Step [BH76, DR75, PB76a]. Stochastic [Ban74a, EGG71, FF75, Gel71, LT70, LF72, LF77, Nie70a, Pag70a, Pag70b, Par72b, Paz70c, Paz71, Pre70, Sil79, Tho75, VN78]. Storage [Anz71, Cho76, GG73, Har76, Hoa76, KLCC77, Mcd72, Nel74, PA78, RC74, SM71b]. Store [HG78b, HPP77, JB76b]. Stored [DEA73, DS74, Dru71a, IC78]. Straight [Ada78, Bar78, BLZ75, Byr71, CY76, DT76, Das78, Fre79a, Ros74b, Sta74, SKK79]. Straight-Line [Ada78, Bar78, DT76, Das78, Sta74]. Strategies [CR72a, FD74a, KR71b, RCG72, WM76]. Stream [KV73, Mil74b, Rub79]. Strings [Wil70]. Strongly [SM78]. Structural [RK71, Yan75b]. Structure [AC71, Bas71b, BR78, Cho74, DS72b, DW73a, Kru71, MN70, NM76a, Weg72, Whi72]. Structured [BS76, BS75, GG73, KD79, Pat76]. Structures [BM77a, Ben77, Ber71, BW76b, Duk74, ET75, Elm76, FE73, Fly72a, Klu77, KC73, Law70, Les71, MAB78, Pat72, Ram73a, Rei74a, RS71a, TJM79, Wan72]. Strum [Pat76]. STT [FG71c, NT79]. Stuck [BCS76, Ios78, KH77c, KR78b, Mei74, Nai79, SS77, Sun79, Wei72]. Stuck-At [Ios78, Mei74, KH77c, Nai79, Sun79]. Stuck-Fault [Wei72]. Stuck-Faults [KR78b]. Stuck-Type [BCS76, SS77]. Students [Fre74a]. Studies [Avi71, Fel79, Nay71, Ros74a, WBRG76]. Study [BCL75, Cha73b, CDL78, DF78, Far75, FL77a, HN73, Ken77a, MM73a, MH70b, Nel74, Red71a, Smi79a, Stu75, WM76]. Style [Gri74]. Stylus [BdB70]. Subcodes [Rao72]. Subdirect [Tur70]. Subfunctions [Reu75]. Subject [Ano71-51]. Submatrices [Ban78]. Suboptimal [Hab77]. Subresultants [Gri78a]. Subroutine [Roc70]. Subroutines [KR71a]. Subscribe [Ano76t, Ano77u, Ano78-42, Ano78-43, Ano79-47, Ano79-48, Ano79-49, Ano79-50]. Subscription [Ano76c, Ano76d]. Subset [NF77]. Subsets [Chi71, MG71b, Rub75b, Tou72a]. Subspace [The75]. substitution [KD79]. substitution-permutation [KD79]. Subsystem [Mal73]. Subtraction [Ban74b]. Successive [JW73]. Sufficient [Smi79d]. Suggested [Hen78]. Suggestion [Bak76, Ste72]. Suitability [GR71b]. Sum [AB78, KO73, Pap79]. Sum-of-Products [AB78]. Summability [GBMC72, Hwa72]. Summation [HK74]. Sums [Cur77, JW73]. Supervised [CF70b]. Support [FSLC78, KS77c, Wor79]. Supporting [De 79]. Surface [Ges71, Mah72]. Surfaces [Gou71, OF73, SF78]. Survey [Age76, Ano77g, Bas73, Gar76, HKD^{+}71]. Survivable [GLW74]. Swap [Kle70]. Swap-Time [Kle70]. Switch [Ast78, IS76, Ogu74, SM73a, SM73b]. Switched [CG74]. Switching [AA71, Bei70, Bei71, BR76a, BR78, BV79, Bro75b, Bro75c, CM77, CDS73, CHSS73, CYNM79, Das71b, DS71, DK72, DW73a, DM75, DM76, Eps73b, HPP79, Han72, Hen71, HE79, Iba71, Kau70a, Kin70, Kin72, Kje72, KS77b, LH76, MN78, Mat71b, Maz71, MS70, MI72, OW72, OY70, Pap77b, Pra78c, Red72a, RNMP77, Sch75, SD74, SM75c, SM70a, SMW71, Smi71b, Spi71, Su70, SC72c, SS72, Sur75, TD73, TD77, Tha78, Tis71, Tom70, UB74, Ung71, VLS70, VW72, Wes78, YT71a, YT71b, ZKK79]. Symbiosis [Arb76]. Symbol [MR72]. Symbolic [Hol72b]. Symmetric [Bis70, Bis73, Bor72, Dah73, HS75, KKK73, KKK74, LL72, LH76, MM78b, RH75b, SK77a, SK77b, YT71b]. Symmetries [EH78]. Symmetry [DS71, KD73, TV79, YT71a]. Symposium [AG73]. Synchronization [Brz70, DW78, GS79, HA76, SH77]. Synchronized [RV74]. Synchronizer [CM73a, Cha79, FD79, Wor77, Wor79]. Synchronizers [CW75, EW77, Pec76]. Synchronizing [LM70, Loo70, Ung77b]. Synchronous [Ano75-28, Chu70, Cur70, Cur73a, DSBC74, FG76, HSP78, HG79, Hu72, Hwa73, LF74, MF76b, NA72, PM77a, Saw75, Ser75, SHR72]. Syntactic [FP75, FB73b, FF75, JK77, KM78, LF78, TG75, Wal76]. Syntax [CR72d, LF72, LF77, Tho75]. Syntax-Directed [CR72d, Tho75]. Synthesis [AAK74, Bar70, BF72, BBP75, BH73, Bre75, CM74, CMS79, CCB70, CK77, CD73, Col77, CYNM79, Dav70a, Edw75, EH78, Fra72, FG76, GGMM71, Ges70, GG75, GL74, Hu72, IM71, Ish77b, Jum70, KH77a, Kje72, Kol72, LV75, Liu75, Liu77a, Liu77b, Lyo70, Mam77, MS77a, Nak79, Pai73, Pap78, Pap72, SAPR70, SA70, SR72b, SR71, SN71, Sur70a, Sur78, Tho74, UB74, Win70a, Yen79, Yoe70b, Yue76]. Synthesizing [KO78]. System [AA72a, Ake76, AKT75, Anz71, Arn73, AKM^{+}71, Avi71, Bar74b, Bas71b, BCA^{+}75, Bha77, BR76b, BdB70, CM76, Cas73, CR78, CDS73, CHSS73, CK79, CW77, CF70a, Cor77, Den76, Dep70, Dos77, DLL^{+}73, Eps73a, EFLR76, Fer77, FSLC78, FK78b, FT79, GSB^{+}76, GSJ^{+}76, GR72b, Gou77, HW76, HP73, HH77, Joh71, KK79a, KF79, Kim70, Kim79, Kin72, KKK73, KKK74, Koc71, Kor76, KF73, Kuc70, LNL72, LTE77, LE77, LE79, LL79, LH79, Low73, Mak71, Mal72, MMB78, Min75, MF76c, O’K73, Pat76, Per78, PB71, RC71a, RCG72, RC72, RSH79, Rau79a, RF78, REFN76, Rei76, RH74, RK75a, RK75b, SNN70, SG74, Sch78c, She70d, Smi79e, Sri76, Sta70b, SA75, Swa79a, TM70, Toi76, TW77, TK74, TJM79, Von75, WN76, WKD70, WG75]. System [YT77, You70, Cam72, Flo72]. System/360 [Joh71, Cam72]. System/370 [Flo72]. Systematic [Cur70, Den74, Rao72]. Systems [Ano72-39, Avi76, Bae76, BB72, Ban74b, Bar75a, BM73, BM74, BGM76, BM78, Bas73, BS74, Bea78, Ber70a, BS75, Bha75b, Boo70, BF75b, BE70b, BW77, Bre71a, Bre73b, BP72, CF78a, Cha74c, CM70, CK75, CS79, CLL78, Coy79, DW78, De 79, Dee72, DF70, Dot70, Eks74, Eks73, FD74a, FS74, Fre74b, FB74, FM71, Fri73, FB73b, Fu74, FK78a, FK78c, Gay78, Gel73a, Gor75, GH72, HA74, Hay76a, Hei79, Hil79, Jos74, JA78, KG72, KO73, KF79, Ker73, Kim74, Kle70, Koh75, KP71, KS79, Kuo74, Kur74, Lan77, LT73, LJ79, Lud74, MH76, MH78, MF76a, MM78a, Mat71c, McL72, MK79, Mil74a, Mor76a, Nay71, NN73, O’K75, Par71, Paz71, Pea75, Pu72, RH75a, Rav72, Red78a, Reg73, SAPR70, Saw77, SM71b, SD79, SM73b, Sie74, Sri79]. Systems [Sri71a, Sri71b, Stu75, Su70, SSR72, Tab74, TM71, TI72b, Tis71, TW76, UB74, VM78, Var79, Vic74, VLS70, Wai70, Wak75c, YL73, Zoh76b]. Systems-Switching [Tis71].

T [GB71b]. T-Partitionable [GB71b]. T. [Wor79]. Table [Cur77, Das71b, DK72, Dia76, MS77b, SD74]. Table-Lookup [MS77b]. Tables [AM79c, Cur71a, DW72a, HS70, VVA78]. Tac [JW76]. Tagged [Feu73]. Tan [Cur73a]. Tan-Like [Cur73a]. Tandem [But78b]. Tangent [Pre78b]. TANT [LD72a, Fra72, Koh71, Lee78b, Vin78]. Tape [Arn70, Har70b, Kam70a, PB76b]. Tapered [Mor71]. Task [GR72b, WN76]. Task-Free [WN76]. Tasks [Koh75, RPL73, RFL76, RL78, RSH79, RV73]. Tausworthe [NM76a]. Teacher [MS76]. Teaching [WBRG76]. Technical [Ano71-52]. Technique [Agr78a, AC70, Bor72, Bre71d, DV78a, GL70, GH72, GP74, Hun72, IR76, JR76, Jol76, KS77a, Koc75, Koh71, KF72b, KF72c, Lan78, LD72a, Maj72c, MJTW75, MBM77, PF72, PH73, Red71b, RNMP77, RB72, SH77, Sic76, Skl78, Sto74a, Sur75, TM76, Toh74, TB74, WNR74]. Techniques [Aar70, CDJ73, CHSS73, Cha74c, CM70, DM78b, FM72a, Gil74a, HKD^{+}71, Kim74, Kri70b, LT70, Lil72, LM78, Mil77, Moh71, MG71b, New75b, PR72, PR76, Pra72, PD77, RPL73, RC70, SF71, Sau72b, STG71, Sie77, Sto70d, SNEL79, Tou72a, Wai70, Weh74, YK76]. Technological [Rem72]. Technology [Ano77b, Ano77c, Fre74a, FLBK76, HHJ^{+}76, Jos72b]. Telecommunications [Cox72]. Teleprocessing [CR72b, IA76, TWB78]. Television [DWT70]. Template [Oza78, VR77]. Temporary [Sri71b]. Ten [Sit74]. Tensor [WS76]. Terminal [Arm73, Har70a, HI72, KR78b, PRD71, SAFV71, Wei72]. Terminals [Hob73, Pat73, TWB78]. Termination [Hor72]. Ternary [AA72b, Bar77, BM77a, BY79, CS70, EI77, FL75, Gra73, HK77, HK78, MJ77, MJ78, SAPR70, SA70, SA74]. Tesselated [MF76d]. Tessellation [SC74, Wag70]. Test [AM79a, AM79b, AA72a, AA75, Ake73, Ake76, Bar70, Bet71, Bre71d, Bre74, BH74b, CSJ74, CDO78, Dan73, DDSM76, HF74, HH77, Hla78, How75, Hus74, Mar71, MST^{+}75, Mut76, OH79, Pap77a, Pra78d, Red73a, RS75, SR75, SLC75, SV74, TV79, Wan75a, Wei72, WA73, Wul74, YT71c]. Test-Input [DDSM76, Mar71, MST^{+}75]. Test-Pattern [SLC75]. Testability [DR74, RSL76, WA73]. Testable [Coy79, Fri73, FNSK75, Kod74, Red72b, RW74, SR74a, SR74b, Sun76]. Tested [RH75b]. Testing [AA76, Agr78b, BK76, Bre73c, CDJ71, GBMC72, GCC74, Hay76c, Kau74b, KH77b, KH77c, LD74, Los78, NTA78, Nai79, PR71, RC72, RS74, Red77b, RS78, Rub75a, SB77a, SB79, Sch79a, Si78, Smi79b, SC74, Sur74, TV76, Yue77e, AGM^{+}71]. Tests [BC71b, Bre72a, FD74b, FD77b, Hay71b, Hay78a, KS71, MSJ72, Red77a, Rob70b, Toi76, Wan75b]. Text [AKM^{+}71, AN74]. Textural [Tho77]. Texture [DM78b, MMB77, RJ72, Ros75, SH72, Tan78, ZRD75]. Texture-Based [DM78b]. Textured [Ehr77, Nev76, Sha78b]. th [AM79c]. Their [Ale71, AP76a, BR76a, BF72, Duk74, Fen74, Fly72b, Har70a, Hay74, HC73, Hwa72, Jos72a, KN70, KMC72, LJ79, Meo75, SK79a, Tha78, The75, Wak74, Wan75c, YT70, FD75, GL74, KT72]. Theorem [Ern76, Fis76, MOW76, Mor76b, Pra71a, Rei76, TL73a, Vil73a, Yue75d]. Theorem-Proving [MOW76]. Theorems [FK78c]. Theoretic [KNF76, Law70, Nus78]. Theoretical [Che71a, CW75, Mil73, Sha78a, Tou72c, Zah71]. Theories [de 70]. Theory [AGM^{+}71, BGM76, BS74, Bei70, Bei71, BA71, Cha74c, CM79d, Fu74, Gar78, Hou72, KS78a, Kar74, Kel74, Lei74, Lew74, LS76b, Moo76, Pra78c, Rei74b, RS71a, Roo73, SP72b, Sto70a, TD73, Tha78, UB74, YK76, Yoe70a]. There [CS70]. Thermoplastic [BL75]. Thought [Mic74]. Three [AA72b, Ano72-39, Ast78, BM77a, BCSM72, Bre72b, CCB70, CC73, Gec77, Koc71, Mor76a, O’K75, Par71, RH75b, Sar71, SF77, UM77, Wei75]. Three-Dimensional [Ano72-39, CC73, Par71, Sar71, UM77, Wei75]. Three-Face [SF77]. Three-Failure-Tolerant [Koc71]. Three-Level [CCB70, RH75b]. Three-Stage [Ast78]. Three-Valued [Bre72b, Mor76a]. Threshold [AA72b, Ama72, Bar70, BW76a, Bre71b, CC72, CM79b, DS72a, Edw75, FD73, GW77, GBMC72, GC73, Ham73, Hen71, HC73, Ish77a, Ken77b, KT72, Lyo70, Mor70a, MTB70, OS76, OY70, SA70, She70b, SR72b, Sli70, Tak76, Wea75, WNR74, Win70a, Win70b, Yue76]. Threshold-Element [Tak76]. Threshold-Logic [Lyo70]. Throughput [BCHL73, Sch79b]. Tic [JW76]. Tic-Tac-Toe [JW76]. Time [AK72, Arn70, AO79, BCKT79, CK75, CK77, Dep70, Der73, DF70, Dot70, Dow73, FB73a, FS74, Gay78, Gel73a, GSJ^{+}76, Goo70, Goy76, HO79, Har70b, HA70, Hun72, Kle70, Koc75, KR78a, Lev70, Mar78, NT78, Nec79, Nie70b, PRD71, Pea75, Pol71, Saw77, Sch79b, Sil77, Spi71, WKD70]. Time-Component [AO79]. Time-Invariant [Pea75]. Time-Shared [Kle70]. Times [Pec76, Spi73]. Timing [Nie70b]. Toe [JW76]. Toeplitz [Eks74]. Toggle [AM79c]. Toggle-Registers [AM79c]. Tolerance [BO70, Ogu74]. Tolerant [AGM^{+}71, Avi76, BCJ^{+}71, Car73, CM76, FM72a, GM76, GCP^{+}77, Hay76a, Hec73, Hop71, HS75, Ios75, Kim75, Koc71, LR72, MS74, Mat71c, Mey71, MR76, OW73b, PR73, PR74b, Pra74, Pra78b, Ram71, Red78b, RC70, SM75a, Sch74a, SG71, Smi75, Sti78b, TA71]. Tombs [Oza78]. Tone [JR76]. Tool [BW77]. TOPI [GGMM71]. Topoglyphs [Gol78]. Topological [Maz71, SV75, TG72b, UB74]. Total [DS71]. Totally [AM73, AR77, Bis70, Bis73, Bor72, Dav78, Usa75, YT71b]. Tours [DS73]. Trace [Ram77a]. Traced [DT70, TD70]. Traces [BBP75]. Tracey [HB70]. Tracing [Weg72]. Track [PB76b].

Tracking [SP72a, SSR72]. Tradeoff [Rus72]. Training [DS72a, LK75]. Transactions [Ano70a, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-37, Ano70-34, Ano70-35, Ano70-36, Ano71a, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72a, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano73a, Ano73m, Ano73n, Ano74a, Ano75a, Ano76b, Ano77a, Ano78h, Ano78i, Ano78j, Ano78b, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano79k, Ano79l, Ano79a, Ano73e, Ano78h, Ano78i, Ano78j, Ano79k, Ano79l]. Transcoder [DG72]. Transfer [ASS72, Bar75a, BEGW72, BT70b, HP75, Low73, Sto70c]. Transferable [Roc70]. Transference [Ull72]. Transfers [Max77, Mcd72]. Transfom [ANR74]. Transform [ARS71b, Cap76, Car71a, CS73, CF79, Des74, Des79, DV78b, Edw75, FA76, GC77, Hen70, Jai77a, Kun75, Lec70a, Lec70b, LD72b, RNR75, RT79, Ree70, Rej72, RRNA74, Sha75, Sha78b, Ulm70, Yip76, Yue75a, Yue76, Yue77e, Ber72, Cai71, Cor71a, Cor71b, CSY75, CL72, Dru71a, Gla70, GW70, Har76, LH75, MS73b, PB76a, SR77, TM78a, WK71]. Transformation [AK72, CL79, Cor77, KF72a, Kri70b, Kri71b, LH76, ND73, Pap76, TM78b, Tur70, VW72, Weh74, YT71b, Zoh73b, Zoh74]. Transformations [Ami77, BD71, DN77, NSR74, PC72, Rat76, Sta70a, WGS78]. Transforms [Bro77, CH72a, CH73b, Dun75, FHTS78, Kun79, Nus78, Pre71b, Rad72, Rad73, RTKH77, Sar78, VR79, Yue72, Goo71, GW75, Hun71, Kar77b]. Transient [Wak75c]. Transistor [HHJ^{+}76]. Transition [BH73, Hay76c, Hay78a, KR78a, NT78, Red77b, Reu72, SR73, Sho74]. Transition-Sensitive [BH73]. Transition-Time [KR78a]. Transitive [Tsi70]. Translation [BB72, IC78, Lew74, Tho75]. Translators [BCH^{+}73, CDJ73]. Transmission [And70, Hor71b, Hor72, Mah74, MMO74, SW70a, SW70b]. Transparent [Sun79]. Transportation [Mah74]. Transposing [All75a, Ari79, DS74, Ekl72, Ram75, Sch73, Van77]. Transposition [TE76]. Trapezoidal [Max77]. Traversing [Klu77]. Treatment [FA76, Hol72b, PM75b, VR79]. Tree [AM79a, AA76, Arb70, BS76, Ber70a, Ber71, CK74, CR72b, Cha70, CE75, FB73b, Gri76, IA76, Ish77b, KH77a, Klu77, KB75, LF78, MH75, MF76c, Pat72, SK77c, Ung77c]. Tree-Searching [Gri76]. Tree-Type [CE75, KH77a]. Trees [BF78, CMS79, Cha74d, Cur79, EF76, GST79, KLCC77, Li77, LS76b, LP76, LP78, MM73b, PM77b, TG78, WC74]. Trends [Hol76, JM74, New76]. Triangular [Fre79a]. Triangulation [LSB77]. Tridiagonal [SK77a, SK77b]. Trigger [HN73]. Triple [AS74, Wak75c]. Truncation [Zoh76b]. Trunk [AKT^{+}77]. Trunk-Division [AKT^{+}77]. Truth [Dia76, Har70a, Pin71]. Truth-Table [Dia76]. Tumors [BS76]. Tuple [Rat76]. Turing [Har70b, Lyn72]. Tutorial [Ano78v, Ano78w, Ano78x, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Hen76a]. Twelve [Car71a, WK71]. Twiddling [Yue73a]. Two [AO79, Bar70, Bar71, BW73, Bla74, BR74, CC72, CC73, CW79c, DM78b, EZS78, ET75, FO71c, GG75, Goo71, Gra71, HND75, Han75, Har70a, HO79, Hwa79, Ios78, Jai77c, JM71, Kri71a, Kro74, LSB77, Mar73b, MG75, MK79, MS78, Moh71, Moo71, Nak79, NA74, OY70, PCM70, Pav75b, PR74b, RR74, Roe78, SR74a, SK78b, SM72a, SM71b, Sil77, Sta74, Tak76, Tom74, VR77, Vic71, Vil73a, Wri73, Yan75a, YRT75, NA76b]. Two-Class [BR74, PCM70]. Two-Dimensional [DM78b, FO71c, HND75, Jai77c, MS78, NA74, Roe78, SR74a, SM72a, Vil73a, NA76b]. Two-Input [SK78b]. Two-Layer [Tak76]. Two-Level [CC72, EZS78, ET75, Ios78, MK79, Nak79, OY70, PR74b, RR74]. Two-Multiplies-per-Cycle [HO79, Sil77]. Two-Point [Vic71]. Two-Space [LSB77, Wri73]. Two-Stage [VR77]. Two-State [CW79c]. Two-Terminal [Har70a]. Two-Way [Bar71, Kri71a, Moo71]. Type [AH76, BCS76, CE75, Cur73b, KH77a, Orc76, PHG77, Pre72, Sik70, SS77]. Types [KF78, LHSM74, Tak76]. Typicality [WL75].

ULM [Voi77b]. Ultimate [Brz70]. Ultrareliable [Mat71c]. Ultrasonic [GW75]. Unate [BPC75, Bau72, Bet71, DSBC74, Mat71b, SK79b, Skl70, TD77]. Unateness [Bar70, Fis74, Pra78a]. Uncertainty [Key77]. Understanding [Dav77b, EFLR76, REFN76, Wal76]. Undetectability [Smi79d]. Unformatted [AKM^{+}71]. Unger [Pra71a]. Unidirectional [Wak75b]. Unified [CM77, DAA79, DB73, FA76, Gel73b, KS77c, LD75, UB74, VR79]. Uniform [CF78b, Dav74, GR71a, Hal74, Hou72, Jum70, Obe72, Wil75]. Unifying [Reu75]. Unimodal [GL70]. Uniprocessor [RFL76, RL78]. Uniqueness [Nie70a]. Unit [Bat70, BT76, Dev79, Jol76, Kau70c, KS77c, Mor70a, SS71, Sri78, Wea75]. Unit-Distance [Dev79, Kau70c]. Unitary [WGS78]. Units [Atk70, GHJ77, Meo75, RS75, RS78, YC76]. Universal [Ake73, Ale71, Alm78, AT70, BB73, But78b, Cha70, CE75, FG71c, GT74, Hua71, Ios75, KH77a, Kan75, Kel74, Kli71, LV75, NT78, NT79, NA72, NK75, OW72, Pat73, Pra78d, Pre71a, Pre72, Smi79e, SB74b, Sur78, TG78, Voi77a, Voi77b, YT70]. Universality [SK79a]. University [AG73]. Unrestricted [MS75a, Ung71].

Unscramble [Swa74a]. Unscrambling [PB76a]. Unspecified [Dav70a]. Unstable [Vic71]. Unsupervised [LD75, MS75b]. Update [GS79]. Upper [Fre79b, Tou71, Vil70, Yue72]. Use [Agr78b, Atk70, Aus70, Ban74c, BEGW72, Ber77, BK71b, Cle73, DT70, EF79, EH76, FB74, Ges71, Hla70, Hun71, Jos72b, Kar70, Mal73, PS71, Sch79a, Sha78a, Tha78, Tri77b, Tri78, TB75, Wak74, RT73a]. Used [Bec71, Len78, O’K73]. User [Har71, Tho74]. Uses [AP76a, Jos72a]. Using [ADMR78, Alm78, ACJ75, Ano79b, BC71a, BPC75, Bar77, Bar70, BT72, BCL75, BH73, Bre75, BB75, Cha72, CD73, CD78, Cur70, Cur71b, DDSM76, DM78b, Dep70, Des74, Elm76, FM72a, Gar72a, GY77, Gel76, Gri73, Gri78a, HLD70, HA77, Hun72, HTD72, Iba71, IC78, JP73, JB76a, Jul78, KF79, KM78, KC71, Kit78, Kon71, KF72a, Kri77, LT70, Lan76, Lin70, Mam77, MS73b, Mar71, Mat71b, Meo75, MST^{+}75, MBSS75, MS78, Mor75, NT70, Oja79, PA78, PM75a, Pra78b, RM71, RGK72, RT70, RE71, RH74, Sau72b, SH77, SM72a, Ser73, SV74, SY75, SR71, SP72a, Tho76, Tjo77, TOS71, Toh74, TJM79, TB74, Tun70b, VD76b, VV74, VVA78, Wak75b, Wil78, WS76, Yen79, You70, BF75a, Rao72]. Utilities [BH76]. Utilization [Cha70, DB74, FGR71, SN79]. Utilizing [Das71b, FK74, SNN70].

V [Tab70b, CDO78]. Valley [KF72c]. Valley-Seeking [KF72c]. Value [Kar70, PR78, Vic71, WK73]. Valued [Bre72b, CM79b, DVS74, DV78a, GLR71, ISN76, KH77a, KT72, Mor76a, Mut76, SV74, Sin72, Tou72d, TG79, VLS70, Vra77]. Values [Wan75b]. Variable [Byr71, CYNM79, Ehr76, Sti78b, Sto74a, Tom74, Wai70]. Variable-Mode [Byr71]. Variables [AA72b, Ano75-28, BCSM72, Bor72, Dah73, MTB70, Mur70, Pav75b, Vin78, Wri73].

Variables-Weighted [Dah73]. Variational [RH71]. Variations [Hoe76]. Various [Bre73b, IS76, MM73a, Sie79]. Varistructured [Lip77]. Vector [Ath75, Eks74, FV79, Hig76, Kit77, Lee76]. Vectorcardiograms [KCLB71]. Vectors [BP72, FS75, Glu71, Kit78, SO79b, Swa74a]. Vehicles [Hop71]. Velocity [Dav72b, SF72]. Verification [BT72, Dia76, DW73a, EH76, HP75, Rot77, ST76]. Versatile [MH70b, TP72, YT77]. Version [WN76]. Versus [Ham71, LR71a, LR72, Pea76]. Vertex [Hwa72]. Vertices [Hwa72, SF77]. Very [BHK79, Des79, RV73]. via [GW75, Rad72, Rad73, RA78, WA73]. Video [GSJ^{+}76, KF73, MCRD72]. Viewed [BG75]. Viewpoint [Car73, FR71]. Views [MA77, UC75]. Visible [Mah72]. Vision [Per78, YT77]. Visual [RT71, UC75]. VLSI [PW79b]. Vol [Ano70a, Ano71a, Ano72a, Ano73a, Ano74a, Ano75a, Ano76b, Ano77a, Ano78b, Ano79a]. Voltage [TS75]. Volume [Cha74c, Lew74]. Voted [Sie75]. Voting [HP73].

W [Tab70b]. Walsh [Car71b, Hen70, KS72, Lec70a, Van73b, Yue76, AK72, Bro77, But72, But78a, CS73, Dav72a, Edw75, FA76, FB74, GC77, Kun79, LM71b, Red71b, RRNA74, Ric74b, TG72a, Yue72, Yue74b, Yue75a, Yue75e, Yue75d, Yue77e]. Washington [Ano74b, Ano74c, Ano74d, Ano74e, Ano74f]. Waveform [PF72, Pav73]. Waveforms [CL71, Dav71, EF76, Tjo77]. Way [Bar71, Du72, Har76, Kri71a, Moo71]. Week [Ano79m, Ano79n, Ano79o, Ano79p, Ano79q]. Weight [Hwa72, TL73b, WL74]. Weight-Preserved [WL74]. Weighted [AA77, Bro75c, Dah73, SLC75, VV74]. Weighting [Bub72, NRS74]. Weights [Mal73]. which [JB76b]. Whirl [Paz71]. White [Jai77a, Mor74, Sok72, Yue74b]. Wide [GSJ^{+}76]. Wide-Band [GSJ^{+}76]. Width [BK79, JB76b]. Wiener [FB74, Hab77, Pra72, PD77]. Window [Cha73b]. Winograd [Wak70b]. Wire [Abe72, AKT^{+}77]. Wire-Routing [AKT^{+}77]. Wiring [Kau74b]. with-Binary [EI77]. Within [WGS78]. Without [BKM73, CF78b, Hla70, MS76, RK75b]. Word [Bak79, CBS71, Far79, RE71, Sch78c]. Word-Level [Far79]. Words [JB76b, KM78, OTK76, Pol71, Rom78]. Work [ES77, Hel72, Hel74, Mas78, Wel74]. Working [GF78, Smi76]. Worst [Bri73, Rub75b, Yue74a]. Writable [LM78].

X [Wei75]. X-Ray [Wei75].

Years [Ano75b, Weg76]. Yield [HB75a].

Zadeh [Tab70b]. Zero [Geo70, NS75]. Zero-Crossing [NS75]. Zero-One [Geo70]. Zoom [Yip76].

References

Ali:1971:RNS

[AA71] M. Ali and R. Ahmed. Realization of nonlinearly separable switching functions. IEEE Transactions on Computers, C-20(6):695– 698, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                                 https://ieeexplore.ieee.org/
                                                 stamp/stamp.jsp?tp=&arnumber=
                                                 1671923.

Agrawal:1972:ATG

[AA72a] Vishwani D. Agrawal and Prathima Agrawal. An automatic test generation system for Illiac IV logic boards. IEEE Transactions on Computers, C-21(9):1015–1017, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009081.

Aibara:1972:ETT

[AA72b] Tsunehiro Aibara and Michihiro Akagi. Enumeration of ternary threshold functions of three variables. IEEE Transactions on Computers, C-21(4):402– 407, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008986.

Agrawal:1975:PAR

[AA75] P. Agrawal and V. D. Agrawal.

Probabilistic analysis of random test generation method for irredundant combinational logic networks. IEEE Transactions on Computers, C-24(7):691– 695, July 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672882.

Agrawal:1976:MCT

[AA76] P. Agrawal and V. D. Agrawal.

On Monte Carlo testing of logic tree networks. IEEE Transactions on Computers, C-25(6):664– 667, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674670.

Al-Alaoui:1977:NWG

[AA77] M. A. Al-Alaoui. A new weighted generalized inverse algorithm for pattern recognition. IEEE Transactions on Computers, C-26(10): 1009–1017, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674736.

Abd-Alla:1974:HSM

[AAK74] A. M. Abd-Alla and D. C. Karlgaard. Heuristic synthesis of microprogrammed computer architecture. IEEE Transactions on Computers, C-23(8):802–807, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672629. See comments [Col77].

Aaron:1970:BAD

[Aar70] M. R. Aaron. B70-3 analog-to-digital and digital-to-analog conversion techniques. IEEE Transactions on Computers, C-19(8):762– 763, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671622.

Agin:1976:CDC

[AB76] Gerald J. Agin and Thomas O. Binford. Computer description of curved objects. IEEE Transactions on Computers, C-25(4):439– 449, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674626.

Arevalo:1978:MSB

[AB78] Z. Arevalo and J. G. Bredeson. A method to simplify a Boolean function into a near minimal sum-of-products for programmable logic arrays. IEEE Transactions on Computers, C-27(11):1028–1039, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674992.

Abramovici:1979:RFD

[AB79a] M. Abramovici and M. A. Breuer.

On redundancy and fault detection in sequential circuits. IEEE Transactions on Computers, C-28(11): 864–865, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675267.

Agrawal:1979:EDR

[AB79b] P. Agrawal and M. A. Breuer. Experiments with a density router for PC cards. IEEE Transactions on Computers, C-28(3):262– 267, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675333.

Abdali:1971:PSM

[Abd71] S. K. Abdali. On proving sequential machine designs. IEEE Transactions on Computers, C-20 (12):1563–1566, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671763.

Abel:1972:OCA

[Abe72] L. C. Abel. On the ordering of connections for automatic wire routing. IEEE Transactions on Computers, C-21(11):1227–1233, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672075.

Aven:1976:SRD

[ABK76] O. I. Aven, L. B. Boguslavsky, and Y. A. Kogan. Some results on distribution-free analysis of paging algorithms. IEEE Transactions on Computers, C-25(7):737– 745, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674682.

Abraham:1975:CSR

[Abr75] J. A. Abraham. A combinatorial solution to the reliability of interwoven redundant logic networks. IEEE Transactions on Computers, C-24 (5):578–584, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672858.

Andrews:1970:GTS

[AC70] Harry C. Andrews and Kenneth L.

Caspari. A generalized technique for spectral analysis. IEEE Transactions on Computers, C-19(1):16– 25, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008895.

Andrews:1971:DFM

[AC71] H. C. Andrews and K. L. Caspari. Degrees of freedom and modular structure in matrix multiplication. IEEE Transactions on Computers, C-20(2):133–141, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671795.

Andrus:1975:DIR

[ACJ75] J. F. Andrus, C. W. Campbell, and R. R. Jayroe. Digital image registration method using boundary maps. IEEE Transactions on Computers, C-24(9):935–940, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672934.

Aggarwal:1975:CAM

[AD75] J. K. Aggarwal and R. O. Duda. Computer analysis of moving polygonal images. IEEE Transactions on Computers, C-24(10):966– 976, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672695.

Adams:1978:CIM

[Ada78] E. N. Adams. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3): 287, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675091. See [DT76, Das78, Bar78].

Ausherman:1972:ECE

[ADL72] D. A. Ausherman, S. J. Dwyer, and G. S. Lodwick. Extraction of connected edges from knee radiographs. IEEE Transactions on Computers, C-21(7): 753–758, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672171.

Ahuja:1978:PAP

[ADMR78] N. Ahuja, L. S. Davis, D. L. Milgram, and A. Rosenfeld. Piecewise approximation of pictures using maximal neighborhoods. IEEE Transactions on Computers, C-27 (4):375–379, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675110.

Atkins:1973:PCA

[AG73] David E. Atkins, III and Harvey L. Garner. Preface: Computer arithmetic: An introduction and overview [Second IEEE Symposium on Computer Arithmetic, May 15–16, 1972, University of Maryland]. IEEE Transactions on Computers, C-22(6):549– 551, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Atkins.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009104.

Agerwala:1976:MOS

[Age76] T. Agerwala. Microprogram optimization: a survey. IEEE Transactions on Computers, C-25(10): 962–973, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674537.

Avizienis:1971:SST

[AGM^{+}71] A. Avizienis, G. C. Gilley, F. P.

Mathur, D. A. Rennels, J. A. Rohr, and D. K. Rubin. The STAR (Self-Testing And Repairing) computer: An investigation of the theory and practice of fault-tolerant computer design. IEEE Transactions on Computers, C-20(11):1312–1321, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671726.

Agrawal:1975:AAN

[Agr75a] D. P. Agrawal. Arithmetic algorithms in a negative base. IEEE Transactions on Computers, C-24(10):998–1000, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672702.

Agrawala:1975:GLP

[Agr75b] A. K. Agrawala. On generating a line “Parallel” to a digital line — a sequential approach. IEEE Transactions on Computers, C-24(10): 983–988, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672698.

Agrawal:1976:SMR

[Agr76] D. P. Agrawal. Signed modified reflected binary code. IEEE Transactions on Computers, C-25(5): 549–552, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674648.

Agrawal:1977:CNB

[Agr77] D. P. Agrawal. Comments on “A Note on Base-2 Arithmetic Logic”. IEEE Transactions on Computers, C-26(5):511, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674869. See [Yue75b].

Agrawal:1978:NTC

[Agr78a] D. P. Agrawal. A novel technique for computing negabinary squares. IEEE Transactions on Computers, C-27(3):266–270, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675084.

Agrawal:1978:WUR

[Agr78b] V. D. Agrawal. When to use random testing. IEEE Transactions on Computers, C-27(11):1054–1055, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674994. See comment [Sch79a].

Agrawal:1979:HSA

[Agr79a] D. P. Agrawal. High-speed arithmetic arrays. IEEE Transactions on Computers, C-28(3):215– 224, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675322.

Agrawal:1979:AR

[Agr79b] V. D. Agrawal. Author’s reply. IEEE Transactions on Computers, C-28(8):581, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675416.

Agrawal:1979:CAH

[Agr79c] V. D. Agrawal. Comments on “An Approach to Highly Integrated Computer-Maintained Cellular Arrays”. IEEE Transactions on Computers, C-28(9):691– 693, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675440. See [Man77].

Acha:1976:INS

[AH76] J. I. Acha and J. L. Huertas. On input and next-state equations of the R–S type M-stable. IEEE Transactions on Computers, C-25 (7):759–763, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674685.

Appelbe:1978:SHM

[AI78] W. F. Appelbe and M. R. Ito. Scheduling heuristics in a multiprogramming environment. IEEE Transactions on Computers, C-27 (7):628–637, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675163.

Alexandridis:1972:RTW

[AK72] Nikitas A. Alexandridis and Allen Klinger. Real-time Walsh–Hadamard transformation. IEEE Transactions on Computers, C-21(3):288– 292, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008951.

Andrews:1975:HCM

[AK75] M. Andrews and G. A. Korn.

Hybrid computer methods for direct functional optimization. IEEE Transactions on Computers, C-24(10):958–965, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672694.

Akers:1970:RMC

[Ake70] S. B. Akers, Jr. R70-40 module clustering to minimize delay in digital networks. IEEE Transactions on Computers, C-19(10):987– 988, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671406.

Akers:1972:RLA

[Ake72] Sheldon B. Akers. A rectangular logic array. IEEE Transactions on Computers, C-21(8):848– 857, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009040.

Akers:1973:UTS

[Ake73] Sheldon B. Akers. Universal test sets for logic networks. IEEE Transactions on Computers, C-22(9): 835–839, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009174.

Akers:1974:FDG

[Ake74] S. B. Akers, Jr. Fault diagnosis as a graph coloring problem. IEEE Transactions on Computers, C-23 (7):706–713, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672611.

Akers:1976:LSF

[Ake76] S. B. Akers, Jr. A logic system for fault test generation. IEEE Transactions on Computers, C-25 (6):620–630, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674662.

Akers:1978:BDD

[Ake78] S. B. Akers. Binary decision diagrams. IEEE Transactions on Computers, C-27(6):509– 516, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675141.

Ascher:1971:ISR

[AKM^{+}71] R. N. Ascher, G. M. Koppelman, M. J. Miller, G. Nagy, and G. L. Shelton, Jr. An interactive system for reading unformatted printed text. IEEE Transactions on Computers, C-20(12):1527–1543, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671759.

Allan:1975:ADA

[AKT75] F. J. Allan, T. Kameda, and S. Toida. An approach to the diagnosability analysis of a system. IEEE Transactions on Computers, C-24(10):1040–1042, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672718.

Asano:1977:WRS

[AKT^{+}77] T. Asano, T. Kitahashi, K. Tanaka, H. Horino, and N. Amano. A wirerouting scheme based on trunkdivision methods. IEEE Transactions on Computers, C-26(8):764– 772, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674914.

Arcelli:1971:PPO

[AL71] C. Arcelli and S. Levialdi. Picture processing and overlapping blobs. IEEE Transactions on Computers, C-20(9):1111–1115, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672007.

Aleksander:1971:CUL

[Ale71] I. Aleksander. Comments on “Universal Logic Modules and Their Applications”. IEEE Transactions on Computers, C-20(5):586– 587, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671892. See [YT70].

Alltop:1975:CAT

[All75a] W. O. Alltop. A computer algorithm for transposing nonsquare matrices. IEEE Transactions on Computers, C-24(10): 1038–1040, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672717. See [Ekl72] and comments [Van77].

Allum:1975:LMS

[All75b] J. H. J. Allum. A least mean squares cubic algorithm for on-line differential of sampled analog signals. IEEE Transactions on Computers, C-24 (6):585–590, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672863. See comments [Wil77b].

Allum:1977:AR

[All77] J. H. J. Allum. Author’s reply.

IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674853.

Almaini:1978:SMI

[Alm78] A. E. A. Almaini. Sequential machine implementations using universal logic modules. IEEE Transactions on Computers, C-27(10):951– 960, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674975.

Anderson:1973:DTS

[AM73] D. A. Anderson and G. Metze. Design of totally self-checking check circuits for m-out-of-n codes. IEEE Transactions on Computers, C-22 (3):263–269, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672298.

Agarwal:1977:ROF

[AM77] V. K. Agarwal and G. M. Masson. Resolution-oriented fault interrelationships in combinational logic networks. IEEE Transactions on Computers, C-26(11):1170–1175, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674771.

Agarwal:1979:FFA

[AM79a] V. K. Agarwal and G. M. Masson. A functional form approach to test set coverage in tree networks. IEEE Transactions on Computers, C-28 (1):50–52, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675221.

Agarwal:1979:RCP

[AM79b] V. K. Agarwal and G. M. Masson. Recursive coverage projection of test sets. IEEE Transactions on Computers, C-28(11):865– 870, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675268.

Arvillias:1979:TRG

[AM79c] A. C. Arvillias and D. G. Maritsas. Toggle-registers generating in parallel k k th decimations of m-sequences x^{P} + x^{k} + 1 design tables. IEEE Transactions on Computers, C-28(2):89– 101, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675298.

Amari:1972:LPP

[Ama72] S.-I. Amari. Learning patterns and pattern sequences by selforganizing nets of threshold elements. IEEE Transactions on Computers, C-21(11):1197–1206, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672070.

Amin:1977:AGL

[Ami77] A. T. Amin. An algorithm for grey-level transformations in digitized images. IEEE Transactions on Computers, C-26(11):1158–1161, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674767.

Ascher:1974:MAH

[AN74] R. N. Ascher and G. Nagy. A means for achieving a high degree of compaction on scan-digitized printed text. IEEE Transactions on Computers, C-23(11):1174–1179, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672419.

Andrews:1970:BTI

[And70] H. C. Andrews. B70-2 transmission of information by orthogonal functions. IEEE Transactions on Computers, C-19(7): 656, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671596.

Andrews:1971:MRF

[And71] H. C. Andrews. Multidimensional rotations in feature selection. IEEE Transactions on Computers, C-20(9):1045–1051, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671993.

Anderson:1972:BCG

[And72] T. F. Anderson. B72-2 A compiler generator. IEEE Transactions on Computers, C-21(1): 109, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672036.

Anderson:1974:MMR

[And74] G. A. Anderson. Multiple match resolvers: a new design method. IEEE Transactions on Computers, C-23(12):1317–1320, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672449.

Andrews:1976:RM

[And76] P. B. Andrews. Refutations by matings. IEEE Transactions on Computers, C-25(8):801– 807, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674698.

Anonymous:1970:IIT

[Ano70a] Anonymous. 1970 index IEEE Transactions on Computers vol. C-19. IEEE Transactions on Computers, C-19(12):1297–b–1297–m, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671463.

Anonymous:1970:ACCa

[Ano70b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(1):85– 94, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008907.

Anonymous:1970:ACCb

[Ano70c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(2):177– 195, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671485.

Anonymous:1970:ACCc

[Ano70d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(3):273– 290, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671502.

Anonymous:1970:ACCd

[Ano70e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(4):365– 385, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671522.

Anonymous:1970:ACCe

[Ano70f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (5):467–482, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671553.

Anonymous:1970:ACCf

[Ano70g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (6):567–582, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671581.

Anonymous:1970:ACCg

[Ano70h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (7):661–678, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671605.

Anonymous:1970:ACCh

[Ano70i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(8):767– 782, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671630.

Anonymous:1970:ACCi

[Ano70j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(9):861– 877, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671660.

Anonymous:1970:ACCj

[Ano70k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19(10): 991–1008, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671412.

Anonymous:1970:ACCk

[Ano70l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (11):1121–1138, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671439.

Anonymous:1970:ACCl

[Ano70m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-19 (12):1229–1297, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671461.

Anonymous:1970:AI

[Ano70n] Anonymous. Author index. IEEE Transactions on Computers, C-19 (1):100, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008910.

Anonymous:1970:CGC

[Ano70o] Anonymous. Comment on “Gray Code Counter”. IEEE Transactions on Computers, C-19(7):653– 654, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671593.

Anonymous:1970:Ca

[Ano70p] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (1):82–84, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008906.

Anonymous:1970:Cb

[Ano70q] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (2):175–176, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cc

[Ano70r] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (5):460–461, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cd

[Ano70s] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (6):561–562, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ce

[Ano70t] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (7):655, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cf

[Ano70u] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (8):760–761, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Cg

[Ano70v] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (9):852–853, September 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ch

[Ano70w] Anonymous. Contributors. IEEE Transactions on Computers, C-19 (10):984–985, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:Ci

[Ano70x] Anonymous. Contributors. IEEE Transactions on Computers, C-19(12):1227–1228, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:CPO

[Ano70y] Anonymous. Correction to “Pictorial Output with a Line Printer”. IEEE Transactions on Computers, C-19(9):851, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=

1671649. See [Mac70].

Anonymous:1970:DCI

[Ano70z] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-19(1):94–99, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008908.

Anonymous:1970:EN

[Ano70-27] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-19(5):387, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671528.

Anonymous:1970:II

[Ano70-28] Anonymous. Identifier index. IEEE Transactions on Computers, C-19(1):99–100, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008909.

Anonymous:1970:ITCa

[Ano70-29] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(2):c1–195–c, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCd

[Ano70-30] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(5):c1–482–b, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCe

[Ano70-31] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(6):c1–582–b, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCf

[Ano70-32] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(7):c1–678–b, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCg

[Ano70-33] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(8):c1–782–b, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCi

[Ano70-34] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(10):c1–1008–d, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCj

[Ano70-35] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(11):c1–1138–d, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCk

[Ano70-36] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-19(12):c1–1297–o, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:ITCh

[Ano70-37] Anonymous. IEEE transactions on computerss. IEEE Transactions on Computers, C-19(9):c1– 877–c, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1970:IAa

[Ano70-38] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(2):195–b, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671487.

Anonymous:1970:IAb

[Ano70-39] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(3):290–c, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671505.

Anonymous:1970:IAc

[Ano70-40] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(4):385–b, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671524.

Anonymous:1970:IAd

[Ano70-41] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(5):482–a, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671554.

Anonymous:1970:IAe

[Ano70-42] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(6):582–a, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671582.

Anonymous:1970:IAf

[Ano70-43] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(7):678–a, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671606.

Anonymous:1970:IAg

[Ano70-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(8):782–a, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671631.

Anonymous:1970:IAh

[Ano70-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(9):877–b, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671662.

Anonymous:1970:IAi

[Ano70-46] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(10):1008–c, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671415.

Anonymous:1970:IAj

[Ano70-47] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(11):1138–c, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671442.

Anonymous:1970:IAk

[Ano70-48] Anonymous. Information for authors. IEEE Transactions on Computers, C-19(12):1297–n, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671464.

Anonymous:1971:IIT

[Ano71a] Anonymous. 1971 index IEEE Transactions on Computers vol. C-20. IEEE Transactions on Computers, C-20(12):1637–b–1637– f, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671789.

Anonymous:1971:ACCa

[Ano71b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(1):115– 132, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671686.

Anonymous:1971:ACCb

[Ano71c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(2):243– 260, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671819.

Anonymous:1971:ACCc

[Ano71d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(3):364– 380, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671841.

Anonymous:1971:ACCd

[Ano71e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(4):478– 492, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671871.

Anonymous:1971:ACCe

[Ano71f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (5):599–616, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671901.

Anonymous:1971:ACCf

[Ano71g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (6):713–726, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671928.

Anonymous:1971:ACCg

[Ano71h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (7):813–830, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671945.

Anonymous:1971:ACCh

[Ano71i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(8):949– 964, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671976.

Anonymous:1971:ACCi

[Ano71j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(9):1121– 1137, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672010.

Anonymous:1971:ACCj

[Ano71k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20(10): 1224–1239, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671708.

Anonymous:1971:ACCk

[Ano71l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (11):1420–1435, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671743.

Anonymous:1971:ACCl

[Ano71m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-20 (12):1623–1637, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671787.

Anonymous:1971:Ca

[Ano71n] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (1):112–114, January 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cb

[Ano71o] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (2):240–242, February 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cc

[Ano71p] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (3):362–363, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cd

[Ano71q] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (4):477, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ce

[Ano71r] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (6):711–712, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cf

[Ano71s] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (8):945–947, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cg

[Ano71t] Anonymous. Contributors. IEEE Transactions on Computers, C-20(9):1118–1120, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ch

[Ano71u] Anonymous. Contributors. IEEE Transactions on Computers, C-20 (10):1220, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Ci

[Ano71v] Anonymous. Contributors. IEEE Transactions on Computers, C-20(11):1415–1419, November 1971.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:Cj

[Ano71w] Anonymous. Contributors. IEEE Transactions on Computers, C-20(12):1619–1621, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ENa

[Ano71x] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-20 (4):381, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671845.

Anonymous:1971:ENb

[Ano71y] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-20 (10):1139, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671690.

Anonymous:1971:ITCa

[Ano71z] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(1):c1–132–b, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCb

[Ano71-27] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(2):c1–260–b, February 1971. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCc

[Ano71-28] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(3):c1–380–b, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCd

[Ano71-29] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(4):c1–492–b, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCe

[Ano71-30] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(5):c1–616–b, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCf

[Ano71-31] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(6):c1–726–d, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCh

[Ano71-32] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(8):c1–964–d, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCi

[Ano71-33] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(9):c1–1137–e, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCj

[Ano71-34] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(10):c1–1239–e, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCk

[Ano71-35] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(11):c1–1435–c, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:ITCl

[Ano71-36] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-20(12):c1–1637–q, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1971:IAa

[Ano71-37] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(1):132–a, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671687.

Anonymous:1971:IAb

[Ano71-38] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(2):260–a, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671820.

Anonymous:1971:IAc

[Ano71-39] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(3):380–a, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671842.

Anonymous:1971:IAd

[Ano71-40] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(4):492–a, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671872.

Anonymous:1971:IAe

[Ano71-41] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(5):616–a, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671902.

Anonymous:1971:IAf

[Ano71-42] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(6):726–c, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671931.

Anonymous:1971:IAg

[Ano71-43] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(7):830–a, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671946.

Anonymous:1971:IAh

[Ano71-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(8):964–c, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671979.

Anonymous:1971:IAi

[Ano71-45] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(9):1137–d, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672014.

Anonymous:1971:IAj

[Ano71-46] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(10):1239–d, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671712.

Anonymous:1971:IAk

[Ano71-47] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(11):1435–b, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671745.

Anonymous:1971:IAl

[Ano71-48] Anonymous. Information for authors. IEEE Transactions on Computers, C-20(12):1637–p, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671792.

Anonymous:1971:NEC

[Ano71-49] Anonymous. New Editor-in-Chief. IEEE Transactions on Computers, C-20(1):1, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671665.

Anonymous:1971:P

[Ano71-50] Anonymous. Preface. IEEE Transactions on Computers, C-20(9):965–966, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671982.

Anonymous:1971:SI

[Ano71-51] Anonymous. Subject index.

IEEE Transactions on Computers, C-20(12):1637–f–1637–n, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671790.

Anonymous:1971:TPR

[Ano71-52] Anonymous. Technical paper referees. IEEE Transactions on Computers, C-20(1):2–6, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:IIT

[Ano72a] Anonymous. 1972 index IEEE Transactions on Computers vol. C-21. IEEE Transactions on Computers, C-21(12):1475–b–1475– o, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672125.

Anonymous:1972:ACCa

[Ano72b] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(1):111– 127, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672039.

Anonymous:1972:ACCb

[Ano72c] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(2):225– 235, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008936.

Anonymous:1972:ACCc

[Ano72d] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(3):323– 331, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008963.

Anonymous:1972:ACCd

[Ano72e] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(4):415– 425, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008995.

Anonymous:1972:ACCe

[Ano72f] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (5):519–536, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672155.

Anonymous:1972:ACCf

[Ano72g] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (6):619–627, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009029.

Anonymous:1972:ACCg

[Ano72h] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (7):821–836, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672181.

Anonymous:1972:ACCh

[Ano72i] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(8):925– 934, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009059.

Anonymous:1972:ACCi

[Ano72j] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(9):1035– 1044, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009094.

Anonymous:1972:ACCj

[Ano72k] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21(10): 1137–1152, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672061.

Anonymous:1972:ACCk

[Ano72l] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (11):1253–1267, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672085.

Anonymous:1972:ACCl

[Ano72m] Anonymous. Abstracts of current computer literature. IEEE Transactions on Computers, C-21 (12):1460–1475, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672123.

Anonymous:1972:AII

[Ano72n] Anonymous. Applications are invited for IEEE/Popov Society Exchange Delegates. IEEE Transactions on Computers, C-21(12): 1475–a, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672124.

Anonymous:1972:AIa

[Ano72o] Anonymous. Author index. IEEE Transactions on Computers, C-21 (2):240, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008939.

Anonymous:1972:AIb

[Ano72p] Anonymous. Author index. IEEE Transactions on Computers, C-21 (3):336, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008966.

Anonymous:1972:AIc

[Ano72q] Anonymous. Author index. IEEE Transactions on Computers, C-21 (4):432, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008998.

Anonymous:1972:AId

[Ano72r] Anonymous. Author index. IEEE Transactions on Computers, C-21 (6):632, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009032.

Anonymous:1972:AIe

[Ano72s] Anonymous. Author index. IEEE Transactions on Computers, C-21 (8):939, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009062.

Anonymous:1972:AIf

[Ano72t] Anonymous. Author index. IEEE Transactions on Computers, C-21 (9):1051, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009097.

Anonymous:1972:BRa

[Ano72u] Anonymous. Book review. IEEE Transactions on Computers, C-21 (2):224, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008935.

Anonymous:1972:BRb

[Ano72v] Anonymous. Book review. IEEE Transactions on Computers, C-21 (3):322, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008962.

Anonymous:1972:BRd

[Ano72w] Anonymous. Book review. IEEE Transactions on Computers, C-21 (6):618, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009028.

Anonymous:1972:BRc

[Ano72x] Anonymous. Book reviews. IEEE Transactions on Computers, C-21 (4):413–414, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008994.

Anonymous:1972:BRe

[Ano72y] Anonymous. Book reviews. IEEE Transactions on Computers, C-21 (8):924, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009058.

Anonymous:1972:BRf

[Ano72z] Anonymous. Book reviews.

IEEE Transactions on Computers, C-21(9):1033–1034, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009093.

Anonymous:1972:Ca

[Ano72-27] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (1):105–107, January 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cb

[Ano72-28] Anonymous. Contributors. IEEE Transactions on Computers, C-21(2):221–223, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5008934.

Anonymous:1972:Cc

[Ano72-29] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (3):321–322, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008961.

Anonymous:1972:Cd

[Ano72-30] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (4):412–413, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008993.

Anonymous:1972:Ce

[Ano72-31] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (5):515–516, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cf

[Ano72-32] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (6):617, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009027.

Anonymous:1972:Cg

[Ano72-33] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (7):819–820, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Ch

[Ano72-34] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (8):922–923, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009057.

Anonymous:1972:Ci

[Ano72-35] Anonymous. Contributors.

IEEE Transactions on Computers, C-21(9):1032–1033, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009092.

Anonymous:1972:Cj

[Ano72-36] Anonymous. Contributors. IEEE Transactions on Computers, C-21 (10):1132–1133, October 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Ck

[Ano72-37] Anonymous. Contributors. IEEE Transactions on Computers, C-21(11):1248–1249, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:Cl

[Ano72-38] Anonymous. Contributors. IEEE Transactions on Computers, C-21(12):1455–1457, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:CBF

[Ano72-39] Anonymous. Correction to “Batch-Fabricated Three-Dimensional Planar Coaxial Interconnections for Microelectronic Systems”. IEEE Transactions on Computers, C-21 (8):922, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009056. See comments [Par71].

Anonymous:1972:CJC

[Ano72-40] Anonymous. Correction to the June 1972 contributors section. IEEE Transactions on Computers, C-21(11):1247, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672079.

Anonymous:1972:DCIa

[Ano72-41] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(2):236– 239, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008937.

Anonymous:1972:DCIb

[Ano72-42] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(3):331– 335, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008964.

Anonymous:1972:DCIc

[Ano72-43] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(4):426– 430, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008996.

Anonymous:1972:DCId

[Ano72-44] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(6):628– 630, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009030.

Anonymous:1972:DCIe

[Ano72-45] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(8):935– 938, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009060.

Anonymous:1972:DCIf

[Ano72-46] Anonymous. Descriptor-in-context index. IEEE Transactions on Computers, C-21(9):1045– 1049, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009095.

Anonymous:1972:ENa

[Ano72-47] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-21(2):129, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5008916.

Anonymous:1972:ENb

[Ano72-48] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-21(6):537, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009005.

Anonymous:1972:ENc

[Ano72-49] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-21(9):941, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009069.

Anonymous:1972:ECR

[Ano72-50] Anonymous. Expanded Computing Reviews now available to IEEE Computer Society members. IEEE Transactions on Computers, C-21(12):1475–p, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672126.

Anonymous:1972:IIa

[Ano72-51] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(2):239–240, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008938.

Anonymous:1972:IIb

[Ano72-52] Anonymous. Identifier index.

IEEE Transactions on Computers, C-21(3):335–336, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5008965.

Anonymous:1972:IIc

[Ano72-53] Anonymous. Identifier index.

IEEE Transactions on Computers, C-21(4):431–432, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5008997.

Anonymous:1972:IId

[Ano72-54] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(6):631, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009031.

Anonymous:1972:IIe

[Ano72-55] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(8):938–939, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009061.

Anonymous:1972:IIf

[Ano72-56] Anonymous. Identifier index. IEEE Transactions on Computers, C-21(9):1050, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009096.

Anonymous:1972:ICSa

[Ano72-57] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-21(11):1268, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672086.

Anonymous:1972:ICSb

[Ano72-58] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-21(12):1475–q, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672127.

Anonymous:1972:ITCa

[Ano72-59] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(1):c1–127–c, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCb

[Ano72-60] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(5):c1–536–b, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCc

[Ano72-61] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(7):c1–836–b, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCd

[Ano72-62] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(10):c1–1152–b, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCe

[Ano72-63] Anonymous. IEEE Transactions On Computers. IEEE Transactions on Computers, C-21(11):c1–1268–b, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:ITCf

[Ano72-64] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-21(12):c1–1475–s, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1972:IAa

[Ano72-65] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(1):127–b, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672041.

Anonymous:1972:IAb

[Ano72-66] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(5):536–a, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672156.

Anonymous:1972:IAc

[Ano72-67] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(7):836–a, July 1972. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672182.

Anonymous:1972:IAd

[Ano72-68] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(10):1152–a, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672062.

Anonymous:1972:IAe

[Ano72-69] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(11):1268–a, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672087.

Anonymous:1972:IAf

[Ano72-70] Anonymous. Information for authors. IEEE Transactions on Computers, C-21(12):1475–r, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672128.

Anonymous:1972:P

[Ano72-71] Anonymous. Preface. IEEE Transactions on Computers, C-21 (7):633–635, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672159.

Anonymous:1973:IIT

[Ano73a] Anonymous. 1973 index IEEE Transactions on Computers vol. C-22. IEEE Transactions on Computers, C-22(12):1144–a–1144– l, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672258.

Anonymous:1973:BRa

[Ano73b] Anonymous. Book reviews. IEEE Transactions on Computers, C-22 (6):628, June 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009120.

Anonymous:1973:BRb

[Ano73c] Anonymous. Book reviews. IEEE Transactions on Computers, C-22 (7):704–708, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009141.

Anonymous:1973:BRc

[Ano73d] Anonymous. Book reviews. IEEE Transactions on Computers, C-22(9):879–884, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009190.

Anonymous:1973:CEP

[Ano73e] Anonymous. A change in editorial policy for Transactions papers and abstracts. IEEE Transactions on Computers, C-22(1): 1, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672185.

Anonymous:1973:CJC

[Ano73f] Anonymous. Correction to a January 1973 contributor entry. IEEE Transactions on Computers, C-22 (4):428, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672329.

Anonymous:1973:EN

[Ano73g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-22 (11):965, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672224.

Anonymous:1973:ECR

[Ano73h] Anonymous. Expanded computing reviews now available to IEEE Computer Society members. IEEE Transactions on Computers, C-22 (2):222–b, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672287.

Anonymous:1973:ICSa

[Ano73i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(2):222–a, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672286.

Anonymous:1973:ICSb

[Ano73j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(3):315–a, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672305.

Anonymous:1973:ICSc

[Ano73k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(4):432, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672334.

Anonymous:1973:ICSd

[Ano73l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-22(5):548, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672360.

Anonymous:1973:ITCd

[Ano73m] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-22(4):c1–432–b, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1973:ITCe

[Ano73n] Anonymous. IEEE Transactions on Computers. IEEE Transactions on Computers, C-22(5):c1–548–b, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1973:IAa

[Ano73o] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(1):112–a, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672202.

Anonymous:1973:IAb

[Ano73p] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(2):222–c, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672288.

Anonymous:1973:IAc

[Ano73q] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(3):315–b, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672306.

Anonymous:1973:IAd

[Ano73r] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(4):432–a, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672335.

Anonymous:1973:IAe

[Ano73s] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(5):548–a, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672361.

Anonymous:1973:IAf

[Ano73t] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(10):964–a, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672220.

Anonymous:1973:IAg

[Ano73u] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(11):1044–a, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672237.

Anonymous:1973:IAh

[Ano73v] Anonymous. Information for authors. IEEE Transactions on Computers, C-22(12):1144–m, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672259.

Anonymous:1973:PCI

[Ano73w] Anonymous. Parallel computation:

An introduction. IEEE Transactions on Computers, C-22(8):709– 710, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009148.

Anonymous:1973:Pa

[Ano73x] Anonymous. Peripherals. IEEE Transactions on Computers, C-22(11):1044–b, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672238.

Anonymous:1973:Pb

[Ano73y] Anonymous. Peripherals. IEEE Transactions on Computers, C-22(12):1144–n, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672260.

Anonymous:1973:RMJ

[Ano73z] Anonymous. Reply by M. James Errico. IEEE Transactions on Computers, C-22(12):1143–1144, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672257.

Anonymous:1974:IIT

[Ano74a] Anonymous. 1974 index IEEE Transactions on Computers vol. C-23. IEEE Transactions on Computers, C-23(12):1330–a–1330– n, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672461.

Anonymous:1974:CGWa

[Ano74b] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(5):560– b, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672585.

Anonymous:1974:CGWb

[Ano74c] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(6):648– b, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672602.

Anonymous:1974:CGWc

[Ano74d] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(7):752– b, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672620.

Anonymous:1974:CGWd

[Ano74e] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(8):880– b, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672641.

Anonymous:1974:CGWe

[Ano74f] Anonymous. Compcon goes to Washington. IEEE Transactions on Computers, C-23(9):992– b, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672669.

Anonymous:1974:ENa

[Ano74g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23 (4):337, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672539.

Anonymous:1974:ENb

[Ano74h] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23 (8):753, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672623.

Anonymous:1974:ENc

[Ano74i] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-23 (12):1217, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672432.

Anonymous:1974:ICSa

[Ano74j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(1):112–b, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672390.

Anonymous:1974:ICSc

[Ano74k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(2):224–b, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672506.

Anonymous:1974:ICSe

[Ano74l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(3):336–b, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672536.

Anonymous:1974:ICSg

[Ano74m] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(4):448–b, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672556.

Anonymous:1974:ICSn

[Ano74n] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(10):1104–b, October 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672407.

Anonymous:1974:ICSp

[Ano74o] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-23(11):1216–b, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672429.

Anonymous:1974:ICSb

[Ano74p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(2):224– a, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672505.

Anonymous:1974:ICSd

[Ano74q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(3): 336–a, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672535.

Anonymous:1974:ICSf

[Ano74r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(4): 448–a, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672555.

Anonymous:1974:ICSh

[Ano74s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(5): 560–a, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672584.

Anonymous:1974:ICSi

[Ano74t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(6): 648–a, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672601.

Anonymous:1974:ICSj

[Ano74u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(7): 752–a, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672619.

Anonymous:1974:ICSk

[Ano74v] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(8):880– a, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672640.

Anonymous:1974:ICSl

[Ano74w] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(9):992– a, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672668.

Anonymous:1974:ICSm

[Ano74x] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(10): 1104–a, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672406.

Anonymous:1974:ICSo

[Ano74y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(11): 1216–a, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672428.

Anonymous:1974:ICSq

[Ano74z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-23(12): 1330–o, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672462.

Anonymous:1974:IA

[Ano74-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-23(1):112–a, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672389.

Anonymous:1974:PGM

[Ano74-28] Anonymous. People are getting more important. IEEE Transactions on Computers, C-23(12): 1330–p, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672463.

Anonymous:1975:IIT

[Ano75a] Anonymous. 1975 index IEEE Transactions on Computers vol. C-24. IEEE Transactions on Computers, C-24(12):1242–a–1242– l, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672766.

Anonymous:1975:CNY

[Ano75b] Anonymous. Computers the next 5 years. IEEE Transactions on Computers, C-24(12):1242–m–1242–n, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672767.

Anonymous:1975:DEMa

[Ano75c] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(6):680–b, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672878.

Anonymous:1975:DEMb

[Ano75d] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(7):760–b, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672897.

Anonymous:1975:DEMc

[Ano75e] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(8):856–b, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672919.

Anonymous:1975:DEMd

[Ano75f] Anonymous. Does easier mean better? IEEE Transactions on Computers, C-24(9):952–b, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672941.

Anonymous:1975:ENa

[Ano75g] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-24 (3):225, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672795.

Anonymous:1975:ENb

[Ano75h] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-24 (9):857, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672922.

Anonymous:1975:ICSd

[Ano75i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(3):336–b, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672816.

Anonymous:1975:ICSf

[Ano75j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(4):456–b, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672836.

Anonymous:1975:ICSh

[Ano75k] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(5):584–b, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672860.

Anonymous:1975:ICSn

[Ano75l] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(10):1048–b, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672722.

Anonymous:1975:ICSp

[Ano75m] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(11):1144–b, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672747.

Anonymous:1975:ICSr

[Ano75n] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-24(12):1242–p, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672769.

Anonymous:1975:ICSa

[Ano75o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(1):112– a, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672689.

Anonymous:1975:ICSb

[Ano75p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(2):224– a, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672791.

Anonymous:1975:ICSc

[Ano75q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(3): 336–a, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672815.

Anonymous:1975:ICSe

[Ano75r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(4): 456–a, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672835.

Anonymous:1975:ICSg

[Ano75s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(5): 584–a, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672859.

Anonymous:1975:ICSi

[Ano75t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(6): 680–a, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672877.

Anonymous:1975:ICSj

[Ano75u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(7): 760–a, July 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672896.

Anonymous:1975:ICSk

[Ano75v] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(8):856– a, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672918.

Anonymous:1975:ICSl

[Ano75w] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(9):952– a, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672940.

Anonymous:1975:ICSm

[Ano75x] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(10): 1048–a, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672721.

Anonymous:1975:ICSo

[Ano75y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(11): 1144–a, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672746.

Anonymous:1975:ICSq

[Ano75z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-24(12): 1242–o, December 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672768.

Anonymous:1975:IA

[Ano75-27] Anonymous. Information for authors. IEEE Transactions on Computers, C-24(1):112, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672688.

Anonymous:1975:MRN

[Ano75-28] Anonymous. A method for reducing the number of input variables to synchronous sequential circuits. IEEE Transactions on Computers, C-24(10):1029–1031, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672713.

Anonymous:1975:PGMa

[Ano75-29] Anonymous. People are getting more important. IEEE Transactions on Computers, C-24(1):112– b, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672690.

Anonymous:1975:PGMb

[Ano75-30] Anonymous. People are getting more important. IEEE Transactions on Computers, C-24(2):224– b, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672792.

Anonymous:1976:LR

[Ano76a] Anonymous. 1975 list of reviewers. IEEE Transactions on Computers, C-25(6):668–672, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1976:IIT

[Ano76b] Anonymous. 1976 index IEEE Transactions on Computers vol. C-25. IEEE Transactions on Computers, C-25(12):1361, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674605.

Anonymous:1976:AICa

[Ano76c] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan. IEEE Transactions on Computers, C-25(5): 552, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674649.

Anonymous:1976:AICb

[Ano76d] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan. IEEE Transactions on Computers, C-25(8): 862, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674705.

Anonymous:1976:CLM

[Ano76e] Anonymous. Computers … by the millions, for the millions. IEEE Transactions on Computers, C-25 (8):862, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674708.

Anonymous:1976:EN

[Ano76f] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-25 (3):209, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009246.

Anonymous:1976:FI

[Ano76g] Anonymous. Foreword and introduction. IEEE Transactions on Computers, C-25(10):961–962, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674536.

Anonymous:1976:GEI

[Ano76h] Anonymous. Guidelines for Editors of the IEEE Computer Society. IEEE Transactions on Computers, C-25(11):1187–1188, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674580.

Anonymous:1976:ICSc

[Ano76i] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-25(5):552, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674651.

Anonymous:1976:ICSe

[Ano76j] Anonymous. IEEE Computer Society membership & publications. IEEE Transactions on Computers, C-25(6):672, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674673.

Anonymous:1976:ICSa

[Ano76k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(4): 456, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674628.

Anonymous:1976:ICSb

[Ano76l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(5): 552, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674650.

Anonymous:1976:ICSd

[Ano76m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(6): 672, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674672.

Anonymous:1976:ICSf

[Ano76n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(7): 768, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674691.

Anonymous:1976:ICSg

[Ano76o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(8): 862, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674707.

Anonymous:1976:ICSh

[Ano76p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(9): 960, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674724.

Anonymous:1976:ICSi

[Ano76q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(10): 1056, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674551.

Anonymous:1976:ICSj

[Ano76r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(11): 1188, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674582.

Anonymous:1976:ICSk

[Ano76s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-25(12): 1361, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674606.

Anonymous:1976:IYC

[Ano76t] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-25 (8):862, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674706.

Anonymous:1976:MAC

[Ano76u] Anonymous. Management auditing of computer operations. IEEE Transactions on Computers, C-25 (4):456, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674629.

Anonymous:1976:PPNa

[Ano76v] Anonymous. Papers to be published in the next issue August 1976. IEEE Transactions on Computers, C-25(7):768, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674692.

Anonymous:1976:PPNd

[Ano76w] Anonymous. Papers to be published in the next issue December 1976. IEEE Transactions on Computers, C-25(11):1188, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674583.

Anonymous:1976:PPNe

[Ano76x] Anonymous. Papers to be published in the next issue January 1977. IEEE Transactions on Computers, C-25(12):1361, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674607.

Anonymous:1976:PPNc

[Ano76y] Anonymous. Papers to be published in the next issue November 1976. IEEE Transactions on Computers, C-25(10):1056, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674552.

Anonymous:1976:PPNb

[Ano76z] Anonymous. Papers to be published in the next issue October 1976. IEEE Transactions on Computers, C-25(9):960, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674725.

Anonymous:1976:P

[Ano76-27] Anonymous. Preface. IEEE Transactions on Computers, C-25 (12):1189–1190, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1674586.

Anonymous:1976:SSA

[Ano76-28] Anonymous. Special sections announcement. IEEE Transactions on Computers, C-25(11):1188, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674581.

Anonymous:1977:IIT

[Ano77a] Anonymous. 1977 index IEEE Transactions on Computers vol. C-26. IEEE Transactions on Computers, C-26(12):1308, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674800.

Anonymous:1977:ACTa

[Ano77b] Anonymous. Announcing computer technology: Status, limits, alternatives. IEEE Transactions on Computers, C-26(11): 1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674775.

Anonymous:1977:ACTb

[Ano77c] Anonymous. Announcing computer technology: Status, limits, alternatives. IEEE Transactions on Computers, C-26(12): 1308, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674804.

Anonymous:1977:BCP

[Ano77d] Anonymous. Books for computer professionals. IEEE Transactions on Computers, C-26(12): 1308, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674802.

Anonymous:1977:CP

[Ano77e] Anonymous. Call for papers. IEEE Transactions on Computers, C-26 (3):320, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1977:CF

[Ano77f] Anonymous. Compcon 77 Fall.

IEEE Transactions on Computers, C-26(5):512, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674873.

Anonymous:1977:CAF

[Ano77g] Anonymous. Computer arithmetic: Foreword and survey. IEEE Transactions on Computers, C-26(7): 609, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674892.

Anonymous:1977:ENc

[Ano77h] Anonymous. Editor’s note. IEEE Transactions on Computers, C-26 (9):837, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674931.

Anonymous:1977:ENa

[Ano77i] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-26 (1):1, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009266.

Anonymous:1977:ENb

[Ano77j] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-26 (6):513, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674876.

Anonymous:1977:ICSa

[Ano77k] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(3): 320, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674837.

Anonymous:1977:ICSb

[Ano77l] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(4): 416, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674854.

Anonymous:1977:ICSc

[Ano77m] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(5): 512, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674872.

Anonymous:1977:ICSd

[Ano77n] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(6): 608, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674888.

Anonymous:1977:ICSe

[Ano77o] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(7): 704, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674904.

Anonymous:1977:ICSf

[Ano77p] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(8): 836, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674927.

Anonymous:1977:ICSg

[Ano77q] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(9): 932, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674943.

Anonymous:1977:ICSh

[Ano77r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(10): 1060, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674748.

Anonymous:1977:ICSi

[Ano77s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(11): 1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674774.

Anonymous:1977:ICSj

[Ano77t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-26(12): 1308, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674803.

Anonymous:1977:IYC

[Ano77u] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-26 (12):1308, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674801.

Anonymous:1977:MDC

[Ano77v] Anonymous. Maxwell’s displacement current. IEEE Transactions on Computers, C-26(3): 318, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674834. See [Cat77].

Anonymous:1977:PPNc

[Ano77w] Anonymous. Papers to be published in the next issue August 1977. IEEE Transactions on Computers, C-26(7):704, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674905.

Anonymous:1977:PPNb

[Ano77x] Anonymous. Papers to be published in the next issue July 1977. IEEE Transactions on Computers, C-26(6):608, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674889.

Anonymous:1977:PPNa

[Ano77y] Anonymous. Papers to be published in the next issue May 1977. IEEE Transactions on Computers, C-26(4):416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674855.

Anonymous:1977:PPNf

[Ano77z] Anonymous. Papers to be published in the next issue November 1977. IEEE Transactions on Computers, C-26(10):1060, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674749.

Anonymous:1977:PPNe

[Ano77-27] Anonymous. Papers to be published in the next issue October 1977. IEEE Transactions on Computers, C-26(9):932, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674944.

Anonymous:1977:PPNd

[Ano77-28] Anonymous. Papers to be published in the next issue September 1977. IEEE Transactions on Computers, C-26(8):836, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674928.

Anonymous:1978:LR

[Ano78a] Anonymous. 1976–1977 list of referees. IEEE Transactions on Computers, C-27(5):472–476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See correction [Ano78s].

Anonymous:1978:IIT

[Ano78b] Anonymous. 1978 index IEEE Transactions on Computers vol. C-27. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675036.

Anonymous:1978:APAa

[Ano78c] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(4): 384, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675113.

Anonymous:1978:APAc

[Ano78d] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(7): 679, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675171.

Anonymous:1978:APAd

[Ano78e] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-27(8): 782, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675194.

Anonymous:1978:APAb

[Ano78f] Anonymous. Advance program announcements. IEEE Transactions on Computers, C-27(6): 582, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675154.

Anonymous:1978:A

[Ano78g] Anonymous. Announcement. IEEE Transactions on Computers, C-27 (7):585, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675156.

Anonymous:1978:ANQa

[Ano78h] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27 (9):879, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675210.

Anonymous:1978:ANQb

[Ano78i] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27 (10):973, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674980.

Anonymous:1978:ANQc

[Ano78j] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27 (11):1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675005.

Anonymous:1978:BCP

[Ano78k] Anonymous. Books for computer professionals. IEEE Transactions on Computers, C-27(5): 476, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675132.

Anonymous:1978:CPa

[Ano78l] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (1):96, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPb

[Ano78m] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (2):192, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPc

[Ano78n] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (8):782, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPd

[Ano78o] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (9):879, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPe

[Ano78p] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (10):973, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPf

[Ano78q] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (11):1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CPg

[Ano78r] Anonymous. Call for papers. IEEE Transactions on Computers, C-27 (12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1978:CLR

[Ano78s] Anonymous. Correction to “1976– 1977 List of Referees”. IEEE Transactions on Computers, C-27(9): 879, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675209. See [Ano78a].

Anonymous:1978:CFC

[Ano78t] Anonymous. Correction to front cover of December 1977 issue. IEEE Transactions on Computers, C-27 (3):288, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675095.

Anonymous:1978:EN

[Ano78u] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-27 (5):385, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675116.

Anonymous:1978:ICSj

[Ano78v] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(10): 973, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674982.

Anonymous:1978:ICSl

[Ano78w] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(11):1089, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675006.

Anonymous:1978:ICSn

[Ano78x] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Computers, C-27(12):1227, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675037.

Anonymous:1978:ICSa

[Ano78y] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(1): 96, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674962.

Anonymous:1978:ICSb

[Ano78z] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(2): 192, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675067.

Anonymous:1978:ICSc

[Ano78-27] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(3): 288, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675096.

Anonymous:1978:ICSd

[Ano78-28] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(4): 384, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675112.

Anonymous:1978:ICSe

[Ano78-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(5): 476, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675133.

Anonymous:1978:ICSf

[Ano78-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(6): 582, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675153.

Anonymous:1978:ICSg

[Ano78-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(7): 679, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675170.

Anonymous:1978:ICSh

[Ano78-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(8): 782, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675195.

Anonymous:1978:ICSi

[Ano78-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(9): 879, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675211.

Anonymous:1978:ICSk

[Ano78-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(10): 973, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674983.

Anonymous:1978:ICSm

[Ano78-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(11): 1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675007.

Anonymous:1978:ICSo

[Ano78-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-27(12): 1227, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675039.

Anonymous:1978:ITPa

[Ano78-37] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27(10): 973, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674981.

Anonymous:1978:ITPb

[Ano78-38] Anonymous. IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-27(11): 1089, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675004.

Anonymous:1978:ITSa

[Ano78-39] Anonymous. IEEE Transactions on Software Engineering. IEEE Transactions on Computers, C-27 (3):288, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675097.

Anonymous:1978:ITSb

[Ano78-40] Anonymous. IEEE Transactions On Software Engineering. IEEE Transactions on Computers, C-27 (7):679, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675169.

Anonymous:1978:ITSc

[Ano78-41] Anonymous. IEEE Transactions On Software Engineering. IEEE Transactions on Computers, C-27 (8):782, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675193.

Anonymous:1978:IYCa

[Ano78-42] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Computers, C-27 (6):582, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675152.

Anonymous:1978:IYCb

[Ano78-43] Anonymous. if your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-27 (12):1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675038.

Anonymous:1978:IA

[Ano78-44] Anonymous. Information for authors. IEEE Transactions on Computers, C-27(5):476, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1675134.

Anonymous:1979:IIT

[Ano79a] Anonymous. 1979 index IEEE Transactions on Computers vol. C-28. IEEE Transactions on Computers, C-28(12):947, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675290.

Anonymous:1979:PCF

[Ano79b] Anonymous. 254: Proceedings: Compcon 79 Fall “Using Microprocessors — Extending our Reach”. IEEE Transactions on Computers, C-28(12):947– o, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675293.

Anonymous:1979:ICPa

[Ano79c] Anonymous. 5th international conference on pattern recognition. IEEE Transactions on Computers, C-28(10):805, October 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:ICPb

[Ano79d] Anonymous. 5th international conference on pattern recognition. IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:APAa

[Ano79e] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(3): 281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675343.

Anonymous:1979:APAb

[Ano79f] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(4): 329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675358.

Anonymous:1979:APAc

[Ano79g] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(5): 378, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675374.

Anonymous:1979:APAd

[Ano79h] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(6): 458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675384.

Anonymous:1979:APAe

[Ano79i] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(7): 543, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675404.

Anonymous:1979:APAf

[Ano79j] Anonymous. Advance program announcement. IEEE Transactions on Computers, C-28(8): 590, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675419.

Anonymous:1979:ANQa

[Ano79k] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-28 (5):378, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675371.

Anonymous:1979:ANQb

[Ano79l] Anonymous. Announcing a new quarterly IEEE Transactions: IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Computers, C-28 (9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675444.

Anonymous:1979:AICa

[Ano79m] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(6): 458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675387.

Anonymous:1979:AICb

[Ano79n] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(7): 543, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675406.

Anonymous:1979:AICc

[Ano79o] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(8): 590, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675422.

Anonymous:1979:AICd

[Ano79p] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675447.

Anonymous:1979:AICe

[Ano79q] Anonymous. Announcing: The IEEE Computer Society’s Tutorial Week 79. IEEE Transactions on Computers, C-28(10): 805, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675257.

Anonymous:1979:CPa

[Ano79r] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (2):175, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPb

[Ano79s] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (2):175, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPc

[Ano79t] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (3):281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPd

[Ano79u] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (3):281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPe

[Ano79v] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (4):329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPf

[Ano79w] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (4):329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPg

[Ano79x] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (6):458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPh

[Ano79y] Anonymous. call for papers. IEEE Transactions on Computers, C-28 (8):590, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPi

[Ano79z] Anonymous. Call for papers. IEEE Transactions on Computers, C-28 (9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPj

[Ano79-27] Anonymous. Call for papers. IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:CPk

[Ano79-28] Anonymous. Call for papers. IEEE Transactions on Computers, C-28(11):875, November 1979. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:ENa

[Ano79-29] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-28(1):1, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675215.

Anonymous:1979:ENb

[Ano79-30] Anonymous. Editor’s notice. IEEE Transactions on Computers, C-28(3):177, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675316.

Anonymous:1979:ENc

[Ano79-31] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(4):285, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675346.

Anonymous:1979:ENd

[Ano79-32] Anonymous. Editor’s notice.

IEEE Transactions on Computers, C-28(8):545, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675409.

Anonymous:1979:ICSa

[Ano79-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(1): 86, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675233.

Anonymous:1979:ICSb

[Ano79-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(2): 175, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675312.

Anonymous:1979:ICSc

[Ano79-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(3): 281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675342.

Anonymous:1979:ICSd

[Ano79-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(4): 329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675359.

Anonymous:1979:ICSe

[Ano79-37] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(5): 378, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675373.

Anonymous:1979:ICSf

[Ano79-38] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(6): 458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675386.

Anonymous:1979:ICSg

[Ano79-39] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(7): 543, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675405.

Anonymous:1979:ICSh

[Ano79-40] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(8): 590, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675421.

Anonymous:1979:ICSi

[Ano79-41] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(9): 693, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675446.

Anonymous:1979:ICSj

[Ano79-42] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(10): 805, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675256.

Anonymous:1979:ICSk

[Ano79-43] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(11): 875, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675274.

Anonymous:1979:ICSl

[Ano79-44] Anonymous. IEEE Computer Society publications. IEEE Transactions on Computers, C-28(12): 947, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675294.

Anonymous:1979:ICFa

[Ano79-45] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-28(1):86, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1675232.

Anonymous:1979:ICFb

[Ano79-46] Anonymous. IEEE copyright form. IEEE Transactions on Computers, C-28(12):947, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675291.

Anonymous:1979:IYCa

[Ano79-47] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28 (3):281, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675340.

Anonymous:1979:IYCb

[Ano79-48] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28 (4):329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675357.

Anonymous:1979:IYCc

[Ano79-49] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28 (5):378, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675372.

Anonymous:1979:IYCd

[Ano79-50] Anonymous. If your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Computers, C-28 (9):693, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675443.

Anonymous:1979:IA

[Ano79-51] Anonymous. Information for authors. IEEE Transactions on Computers, C-28(1):86, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675234.

Anonymous:1979:LR

[Ano79-52] Anonymous. List of referees. IEEE Transactions on Computers, C-28(12):945–947, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:1979:NAI

[Ano79-53] Anonymous. Now available from the IEEE Computer Society. IEEE Transactions on Computers, C-28(12):947–q, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675295.

Anonymous:1979:PCS

[Ano79-54] Anonymous. Proceedings of the Conference on Specifications of Reliable Software. IEEE Transactions on Computers, C-28(12):947– n, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675292.

Ahmed:1974:DCT

[ANR74] N. Ahmed, T. Natarajan, and K. R. Rao. Discrete Cosine Transfom. IEEE Transactions on Computers, C-23(1):90– 93, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672377. See comments [Sha75].

Anzelmo:1971:DSF

[Anz71] F. D. Anzelmo. A datastorage format for information system files. IEEE Transactions on Computers, C-20(1):39– 43, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671672.

Atkins:1979:TCC

[AO79] D. E. Atkins and Shau-Chi Ong.

Time-component complexity of two approaches to multioperand binary addition. IEEE Transactions on Computers, C-28(12):918– 926, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675281.

Aminoff:1974:CRA

[AOR74] C. G. Aminoff, L. Ojala, and E. T.

Rautanen. On a class of recursive algorithms for continuous estimation of the mean. IEEE Transactions on Computers, C-23(2):191– 194, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672479.

Andrews:1976:OPE

[AP76a] Harry C. Andrews and Claude L. Patterson. Outer product expansions and their uses in digital image processing. IEEE Transactions on Computers, C-25(2):140– 148, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009228.

Andrews:1976:DID

[AP76b] Harry C. Andrews and Claude L. Patterson III. Digital interpolation of discrete images. IEEE Transactions on Computers, C-25(2):196– 202, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009235.

Agrawala:1974:MPS

[AR74] A. K. Agrawala and T. G. Rauscher. Microprogramming: Perspective and status. IEEE Transactions on Computers, C-23(8):817– 837, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672631.

Ashjaee:1977:TSC

[AR77] M. J. Ashjaee and S. M. Reddy.

On totally self-checking checkers for separable codes. IEEE Transactions on Computers, C-26(8):737– 744, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674911.

Agrawal:1978:MOA

[AR78] D. P. Agrawal and T. R. N.

Rao. On multiple operand addition of signed binary numbers. IEEE Transactions on Computers, C-27(11):1068–1070, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674998.

Arbib:1970:RMT

[Arb70] M. A. Arbib. R70-10 the minimalization of tree automata. IEEE Transactions on Computers, C-19 (6):563, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671574.

Arbib:1976:AIC

[Arb76] M. A. Arbib. Artificial intelligence: Cooperative computation and man–machine symbiosis. IEEE Transactions on Computers, C-25(12):1346–1352, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674603.

Ari:1979:TLM

[Ari79] M. B. Ari. On transposing large 2^{n} × 2^{n} matrices. IEEE Transactions on Computers, C-28(1):72– 75, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675228.

Armstrong:1972:DMS

[Arm72] D. B. Armstrong. A deductive method for simulating faults in logic circuits. IEEE Transactions on Computers, C-21(5):464– 471, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672135.

Armstrong:1973:DGG

[Arm73] J. R. Armstrong. Design of a graphic generator for remote terminal application. IEEE Transactions on Computers, C-22(5): 464–468, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672342.

Arnold:1970:RTT

[Arn70] T. F. Arnold. R70-21 time and tape complexity of pushdown automaton languages. IEEE Transactions on Computers, C-19(7): 658–659, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671601.

Arnold:1973:CCE

[Arn73] T. F. Arnold. The concept of coverage and its effect on the reliability model of a repairable system. IEEE Transactions on Computers, C-22 (3):251–254, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672296.

Ahlgren:1971:CRA

[ARS71a] R. C. Ahlgren, H. F. Ryan, and C. W. Swonger. A character recognition application of an iterative procedure for feature selection. IEEE Transactions on Computers, C-20(9):1067– 1075, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671998.

Ahmed:1971:FCB

[ARS71b] N. Ahmed, K. R. Rao, and R. B. Schultz. Fast complex BIFORE transform by matrix partitioning. IEEE Transactions on Computers, C-20(6):707– 710, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671926.

Abraham:1974:AAR

[AS74] J. A. Abraham and D. P. Siewiorek. An algorithm for the accurate reliability evaluation of triple modular redundancy networks. IEEE Transactions on Computers, C-23 (7):682–692, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672609.

Arguello:1972:TFC

[ASS72] R. J. Arguello, H. R. Sellner, and J. A. Stuller. Transfer function compensation of sampled imagery. IEEE Transactions on Computers, C-21(7):812– 818, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672179.

Asthana:1978:DCT

[Ast78] A. Asthana. Design and control of a three-stage switch matrix in the presence of fan-out. IEEE Transactions on Computers, C-27(10): 886–895, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674967.

Avizienis:1970:UAB

[AT70] Algirdas Avižienis and Chin Tung.

A universal arithmetic building element (ABE) and design methods for arithmetic processors. IEEE Transactions on Computers, C-19(8):733–745, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Avizienis.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671616.

Athani:1975:VAC

[Ath75] V. V. Athani. The 16-vector algorithm for computer controlled digital X–Y plotter. IEEE Transactions on Computers, C-24(8):831– 835, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672909.

Atkins:1970:DAU

[Atk70] D. E. Atkins. Design of the arithmetic units of ILLIAC III: Use of redundancy and higher radix methods. IEEE Transactions on Computers, C-19(8):720–733, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Atkins.
pdf;   https://ieeexplore.ieee.

                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671615.

Arnold:1970:IRS

[ATN70] Thomas F. Arnold, Chung-Jen Tan, and Monroe M. Newborn. Iteratively realized sequential circuits. IEEE Transactions on Computers, C-19(1):54– 66, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008900.

Aho:1974:DMR

[AU74] A. V. Aho and J. D. Ullman. Dynamic memories with rapid random and sequential access. IEEE Transactions on Computers, C-23 (3):272–276, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672515.

Austin:1970:UMP

[Aus70] B. J. Austin. Use of a macro processor in logical design. IEEE Transactions on Computers, C-19 (11):1085–1089, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671429.

Avizienis:1971:AEC

[Avi71] A. Avizienis. Arithmetic error codes: Cost and effectiveness studies for application in digital system design. IEEE Transactions on Computers, C-20(11):1322–1331, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671727.

Avizienis:1973:AAE

[Avi73] Algirdas Avižienis. Arithmetic algorithms for error-coded operands. IEEE Transactions on Computers, C-22(6):567–572, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Avizienis.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009108.

Aviziens:1976:FTS

[Avi76] A. Aviziens. Fault-tolerant systems. IEEE Transactions on Computers, C-25(12):1304–1312, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674598.

Bruno:1971:TAC

[BA71] J. Bruno and S. M. Altman.

A theory of asynchronous control networks. IEEE Transactions on Computers, C-20(6):629– 638, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671907.

Babu:1972:EPF

[Bab72] C. Chitti Babu. On the extraction of pattern features from imperfectly identified samples. IEEE Transactions on Computers, C-21 (4):410–411, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

5008992. See comments [Tou74].

Babu:1974:AR

[Bab74] C. C. Babu. Author’s reply. IEEE Transactions on Computers, C-23 (9):989, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672663.

Baer:1974:PPD

[Bae74] J.-L. Baer. On program placement in a directly executable hierarchy of memories. IEEE Transactions on Computers, C-23(8):838– 849, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672632.

Baer:1976:MS

[Bae76] J.-L. Baer. Multiprocessing systems. IEEE Transactions on Computers, C-25(12):1271–1277, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674594.

Baker:1975:MER

[Bak75a] P. W. Baker. More efficient radix-2 algorithms for some elementary functions. IEEE Transactions on Computers, C-24(11):1049–1054, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672725.

Baker:1975:PMA

[Bak75b] P. W. Baker. Parallel multiplicative algorithms for some elementary functions. IEEE Transactions on Computers, C-24(3):322– 325, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672808.

Baker:1976:SFB

[Bak76] P. W. Baker. Suggestion for a fast binary sine/cosine generator. IEEE Transactions on Computers, C-25(11):1134–1136, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674566.

Baker:1979:SDE

[Bak79] P. W. Baker. The solution of differential equations on short-word-length computing devices. IEEE Transactions on Computers, C-28 (3):205–214, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675321.

Bancilhon:1974:GMS

[Ban74a] F. Bancilhon. A geometric model for stochastic automata. IEEE Transactions on Computers, C-23(12):1290–1299, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672441.

Banerji:1974:NIM

[Ban74b] D. K. Banerji. A novel implementation method for addition and subtraction in residue number systems. IEEE Transactions on Computers, C-23(1):106– 109, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672383;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35077.

Banerji:1974:URA

[Ban74c] D. K. Banerji. On the use of residue arithmetic for computation. IEEE Transactions on Computers, C-23(12):1315–1317, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672448.

Banerji:1978:CBM

[Ban78] R. B. Banerji. The construction of binary matrices with distinct submatrices. IEEE Transactions on Computers, C-27(2):162– 164, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675050.

Bargainer:1970:RUT

[Bar70] J. D. Bargainer. R70-34 unateness test of a Boolean function and two general synthesis methods using threshold logic. IEEE Transactions on Computers, C-19 (9):858, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671655.

Barnes:1971:TWA

[Bar71] B. H. Barnes. A two-way automaton with fewer states than any equivalent one-way automaton. IEEE Transactions on Computers, C-20 (4):474–475, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671866.

Barnard:1974:BPD

[Bar74a] A. C. L. Barnard. B 74-33 peripheral devices. IEEE Transactions on Computers, C-23(5):

559, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672582. See comments [Flo75].

Barnard:1974:BCS

[Bar74b] A. C. L. Barnard. B74-18 computer system organization. IEEE Transactions on Computers, C-23 (2):220, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672500.

Barnes:1974:BIM

[Bar74c] E. R. Barnes. B74-15 iterative methods for nonlinear optimization problems. IEEE Transactions on Computers, C-23(2): 218, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672497.

Barbacci:1975:CRT

[Bar75a] M. R. Barbacci. A comparison of register transfer languages for describing computers and digital systems. IEEE Transactions on Computers, C-24(2):137– 150, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672774.

Barnard:1975:RR

[Bar75b] A. C. L. Barnard. Reviewer’s reply. IEEE Transactions on Computers, C-24(2):224, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672790.

Barak:1977:MAT

[Bar77] A. B. Barak. Multiplicative algorithms for ternary arithmetic using binary logic. IEEE Transactions on Computers, C-26(8):823– 826, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674922.

Barnes:1978:CIM

[Bar78] G. E. Barnes. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3):286– 287, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675090. See [DT76, Das78, Ada78].

Bashkow:1971:BIC

[Bas71a] T. R. Bashkow. B71-2 introduction to computer organization. IEEE Transactions on Computers, C-20 (10):1221, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671704.

Bashkow:1971:BSS

[Bas71b] T. R. Bashkow. B71-6 system structure in data, programs, and computers. IEEE Transactions on Computers, C-20(12):1622, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671786.

Bashkow:1973:CRO

[Bas73] T. R. Bashkow. Comment on review of operating systems survey. IEEE Transactions on Computers, C-22 (12):1143, December 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672256.

Batra:1970:DAU

[Bat70] V. Batra. Design of asynchronous unit delays. IEEE Transactions on Computers, C-19(10):896– 902, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671389.

Batcher:1977:MAM

[Bat77] Kenneth E. Batcher. The multidimensional access memory in STARAN. IEEE Transactions on Computers, C-26(2):174–177, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009297.

Baugh:1971:BNP

[Bau71] C. R. Baugh. Bounds on the number of pseudothreshold functions. IEEE Transactions on Computers, C-20(12):1602–1605, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671774.

Baugh:1972:GRF

[Bau72] C. R. Baugh. Generation of representative functions of the NPN equivalence classes of unate Boolean functions. IEEE Transactions on Computers, C-21(12):1373–1379, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672102.

Bauer:1977:CFP

[Bau77] Henry R. Bauer. Comments on “On the Floating Point Representation of Complex Numbers”. IEEE Transactions on Computers, C-26 (2):191, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009301. See [Yue75c].

Banerji:1972:TAR

[BB72] D. K. Banerji and J. A. Brzozowski. On translation algorithms in residue number systems. IEEE Transactions on Computers, C-21(12):1281–1285, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672092;   https://ieeexplore.
                                        ieee.org/xpl/tocresult.jsp?isnumber=
                                        35066.

Butler:1973:SCU

[BB73] J. T. Butler and K. J. Breeding. Some characteristics of universal cell nets. IEEE Transactions on Computers, C-22(10):897– 903, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672208.

Bracha-Barak:1974:ACF

[BB74a] A. Bracha-Barak. Application of continued fractions for fast evaluation of certain functions on a digital computer. IEEE Transactions on Computers, C-23(3):301– 309, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672519.

Bracha-Barak:1974:MSP

[BB74b] A. Bracha-Barak. A method for solving polynomial equations by continued fractions. IEEE Transactions on Computers, C-23(10): 1093–1097, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672402.

Brubaker:1975:MUL

[BB75] T. A. Brubaker and J. C.

Becker. Multiplication using logarithms implemented with readonly memory. IEEE Transactions on Computers, C-24(8):761– 765, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672900.

Ben-Bassat:1978:IFP

[BB78a] M. Ben-Bassat. Irrelevant features in pattern recognition. IEEE Transactions on Computers, C-27 (8):746–749, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675182.

Ben-Bassat:1978:MPS

[BB78b] M. Ben-Bassat. Myopic policies in sequential classification. IEEE Transactions on Computers, C-27(2):170–174, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675054.

Bender:1978:AAN

[BB78c] E. A. Bender and J. T. Butler. Asymptotic approximations for the number of fanout-free functions. IEEE Transactions on Computers, C-27(12):1180–1183, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675021.

Bandyopadhyay:1972:CPA

[BBC72a] S. Bandyopadhyay, S. Basu, and A. K. Choudhury. A cellular permuter array. IEEE Transactions on Computers, C-21(10): 1116–1119, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672052.

Bandyopadhyay:1972:IAM

[BBC72b] S. Bandyopadhyay, S. Basu, and A. K. Choudhury. An iterative array for multiplication of signed binary numbers. IEEE Transactions on Computers, C-21(8):921– 922, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009055.

Barr:1973:ROD

[BBLT73] R. G. Barr, J. A. Becker, W. P. Lidinsky, and V. V. Tantillo. A research-oriented dynamic microprocessor. IEEE Transactions on Computers, C-22(11):976–985, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672227.

Biermann:1975:SSP

[BBP75] A. W. Biermann, R. I. Baum, and F. E. Petry. Speeding up the synthesis of programs from traces. IEEE Transactions on Computers, C-24(2):122–136, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672773.

Babu:1971:APC

[BC71a] C. C. Babu and Wah-Chun Chan.

An algorithm for pattern classification using eigenvectors. IEEE Transactions on Computers, C-20 (5):575–578, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671886.

Bearnson:1971:DML

[BC71b] L. W. Bearnson and C. C. Carroll. On the design of minimum length fault tests for combinational circuits. IEEE Transactions on Computers, C-20(11):1353–1356, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671730.

Benowitz:1975:AFI

[BCA^{+}75] N. Benowitz, D. F. Calhoun, G. E. Alderson, J. E. Bauer, and C. T. Joeckel. An advanced fault isolation system for digital logic. IEEE Transactions on Computers, C-24 (5):489–497, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672844.

Bell:1974:IAC

[BCB74] J. Bell, D. Casasent, and C. G. Bell. An investigation of alternative cache organizations. IEEE Transactions on Computers, C-23 (4):346–351, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672541.

Bossen:1978:MGE

[BCC78] D. C. Bossen, L. C. Chang, and Chin-Long Chen. Measurement and generation of error correcting codes for package failures. IEEE Transactions on Computers, C-27(3):201– 204, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675072.

Bouricius:1973:MBM

[BCH^{+}73] W. G. Bouricius, W. C. Carter, E. P. Hsieh, D. C. Jessep, Jr., and A. B. Wadia. Modeling of a bubble-memory organization with self-checking translators to achieve high reliability. IEEE Transactions on Computers, C-22(3):269– 275, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672299.

Browne:1973:ETM

[BCHL73] James C. Browne, K. M. Chandy, John Hogarth, and Chester C.-A. Lee. The effect on throughput of multiprocessing in a multiprogramming environment. IEEE Transactions on Computers, C-22(8):728– 735, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009151.

Bouricius:1971:RMF

[BCJ^{+}71] W. G. Bouricius, W. C. Carter, D. C. Jessep, P. R. Schneider, and A. B. Wadia. Reliability modeling for fault-tolerant computers. IEEE Transactions on Computers, C-20(11):1306–1311, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671725.

Banerjee:1979:TPP

[BCKT79] Utrpal Banerjee, Shyh-Ching Chen, David J. Kuck, and Ross A. Towle. Time and parallel processor bounds for FORTRANlike loops. IEEE Transactions on Computers, C-28(9):660–670, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675434.

Borucki:1975:MUC

[BCL75] W. J. Borucki, D. H. Card, and G. C. Lyle. A method of using cluster analysis to study statistical dependence in multivariate data. IEEE Transactions on Computers, C-24(12):1183–1191, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672755.

Breuer:1976:IMS

[BCS76] Melvin A. Breuer, Shih-Jeh Chang, and Stephen Y. H. Su. Identification of multiple stuck-type faults in combinational networks. IEEE Transactions on Computers, C-25 (1):44–54, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009204.

Baugh:1972:ONN

[BCSM72] Charles R. Baugh, C. S. Chandersekaran, Richard S. Swee, and Saburo Muroga. Optimal networks of NOR–OR gates for functions of three variables. IEEE Transactions on Computers, C-21(2):153– 160, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008920.

Biss:1976:SMD

[BCSW76] K. Biss, R. T. Chien, F. Stahl, and S. J. Weissman. Semantic modeling for deductive question-answering. IEEE Transactions on Computers, C-25(4):358– 366, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674618.

Bastien:1971:GTP

[BD71] P. L. Bastien and L. A. Dunn.

Global transformations in pattern recognition of bubble chamber photographs. IEEE Transactions on Computers, C-20(9):995– 1001, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671987.

Bennett:1975:MCQ

[BD75a] J. R. Bennett and J. S. Mac Donald.

On the measurement of curvature in a quantized environment. IEEE Transactions on Computers, C-24 (8):803–820, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672905.

Bezdek:1975:OFP

[BD75b] J. C. Bezdek and J. C. Dunn. Optimal fuzzy partitions: a heuristic for estimating the parameters in a mixture of normal distributions. IEEE Transactions on Computers, C-24 (8):835–838, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672910.

Briggs:1977:OSM

[BD77] Faye A. Briggs and Edward S. Davidson. Organization of semiconductor memories for parallelpipelined processors. IEEE Transactions on Computers, C-26(2): 162–169, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009295.

Brenner:1970:SPD

[BdB70] A. E. Brenner and P. de Bruyne. A sonic pen: a digital stylus system. IEEE Transactions on Computers, C-19(6):546– 548, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671566.

Blumenthal:1977:DNP

[BDR77] Andrew F. Blumenthal, Larry S. Davis, and Azriel Rosenfeld. Detecting natural “Plateaus” in onedimensional patterns. IEEE Transactions on Computers, C-26(2): 178–179, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009298.

Bovet:1970:DMA

[BE70a] D. P. Bovet and G. Estrin. A dynamic memory allocation algorithm. IEEE Transactions on Computers, C-19(5):403–411, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671531.

Bovet:1970:SMA

[BE70b] D. P. Bovet and G. Estrin. On static memory allocation in computer systems. IEEE Transactions on Computers, C-19(6):492– 503, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671559.

Beaudry:1978:PRR

[Bea78] M. D. Beaudry. Performancerelated reliability measures for computing systems. IEEE Transactions on Computers, C-27(6):540– 547, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675145.

Beck:1971:SFP

[Bec71] F. Beck. The special features of POLLY: How they are used for automatic scanning. IEEE Transactions on Computers, C-20(9):1002– 1006, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671988.

Book:1971:AGE

[BEGO71] R. Book, S. Even, S. Greibach, and G. Ott. Ambiguity in graphs and expressions. IEEE Transactions on Computers, C-20(2):149– 153, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671797.

Bell:1972:DUR

[BEGW72] C. G. Bell, J. L. Eggert, J. Grason, and P. Williams. The description and use of register-transfer modules (RTM’s)(R). IEEE Transactions on Computers, C-21(5): 495–500, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672141.

Beizer:1970:TNT

[Bei70] B. Beizer. Towards a new theory of sequential switching networks. IEEE Transactions on Computers, C-19(10):939–956, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671394. See correction [Bei71].

Beizer:1971:CTN

[Bei71] B. Beizer. Correction to “Towards a New Theory of Sequential Switching Networks”. IEEE Transactions on Computers, C-20 (8):943, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671972. See [Bei70].

Belser:1976:CIA

[Bel76] Karl Belser. Comment on “An improved algorithm for the generation of nonparametric curves”. IEEE Transactions on Computers, C-25 (1):103, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009217. See [JLH73].

Benyon:1970:RCM

[Ben70] P. R. Benyon. R70-25 construction of multistep integration formulas for simulation purposes. IEEE Transactions on Computers, C-19 (8):763–764, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671623.

Bennetts:1971:IMP

[Ben71] R. G. Bennetts. An improved method of prime C-class derivation in the state reduction of sequential networks. IEEE Transactions on Computers, C-20(2):229– 231, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671814.

Benedek:1977:DLB

[Ben77] M. Benedek. Developing large binary to BCD conversion structures. IEEE Transactions on Computers, C-26(7):688–700, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1674902.

Bergman:1970:RTG

[Ber70a] S. Bergman. R70-27 tree generating regular systems. IEEE Transactions on Computers, C-19(8): 764, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671625.

Berndt:1970:FML

[Ber70b] H. Berndt. Functional microprogramming as a logic design aid. IEEE Transactions on Computers, C-19(10):902–907, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671390.

Berkling:1971:CMB

[Ber71] K. J. Berkling. A computing machine based on tree structures. IEEE Transactions on Computers, C-20(4):404–418, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671849.

Bergland:1972:PIF

[Ber72] G. D. Bergland. A parallel implementation of the Fast Fourier Transform algorithm. IEEE Transactions on Computers, C-21(4): 366–370, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008977.

Berges:1977:UPN

[Ber77] George A. Berges. Use of postfix notation for the description of microcellular arrays. IEEE Transactions on Computers, C-26(2):185– 191, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009300.

Betancourt:1971:DMT

[Bet71] R. Betancourt. Derivation of minimum test sets for unate logical circuits. IEEE Transactions on Computers, C-20(11):1264–1269, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671719.

Biermann:1972:SFS

[BF72] A. W. Biermann and J. A. Feldman. On the synthesis of finitestate machines from samples of their behavior. IEEE Transactions on Computers, C-21(6):592– 597, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009015.

Birk:1975:AMD

[BF75a] J. E. Birk and D. E. Farmer. An algebraic method for designing multivalued logic circuits using principally binary components. IEEE Transactions on Computers, C-24(11):1101–1104, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672731.

Borgerson:1975:RMG

[BF75b] B. R. Borgerson and R. F. Freitas. A reliability model for gracefully degrading and standbysparing systems. IEEE Transactions on Computers, C-24(5):517– 525, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672848.

Bentley:1978:FAC

[BF78] Jon Louis Bentley and Jerome H.

Friedman. Fast algorithms for constructing minimal spanning trees in coordinate spaces. IEEE Transactions on Computers, C-27(2):97– 105, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675043.

Bates:1975:NOP

[BG75] R. H. T. Bates and P. T.

Gough. New outlook on processing radiation received from objects viewed through randomly fluctuating media. IEEE Transactions on Computers, C-24(4):449– 456, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672834.

Breitbart:1978:AAE

[BG78] Y. Breitbart and S. Gal. Analysis of algorithms for the evaluation of monotonic Boolean functions. IEEE Transactions on Computers, C-27(11):1083–1087, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675002.

Braun:1979:IAD

[BG79] R. D. Braun and D. D. Givone. An improved algorithm for deriving checking experiments. IEEE Transactions on Computers, C-28(2): 153–156, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675306.

Barsi:1976:TDD

[BGM76] F. Barsi, F. Grandoni, and P. Maestrini. A theory of diagnosability of digital systems. IEEE Transactions on Computers, C-25 (6):585–593, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674658.

Bossen:1971:CEA

[BH71a] D. C. Bossen and Se June Hong. Cause-effect analysis for multiple fault detection in combinational networks. IEEE Transactions on Computers, C-20(11):1252–1257, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671717. See comments [LB74].

Bredeson:1971:GCP

[BH71b] J. G. Bredeson and P. T. Hulina. Generation of a clock pulse for asynchronous sequential machines to eliminate critical races.

IEEE Transactions on Computers, C-20(2):225–226, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671812.

Bredeson:1971:GPI

[BH71c] J. G. Bredeson and P. T. Hultina.

Generation of prime implicants by direct multiplication. IEEE Transactions on Computers, C-20(4): 475–476, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671867.

Bredeson:1973:SMI

[BH73] J. G. Bredeson and P. T. Hulina.

Synthesis of multiple-input change asynchronous circuits using transitionsensitive flip-flops. IEEE Transactions on Computers, C-22(5): 524–531, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672349. See comments [Bre75].

Bossen:1974:AR

[BH74a] D. C. Bossen and S. J. Hong. Authors’ reply. IEEE Transactions on Computers, C-23(12):1326–1327, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672453.

Breuer:1974:PES

[BH74b] M. A. Breuer and R. L. Harrison. Procedures for eliminating static and dynamic hazards in test generation. IEEE Transactions on Computers, C-23(10): 1069–1078, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672400. See comments [Hla78].

Baum:1976:DCS

[BH76] R. I. Baum and D. K. Hsiao. Database computers — a step towards data utilities. IEEE Transactions on Computers, C-25 (12):1254–1259, December 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1674592.

Bhandarkar:1975:AMI

[Bha75a] D. P. Bhandarkar. Analysis of memory interference in multiprocessors. IEEE Transactions on Computers, C-24(9):897–908, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672928.

Bhandarkar:1975:PMB

[Bha75b] D. P. Bhandarkar. On the performance of magnetic bubble memories in computer systems. IEEE Transactions on Computers, C-24(11):1125–1129, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672738.

Bhandarkar:1977:SPI

[Bha77] D. P. Bhandarkar. Some performance issues in multiprocessor system design. IEEE Transactions on Computers, C-26(5):506– 511, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674868.

Bhargava:1978:SCR

[Bha78] V. K. Bhargava. Some codes of Rahman and Blake for computer applications. IEEE Transactions on Computers, C-27(8):765– 767, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675188.

Banerjee:1979:DDC

[BHK79] J. Banerjee, D. K. Hsiao, and K. Kannan. DBC — a database computer for very large databases. IEEE Transactions on Computers, C-28(6):414–429, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675381.

Bouricius:1971:ADF

[BHP^{+}71] W. G. Bouricius, E. P. Hsieh, G. R. Putzolu, J. P. Roth, P. R. Schneider, and C. Tan. Algorithms for detection of faults in logic circuits. IEEE Transactions on Computers, C-20(11):1258–1264, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671718.

Beltrami:1972:ARS

[BI72] E. J. Beltrami and J. P. Indusi. An adaptive random search algorithm for constrained minimization. IEEE Transactions on Computers, C-21(9):1004–1008, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009077.

Biberman:1974:BPI

[Bib74] L. M. Biberman. B74-12 principles of interactive computer graphics. IEEE Transactions on Computers, C-23(2):216, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672494.

See comments [New75a].

Biberman:1975:RR

[Bib75] L. M. Biberman. Reviewer’s reply. IEEE Transactions on Computers, C-24(3):336, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672813.

Biswas:1970:ITS

[Bis70] N. N. Biswas. On identification of totally symmetric Boolean functions. IEEE Transactions on Computers, C-19(7): 645–648, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671590. See comments [Bis73].

Biswas:1971:MBF

[Bis71] N. N. Biswas. Minimization of Boolean functions. IEEE Transactions on Computers, C-20(8):925– 929, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671966.

Biswas:1973:CIT

[Bis73] Nripendra N. Biswas. Comments on “Identification of Totally Symmetric Boolean Functions”. IEEE Transactions on Computers, C-22(9):863–864, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009180.

See [Bis70].

Biswas:1974:SMI

[Bis74] N. N. Biswas. State minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-23(1):80– 84, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672373.

Bitner:1977:CED

[Bit77] J. R. Bitner. Capacity and efficiency of decision functions. IEEE Transactions on Computers, C-26(11):1147–1151, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674763.

Basu:1975:SPN

[BJ75] D. Basu and T. Jayashree. On a simple postcorrection for nonrestoring division. IEEE Transactions on Computers, C-24(10): 1019–1020, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672709.

Berg:1971:SAO

[BK71a] R. O. Berg and L. L. Kinney. Serial adders with overflow correction. IEEE Transactions on Computers, C-20(6):668– 671, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671914.

Budnik:1971:OUP

[BK71b] P. Budnik and D. J. Kuck. The organization and use of parallel memories. IEEE Transactions on Computers, C-20(12):1566–1569, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671764.

Bassett:1972:IPD

[BK72] James C. Bassett and Charles R.

Kime. Improved procedures for determining diagnostic resolution. IEEE Transactions on Computers, C-21(4):385–388, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008980.

Berger:1973:FDF

[BK73] I. Berger and Z. Kohavi. Fault detection in fanout-free combinational networks. IEEE Transactions on Computers, C-22(10):908– 914, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672210.

Batni:1976:MLT

[BK76] R. P. Batni and C. R. Kime. A module-level testing approach for combinational networks. IEEE Transactions on Computers, C-25 (6):594–604, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674659.

Baer:1979:MWC

[BK79] J.-L. Baer and B. Koyama. On the minimization of the width of the control memory of microprogrammed processors. IEEE Transactions on Computers, C-28(4): 310–316, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675352.

Brent:1973:PEA

[BKM73] R. Brent, D. Kuck, and K. Maruyama. The parallel evaluation of arithmetic expressions without division. IEEE Transactions on Computers, C-22(5):532–534, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672350.

Butter:1975:THR

[BL75] C. D. Butter and T. C. Lee. Thermoplastic holographic recording of binary patterns in PLZT line composer. IEEE Transactions on Computers, C-24(4):402– 406, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672827.

Blankenship:1974:CTC

[Bla74] P. E. Blankenship. Comments on “A Two’s Complement Parallel Array Multiplication Algorithm”. IEEE Transactions on Computers, C-23(12):1327, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672455. See [BW73, Kro74].

Blankenbaker:1979:CIP

[Bla79] J. V. Blankenbaker. Comments on “Inner Product Computers”. IEEE Transactions on Computers, C-28 (12):944, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675288. See [SGR78].

Bongiovanni:1975:DES

[BLZ75] G. Bongiovanni, F. Luccio, and A. Zorat. The discrete equation of the straight line. IEEE Transactions on Computers, C-24(3):310– 313, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672805.

Bowman:1970:MFA

[BM70] R. M. Bowman and E. S. McVey.

A method for the fast approximate solution of large prime implicant charts. IEEE Transactions on Computers, C-19(2):169– 173, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671481.

Barsi:1973:ECP

[BM73] F. Barsi and P. Maestrini. Error correcting properties of redundant residue number systems. IEEE Transactions on Computers, C-22 (3):307–315, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672304;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35074.

Barsi:1974:EDC

[BM74] F. Barsi and P. Maestrini. Error detection and correction by product codes in residue number systems. IEEE Transactions on Computers, C-23(9):915–924, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672648;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35088.

Bate:1977:TCS

[BM77a] J. A. Bate and J. C. Muzio. Three cell structures for ternary cellular arrays. IEEE Transactions on Computers, C-26(12):1191–1202, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674780.

Bivins:1977:SAA

[BM77b] Robert L. Bivins and Nicholas C.

Metropolis. Significance arithmetic: Application to a partial differential equation. IEEE Transactions on Computers, C-26(7):639– 642, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

/ieeexplore.ieee.org/iel5/12/
35159/01674896.pdf?tp=&isnumber=
35159&arnumber=1674896&punumber=
12;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674896.

Barsi:1978:ACR

[BM78] F. Barsi and P. Maestrini. Arithmetic codes in residue number systems with magnitude index. IEEE Transactions on Computers, C-27(12):1185–1188, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675023;   https://ieeexplore.
ieee.org/xpl/RecentIssue.jsp?
punumber=12;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35165.

Bohling:1970:ICA

[BO70] Dorothea M. Bohling and Lawrence A.

O’Neill. An interactive computer approach to tolerance analysis. IEEE Transactions on Computers, C-19(1):10–16, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5008894.

Bentley:1979:ARC

[BO79] J. L. Bentley and T. A. Ottmann. Algorithms for reporting and counting geometric intersections. IEEE Transactions on Computers, C-28(9):643–647, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675432.

Boehm:1976:SE

[Boe76] B. W. Boehm. Software engineering. IEEE Transactions on Computers, C-25(12):1226–1241, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674590.

Bohlender:1977:FPC

[Boh77] Gerd Bohlender. Floating-point computation of functions with maximum accuracy. IEEE Transactions on Computers, C-26(7):621– 632, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674894.

Bonn:1974:BMC

[Bon74] T. H. Bonn. B74-2 man and the computer. IEEE Transactions on Computers, C-23(1): 110, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672385.

Booth:1970:EPS

[Boo70] T. L. Booth. Estimation, prediction, and smoothing in discrete parameter systems. IEEE Transactions on Computers, C-19 (12):1193–1203, December 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671451.

Born:1972:ITD

[Bor72] R. C. Born. An iterative technique for determining the minimal number of variables for a totally symmetric function with repeated variables. IEEE Transactions on Computers, C-21(10): 1129–1131, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672055.

Bose:1979:CMF

[Bos79] B. Bose. Comments on “Multiple Fault Detection in Combinational Network”. IEEE Transactions on Computers, C-28(10):804– 805, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675254. See [DW73b, GS74].

Boute:1974:DSO

[Bou74a] R. T. Boute. Distinguishing sets for optimal state identification in checking experiments. IEEE Transactions on Computers, C-23(8): 874–877, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672636.

Boute:1974:ONO

[Bou74b] R. T. Boute. Optimal and nearoptimal checking experiments for output faults in sequential machines. IEEE Transactions on Computers, C-23(11):1207–1213, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672425.

Boyes:1974:BNS

[Boy74] J. D. Boyes. Binary noise sources incorporating modulo-N dividers. IEEE Transactions on Computers, C-23(5):550– 552, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672573.

Brown:1972:MPR

[BP72] W. G. S. Brown and E. A. Parrish. A model for pattern recognition systems with binary pattern vectors. IEEE Transactions on Computers, C-21(2):219, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008931.

Bandyopadhyay:1975:CUC

[BPC75] S. Bandyopadhyay, A. Pal, and A. K. Choudhury. Characterization of unate cascade realizability using parameters. IEEE Transactions on Computers, C-24(2):218– 219, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672784.

Breeding:1971:BRC

[BR71] K. J. Breeding and R. G. Rich.

BCD and radix-2 conversions. IEEE Transactions on Computers, C-20 (3):360–361, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671838.

Butler:1974:EMI

[BR74] G. A. Butler and H. B. Ritea. Estimation of mutual information in two-class pattern recognition. IEEE Transactions on Computers, C-23 (4):410–420, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672549.

Benjauthrit:1976:GSF

[BR76a] Boonsieng Benjauthrit and Irving S. Reed. Galois switching functions and their applications. IEEE Transactions on Computers, C-25(1):78– 86, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009207.

Blevins:1976:ADA

[BR76b] P. R. Blevins and C. V. Ramamoorthy. Aspects of a dynamically adaptive operating system. IEEE Transactions on Computers, C-25 (7):713–725, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674680.

Brafmen:1977:IC

[BR77] H. Brafmen and B. J. Reuter. An incremental computer. IEEE Transactions on Computers, C-26 (11):1072–1081, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1674753.

Benjauthrit:1978:FSG

[BR78] B. Benjauthrit and I. S. Reed. On the fundamental structure of Galois switching functions. IEEE Transactions on Computers, C-27(8):757– 762, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675185.

Brent:1970:ABN

[Bre70a] R. Brent. On the addition of binary numbers. IEEE Transactions on Computers, C-19(8):758–759, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Brent.pdf;
 https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671620.

Breuer:1970:FPS

[Bre70b] M. A. Breuer. Functional partitioning and simulation of digital circuits. IEEE Transactions on Computers, C-19(11):1038–1046, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671423.

Bredt:1971:APS

[Bre71a] T. H. Bredt. Analysis of parallel systems. IEEE Transactions on Computers, C-20(11):1403–1407, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671739.

Breeding:1971:SPM

[Bre71b] K. J. Breeding. Some properties of minimal threshold approximations. IEEE Transactions on Computers, C-20(12):1544–1551, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671760.

Brenner:1971:BMP

[Bre71c] M. Brenner. B71-4 mathematical programming in practice. IEEE Transactions on Computers, C-20 (10):1222, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671706.

Breuer:1971:RAT

[Bre71d] M. A. Breuer. A random and an algorithmic technique for fault detection test generation for sequential circuits. IEEE Transactions on Computers, C-20(11):1364–1370, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671733.

Breuer:1972:GFT

[Bre72a] M. A. Breuer. Generation of fault tests for linear logic networks. IEEE Transactions on Computers, C-21 (1):79–83, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672026.

Breuer:1972:NTV

[Bre72b] Melvin A. Breuer. A note on three-valued logic simulation. IEEE Transactions on Computers, C-21 (4):399–402, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008985.

Breeding:1973:BCO

[Bre73a] K. J. Breeding. B73-5 computeroriented approaches to pattern recognition. IEEE Transactions on Computers, C-22(4): 429, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672330.

Brent:1973:PAV

[Bre73b] Richard P. Brent. On the precision attainable with various floatingpoint number systems. IEEE Transactions on Computers, C-22(6):601–607, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Brent.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009113.

Breuer:1973:TIF

[Bre73c] M. A. Breuer. Testing for intermittent faults in digital circuits. IEEE Transactions on Computers, C-22 (3):241–246, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672294.

Breuer:1974:ERD

[Bre74] M. A. Breuer. The effects of races, delays, and delay faults on test generation. IEEE Transactions on Computers, C-23(10): 1078–1092, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672401.

Bredeson:1975:CSM

[Bre75] J. Bredeson. Comments on “Synthesis of Multiple-Input Change Asynchronous Machines Using Controlled Excitation and Flip-Flops”. IEEE Transactions on Computers, C-24(11):1142–1143, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672744. See [BH73].

Briley:1973:SNR

[Bri73] B. E. Briley. Some new results on average worst case carry. IEEE Transactions on Computers, C-22 (5):459–463, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672341. See comment [Yue74a].

Brown:1970:CDM

[Bro70a] F. M. Brown. Comment on “The Determination of the Maximum Compatibility Classes”. IEEE Transactions on Computers, C-19 (5):459, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671541.

Brown:1970:RSB

[Bro70b] F. M. Brown. Reduced solutions of Boolean equations. IEEE Transactions on Computers, C-19(10): 976–981, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671399.

Brown:1971:SPS

[Bro71] F. M. Brown. Single-parameter solutions for flip-flop equations. IEEE Transactions on Computers, C-20 (4):452–454, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671856.

Brown:1974:EL

[Bro74] F. M. Brown. Equational logic. IEEE Transactions on Computers, C-23(12):1228–1237, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=

1672434. See comments [Cer76].

Brown:1975:CIP

[Bro75a] F. M. Brown. The constrainedinput problem. IEEE Transactions on Computers, C-24(1):102– 106, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672682.

Brown:1975:ERS

[Bro75b] F. M. Brown. Equational realizations of switching functions. IEEE Transactions on Computers, C-24(11):1054–1066, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672726.

Brown:1975:WRS

[Bro75c] F. M. Brown. Weighted realizations of switching functions. IEEE Transactions on Computers, C-24(12):1217–1221, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672759.

Brown:1977:RAS

[Bro77] R. D. Brown. A recursive algorithm for sequency-ordered fast Walsh transforms. IEEE Transactions on Computers, C-26(8):819– 822, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674921.

Brzozowski:1970:RSG

[Brz70] J. A. Brzozowski. R70-44 synchronization and general repetitive machines, with applications to ultimate definite automata. IEEE Transactions on Computers, C-19(10):989–990, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671410.

Brzozowski:1971:AFS

[Brz71] J. A. Brzozowski. About feedback and SR flip-flops. IEEE Transactions on Computers, C-20 (4):476, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671868.

Barnea:1972:CAF

[BS72] Daniel I. Barnea and Harvey F.

Silverman. A class of algorithms for fast digital image registration. IEEE Transactions on Computers, C-21(2):179–186, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008923.

Baskett:1974:BOS

[BS74] F. Baskett and R. Sedgewick. B74-38 operating systems theory. IEEE Transactions on Computers, C-23 (8):879–880, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672639.

Bernstein:1975:CAL

[BS75] A. J. Bernstein and P. Siegel. A computer architecture for level structured systems. IEEE Transactions on Computers, C-24(8):785– 793, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672903.

Ballard:1976:LSD

[BS76] D. H. Ballard and J. Sklansky. A ladder-structured decision tree for recognizing tumors in chest radiographs. IEEE Transactions on Computers, C-25(5): 503–513, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674638.

Birta:1977:MIA

[BS77] Louis G. Birta and Ratilal R. Shah. The minimization of inverters in analog computer programming. IEEE Transactions on Computers, C-26(1):2– 12, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009267.

Baudet:1978:OSA

[BS78] G. Baudet and D. Stevenson. Optimal sorting algorithms for parallel computers. IEEE Transactions on Computers, C-27(1):84– 87, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674957.

Beakley:1970:CNP

[BT70a] G. W. Beakley and F. B. Tuteur.

Comments on “A Nonparametric Partitioning Procedure for Pattern Classification”. IEEE Transactions on Computers, C-19(4):362– 363, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671520.

Beuscher:1970:CSI

[BT70b] H. J. Beuscher and W. N.

Toy. Check schemes for integrated microprogrammed control and data transfer circuitry. IEEE Transactions on Computers, C-19(12):1153–1159, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671447.

Beakley:1972:DFP

[BT72] G. W. Beakley and F. B. Tuteur. Distribution-free pattern verification using statistically equivalent blocks. IEEE Transactions on Computers, C-21(12):1337–1347, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672098.

Booth:1973:APM

[BT73] T. L. Booth and R. A. Thompson. Applying probability measures to abstract languages. IEEE Transactions on Computers, C-22 (5):442–450, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672339.

Bednar:1974:ACD

[BT74] G. M. Bednar and J. H. Tracey. An Asynchronous Circuit Design Language (ACDL). IEEE Transactions on Computers, C-23(9):971– 976, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672657.

Bowra:1976:MDM

[BT76] James W. Bowra and H. C. Torng. The modeling and design of multiple function-unit processors. IEEE Transactions on Computers, C-25 (3):210–221, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009247.

Bubenik:1972:WMD

[Bub72] V. Bubenik. Weighting method for the determination of the irredundant set of prime implicants. IEEE Transactions on Computers, C-21(12):1449–1451, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672116. See comments [NRS74].

Burhardt:1974:ASO

[Bur74] K. K. Burhardt. An adaptive search optimization algorithm. IEEE Transactions on Computers, C-23(9):890–897, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672645.

Butz:1971:AAH

[But71] A. R. Butz. Alternative algorithm for Hilbert’s space-filling curve. IEEE Transactions on Computers, C-20(4):424–426, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671851.

Butin:1972:CDW

[But72] H. Butin. A compact definition of Walsh functions. IEEE Transactions on Computers, C-21(6): 590–592, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009014.

Butler:1975:NFR

[But75] J. T. Butler. On the number of functions realized by cascades and disjunctive networks. IEEE Transactions on Computers, C-24 (7):681–690, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672881.

Butler:1976:RCN

[But76] J. T. Butler. Restricted cellular networks. IEEE Transactions on Computers, C-25(11):1139–1142, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674568.

Butin:1978:OWF

[But78a] H. Butin. On an ordering of Walsh functions. IEEE Transactions on Computers, C-27(1):87– 90, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674958.

Butler:1978:TNU

[But78b] J. T. Butler. Tandem networks of universal cells. IEEE Transactions on Computers, C-27(9):785– 799, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675199.

Butz:1979:FLD

[But79] A. R. Butz. FFT length in digital filtering. IEEE Transactions on Computers, C-28(8):577– 580, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675414.

Buzbee:1973:FPS

[Buz73] Billy L. Buzbee. A fast Poisson solver amenable to parallel computation. IEEE Transactions on Computers, C-22(8):793– 796, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009160.

Bennett:1978:CSA

[BV78] J. R. Bennett and W. Vogel. On the characteristics of a simple architecture for finite impulse response digital filtering. IEEE Transactions on Computers, C-27(12):1197–1202, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675027.

Breitbart:1979:CCC

[BV79] Y. Breitbart and K. Vairavan. The computational complexity of a class of minimization algorithms for switching functions. IEEE Transactions on Computers, C-28(12): 941–943, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675286.

Burton:1970:CDF

[BW70] D. P. Burton and M. C. Waters. Comment on “Delay-Free Asynchronous Circuits with Constrained Line Delays”. IEEE Transactions on Computers, C-19(10): 982, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671400.

Baugh:1973:TCP

[BW73] C. R. Baugh and B. A. Wooley. A two’s complement parallel array multiplication algorithm. IEEE Transactions on Computers, C-22(12):1045–1047, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672241. See comments [Bla74, Kro74].

Baugh:1976:SAD

[BW76a] C. R. Baugh and B. A. Wooley. Statistical analysis of a differential threshold logic circuit configuration. IEEE Transactions on Computers, C-25(7):745– 754, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674683.

Bobrow:1976:MCS

[BW76b] D. G. Bobrow and B. Wegbreit.

A model for control structures for artificial intelligence programming languages. IEEE Transactions on Computers, C-25(4):347– 353, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674616.

Brandsma:1977:HST

[BW77] J. R. Brandsma and B. L. A.

Waumans. The hardware simulator: a tool for evaluating computer systems. IEEE Transactions on Computers, C-26(1):68– 72, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009275.

Brzozowski:1979:TMG

[BY79] J. A. Brzozowski and M. Yoeli.

On a ternary model of gate networks. IEEE Transactions on Computers, C-28(3):178–184, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675317.

Byrd:1971:VMC

[Byr71] J. S. Byrd. Variable-mode counting with straight binary counters. IEEE Transactions on Computers, C-20 (1):97–98, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671679.

Chen:1974:NID

[CA74] Chiu Kuan Chen and H. C. Andrews. Nonlinear intrinsic dimensionality computations. IEEE Transactions on Computers, C-23(2):178–184, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672475.

Chow:1977:CAP

[CA77] W. K. Chow and J. K. Aggarwal. Computer analysis of planar curvilinear moving images. IEEE Transactions on Computers, C-26(2): 179–185, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009299.

Cairns:1971:FFT

[Cai71] T. W. Cairns. On the Fast Fourier Transform on finite Abelian groups. IEEE Transactions on Computers, C-20(5):569–571, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671884.

Calvert:1970:NPF

[Cal70] T. W. Calvert. Nonorthogonal projections for feature extraction in pattern recognition. IEEE Transactions on Computers, C-19(5): 447–452, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671536.

Campbell:1972:BSJ

[Cam72] S. Campbell. B72-4 System/360 Job Control Language. IEEE Transactions on Computers, C-21(1):108– 109, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672035.

Cantoni:1971:OCF

[Can71] A. Cantoni. Optimal curve fitting with piecewise linear functions. IEEE Transactions on Computers, C-20(1):59–67, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671675.

Capowski:1976:MTP

[Cap76] J. J. Capowski. The matrix transform processor. IEEE Transactions on Computers, C-25(7):703– 712, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674679.

Carl:1971:CAF

[Car71a] J. W. Carl. Comment on “An Algorithm for a Fast Hadamard Matrix Transform of Order Twelve”. IEEE Transactions on Computers, C-20 (5):587–588, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671896.

Carl:1971:CSD

[Car71b] J. W. Carl. Comments on “A Simplified Definition of Walsh Functions”. IEEE Transactions on Computers, C-20(12):1617, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671782. See [LM71b, KS72].

Carter:1973:FTC

[Car73] W. C. Carter. Fault-tolerant computing: An introduction and a viewpoint. IEEE Transactions on Computers, C-22(3):225– 229, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672291.

Casasent:1973:HDO

[Cas73] David P. Casasent. A hybrid digital/optical computer system. IEEE Transactions on Computers, C-22(9):852–858, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009178.

Catt:1977:CMD

[Cat77] I. Catt. Correction to “Maxwell’s Displacement Current”. IEEE Transactions on Computers, C-26 (11):1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674773. See [Ano77v].

Carter:1972:SDE

[CB72] W. H. Carter and V. D. Brown.

A signal-dependent error arising in digitally processed images due to quantization. IEEE Transactions on Computers, C-21(12):1380–1385, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672103.

Choudhury:1970:AR

[CBD70] A. K. Choudhury, A. K. Basu, and S. C. DeSarkar. Authors’ reply. IEEE Transactions on Computers, C-19(5):459, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671542.

Coffman:1971:PIM

[CBS71] E. G. Coffman, Jr., G. J. Burnett, and R. A. Snowdon. On the performance of interleaved memories with multiple-word bandwidths. IEEE Transactions on Computers, C-20(12):1570–1573, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671765.

Carroll:1972:MTL

[CC72] B. D. Carroll and C. L. Coates, Jr. Minimum two-level threshold gate realizations. IEEE Transactions on Computers, C-21(10): 1086–1098, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672048.

Chang:1973:RTD

[CC73] Shi-Kuo Chang and C. K. Chow. The reconstruction of three-dimensional objects from two orthogonal projections and its application to cardiac cineangiography. IEEE Transactions on Computers, C-22(1):18– 28, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672189.

Chakrabarti:1970:CFA

[CCB70] K. K. Chakrabarti, A. K. Choudhury, and M. S. Basu. Complementary function approach to the synthesis of three-level NAND network. IEEE Transactions on Computers, C-19(6):509– 514, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671561.

Chang:1974:CPD

[CCES74] H. Y.-P. Chang, S. G. Chappell, C. H. Elmendorf, and L. D. Schmidt. Comparison of parallel and deductive fault simulation methods. IEEE Transactions on Computers, C-23(11):1132–1138, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672413.

Chen:1979:APN

[CCF79] I-Ngo Chen, P. Y. Chen, and Tse-Yun Feng. Associative processing of network flow problems. IEEE Transactions on Computers, C-28 (3):184–190, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675318.

Chuang:1973:SMI

[CD73] H. Y. H. Chuang and S. Das. Synthesis of multiple-input change asynchronous machines using controlled excitation and flip-flops. IEEE Transactions on Computers, C-22(12):1103–1109, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672249.

Chuang:1975:AR

[CD75] H. Y. H. Chuang and S. Das. Authors’ reply. IEEE Transactions on Computers, C-24(11):1143–1144, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672745.

Chuang:1978:DFS

[CD78] H. Y. H. Chuang and S. Das. Design of fail-safe sequential machines using separable codes. IEEE Transactions on Computers, C-27(3):249– 252, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675078.

Carter:1971:SST

[CDJ71] W. C. Carter, K. A. Duke, and D. C. Jessep. A simple self-testing decoder checking circuit. IEEE Transactions on Computers, C-20(11):1413–1414, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671741.

Carter:1973:LTM

[CDJ73] W. C. Carter, K. A. Duke, and D. C. Jessep, Jr. Lookaside techniques for minimum circuit memory translators. IEEE Transactions on Computers, C-22(3):283– 289, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672301.

Cordella:1978:ACC

[CDL78] L. P. Cordella, M. J. B. Duff, and S. Levialdi. An analysis of computational cost in image processing: a case study. IEEE Transactions on Computers, C-27(10): 904–910, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674969.

Cha:1978:VAT

[CDO78] C. W. Cha, W. E. Donath, and F. Ozguner. 9-V algorithm for test pattern generation of combinational digital circuits. IEEE Transactions on Computers, C-27(3):193– 200, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675071.

Chang:1973:DMS

[CDS73] H. Y.-P. Chang, R. C. Dorr, and D. J. Senese. The design of a microprogrammed selfchecking processor of an electronic switching system. IEEE Transactions on Computers, C-22(5): 489–500, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672345.

Cheung:1975:MTT

[CE75] C. C. Cheung and R. W. Ehrich. On the minimization of tree-type universal logic circuits. IEEE Transactions on Computers, C-24 (11):1110–1113, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672734.

Cerny:1976:CEL

[Cer76] E. Cerny. Comments on “Equational Logic”. IEEE Transactions on Computers, C-25(1):102– 103, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

5009215. See [Bro74].

Cerny:1978:CFO

[Cer78] E. Cerny. Controllability and fault observability in modular combinational circuits. IEEE Transactions on Computers, C-27(10):896– 903, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674968.

Cook:1970:SDD

[CF70a] R. W. Cook and M. J. Flynn.

System design of a dynamic microprocessor. IEEE Transactions on Computers, C-19(3):213– 222, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671492.

Cooper:1970:AIO

[CF70b] D. B. Cooper and J. H. Freeman.

On the asymptotic improvement in the outcome of supervised learning provided by additional nonsupervised learning. IEEE Transactions on Computers, C-19(11):1055–1063, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671425.

Cook:1973:LNC

[CF73] Robert W. Cook and Michael J.

Flynn. Logical network cost and entropy. IEEE Transactions on Computers, C-22(9):823–826, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009171. See comments [Mas78].

Censier:1978:NSC

[CF78a] L. M. Censier and P. Feautrier. A new solution to coherence problems in multicache systems. IEEE Transactions on Computers, C-27(12):1112–1118, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675013.

Corsini:1978:USM

[CF78b] P. Corsini and G. Frosini. Uniform shift multiplication algorithms without overflow. IEEE Transactions on Computers, C-27(3):256– 258, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675081.

Corsini:1979:PMG

[CF79] P. Corsini and G. Frosini. Properties of the multidimensional generalized discrete Fourier transform. IEEE Transactions on Computers, C-28(11):819–830, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675262.

Cantor:1974:ORP

[CG74] D. G. Cantor and M. Gerla. Optimal routing in a packet-switched computer network. IEEE Transactions on Computers, C-23(10): 1062–1069, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672399.

Carl:1972:AFT

[CH72a] J. W. Carl and C. F. Hall.

The application of filtered transforms to the general classification problem. IEEE Transactions on Computers, C-21(7):785– 790, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672175. See comments [CH73b].

Chien:1972:ECH

[CH72b] R. T. Chien and Se June Hong.

Error correction in high-speed arithmetic. IEEE Transactions on Computers, C-21(5):433– 438, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672131.

Cappa:1973:AIA

[CH73a] M. Cappa and V. C. Hamacher. An augmented iterative array for highspeed binary division. IEEE Transactions on Computers, C-22(2): 172–175, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672273. See comments [Gar74].

Carl:1973:CAF

[CH73b] J. W. Carl and C. F. Hall. Comments on “The Application of Filtered Transforms to the General Classification Problem”. IEEE Transactions on Computers, C-22 (1):111, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672198. See [CH72a].

Casasent:1975:ISI

[CH75] D. Casasent and S. Horvitz. Introduction to the special issue on optical computing. IEEE Transactions on Computers, C-24(4):337– 339, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672819.

Chang:1970:RDU

[Cha70] H. Y. Chang. R70-41 diagnosis and utilization of faulty universal tree circuits. IEEE Transactions on Computers, C-19(10):988– 989, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671407.

Chang:1971:ARM

[Cha71] Chin Liang Chang. The accelerated relaxation method for linear inequalities. IEEE Transactions on Computers, C-20(2):222– 225, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671811.

Chang:1972:EFP

[Cha72] Shi-Kuo Chang. On the execution of fuzzy programs using finitestate machines. IEEE Transactions on Computers, C-21(3):241– 253, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008946.

Chang:1973:PRP

[Cha73a] Chin-Liang Chang. Pattern recognition by piecewise linear discriminant functions. IEEE Transactions on Computers, C-22(9):859– 862, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009179.

Chang:1973:CWO

[Cha73b] Shi-Kuo Chang. The computation of window operations on a parallel organized computer — a case study. IEEE Transactions on Computers, C-22(1):34– 40, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672191.

Chakrabarti:1974:DCA

[Cha74a] S. Chakrabarti. Divide-and-correct algorithm for division in a negative base. IEEE Transactions on Computers, C-23(9):981–983, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672659.

Chang:1974:FPN

[Cha74b] Chin-Liang Chang. Finding prototypes for nearest neighbor classifiers. IEEE Transactions on Computers, C-23(11):1179–1184, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672420.

Chang:1974:BDA

[Cha74c] H. Y. Chang. B74-3 design automation of digital systems — volume 1, theory and techniques. IEEE Transactions on Computers, C-23(1): 110–111, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672386.

Chang:1974:PBB

[Cha74d] Shi-Kuo Chang. Parallel balancing of binary search trees. IEEE Transactions on Computers, C-23 (4):441–444, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672552.

Chaney:1979:CNS

[Cha79] T. J. Chaney. Comments on “A Note on Synchronizer or Interlock Maloperation”. IEEE Transactions on Computers, C-28(10):802– 804, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675252. See [Wor77, Wor79].

Chen:1971:TCC

[Che71a] C. H. Chen. Theoretical comparison of a class of feature selection criteria in pattern recognition. IEEE Transactions on Computers, C-20(9):1054–1056, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671995. See comments [Tou72c].

Chen:1971:BMS

[Che71b] Tien Chi Chen. A binary multiplication scheme based on squaring. IEEE Transactions on Computers, C-20(6):678–680, June 1971. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671918.

Chen:1972:AR

[Che72] C. H. Chen. Author’s reply. IEEE Transactions on Computers, C-21 (6):616, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009025.

Chen:1973:NMI

[Che73] R. M.-M. Chen. New matrix inversion algorithms based on exchange method. IEEE Transactions on Computers, C-22(10):885– 890, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672206.

Chen:1974:MLS

[Che74] Chi-Tsong Chen. Minimization of linear sequential machines. IEEE Transactions on Computers, C-23 (1):93–95, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672378.

Chen:1978:GPM

[Che78] C. L. Chen. On the generation of permutations in magnetic bubble memories. IEEE Transactions on Computers, C-27(12):1183–1185, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675022.

Chien:1971:SDM

[Chi71] Yi-Tzuu Chien. A sequential decision model for selecting feature subsets in pattern recognition. IEEE Transactions on Computers, C-20 (3):282–290, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671825.

Chinal:1972:SCP

[Chi72] J. P. Chinal. Some comments on postcorrections for nonrestoring division. IEEE Transactions on Computers, C-21(12):1385–1394, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672104.

Chinal:1977:BFS

[Chi77] Jean P. Chinal. Boolean features of sign-flipping and signed-carry logic in base-2. IEEE Transactions on Computers, C-26(1):82– 86, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009278.

Chiba:1978:ILH

[Chi78] T. Chiba. Impact of the LSI on high-speed computer packaging. IEEE Transactions on Computers, C-27(4):319–325, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1675104.

Choueka:1974:SA

[Cho74] Y. A. Choueka. Structure automata. IEEE Transactions on Computers, C-23(12):1218–1227, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672433.

Chow:1976:DCC

[Cho76] C. K. Chow. Determination of cache’s capacity and its matching storage hierarchy. IEEE Transactions on Computers, C-25(2):157– 164, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009230.

Chang:1973:MTM

[CHSS73] H. Y. P. Chang, G. W. Heimbigner, D. J. Senese, and T. L. Smith. Maintenance techniques of a microprogrammed self-checking control complex of an electronic switching system. IEEE Transactions on Computers, C-22(5): 501–512, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672346.

Chu:1970:BBP

[Chu70] W. W. Chu. Buffer behavior for Poisson arrivals and multiple synchronous constant outputs. IEEE Transactions on Computers, C-19 (6):530–534, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671563.

Chu:1971:EBC

[Chu71] J. T. Chu. Error bounds for a contextual recognition procedure. IEEE Transactions on Computers, C-20(10):1203–1207, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671699. See comments [Tou72b].

Chu:1972:AR

[Chu72] John T. Chu. Author’s reply. IEEE Transactions on Computers, C-21(9):1027–1028, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009088.

Chu:1974:SNE

[Chu74] J. T. Chu. Some new error bounds and approximations for pattern recognition. IEEE Transactions on Computers, C-23(2):194– 199, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672480.

Chuang:1976:FSA

[Chu76] H. Y. H. Chuang. Fail-safe asynchronous machines with multipleinput changes. IEEE Transactions on Computers, C-25(6):637– 642, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674664.

Chu:1979:AHD

[Chu79] Yaohan Chu. Architecture of a hardware data interpreter. IEEE Transactions on Computers, C-28(2):101–109, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675299.

Cioffi:1978:ASI

[Cio78] G. Cioffi. Autotesting speedindependent sequential circuits. IEEE Transactions on Computers, C-27(1):90–94, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674959.

Chandrasekaran:1974:QCI

[CJ74] B. Chandrasekaran and A. K.

Jain. Quantization complexity and independent measurements. IEEE Transactions on Computers, C-23(1):102–106, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672382.

Carter:1971:LDD

[CJW^{+}71] W. C. Carter, D. C. Jessep, A. B. Wadia, P. R. Schneider, and W. G. Bouricius. Logic design for dynamic and interactive recovery. IEEE Transactions on Computers, C-20(11):1300–1305, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671724.

Chakrabarti:1974:FCT

[CK74] K. Chakrabarti and O. Kolp. Fanin constrained tree networks of flexible cells. IEEE Transactions on Computers, C-23(12):1238–1249, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672435.

Chen:1975:TPP

[CK75] Shyh-Ching Chen and D. J. Kuck.

Time and parallel processor bounds for linear recurrence systems. IEEE Transactions on Computers, C-24 (7):701–717, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672884.

Chen:1977:CCS

[CK77] Shyh-Ching Chen and D. J. Kuck. Combinational circuit synthesis with time and component bounds. IEEE Transactions on Computers, C-26(8):712– 726, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674909.

Chow:1979:MDL

[CK79] Yuan-Chieh Chow and W. H. Kohler. Models for dynamic load balancing in a heterogeneous multiple processor system. IEEE Transactions on Computers, C-28(5): 354–361, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675365.

Chang:1977:EBP

[CKL77] D. Y. Chang, D. J. Kuck, and D. H. Lawrie. On the effective bandwidth of parallel memories. IEEE Transactions on Computers, C-26 (5):480–490, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674865.

Cohn:1971:SPW

[CL71] M. Cohn and A. J. Lincoln. Spectral properties of waveforms generated by smooth sequences. IEEE Transactions on Computers, C-20 (3):330–332, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671833.

Cyre:1972:GMC

[CL72] W. R. Cyre and G. J. Lipovski.

On generating multipliers for a cellular Fast Fourier Transform processor. IEEE Transactions on Computers, C-21(1):83– 87, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672027.

Chao:1975:HRR

[CL75] Chin Chao and H. H. Loomis, Jr. High rate realization of finitestate machines. IEEE Transactions on Computers, C-24(7):729– 740, July 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672886.

Cheng:1977:ASO

[CL77] D. Cheng and J. J. Liu. An algorithm for sequency ordering of Hadamard functions. IEEE Transactions on Computers, C-26(3): 308–309, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674828.

Cheng:1979:GOT

[CL79] D. K. Cheng and J. J. Liu. A generalized orthogonal transformation matrix. IEEE Transactions on Computers, C-28(2):147– 150, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675304.

Clegg:1973:USA

[Cle73] F. W. Clegg. Use of SPOOF’s in the analysis of faulty logic networks. IEEE Transactions on Computers, C-22(3):229–234, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672292.

Costes:1978:RAM

[CLL78] A. Costes, C. Landrault, and J. C. Laprie. Reliability and availability models for maintained systems featuring hardware failures and design faults. IEEE Transactions on Computers, C-27(6): 548–560, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675146.

Chen:1970:ISL

[CM70] Wei Chen and L. P. McNamee. Iterative solution of large-scale systems by hybrid techniques. IEEE Transactions on Computers, C-19(10): 879–889, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671387.

Chaney:1973:ABS

[CM73a] T. J. Chaney and C. E. Molnar.

Anomalous behavior of synchronizer and arbiter circuits. IEEE Transactions on Computers, C-22 (4):421–422, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672323.

Chien:1973:PAR

[CM73b] Robert T. Chien and Edward A.

Mark. On performance analysis of retrieval functions. IEEE Transactions on Computers, C-22(9):840– 844, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009175.

Cerny:1974:CAS

[CM74] E. Cerny and M. A. Marin. A computer algorithm for the synthesis of memoryless logic circuits. IEEE Transactions on Computers, C-23 (5):455–465, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672560.

Carter:1976:IEF

[CM76] W. C. Carter and C. E. Mc-Carthy. Implementation of an experimental fault-tolerant memory system. IEEE Transactions on Computers, C-25(6):557– 568, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674655.

Cerny:1977:AUM

[CM77] E. Cerny and M. A. Marin. An approach to unified methodology of combinational switching circuits. IEEE Transactions on Computers, C-26(8):745– 756, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674912.

Chang:1979:AR

[CM79a] D. M. Chang and T. H. Mott, Jr. Author’s reply. IEEE Transactions on Computers, C-28(11): 875, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675271.

Current:1979:IPC

[CM79b] K. W. Current and D. A. Mow. Implementing parallel counters with four-valued threshold logic. IEEE Transactions on Computers, C-28 (3):200–204, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675320.

Cutler:1979:CCI

[CM79c] R. B. Cutler and S. Muroga. Comments on “Computing Irredundant Normal Forms from Abbreviated Presence Functions”. IEEE Transactions on Computers, C-28(11): 874–875, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675270.

Cutler:1979:CGC

[CM79d] R. B. Cutler and S. Muroga.

Comments on “Generalization of Consensus Theory and Application to the Minimization of Boolean Functions”. IEEE Transactions on Computers, C-28(7): 542–543, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675403.

Cerny:1979:SMB

[CMS79] E. Cerny, D. Mange, and E. Sanchez.

Synthesis of minimal binary decision trees. IEEE Transactions on Computers, C-28(7):472– 482, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675391.

Cody:1973:SDN

[Cod73] William J. Cody, Jr. Static and dynamic numerical characteristics of floating-point arithmetic. IEEE Transactions on Computers, C-22(6):598–601, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Cody.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009112.

Cohn:1970:AR

[Coh70] M. Cohn. Authors’ reply. IEEE Transactions on Computers, C-19 (7):654, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671594.

Coles:1977:CHS

[Col77] C. J. Coles. Comments on “Heuristic Synthesis of Microprogrammed Computer Architecture”. IEEE Transactions on Computers, C-26(9):930–932, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674941. See [AAK74].

Cooper:1972:BAI

[Coo72] D. W. Cooper. B72-25 artificial intelligence: The heuristic programming approach. IEEE Transactions on Computers, C-21(12): 1458, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672120.

Cooper:1977:DP

[Coo77] R. G. Cooper. The distributed pipeline. IEEE Transactions on Computers, C-26(11):1123–1132, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674758.

Corinthios:1971:DCF

[Cor71a] M. J. Corinthios. The design of a class of Fast Fourier Transform computers. IEEE Transactions on Computers, C-20(6):617– 623, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671905.

Corinthios:1971:FFT

[Cor71b] M. J. Corinthios. A Fast Fourier Transform for high-speed signal processing. IEEE Transactions on Computers, C-20(8):843– 846, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671952.

Corinthios:1977:FTA

[Cor77] Michael J. Corinthios. A fast Z transformation algorithm for system identification. IEEE Transactions on Computers, C-26(1):55– 67, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009274.

Covo:1974:AMC

[Cov74] A. A. Covo. Analysis of multiprocessor control organizations with partial program memory replication. IEEE Transactions on Computers, C-23(2):113–120, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672466.

Cox:1972:BFD

[Cox72] J. E. Cox. B72-19 future developments in telecommunications. IEEE Transactions on Computers, C-21(10):1135–1136, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672059.

Coy:1979:DET

[Coy79] W. Coy. On the design of easily testable iterative systems of combinational cells. IEEE Transactions on Computers, C-28(5):367– 371, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675368.

Chandy:1972:RRS

[CR72a] K. M. Chandy and C. V. Ramamoorthy. Rollback and recovery strategies for computer programs. IEEE Transactions on Computers, C-21(6):546–556, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009007.

Chandy:1972:DML

[CR72b] K. M. Chandy and R. A. Russell. The design or multipoint linkages in a teleprocessing tree network. IEEE Transactions on Computers, C-21(10):1062–1066, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672045.

Chiang:1972:NAB

[CR72c] A. C. L. Chiang and I. S. Reed.

Notes on the arithmetic BN modulo A codes. IEEE Transactions on Computers, C-21(8):891– 894, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009046.

Chien:1972:NDB

[CR72d] Y. T. Chien and R. Ribak. A new data base for syntax-directed pattern analysis and recognition. IEEE Transactions on Computers, C-21 (7):790–801, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672176.

Casasent:1978:SFO

[CR78] D. Casasent and P. D. Rapp. System functions for an optical/ digital processor. IEEE Transactions on Computers, C-27(8):732– 738, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675180.

Chiang:1972:PSP

[CRB72] A. C. L. Chiang, I. S. Reed, and A. V. Banes. Path sensitization, partial Boolean difference, and automated fault diagnosis. IEEE Transactions on Computers, C-21(2):189–195, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008925.

Cassee:1970:TAT

[CS70] P. R. Cassee and M. J. O. Strutt. Is there any advantage of ternary logic as compared with binary? IEEE Transactions on Computers, C-19 (6):559, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671569.

Carl:1973:HWT

[CS73] Joseph W. Carl and Richard V.

Swartwood. A hybrid Walsh transform computer. IEEE Transactions on Computers, C-22(7):669– 672, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009133.

Casasent:1975:ODP

[CS75] D. Casasent and W. M. Sterling. An optical/digital processor: Hardware and applications. IEEE Transactions on Computers, C-24(4):348– 358, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672821.

Ciompi:1979:AOD

[CS79] P. Ciompi and L. Simoncini. Analysis and optimal design of selfdiagnosable systems with repair. IEEE Transactions on Computers, C-28(5):362–365, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675366.

Carroll:1974:EAT

[CSJ74] B. D. Carroll, H. G. Shah, and D. M. Jones. An examination of algebraic test generation methods for multiple faults. IEEE Transactions on Computers, C-23(7): 743–745, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672617.

Cook:1973:DSC

[CSST73] R. W. Cook, W. H. Sisson, T. F. Storey, and W. N. Toy. Design of a self-checking microprogram control. IEEE Transactions on Computers, C-22(3):255– 262, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672297.

Corinthios:1975:PRF

[CSY75] M. J. Corinthios, K. C. Smith, and J. L. Yen. A parallel radix-4 Fast Fourier Transform computer. IEEE Transactions on Computers, C-24 (1):80–92, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672678.

Curtis:1970:SPR

[Cur70] H. Allen Curtis. Systematic procedures for realizing synchronous sequential machines using flip-flop memory: Part II. IEEE Transactions on Computers, C-19(1):66– 73, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008901.

Curtis:1971:FRC

[Cur71a] H. A. Curtis. The further reduction of CC-tables. IEEE Transactions on Computers, C-20(4):454– 456, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671857.

Curtis:1971:RPS

[Cur71b] H. A. Curtis. The realization of polylinear sequential circuits using flip-flop memory. IEEE Transactions on Computers, C-20(1):87– 94, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671677.

Curtis:1973:TLS

[Cur73a] H. A. Curtis. Tan-like state assignments for synchronous sequential machines. IEEE Transactions on Computers, C-22(2):181– 187, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672275.

Curtis:1973:NTD

[Cur73b] H. Allen Curtis. A new type doublerank sequential machine. IEEE Transactions on Computers, C-22 (8):796–803, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009161.

Curtis:1976:SDB

[Cur76] H. A. Curtis. Simplified decomposition of Boolean functions. IEEE Transactions on Computers, C-25(10):1033–1044, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674544.

Curts:1977:ATM

[Cur77] H. A. Curts. Adjacency table method of deriving minimal sums. IEEE Transactions on Computers, C-26(11):1136–1141, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674760.

Curtis:1979:SCM

[Cur79] H. A. Curtis. Short-cut method of deriving nearly optimal arrays of NAND trees. IEEE Transactions on Computers, C-28(7):521– 528, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675398.

Cutlip:1970:RSR

[Cut70] W. F. Cutlip. R70-14 some results on cascade decomposition of automata. IEEE Transactions on Computers, C-19(6): 565, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671578.

Couranz:1975:TEB

[CW75] G. R. Couranz and D. F. Wann.

Theoretical and experimental behavior of synchronizers operating in the metastable region. IEEE Transactions on Computers, C-24 (6):604–616, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672866.

Conrad:1977:ISL

[CW77] V. Conrad and Y. Wallach. Iterative solution of linear equations on a parallel processor system. IEEE Transactions on Computers, C-26(9):838–847, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674932.

Chandra:1979:MPC

[CW79a] A. K. Chandra and C. K. Wong. The movement and permutation of columns in magnetic bubble lattice files. IEEE Transactions on Computers, C-28(1):8– 15, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675217.

Chen:1979:PMB

[CW79b] I-Ngo Chen and R. Willoner. An O(n) parallel multiplier with bit-sequential input and output. IEEE Transactions on Computers, C-28(10):721–727, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675239.

Chung:1979:AOI

[CW79c] Kin-Man Chung and C. K. Wong. Asymptotically optimal interconnection networks from two-state cells. IEEE Transactions on Computers, C-28(7):500– 505, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675395.

Cooper:1976:CCA

[CY76] D. B. Cooper and N. Yalabik. On the computational cost of approximating and recognizing noise-perturbed straight lines and quadratic arcs in the plane. IEEE Transactions on Computers, C-25(10):1020–1032, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674543.

Culliney:1979:RSO

[CYNM79] J. N. Culliney, M. H. Young, T. Nakagawa, and S. Muroga. Results of the synthesis of optimal networks of AND and OR gates for four-variable switching functions. IEEE Transactions on Computers, C-28(1):76– 85, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675229.

Diaz:1979:UDS

[DAA79] M. Diaz, P. Azema, and J. M. Ayache. Unified design of self-checking and fail-safe combinational circuits and sequential machines. IEEE Transactions on Computers, C-28 (3):276–281, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675338.

Dahlberg:1973:SFR

[Dah73] B. Dahlberg. On symmetric functions with redundant variablesweighted functions. IEEE Transactions on Computers, C-22(5): 450–458, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672340.

Danielsson:1970:ICG

[Dan70] P. E. Danielsson. Incremental curve generation. IEEE Transactions on Computers, C-19(9):783– 793, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671634.

Dandapani:1973:DMT

[Dan73] Ramaswami Dandapani. Derivation of minimal test sets for monotonic logic circuits. IEEE Transactions on Computers, C-22(7): 657–661, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009131.

Das:1971:FSL

[Das71a] S. K. Das. Feature selection with a linear dependence measure. IEEE Transactions on Computers, C-20(9):1106–1109, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672005. See comments [TV72].

Das:1971:ASS

[Das71b] S. R. Das. An approach for simplifying switching functions by utilizing the cover table representation. IEEE Transactions on Computers, C-20(3):355– 359, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671837.

Das:1971:CNA

[Das71c] S. R. Das. Comments on “A New Algorithm for Generating Prime Implicants”. IEEE Transactions on Computers, C-20(12):1614–1615, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671778. See [SCL70].

Das:1972:AR

[Das72] Subrata K. Das. Author’s reply.

IEEE Transactions on Computers, C-21(4):409, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008990.

Das:1973:NAF

[Das73] S. R. Das. On a new approach for finding all the modified cut-sets in an incompatibility graph. IEEE Transactions on Computers, C-22(2):187–193, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672276.

Dasgupta:1978:CIM

[Das78] S. Dasgupta. Comments on “The Identification of Maximal Parallelism in Straight-Line Microprograms”. IEEE Transactions on Computers, C-27(3):285– 286, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675089. See [DT76, Bar78, Ada78].

Davis:1970:RSL

[Dav70a] W. A. Davis. R70-42 synthesis of linear sequential machines with unspecified outputs. IEEE Transactions on Computers, C-19(10): 989, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671408.

Davis:1970:SMR

[Dav70b] W. A. Davis. Sequential machines realizable with delay elements only. IEEE Transactions on Computers, C-19(4):353– 355, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671515.

Davies:1971:PWO

[Dav71] A. C. Davies. Properties of waveforms obtained by nonrecursive digital filtering of pseudorandom binary sequences. IEEE Transactions on Computers, C-20(3):270– 281, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671824.

Davies:1972:DGW

[Dav72a] Anthony C. Davies. On the definition and generation of Walsh functions. IEEE Transactions on Computers, C-21(2):187–189, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008924. See comments [Van73b].

Davis:1972:VAA

[Dav72b] J. M. Davis. Velocity analysis: An application of deterministic estimation to reflection seismology. IEEE Transactions on Computers, C-21 (7):730–738, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672168.

Davis:1974:USN

[Dav74] R. L. Davis. Uniform shift networks.

????, C-24(3):317–322, March 1974. CODEN ???? ISSN ????

David:1977:MDA

[Dav77a] R. David. Modular design of asynchronous circuits defined by graphs. IEEE Transactions on Computers, C-26(8):727–737, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674910.

Davis:1977:USA

[Dav77b] L. S. Davis. Understanding shape: Angles and sides. IEEE Transactions on Computers, C-26(3):236– 242, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674812.

David:1978:TSC

[Dav78] R. David. A totally self-checking 1-out-of-3 checker. IEEE Transactions on Computers, C-27(6): 570–572, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675149.

Durling:1973:UMR

[DB73] A. E. Durling and T. E. Bullock. A unified method for the reconstruction of sampled data. IEEE Transactions on Computers, C-22 (4):388–396, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672319.

DeFiore:1974:QAU

[DB74] C. R. DeFiore and P. B. Berra. A quantitative analysis of the utilization of associative memories in data management. IEEE Transactions on Computers, C-23(2):121– 133, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672467.

David:1976:ARF

[DB76] R. David and G. Blanchet. About random fault detection of combinational networks. IEEE Transactions on Computers, C-25(6): 659–664, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674669.

DeSarkar:1971:DIP

[DBC71] S. C. De Sarkar, A. K. Basu, and A. K. Choudhury. On the determination of irredundant prime closed sets. IEEE Transactions on Computers, C-20(8):933– 938, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671968.

Das:1973:CMM

[DBC73] S. R. Das, D. K. Banerji, and A. Chattopadhyay. On control memory minimization in microprogrammed digital computers. IEEE Transactions on Computers, C-22(9):845–848, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009176.

Dudani:1977:AIM

[DBM77] Sahibsingh A. Dudani, Kenneth J. Breeding, and Robert B. McGhee. Aircraft identification by moment invariants. IEEE Transactions on Computers, C-26(1):39– 46, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009272.

Das:1976:CDM

[DDSM76] S. R. Das, C. R. Datta, P. K. Srimani, and K. Mandal. Comments on “Derivation of Minimal Complete Sets of Test-Input Sequences Using Boolean Differences”. IEEE Transactions on Computers, C-25(10):1053–1056, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1674548. See [MST^{+}75].

deRennaeSouza:1970:BTA

[de 70] Celso de Renna e Souza. B70-5 theories of abstract automata. IEEE Transactions on Computers, C-19(9):855–857, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671652.

deRennaeSouza:1971:PAM

[de 71] Celso de Renna e Souza. Probabilistic automata with monitored final state sets. IEEE Transactions on Computers, C-20(4):448– 452, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671855.

DeWitt:1979:DMO

[De 79] D. J. De Witt. Direct — a multiprocessor organization for supporting relational database management systems. IEEE Transactions on Computers, C-28(6):395– 406, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675379.

Dawoud:1973:PDD

[DEA73] D. Sh. Dawoud and N. Z. El-Araby. Parallel digital differential analyzer with arbitrary stored interconnections. IEEE Transactions on Computers, C-22(1):41– 46, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672192.

Deekshatulu:1972:BAL

[Dee72] B. L. Deekshatulu. B72-24 adaptation and learning in automatic systems. IEEE Transactions on Computers, C-21(11):

1252, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672084.

Denning:1974:BSP

[Den74] P. J. Denning. B74-20 systematic programming: An introduction. IEEE Transactions on Computers, C-23(2):222–223, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672502.

Denman:1976:AAS

[Den76] E. D. Denman. An additional algorithm for a system of coupled algebraic matrix Riccati equations. IEEE Transactions on Computers, C-25(1):91– 93, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009210. See comments [Gou77].

Depeyrot:1970:LSI

[Dep70] M. Depeyrot. Linear system identification using real-time deconvolution. IEEE Transactions on Computers, C-19(12):1139–1145, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671445.

Dertouzos:1973:TBS

[Der73] M. L. Dertouzos. Time bounds on space computations. IEEE Transactions on Computers, C-22(1):12– 17, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672188.

Despain:1974:FTC

[Des74] Alvin M. Despain. Fourier transform computers using CORDIC iterations. IEEE Transactions on Computers, C-23(10):993–1001, October 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672393.

Despain:1979:VFF

[Des79] A. M. Despain. Very fast Fourier transform algorithms hardware for implementation. IEEE Transactions on Computers, C-28(5):333– 341, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675363.

Deutsch:1970:POH

[Deu70] E. S. Deutsch. On parallel operations on hexagonal arrays. IEEE Transactions on Computers, C-19(10):982–983, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671401.

Deuser:1971:HMF

[Deu71] L. M. Deuser. A hybrid multispectral feature selection criterion. IEEE Transactions on Computers, C-20(9):1116–1117, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672008.

Devijver:1974:NCB

[Dev74] P. A. Devijver. On a new class of bounds on Bayes risk in multihypothesis pattern recognition. IEEE Transactions on Computers, C-23 (1):70–80, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672372. See comments [Tou75].

Deverell:1975:PIA

[Dev75] J. Deverell. Pipeline iterative arithmetic arrays. IEEE Transactions on Computers, C-24(3):317– 322, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672807.

Devries:1979:CRI

[Dev79] R. C. Devries. Comments on “A Readily Implemented Single-Error-Correcting Unit-Distance Counting Code”. IEEE Transactions on Computers, C-28(3):253– 255, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675328. See [Kau70c].

Doty:1970:DTS

[DF70] K. L. Doty and H. Frank. Discretetime systems with the decomposition property. IEEE Transactions on Computers, C-19(5):412– 420, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671532.

Das:1975:FDE

[DF75] P. Das and D. E. Farmer.

Fault-detection experiments for parallel-decomposable sequential machines. IEEE Transactions on Computers, C-24(11):1104–1109, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672732.

Deutsch:1978:QSO

[DF78] E. S. Deutsch and J. R. Fram.

A quantitative study of the orientation bias of some edge detector schemes. IEEE Transactions on Computers, C-27(3):205– 213, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675073.

Degryse:1972:LT

[DG72] D. Degryse and B. Guerin. A logarithmic transcoder. IEEE Transactions on Computers, C-21 (11):1165–1168, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672066.

Diaz:1974:SRF

[DGC74] M. Diaz, J. C. Geffroy, and M. Courvoisier. On-set realization of fail-safe sequential machines. IEEE Transactions on Computers, C-23(2):133–138, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672468.

Dunworth:1979:ESM

[DH79] A. Dunworth and H. V. Hartog. An efficient state minimization algorithm for some special classes of incompletely specified sequential machines. IEEE Transactions on Computers, C-28(7):531– 535, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675400.

Daniel:1971:DRM

[DI71] J. M. Daniel and J. D. Irwin. Dynamic resolution of memory access conflicts. IEEE Transactions on Computers, C-20(12):1583–1586, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671769.

Dias:1975:FMC

[Dia75] F. J. O. Dias. Fault masking in combinational logic circuits. IEEE Transactions on Computers, C-24 (5):476–482, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672842.

Dias:1976:TTV

[Dia76] F. J. O. Dias. Truth-table verification of an iterative logic array. IEEE Transactions on Computers, C-25 (6):605–613, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674660.

Digby:1973:SMM

[Dig73] David W. Digby. A search memory for many-to-many comparisons. IEEE Transactions on Computers, C-22(8):768–772, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009156.

Dirilten:1972:MMC

[Dir72] Hudai Dirilten. On the mathematical models characterizing faulty four-phase MOS logic arrays. IEEE Transactions on Computers, C-21 (3):301–305, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008954.

Das:1972:CCT

[DK72] S. R. Das and N. S. Khabra. Clause-column table approach for generating all the prime implicants of switching functions. IEEE Transactions on Computers, C-21(11):1239–1246, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672077.

Dowd:1973:SGH

[DLL^{+}73] P. G. Dowd, D. J. Ladd, G. W. R. Luderer, P. E. Rosenfeld, and R. L. Salmon. Stare 2 — a graphical hard-copy system. IEEE Transactions on Computers, C-22(4):359– 367, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672315.

Ducasse:1975:CRB

[DM75] E. Ducasse and G. Metze. Comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra under Different Definitions of Complement”. IEEE Transactions on Computers, C-24(2):220, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672785.

See [DM76].

Ducasse:1976:RAR

[DM76] Edgar Ducasse and Gernot Metze.

Rebuttal to the authors’ reply to our comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra Under Different Definitions of Complement”. IEEE Transactions on Computers, C-25(3):311– 312, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009259. See [Eps73b, DM75].

DeSousa:1978:SMR

[DM78a] P. T. De Sousa and F. P. Mathur.

Sift-out modular redundancy. IEEE Transactions on Computers, C-27 (7):624–627, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675162.

Deguchi:1978:TCT

[DM78b] K. Deguchi and I. Morishita. Texture characterization and texturebased image partitioning using two-dimensional linear estimation techniques. IEEE Transactions on Computers, C-27(8):739– 745, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675181.

Dao:1977:MII

[DMR77] T. T. Dao, E. J. Mccluskey, and L. K. Russell. Multivalued integrated injection logic. IEEE Transactions on Computers, C-26(12):1233–1241, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674784.

Dirilten:1977:PMU

[DN77] H. Dirilten and T. G. Newman. Pattern matching under affine transformations. IEEE Transactions on Computers, C-26(3):314– 317, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674832.

Dorr:1972:SCC

[Dor72] R. C. Dorr. Self-checking combinational logic binary counters. IEEE Transactions on Computers, C-21(12):1426–1430, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672111.

Doster:1977:CPS

[Dos77] W. Doster. Contextual postprocessing system for cooperation with a multiple-choice characterrecognition system. IEEE Transactions on Computers, C-26(11): 1090–1101, November 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674755.

Doty:1970:ILD

[Dot70] K. L. Doty. On information-lossless discrete-time systems. IEEE Transactions on Computers, C-19(6): 548–551, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671567.

Downs:1973:RTA

[Dow73] H. Robert Downs. Real-time algorithms and data management on Illiac IV. IEEE Transactions on Computers, C-22(8):773– 777, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009157.

Dandapani:1974:DLN

[DR74] R. Dandapani and S. M. Reddy.

On the design of logic networks with redundancy and testability considerations. IEEE Transactions on Computers, C-23(11):1139–1149, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672414.

Davis:1975:DSE

[DR75] L. S. Davis and A. Rosenfeld. Detection of step edges in noisy onedimensional data. IEEE Transactions on Computers, C-24(10): 1006–1010, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672704.

Davis:1977:CSR

[DR77] L. S. Davis and A. Rosenfeld. Curve segmentation by relaxation labeling. IEEE Transactions on Computers, C-26(10):1053–1057, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674746.

DeMori:1975:SPC

[DRS75] R. De Mori, S. Rivoira, and A. Serra. A special-purpose computer for digital signal processing. IEEE Transactions on Computers, C-24(12):1202–1211, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672757.

Drubin:1971:CFF

[Dru71a] M. Drubin. Computation of the Fast Fourier Transform from data stored in external auxiliary memory for any general radix r = 2^{n}, n ≥ 1. IEEE Transactions on Computers, C-20(12):1552–1558, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671761.

Drubin:1971:KPF

[Dru71b] M. Drubin. Kronecker product factorization of the FFT matrix. IEEE Transactions on Computers, C-20 (5):590–593, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671899.

Dere:1970:BAP

[DS70] W. Y. Dere and D. J. Sakrison.

Berkeley array processor. IEEE Transactions on Computers, C-19 (5):444–447, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671535.

Das:1971:DTP

[DS71] S. R. Das and C. L. Sheng. On detecting total or partial symmetry of switching functions. IEEE Transactions on Computers, C-20 (3):352–355, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671836.

Das:1972:MTP

[DS72a] S. K. Das and D. F. Stanat. A modified training procedure for linear threshold devices. IEEE Transactions on Computers, C-21(4):396– 397, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008983.

DeMori:1972:PSS

[DS72b] R. De Mori and A. Serra. A parallel structure for signednumber multiplication and addition. IEEE Transactions on Computers, C-21(12):1453–1454, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672118.

Dewdney:1973:TMD

[DS73] A. K. Dewdney and A. L. Szilard. Tours in machines and digraphs. IEEE Transactions on Computers, C-22(7):635– 639, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009128.

Delcaro:1974:MTE

[DS74] L. G. Delcaro and G. L. Sicuranza.

A method for transposing externally stored matrices. IEEE Transactions on Computers, C-23(9): 967–970, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672655.

DeVries:1975:MOO

[DS75] R. C. De Vries and A. Svoboda.

Multiple-output optimization with mosaics of Boolean functions. IEEE Transactions on Computers, C-24 (8):777–785, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672902.

De-Sarkar:1974:UCR

[DSBC74] S. C. De-Sarkar, S. Bandyopadhyay, and A. K. Choudhury. Unate cascade realization of synchronous sequential machines. IEEE Transactions on Computers, C-23(10): 1008–1019, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672395.

Donaldson:1970:UCC

[DT70] R. W. Donaldson and G. T.

Toussaint. Use of contextual constraints in recognition of contour-traced handprinted characters. IEEE Transactions on Computers, C-19(11):1096–1099, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671432.

Dasgupta:1976:IMP

[DT76] S. Dasgupta and J. Tartar. The identification of maximal parallelism in straight-line microprograms. IEEE Transactions on Computers, C-25(10):986–992, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674539. See comments [Das78].

David:1979:CEL

[DTG79] R. David and R. Tellez-Giron. Comments on “The Error Latency of a Fault in a Sequential Digital Circuit”. IEEE Transactions on Computers, C-28(1):85– 86, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675230. See [SM76].

Do-Tu:1978:LAN

[DTI78] Hai Do-Tu and M. Installe. Learning algorithms for nonparametric solution to the minimum error classification problem. IEEE Transactions on Computers, C-27(7): 648–659, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675165.

Du:1971:CNS

[Du71] Min-Wen Du. Comments on “Nonlinear Sequential Circuits”. IEEE Transactions on Computers, C-20 (5):585–586, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671891. See [Kai70].

Du:1972:WFL

[Du72] Min-Wen Du. A way to find a lower bound for the minimal solution of the covering problem. IEEE Transactions on Computers, C-21 (3):317–318, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008958.

Dube:1979:PSS

[Dub79] R. P. Dube. Preliminary specification of spline curves. IEEE Transactions on Computers, C-28 (4):286–290, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675347.

Duin:1976:CSP

[Dui76] R. P. W. Duin. On the choice of smoothing parameters for Parzen estimators of probability density functions. IEEE Transactions on Computers, C-25(11):1175–1179, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674577.

Duke:1974:BDM

[Duk74] K. A. Duke. B 74-26 discrete mathematical structures and their applications. IEEE Transactions on Computers, C-23(5): 555, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672575.

Dunn:1975:GAL

[Dun75] J. C. Dunn. Group averaged linear transforms that detect corners and edges. IEEE Transactions on Computers, C-24(12):1191–1201, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672756.

Dutta:1974:EPH

[Dut74] A. K. Dutta. An experimental procedure for handwritten character recognition. IEEE Transactions on Computers, C-23(5): 536–545, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672571.

Druzeta:1978:HRT

[DV78a] A. Druzeta and Z. G. Vranesic. A higher radix technique for fault detection in many-valued multithreshold networks. IEEE Transactions on Computers, C-27(11):1070–1073, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674999.

Dubois:1978:DFT

[DV78b] E. Dubois and A. N. Venetsanopoulos. The discrete Fourier transform over finite rings with application to fast convolution. IEEE Transactions on Computers, C-27 (7):586–593, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675158.

Druzeta:1974:AME

[DVS74] A. Druzeta, Z. G. Vranesic, and A. S. Sedra. Application of multithreshold elements in the realization of many-valued logic networks. IEEE Transactions on Computers, C-23(11):1194–1198, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672422.

Dollhoff:1972:RSE

[DW72a] Terry L. Dollhoff and Bernard L.

Weinberg. A result on set extraction and application to coveringclosure tables. IEEE Transactions on Computers, C-21(6):603– 606, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009018.

Drilman:1972:MNN

[DW72b] J. Drilman and P. Weiner. Modular networks and nondeterministic sequential machines. IEEE Transactions on Computers, C-21(10): 1124–1129, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672054.

Du:1973:CSS

[DW73a] Min-Wen Du and C. Dennis Weiss. Circuit structure and switching function verification. IEEE Transactions on Computers, C-22(6):618–625, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Du.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009116.

Du:1973:MFD

[DW73b] Min-Wen Du and C. Dennis Weiss. Multiple fault detection in combinational circuits: Algorithms and computational results. IEEE Transactions on Computers, C-22(3): 235–240, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672293. See comments [GS74, Bos79].

Davies:1978:SMR

[DW78] D. Davies and J. F. Wakerly. Synchronization and matching in redundant systems. IEEE Transactions on Computers, C-27(6):531– 539, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675144.

Dinn:1970:CCI

[DWT70] D. F. Dinn, D. A. Winter, and B. G. Trenholm. CINTEL — Computer Interface for Television. IEEE Transactions on Computers, C-19(11):1091–1095, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671431.

Easton:1978:CCS

[Eas78] M. C. Easton. Computation of cold-start miss ratios. IEEE Transactions on Computers, C-27(5): 404–408, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675119.

Edwards:1975:ARW

[Edw75] C. R. Edwards. The application of the Rademacher–Walsh transform to Boolean function classification and threshold logic synthesis. IEEE Transactions on Computers, C-24 (1):48–62, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672675. See comments [Yue76].

Edwards:1976:AR

[Edw76] C. R. Edwards. Author’s reply. IEEE Transactions on Computers, C-25(7):767, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674689.

Ehrich:1976:RRW

[EF76] R. W. Ehrich and J. P. Foith. Representation of random waveforms by relational trees. IEEE Transactions on Computers, C-25 (7):725–736, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674681.

Easton:1979:UBS

[EF79] M. C. Easton and P. A. Franaszek. Use bit scanning in replacement decisions. IEEE Transactions on Computers, C-28(2):133–141, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675302.

Erman:1976:SOS

[EFLR76] L. D. Erman, R. D. Fennell, V. R.

Lesser, and D. R. Reddy. System organizations for speech understanding: Implications of network and multiprocessor computer architectures for AI. IEEE Transactions on Computers, C-25(4):414– 421, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674623.

El-Ghoroury:1971:RSA

[EGG71] H. N. El-Ghoroury and S. C.

Gupta. Realization of stochastic automata. IEEE Transactions on Computers, C-20(8):889– 893, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671959.

El-Ghoroury:1972:ABN

[EGG72] H. N. El-Ghoroury and S. C. Gupta.

Additive Bernoulli noise linear sequential circuits. IEEE Transactions on Computers, C-21(10): 1119–1124, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672053.

Ernst:1976:UHO

[EH76] G. W. Ernst and R. J. Hookway. The use of higher order logic in program verification. IEEE Transactions on Computers, C-25(8):844– 851, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674703.

Edwards:1978:DSP

[EH78] C. R. Edwards and S. L. Hurst. A digital synthesis procedure under function symmetries and mapping methods. IEEE Transactions on Computers, C-27(11):985– 997, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674988.

Ehrich:1972:NSM

[Ehr72] H.-D. Ehrich. A note on state minimization of a special class of incomplete sequential machines. IEEE Transactions on Computers, C-21 (5):500–502, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672142.

Ehrich:1976:RCD

[Ehr76] R. W. Ehrich. A readily computable decision rule with variable dimensionality. IEEE Transactions on Computers, C-25(5): 539–542, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674644.

Ehrich:1977:DGE

[Ehr77] R. W. Ehrich. Detection of global edges in textured images. IEEE Transactions on Computers, C-26 (6):589–603, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674884.

Etiemble:1977:ITC

[EI77] D. Etiemble and M. Isreal. Implementation of ternary circuits with-binary integrated circuits. IEEE Transactions on Computers, C-26(12):1222–1233, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674783.

Ehrich:1975:ECR

[EK75] R. W. Ehrich and K. J. Koehler. Experiments in the contextual recognition of cursive script. IEEE Transactions on Computers, C-24(2): 182–194, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672777.

Eklundh:1972:FCM

[Ekl72] J. O. Eklundh. A fast computer method for matrix transposing. IEEE Transactions on Computers, C-21(7):801–803, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672177.

See also [Ekl73, Sch73, All75a, Ram75, TE76, Ari79].

Eklundh:1973:AR

[Ekl73] J. O. Eklundh. Author’s reply. IEEE Transactions on Computers, C-22(5):543–544, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672354.

Ekstrom:1973:NAI

[Eks73] M. P. Ekstrom. A numerical algorithm for identifying spread functions of shift-invariant imaging systems. IEEE Transactions on Computers, C-22(4):322– 328, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672311.

Ekstorm:1974:IIA

[Eks74] M. P. Ekstorm. An iterativeimprovement approach to the numerical solution of vector Toeplitz systems. IEEE Transactions on Computers, C-23(3):320–325, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672522.

Elmasry:1976:LDU

[Elm76] M. I. Elmasry. Logic design using EFL structures. IEEE Transactions on Computers, C-25(9):952– 956, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674722.

Elschlager:1974:BDM

[Els74] R. Elschlager. B74-8 design of mancomputer dialogues. IEEE Transactions on Computers, C-23(2):212– 213, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672490.

En:1977:NRE

[En77] John En. A new random-error-correction code. IEEE Transactions on Computers, C-26(1):86– 88, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009279. See comments [Paa79].

En:1979:AR

[En79] J. En. Author’s reply. IEEE Transactions on Computers, C-28 (3):257–258, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675330.

Eschenbach:1978:ECR

[EO78] R. F. Eschenbach and B. M. Oliver. An efficient coordinate rotation algorithm. IEEE Transactions on Computers, C-27(12):1178–1180, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675020.

Epstein:1973:EAD

[Eps73a] G. Epstein. An equational axiomatization for the disjoint system of post algebras. IEEE Transactions on Computers, C-22(4):422– 423, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672324.

Epstein:1973:CRB

[Eps73b] George Epstein. Comments on “The Relationship Between Multivalued Switching Algebra and Boolean Algebra Under Different Definitions of Complement”. IEEE Transactions on Computers, C-22(9):864, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009181. See [SS72, DM76].

Ercegovac:1973:REC

[Erc73] Miloš D. Ercegovac. Radix-16 evaluation of certain elementary functions. IEEE Transactions on Computers, C-22(6):561–566, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Ercegovac.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009107.

Ercegovac:1977:GHO

[Erc77] Miloš D. Ercegovac. A general hardware-oriented method for evaluation of functions and computations in a digital computer. IEEE Transactions on Computers, C-26 (7):667–680, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674900.

Ernbo:1972:AIM

[Ern72] Arne Ernbo. Application of intensity-modulated ink jets to alphanumeric printing devices. IEEE Transactions on Computers, C-21(9):942–947, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009070.

Ernst:1976:DDT

[Ern76] G. W. Ernst. A definition-driven theorem prover. IEEE Transactions on Computers, C-25(4):317– 322, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674611.

Ecker:1977:CWC

[ES77] K. Ecker and D. Schutt. Computational work of combinational machines. IEEE Transactions on Computers, C-26(11):1161–1163, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674768.

Elmasry:1975:TLE

[ET75] M. I. Elmasry and P. M. Thompson. Two-level emitter-function logic structures for logic-in-memory computers. IEEE Transactions on Computers, C-24(3):250–258, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672799.

Ejiri:1972:PIR

[EUY^{+}72] Masakazu Ejiri, Takeshi Uno, Haruo Yoda, Tatsuo Goto, and Kiyoo Takeyasu. A prototype intelligent robot that assembles objects from plan drawings. IEEE Transactions on Computers, C-21(2):161– 170, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
5008921.

Elineau:1977:NJK

[EW77] G. Elineau and W. Wiesbeck. A new J–K flip-flop for synchronizers. IEEE Transactions on Computers, C-26(12):1277–1279, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674789.

El-Ziq:1978:CAL

[EZS78] Y. M. El-Ziq and S. Y. H.

Su. Computer-aided logic design of two-level MOS combinational networks with statistical results. IEEE Transactions on Computers, C-27(10):911–923, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674970.

Fino:1976:UMT

[FA76] B. J. Fino and V. R. Algazi. Unified matrix treatment of the fast Walsh–Hadamard transform. IEEE Transactions on Computers, C-25(11):1142–1146, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674569.

Faiman:1973:BDC

[Fai73] M. Faiman. B73-11 digital circuits and devices. IEEE Transactions on Computers, C-22(5): 546–547, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672358.

Farmer:1973:ADF

[Far73] D. E. Farmer. Algorithms for designing fault-detection experiments for sequential machines. IEEE Transactions on Computers, C-22(2):159–167, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672271.

Farhat:1975:NOD

[Far75] N. H. Farhat. Nonlinear optical data processing and filtering: a feasibility study. IEEE Transactions on Computers, C-24(4):443– 448, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672833.

Farag:1979:WLR

[Far79] R. F. H. Farag. Word-level recognition of cursive script. IEEE Transactions on Computers, C-28(2): 172–175, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675310.

Fernandez:1973:BNP

[FB73a] Eduardo B. Fernandez and Bertram Bussell. Bounds on the number of processors and time for multiprocessor optimal schedules. IEEE Transactions on Computers, C-22 (8):745–751, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009153.

Fu:1973:TSS

[FB73b] King-Sun Fu and B. K. Bhargava.

Tree systems for syntactic pattern recognition. IEEE Transactions on Computers, C-22(12):1087–1099, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672247.

French:1974:UWF

[FB74] A. S. French and E. G. Butz.

The use of Walsh functions in the Wiener analysis of nonlinear systems. IEEE Transactions on Computers, C-23(3):225– 232, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672509.

Friedman:1975:AFN

[FBS75] J. H. Friedman, F. Baskett, and L. J. Shustek. An algorithm for finding nearest neighbors. IEEE Transactions on Computers, C-24(10): 1000–1006, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672703.

Frei:1977:FBD

[FC77] W. Frei and Chung-Ching Chen. Fast boundary detection: a generalization and a new algorithm. IEEE Transactions on Computers, C-26(10):988–998, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674733.

Fu:1972:AR

[FCC72] K. S. Fu, Y. T. Chien, and G. P. Cardillo. Authors’ reply. IEEE Transactions on Computers, C-21 (3):320, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008960.

Fisher:1973:BPT

[FD73] L. T. Fisher and D. W. Dearholt. Boundary points of threshold functions. IEEE Transactions on Computers, C-22(12):1132–1139, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672252.

Filene:1974:RIM

[FD74a] R. J. Filene and W. M. Daly.

The reliability impact of mission abort strategies on redundant flight computer systems. IEEE Transactions on Computers, C-23(7): 739–743, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672616.

Fridrich:1974:MFT

[FD74b] M. Fridrich and W. A. Davis.

Minimal fault tests for combinational networks. IEEE Transactions on Computers, C-23(8):850– 859, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672633. See comments [Red77a].

Fram:1975:QEE

[FD75] J. R. Fram and E. S. Deutsch.

On the quantitative evaluation of edge detection schemes and their comparison with human performance. IEEE Transactions on Computers, C-24(6):616– 628, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672867.

Freeman:1977:CFA

[FD77a] H. Freeman and L. S. Davis. A corner-finding algorithm for chaincoded curves. IEEE Transactions on Computers, C-26(3):297– 303, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674825.

Fridrich:1977:MFT

[FD77b] M. Fridrich and W. A. Davis. Minimal fault tests for redundant combinational networks. IEEE Transactions on Computers, C-26(10): 1057–1060, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674747.

Fleischhammer:1979:ABF

[FD79] W. Fleischhammer and O. Dortok. The anomalous behavior of flipflops in synchronizer circuits. IEEE Transactions on Computers, C-28 (3):273–276, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675337.

Fischler:1973:RMP

[FE73] M. A. Fischler and R. A. Elschlager. The representation and matching of pictorial structures. IEEE Transactions on Computers, C-22(1):67– 92, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672195.

Fehlauer:1978:DCF

[FE78] J. Fehlauer and B. A. Eisenstein. A declustering criterion for feature extraction in pattern recognition. IEEE Transactions on Computers, C-27(3):261– 266, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675083.

Feldman:1979:AOP

[Fel79] M. B. Feldman. An applicationoriented programming language for sequential machine studies. IEEE Transactions on Computers, C-28 (8):582–586, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675417.

Feng:1974:DMF

[Fen74] Tse-Yun Feng. Data manipulating functions in parallel processors and their implementations. IEEE Transactions on Computers, C-23 (3):309–318, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672520.

Feng:1977:PPP

[Fen77] Tse-Yun Feng. Parallel processors and processing: An introduction. IEEE Transactions on Computers, C-26(2):97–98, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009288.

Ferguson:1977:ANG

[Fer77] J. C. Ferguson. Algorithms in numerical geometry system. IEEE Transactions on Computers, C-26(12):1290–1292, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674794.

Feustel:1972:BCC

[Feu72] E. A. Feustel. B72-9 compiler construction for digital computers. IEEE Transactions on Computers, C-21(5):517, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672152.

Feustel:1973:ATA

[Feu73] Edward A. Feustel. On the advantages of tagged architecture. IEEE Transactions on Computers, C-22 (7):644–656, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
5009130.

Fung:1975:SSD

[FF75] Lai-Wo Fung and King-Sun Fu.

Stochastic syntactic decoding for pattern classification. IEEE Transactions on Computers, C-24(6): 662–667, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672871.

Foster:1971:CIO

[FG71a] C. C. Foster and R. Gonter.

Conditional interpretation of operation codes. IEEE Transactions on Computers, C-20(1):108– 111, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671684.

Frosini:1971:PIA

[FG71b] G. Frosini and G. B. Gerace.

Pulse input asynchronous sequential circuits. IEEE Transactions on Computers, C-20(4):436– 442, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671853.

Frosini:1971:USS

[FG71c] G. Frosini and G. B. Gerace.

A universal STT state assignment method for pulse input asynchronous sequential circuits. IEEE Transactions on Computers, C-20 (8):856–861, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671954.

Frosini:1976:SAP

[FG76] G. Frosini and G. B. Gerace. Synchronous and asynchronous procedures for the synthesis of masterslave circuits. IEEE Transactions on Computers, C-25(11):1179–1184, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674578.

Foster:1971:MOC

[FGR71] C. C. Foster, R. H. Gonter, and E. M. Riseman. Measures of opcode utilization. IEEE Transactions on Computers, C-20(5):582– 584, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671889.

Felstead:1978:CDC

[FHTS78] E. B. Felstead, K. O. Hill, and A. U. Tenne-Sens. Comments on “Direct Coherent Optical Fourier Transforms of Curves”. IEEE Transactions on Computers, C-27(2):190– 191, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=

1675063. See [LH75].

Fisher:1974:UPE

[Fis74] L. T. Fisher. Unateness properties of and-exclusive-or logic circuits. IEEE Transactions on Computers, C-23(2):166–172, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672472.

Fishman:1976:POS

[Fis76] D. H. Fishman. A problemoriented search procedure for theorem proving. IEEE Transactions on Computers, C-25(8):807– 815, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674699.

Fukunaga:1970:AKL

[FK70a] K. Fukunaga and W. L. G. Koontz.

Application of the Karhunen– Loève expansion to feature selection and ordering. IEEE Transactions on Computers, C-19(4):311– 318, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671511.

Fukunaga:1970:CAG

[FK70b] K. Fukunaga and W. L. G. Koontz.

A criterion and an algorithm for grouping data. IEEE Transactions on Computers, C-19(10):917– 923, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671392.

Fukunaga:1971:ECE

[FK71] K. Fukunaga and D. L. Kessell. Estimation of classification error. IEEE Transactions on Computers, C-20(12):1521–1527, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671758.

Fujiwara:1974:DDS

[FK74] H. Fujiwara and K. Kinoshita. Design of diagnosable sequential machines utilizing extra outputs. IEEE Transactions on Computers, C-23(2):138–145, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672469.

Fujiwara:1978:CAP

[FK78a] H. Fujiwara and K. Kinoshita. Connection assignments for probabilistically diagnosable systems. IEEE Transactions on Computers, C-27 (3):280–283, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675087.

Fujiwara:1978:CCS

[FK78b] H. Fujiwara and K. Kinoshita. On the computational complexity of system diagnosis. IEEE Transactions on Computers, C-27(10):881– 885, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674966.

Fujiwara:1978:SET

[FK78c] H. Fujiwara and K. Kinoshita. Some existence theorems for probabilistically diagnosable systems. IEEE Transactions on Computers, C-27 (4):379–384, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675111.

Frieder:1975:ABC

[FL75] G. Frieder and C. Luk. Algorithms for binary coded balanced and ordinary ternary operations. IEEE Transactions on Computers, C-24(2):212–215, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672781.

Fennell:1977:PAI

[FL77a] Richard D. Fennell and Victor R.

Lesser. Parallelism in artificial intelligence problem solving: a case study of Hearsay II. IEEE Transactions on Computers, C-26(2):98– 111, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009289.

Finnila:1977:ALA

[FL77b] Charles A. Finnila and Hubert H. Love. The associative linear array processor. IEEE Transactions on Computers, C-26(2):112– 125, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009290.

Fuller:1976:EET

[FLBK76] S. H. Fuller, V. R. Lesser, C. G. Bell, and C. H. Kaman. The effects-of emerging technology and emulation requirements on microprogramming. IEEE Transactions on Computers, C-25(10): 1000–1009, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674541.

Fleisher:1979:ISS

[Fle79] H. Fleisher. Introduction to special section on programmable logic arrays. IEEE Transactions on Computers, C-28(9):593, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675425.

Flores:1972:BCO

[Flo72] I. Flores. B72-26 computer organization and the System/ 370. IEEE Transactions on Computers, C-21(12):1458–1459, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672121.

Flores:1975:CRP

[Flo75] I. Flores. Comments on the review of “Peripheral Devices”. IEEE Transactions on Computers, C-24(2):223–224, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672789.

See [Bar74a].

Fusco:1978:QFM

[FLR78] F. Fusco, S. Levialdi, and L. Ritondale. A quality factor for measuring program efficiency. IEEE Transactions on Computers, C-27 (3):275–280, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675086.

Flynn:1970:DFI

[Fly70] M. J. Flynn. On division by functional iteration. IEEE Transactions on Computers, C-19(8):702–706, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        www.acsel-lab.com/arithmetic/
                                        arith1/papers/ARITH1_Flynn.pdf;
                                         https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671612. Reprinted in [Swa90].

Flynn:1972:BCS

[Fly72a] M. J. Flynn. B72-3 computer structures, readings and examples. IEEE Transactions on Computers, C-21(1):109–110, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672037.

Flynn:1972:SCO

[Fly72b] Michael J. Flynn. Some computer organizations and their effectiveness. IEEE Transactions on Computers, C-21(9):948–960, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009071.

Friedman:1971:SAO

[FM71] A. D. Friedman and P. R. Menon. Systems of asynchronously operating modules. IEEE Transactions on Computers, C-20(1):100– 104, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671681.

Freeman:1972:FTC

[FM72a] Harvey A. Freeman and Gernot Metze. Fault-tolerant computers using “Dotted Logic” redundancy techniques. IEEE Transactions on Computers, C-21(8):867– 871, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009042.

Friedman:1972:CDD

[FM72b] A. D. Friedman and P. R. Menon.

Comments on “Design of Diagnosable Iterative Arrays”. IEEE Transactions on Computers, C-21 (5):511, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672146. See [LY71].

Friedman:1973:RCS

[FM73] A. D. Friedman and P. R. Menon.

Restricted checking sequences for sequential machines. IEEE Transactions on Computers, C-22(4): 397–399, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672320.

Fantauzzi:1974:MFD

[FM74] G. Fantauzzi and A. Marsella.

Multiple-fault detection and location in fan-out free combinational circuits. IEEE Transactions on Computers, C-23(1):48– 55, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672369.

Fuller:1978:MLC

[FM78] S. H. Fuller and P. F. McGehearty. Minimizing latency in CCD memories. IEEE Transactions on Computers, C-27(3):252– 254, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675079.

Fukunage:1975:BBA

[FN75] K. Fukunage and P. M. Narendra. A branch and bound algorithm for computing k-nearest neighbors. IEEE Transactions on Computers, C-24(7):750–753, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672890.

Fujiwara:1975:ETS

[FNSK75] H. Fujiwara, Y. Nagao, T. Sasao, and K. Kinoshita. Easily testable sequential machines with extra inputs. IEEE Transactions on Computers, C-24(8):821– 826, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672906.

Fukunaga:1971:AFI

[FO71a] K. Fukunaga and D. R. Olsen. An algorithm for finding intrinsic dimensionality of data. IEEE Transactions on Computers, C-20(2): 176–183, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671801. See comments [Tru71].

Fukunaga:1971:AR

[FO71b] K. Fukunaga and D. R. Olsen. Authors’ reply. IEEE Transactions on Computers, C-20(12):1615–1616, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671780.

Fukunaga:1971:TDD

[FO71c] K. Fukunaga and D. R. Olsen.

A two-dimensional display for the classification of multivariate data. IEEE Transactions on Computers, C-20(8):917–923, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671964.

Fontao:1971:GMC

[Fon71] R. O. Fontao. A graphical method for checking complete monotonicity. IEEE Transactions on Computers, C-20(4):461– 464, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671860.

Fosdick:1973:BAC

[Fos73] L. D. Fosdick. B73-2 approximate calculation of multiple integrals. IEEE Transactions on Computers, C-22(1):112, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672201.

Feng:1975:DPS

[FP75] H.-Y. F. Feng and T. Pavlidis. Decomposition of polygons into simpler components: Feature generation for syntactic pattern recognition. IEEE Transactions on Computers, C-24(6):636– 650, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672869.

Flynn:1971:MIV

[FR71] M. J. Flynn and R. F. Rosin. Microprogramming: An introduction and a viewpoint. IEEE Transactions on Computers, C-20(7): 727–731, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671934.

Foster:1972:PCE

[FR72] C. C. Foster and E. M. Riseman. Percolation of code to enhance parallel dispatching and execution. IEEE Transactions on Computers, C-21(12):1411–1415, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672108.

Fu:1976:PRI

[FR76] King-Sun Fu and A. Rosenfeld. Pattern recognition and image processing. IEEE Transactions on Computers, C-25(12):1336–1346, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674602.

Frackowiak:1972:SMH

[Fra72] J. Frackowiak. The synthesis of minimal hazardless TANT networks. IEEE Transactions on Computers, C-21(10):1099–1108, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672049.

Franklin:1978:PSO

[Fra78] M. A. Franklin. Parallel solution of ordinary differential equations. IEEE Transactions on Computers, C-27(5):413–420, May 1978. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675121.

Freeman:1974:BFA

[Fre74a] P. Freeman. B 74-34 Fortran and Algol: a programmed course for students of science and technology. IEEE Transactions on Computers, C-23(5):559– 560, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672583.

Freeman:1974:BOS

[Fre74b] P. Freeman. B74-13 operating systems — Infotech state of the art report #14. IEEE Transactions on Computers, C-23(2):216– 217, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672495.

French:1977:CSS

[Fre77] A. S. French. Computer simulation of space-filling molecular models. IEEE Transactions on Computers, C-26(10):1026–1028, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674738.

Freeman:1979:AGD

[Fre79a] H. Freeman. Algorithm for generating a digital straight line on a triangular grid. IEEE Transactions on Computers, C-28(2):150– 152, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675305.

Freeman:1979:AUB

[Fre79b] M. Freeman. Aspects of the upper bounds of finite input-memory and finite output-memory sequential machines. IEEE Transactions on Computers, C-28(3):249– 253, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675327.

Friedman:1970:RAS

[Fri70] A. D. Friedman. R70-17 asynchronous sequential circuits with feedback. IEEE Transactions on Computers, C-19(7):657, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671597.

Friedman:1973:ETI

[Fri73] A. D. Friedman. Easily testable iterative systems. IEEE Transactions on Computers, C-22(12):1061–1064, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672244.

Friedman:1974:DSC

[Fri74] A. D. Friedman. Diagnosis of short-circuit faults in combinational circuits. IEEE Transactions on Computers, C-23(7):746– 752, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672618.

Friedman:1977:RPD

[Fri77] J. H. Friedman. A recursive partitioning decision rule for nonparametric classification. IEEE Transactions on Computers, C-26(4): 404–408, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674849.

Farber:1970:MLB

[FS70] A. S. Farber and E. S. Schlig. Mathematical “Lower Bounds” and the logic circuit designer. IEEE Transactions on Computers, C-19(1):80– 81, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008905.

Foster:1971:CRA

[FS71] C. C. Foster and F. D. Stockton. Counting responders in an associative memory. IEEE Transactions on Computers, C-20(12):1580–1583, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671768.

Franklin:1974:ART

[FS74] M. A. Franklin and A. Sen. An analytic response time model for single-and dual-density disk systems. IEEE Transactions on Computers, C-23(12):1269–1276, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672439.

Foley:1975:OSD

[FS75] D. H. Foley and J. W. Sammon, Jr. An optimal set of discriminant vectors. IEEE Transactions on Computers, C-24(3):281– 289, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672801.

Fukunaga:1978:GCP

[FS78] K. Fukunaga and R. D. Short.

Generalized clustering for problem localization. IEEE Transactions on Computers, C-27(2):176– 181, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675056.

Fernandez:1978:ASS

[FSLC78] E. B. Fernandez, R. C. Summers, T. Lang, and C. D. Coleman. Architectural support for system protection and database security. IEEE Transactions on Computers, C-27 (8):767–771, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675189.

Friedman:1974:PPA

[FT74] J. H. Friedman and J. W. Tukey. A projection pursuit algorithm for exploratory data analysis. IEEE Transactions on Computers, C-23(9):881–890, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672644.

Fraley:1976:RAM

[FT76] D. Fraley and J. Talavage. A realization algorithm for multiple preset experiments. IEEE Transactions on Computers, C-25(11):1130–1132, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674564.

Fung:1979:AMC

[FT79] Kwok-Tung Fung and H. C. Torng. On the analysis of memory conflicts and bus contentions in a multiplemicroprocessor system. IEEE Transactions on Computers, C-28 (1):28–37, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675219.

Fu:1974:BFT

[Fu74] K. S. Fu. B 74-30 foundation of the theory of learning systems. IEEE Transactions on Computers, C-23 (5):557, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672579.

Fuller:1972:ODS

[Ful72] S. H. Fuller. An optimal drum scheduling algorithm. IEEE Transactions on Computers, C-21(11): 1153–1165, November 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672065.

Frosini:1979:AEF

[FV79] G. Frosini and F. M. Viterbo. An algorithm for evaluating the frequency of a rotating vector. IEEE Transactions on Computers, C-28 (8):560–566, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675411.

Friedman:1978:AAP

[FW78] D. P. Friedman and D. S. Wise. Aspects of applicative programming for parallel processing. IEEE Transactions on Computers, C-27 (4):289–296, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675100.

Gaddess:1970:EDB

[Gad70] Terry G. Gaddess. An errordetecting binary adder: a hardwareshared implementation. IEEE Transactions on Computers, C-19 (1):34–38, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008897.

Gagliardi:1972:BFG

[Gag72] U. O. Gagliardi. B72-27 the Fourth Generation. IEEE Transactions on Computers, C-21(12): 1459, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672122.

Gaines:1979:MGI

[Gai79] B. R. Gaines. Maryanski’s grammatical inferencer. IEEE Transactions on Computers, C-28(1):62– 64, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675224.

Gajewski:1975:PEF

[Gaj75] T. Gajewski. On the periodic equivalents of finite automata. IEEE Transactions on Computers, C-24(10):991–994, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672700.

Galperin:1978:LPC

[Gal78] E. A. Galperin. Loop properties and controllability of linear congruential sequences. IEEE Transactions on Computers, C-27(1):68– 76, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674954.

Gardner:1971:FMM

[Gar71] P. L. Gardner. Functional memory and its microprogramming implications. IEEE Transactions on Computers, C-20(7):764– 775, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671938.

Garey:1972:RBC

[Gar72a] M. R. Garey. Resident-bubble cellular logic using magnetic domains. IEEE Transactions on Computers, C-21(4):392–396, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5008982.

Garey:1972:SBI

[Gar72b] M. R. Garey. Simple binary identification problems. IEEE Transactions on Computers, C-21(6): 588–590, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009013.

Gardiner:1974:CAI

[Gar74] A. B. Gardiner. Comments on “An Augmented Iterative Array for High Speed Division”. IEEE Transactions on Computers, C-23(3):326– 327, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672524. See [CH73a].

Garner:1976:SSR

[Gar76] H. L. Garner. A survey of some recent contributions to computer arithmetic. IEEE Transactions on Computers, C-25(12):1277–1282, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674595.

Garner:1978:TCA

[Gar78] H. L. Garner. Theory of computer addition and overflows. IEEE Transactions on Computers, C-27 (4):297–301, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675101.

Gay:1978:EMS

[Gay78] F. A. Gay. Evaluation of maintenance software in realtime systems. IEEE Transactions on Computers, C-27(6):576– 582, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675151.

Grzymala-Busse:1971:DPR

[GB71a] J. W. Grzymala-Busse. On the decomposition of periodic representations of sequential machines. IEEE Transactions on Computers, C-20 (8):929–933, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671967.

Grzymala-Busse:1971:PRP

[GB71b] J. W. Grzymala-Busse. Periodic representations and Tpartitionable equivalents of sequential machines. IEEE Transactions on Computers, C-20(2):190– 198, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671803.

Ghosh:1972:SMT

[GBMC72] S. Ghosh, S. Bandyopadhyay, S. K. Mitra, and A. K. Choudhury. Simple methods for the testing of 2-summability of Boolean functions and isobaricity of threshold functions. IEEE Transactions on Computers, C-21(5):503– 507, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672144.

Ghosh:1971:CMN

[GC71a] S. Ghosh and A. K. Choudhury. Cascaded multithreshold networks. IEEE Transactions on Computers, C-20(6):655– 662, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671910.

Griswold:1971:LSC

[GC71b] V. M. Griswold and C. C. Carroll. Large-scale circuit interconnection for Boolean function implementation. IEEE Transactions on Computers, C-20(5):572– 575, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671885.

Ghosh:1973:PBF

[GC73] S. Ghosh and A. K. Choudhury. Partition of Boolean functions for realization with multithreshold threshold logic elements. IEEE Transactions on Computers, C-22(2):204–215, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672279.

Geadah:1977:NDS

[GC77] Y. A. Geadah and M. J. G.

Corinthios. Natural, dyadic, and sequency order algorithms and processors for the Walsh– Hadamard transform. IEEE Transactions on Computers, C-26(5): 435–442, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674860.

Gupta:1974:LTR

[GCC74] A. S. Gupta, D. K. Chattopadhyay, and A. K. Choudhury. Linearity testing and realization of sequential machines. IEEE Transactions on Computers, C-23(5): 524–528, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672567.

Gupta:1977:RFT

[GCP^{+}77] A. Sen Gupta, D. K. Chattopadhyay, A. Palit, A. K. Bandyopadhyay, M. S. Basu, and A. K. Choudhury. Realization of fault-tolerant and fail-safe sequential machines. IEEE Transactions on Computers, C-26(1):91– 96, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009281.

Goke:1972:DRP

[GD72] L. R. Goke and K. L. Doty. Design of a random-pulse computer for classifying binary patterns. IEEE Transactions on Computers, C-21(12):1347–1354, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672099.

Gupta:1973:FSM

[GD73] P. K. Gupta and D. L. Dietmeyer. Fast state minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-22(2):215– 217, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672280.

Gecsei:1977:INT

[Gec77] J. Gecsei. Interconnection networks from three state cells. IEEE Transactions on Computers, C-26 (8):705–711, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674908.

Gelenbe:1971:RMS

[Gel71] S. E. Gelenbe. A realizable model for stochastic sequential machines. IEEE Transactions on Computers, C-20(2):199–204, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671804.

Gelenbe:1973:BTS

[Gel73a] E. Gelenbe. B73-10 time sharing systems. IEEE Transactions on Computers, C-22(5): 546, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672357.

Gelenbe:1973:UAE

[Gel73b] Erol Gelenbe. A unified approach to the evaluation of a class of replacement algorithms. IEEE Transactions on Computers, C-22(6):611– 618, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Gelenbe.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009115.

Gelperin:1976:RBP

[Gel76] D. Gelperin. A resolution-based proof procedure using deletiondirected search. IEEE Transactions on Computers, C-25(4):323– 327, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674612.

Geoffrion:1970:RAA

[Geo70] A. M. Geoffrion. R70-4 analysis of algorithms for the zeroone programming problem. IEEE Transactions on Computers, C-19 (5):464, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671548.

Gestri:1970:SMS

[Ges70] G. Gestri. Synthesis of multiple sequential machines having different inputs. IEEE Transactions on Computers, C-19(11):1105–1108, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671434.

Geselowitz:1971:UME

[Ges71] D. B. Geselowitz. Use of the multipole expansion to extract significant features of the surface electrocardiogram. IEEE Transactions on Computers, C-20(9):1086– 1089, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672000.

Gupta:1978:WSP

[GF78] R. K. Gupta and M. A. Franklin.

Working set and page fault frequency paging algorithms: a performance comparison. IEEE Transactions on Computers, C-27(8):706– 712, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675177.

Gabrielian:1973:SSA

[GG73] A. Gabrielian and S. Ginsburg.

Structured storage AFA. IEEE Transactions on Computers, C-22 (5):534–537, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672351.

Gibson:1975:SCT

[GG75] J. A. Gibson and R. W. Gibbard. Synthesis and comparison of two’s complement parallel multipliers. IEEE Transactions on Computers, C-24(10): 1020–1027, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672710.

Gerace:1971:TSP

[GGMM71] G. B. Gerace, L. Gilli, P. Maestrini, and A. R. Meo. TOPI — a specialpurpose computer for Boolean analysis and synthesis. IEEE Transactions on Computers, C-20(8):837– 842, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671951.

Govindan:1972:CIT

[GH72] G. N. Govindan and R. A. Higgins. Convergence of an identification technique for nonlinear systems. IEEE Transactions on Computers, C-21(11):1216–1219, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672072.

Gaitanis:1978:NDR

[GH78] N. Gaitanis and C. Halatsis. A new double-rank realization of sequential machines. IEEE Transactions on Computers, C-27(12):1137–1143, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675016.

Garey:1977:ASP

[GHJ77] M. R. Garey, F. K. Hwang, and D. S. Johnson. Algorithms for a set partitioning problem arising in the design of multipurpose units. IEEE Transactions on Computers, C-26 (4):321–328, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674841.

Gill:1970:SCM

[Gil70a] A. Gill. Single-channel and multichannel finite-state machines. IEEE Transactions on Computers, C-19(11):1073–1078, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671427.

See comments [Ric72].

Gilliland:1970:RSC

[Gil70b] M. C. Gilliland. R70-7 stability controls for the analysis of analog/digital hybrid loops. IEEE Transactions on Computers, C-19 (5):465, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671551.

Gillespie:1974:BPR

[Gil74a] R. C. Gillespie. B74-19 pattern recognition techniques. IEEE Transactions on Computers, C-23 (2):220–222, February 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672501.

Gillespie:1974:BCO

[Gil74b] R. C. Gillespie. B74-5 computeroriented approaches to pattern recognition. IEEE Transactions on Computers, C-23(1):112, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672388.

Gitman:1970:ADU

[GL70] I. Gitman and M. D. Levine. An algorithm for detecting unimodal fuzzy sets and its application as a clustering technique. IEEE Transactions on Computers, C-19(7): 583–593, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671585.

Giloi:1974:FDS

[GL74] W. K. Giloi and H. Liebig. A formalism for description and synthesis of logical algorithms and their hardware implementation. IEEE Transactions on Computers, C-23(9):897–906, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672646.

Glassman:1970:GFF

[Gla70] J. A. Glassman. A generalization of the Fast Fourier Transform. IEEE Transactions on Computers, C-19(2):105–116, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671468.

Gupta:1979:OSC

[GLL79] U. I. Gupta, D. T. Lee, and J. Y.-T. Leung. An optimal solution for the channel-assignment problem. IEEE Transactions on Computers, C-28(11):807–810, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675260.

Gold:1971:FFP

[GLMR71] B. Gold, I. L. Lebow, P. G.

McHugh, and C. M. Rader. The FDP, a fast programmable signal processor. IEEE Transactions on Computers, C-20(1):33– 38, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671671.

Givone:1971:MSM

[GLR71] D. D. Givone, M. E. Liebler, and R. P. Roesser. A method of solution for multiple-valued logic expressions. IEEE Transactions on Computers, C-20(4):464– 467, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671861.

Glucksman:1971:MCP

[Glu71] H. A. Glucksman. Multicategory classification of patterns represented by high-order vectors of multilevel measurements. IEEE Transactions on Computers, C-20(12):1593–1598, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671772.

Goldberg:1974:OHS

[GLW74] J. Goldberg, K. N. Levitt, and J. H. Wensley. An organization for a highly survivable memory. IEEE Transactions on Computers, C-23 (7):693–705, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672610.

Grasselli:1970:MRO

[GM70] A. Grasselli and U. Montanari. On the minimization of READ-ONLY memories in microprogrammed digital computers. IEEE Transactions on Computers, C-19(11):1111–1114, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671436.

Gray:1971:LFC

[GM71] F. G. Gray and J. F. Meyer. Locatability of faults in combinational networks. IEEE Transactions on Computers, C-20(11):1407–1412, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671740.

Gibson:1972:FHC

[GM72a] John A. Gibson and Terence W.

Marks. Fast hybrid computer implementation of the Dynostat algorithm. IEEE Transactions on Computers, C-21(8):872– 880, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009043.

Gose:1972:AR

[GM72b] Earl E. Gose and Anthony N. Mucciardi. Authors’ reply. IEEE Transactions on Computers, C-21 (9):1029, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009090.

Gray:1976:APF

[GM76] F. G. Gray and J. F. Meyer. Algebraic properties of functions affecting optimum fault-tolerant realizations. IEEE Transactions on Computers, C-25(11):1078–1088, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674558.

Godbole:1971:AMC

[God71] S. S. Godbole. An algorithm for the machine computation of partialfractions expansion of functions having multiple poles. IEEE Transactions on Computers, C-20(10): 1147–1152, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671692.

Godbole:1973:ECM

[God73] Sadashiva S. Godbole. On efficient computation of matrix chain products. IEEE Transactions on Computers, C-22(9):864–866, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009182.

Golay:1972:SDL

[Gol72] Marcel J. E. Golay. Smoothing of data by least squares procedures and by filtering. IEEE Transactions on Computers, C-21(3):299– 301, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008953.

Golumbic:1976:CM

[Gol76] M. C. Golumbic. Combinatorial merging. IEEE Transactions on Computers, C-25(11):1164–1167, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674574.

Golay:1978:T

[Gol78] M. J. E. Golay. Topoglyphs.

IEEE Transactions on Computers, C-27(2):164–167, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675051.

Gonenc:1970:MDF

[Gon70] G. Gonenc. A method for the design of fault detection experiments. IEEE Transactions on Computers, C-19(6):551– 558, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671568.

Gonzalez:1979:NOS

[Gon79] T. Gonzalez. A note on open shop preemptive schedules. IEEE Transactions on Computers, C-28(10): 782–786, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675246.

Goodman:1970:RRT

[Goo70] E. D. Goodman. R70-18 realtime computation by n-dimensional iterative arrays of finite-state machines. IEEE Transactions on Computers, C-19(7):657–658, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671598.

Good:1971:RBT

[Goo71] I. J. Good. The relationship between two Fast Fourier Transforms. IEEE Transactions on Computers, C-20(3):310– 317, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671829.

Gorenstein:1975:BSS

[Gor75] S. Gorenstein. B75-2 systems simulation. IEEE Transactions on Computers, C-24(1):110–111, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672687.

Gouraud:1971:CDC

[Gou71] Henri Gouraud. Continuous shading of curved surfaces. IEEE Transactions on Computers, C-20(6): 623–629, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671906.

Gourishankar:1977:CAA

[Gou77] V. Gourishankar. Comments on “An Additional Algorithm for a System of Coupled Algebraic Matrix Riccati Equations”. IEEE Transactions on Computers, C-26 (3):320, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674836. See [Den76].

Goyal:1976:NAR

[Goy76] L. N. Goyal. A note on Atrubin’s real-time iterative multiplier. IEEE Transactions on Computers, C-25 (5):546–548, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674646.

Grewal:1974:BFE

[GP74] K. S. Grewal and J. D. Patterson. A binary feature extraction technique. IEEE Transactions on Computers, C-23(5):545– 549, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672572.

Gelenbe:1971:UMR

[GR71a] S. E. Gelenbe and N. Rossi. Uniform modular realizations and linear machines. IEEE Transactions on Computers, C-20(12):1616–1617, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671781.

Gonzalez:1971:PSP

[GR71b] M. J. Gonzalez, Jr. and C. V. Ramamoorthy. Program suitability for parallel processing. IEEE Transactions on Computers, C-20 (6):647–654, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671909.

Givone:1972:MLI

[GR72a] D. D. Givone and R. P. Roesser. Multidimensional linear iterative circuits-general properties. IEEE Transactions on Computers, C-21(10):1067–1073, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672046.

Gonzalez:1972:PTE

[GR72b] M. J. Gonzalez, Jr. and C. V. Ramamoorthy. Parallel task execution in a decentralized system. IEEE Transactions on Computers, C-21(12):1310–1322, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672095.

Givone:1973:AR

[GR73a] D. D. Givone and R. P. Roesser. Authors’ reply. IEEE Transactions on Computers, C-22(12):1140, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672254.

Givone:1973:MML

[GR73b] Donald D. Givone and Robert P.

Roesser. Minimization of multidimensional linear iterative circuits. IEEE Transactions on Computers, C-22(7):673–678, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009134.

Grado:1970:REI

[Gra70] G. Grado. R70-5 the effect of IC’s on the design of analog computers. IEEE Transactions on Computers, C-19(5):464– 465, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671549.

Gray:1971:LPB

[Gra71] S. B. Gray. Local properties of binary images in two dimensions. IEEE Transactions on Computers, C-20(5):551– 561, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671882.

Granlund:1972:FPH

[Gra72] G. H. Granlund. Fourier preprocessing for hand print character recognition. IEEE Transactions on Computers, C-21(2):195– 201, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008926.

Gravitz:1973:NMT

[Gra73] M. E. Gravitz. A note on a modified ternary simulator capable of initializing all fault machine memory elements. IEEE Transactions on Computers, C-22(11):1042–1044, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672236.

Greibach:1975:RCS

[Gre75] S. A. Greibach. A remark on code sets and context-free languages. IEEE Transactions on Computers, C-24(7):741–742, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672887.

Griffith:1973:EDS

[Gri73] A. K. Griffith. Edge detection in simple scenes using a priori information. IEEE Transactions on Computers, C-22(4):371– 381, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672317.

Griswold:1974:BEP

[Gri74] R. E. Griswold. B74-44 the elements of programming style. IEEE Transactions on Computers, C-23 (12):1329, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672458.

Griffith:1976:EEP

[Gri76] Arnold K. Griffith. Empirical exploration of the performance of the alpha beta treesearching heuristic. IEEE Transactions on Computers, C-25(1):6– 11, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009198.

Griss:1978:UES

[Gri78a] Martin L. Griss. Using an efficient sparse minor expansion algorithm to compute polynomial subresultants and the greatest common denominator. IEEE Transactions on Computers, C-27(10): 945–950, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674974.

Gritsenko:1978:CEC

[Gri78b] V. M. Gritsenko. Comments on “Error-Correcting Codes for Byte-Organized Arithmetic Processors”. IEEE Transactions on Computers, C-27(2):189–190, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675061. See [NR75].

Gault:1972:MFD

[GRR72] J. W. Gault, J. P. Robinson, and S. M. Reddy. Multiple fault detection in combinational networks. IEEE Transactions on Computers, C-21(1):31– 36, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672021.

Gault:1974:AR

[GRR74] J. W. Gault, J. P. Robinson, and S. M. Reddy. Authors’ reply. IEEE Transactions on Computers, C-23 (3):334, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672531.

Goel:1974:CMF

[GS74] P. Goel and D. P. Siewiorek. Comments on “Multiple Fault Detection in Combinational Networks”. IEEE Transactions on Computers, C-23 (3):333–334, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672530. See [DW73b].

Gelenbe:1979:AUS

[GS79] E. Gelenbe and K. Sevcik. Analysis of update synchronization for multiple copy data bases. IEEE Transactions on Computers, C-28(10): 737–747, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675241.

Gilbert:1976:PDM

[GSB^{+}76] B. K. Gilbert, M. T. Storma, K. C. Ballard, L. W. Hobrock, C. E. James, and E. H. Wood. A programmable dynamic memory allocation system for input/ output of digital data into standard computer memories at 40 megasamples/s. IEEE Transactions on Computers, C-25(11):1101–1109, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674560.

Gilbert:1976:RTH

[GSJ^{+}76] B. K. Gilbert, M. T. Storma, C. E. James, L. W. Hobrock, E. S. Yang, K. C. Ballard, and E. H. Wood. A real-time hardware system for digital processing of wide-band video images. IEEE Transactions on Computers, C-25(11):1089–1100, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674559.

Gray:1979:DFM

[GST79] F. G. Gray, L. C. C. Shih, and R. A. Thompson. Diagnosis of faults in modular trees. IEEE Transactions on Computers, C-28 (5):342–353, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675364.

Gray:1974:RRC

[GT74] F. G. Gray and R. A. Thompson. Reconfiguration for repair in a class of universal logic modules. IEEE Transactions on Computers, C-23(11):1185–1194, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672421.

Gray:1978:FDB

[GT78a] F. G. Gray and R. A. Thompson. Fault detection in bilateral arrays of combinational cells. IEEE Transactions on Computers, C-27(12):1206–1213, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675029.

Greer:1978:CLD

[GT78b] C. R. Greer and R. A. Thompson. Combinational logic design with decoders. IEEE Transactions on Computers, C-27(9):869– 875, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675207.

Guffin:1971:MSC

[Guf71] R. M. Guffin. Microdiagnostics for the Standard Computer MLP-900 processor. IEEE Transactions on Computers, C-20(7):803– 808, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671943.

Groginsky:1970:PFF

[GW70] H. L. Groginsky and G. A. Works.

A pipeline Fast Fourier Transform. IEEE Transactions on Computers, C-19(11):1015–1019, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671419.

Greguss:1975:UHF

[GW75] P. Greguss and W. Waigelich.

Ultrasonic holographic Fourier spectroscopy via Optical Fourier Transforms. IEEE Transactions on Computers, C-24(4):412– 418, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672829.

Gaines:1977:SAA

[GW77] B. R. Gaines and I. H. Witten. Stability and admissibility of adaptive threshold logic convergence. IEEE Transactions on Computers, C-26 (1):88–91, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009280.

Garnett:1977:NEB

[GY77] James M. Garnett and Stephen S. Yau. Nonparametric estimation of the Bayes error of feature extractors using ordered nearest neighbor sets. IEEE Transactions on Computers, C-26(1):46– 54, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009273.

Hornbuckle:1970:LMA

[HA70] G. D. Hornbuckle and E. I. Ancona. The LX-1 microprocessor and its application to real-time signal processing. IEEE Transactions on Computers, C-19(8):710–720, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        www.acsel-lab.com/arithmetic/
                                        arith1/papers/ARITH1_Hornbuckle.
                                        pdf;   https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671614.

Hakimi:1974:CCA

[HA74] S. L. Hakimi and A. T. Amin. Characterization of connection assignment of diagnosable systems. IEEE Transactions on Computers, C-23 (1):86–88, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672375.

Huertas:1976:SSA

[HA76] J. L. Huertas and J. I. Acha. Selfsynchronization of asynchronous sequential circuits employing a general clock function. IEEE Transactions on Computers, C-25(3):297– 300, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009256.

Hou:1977:LSI

[HA77] Hsieh S. Hou and Harry C.

Andrews. Least squares image restoration using spline basis functions. IEEE Transactions on Computers, C-26(9):856–873, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674934.

Habibi:1977:FSW

[Hab77] A. Habibi. Fast suboptimal Wiener filtering of Markov sequences. IEEE Transactions on Computers, C-26 (5):443–449, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674861.

Hall:1974:AUD

[Hal74] E. L. Hall. Almost uniform distributions for computer image enhancement. IEEE Transactions on Computers, C-23(2):207–208, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672485.

Hamacher:1971:MCV

[Ham71] V. C. Hamacher. Machine complexity versus interconnection complexity in iterative arrays. IEEE Transactions on Computers, C-20 (3):321–323, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671831.

Hampel:1973:MTG

[Ham73] D. Hampel. MultiFunction threshold gates. IEEE Transactions on Computers, C-22(2):197–203, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672278.

Hamacher:1974:BCH

[Ham74] V. C. Hamacher. B 74-28 computer hardware and software: An interdisciplinary introduction. IEEE Transactions on Computers, C-23 (5):556, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672577.

Hansalik:1972:LBC

[Han72] W. E. Hansalik. A lower bound on the complexity of arbitrary switching function realizers. IEEE Transactions on Computers, C-21(5): 507–510, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672145.

Handoko:1975:DTA

[Han75] F. Handoko. A discussion on two algorithms for determining maximum compatibles. IEEE Transactions on Computers, C-24(8):838– 840, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672911.

Harrison:1970:BTF

[Har70a] M. A. Harrison. B70-1 truth functions and the problem of their realization by two-terminal graphs. IEEE Transactions on Computers, C-19(5):462, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671544.

Hartmanis:1970:RNC

[Har70b] J. Hartmanis. R70-1 A note on computing time for the recognition of context-free languages by a single-tape Turing machine. IEEE Transactions on Computers, C-19 (5):463, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671545.

Harr:1971:BAU

[Har71] J. A. Harr. B715 APL users’ guide. IEEE Transactions on Computers, C-20(10):1222–1223, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671707.

Harada:1972:SPN

[Har72] K. Harada. Sequential permutation networks. IEEE Transactions on Computers, C-21(5):472– 479, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672136.

Harrison:1973:NCB

[Har73] M. A. Harrison. On the number of classes of binary matrices. IEEE Transactions on Computers, C-22(12):1048–1052, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672242. See comments [Sag77].

Harrison:1974:BNC

[Har74a] T. J. Harrison. B74-23 numerical control — mathematics and applications. IEEE Transactions on Computers, C-23(3): 335, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672532.

Hart:1974:CPA

[Har74b] J. M. Hart. A class of push-down acceptors for adjunct languages. IEEE Transactions on Computers, C-23 (5):449–455, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672559.

Haralick:1976:SEW

[Har76] R. M. Haralick. A storage efficient way to implement the Discrete Cosine Transform. IEEE Transactions on Computers, C-25 (7):764–765, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674687.

Hassab:1973:SSC

[Has73] Joseph C. Hassab. A simply solvable class of simultaneous linear equations. IEEE Transactions on Computers, C-22(9):866– 867, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009183.

Hassab:1974:CPC

[Has74] J. C. Hassab. The characteristic polynomials for a class of eigenvalue problems. IEEE Transactions on Computers, C-23(9):967, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672654.

Hayes:1971:NMF

[Hay71a] J. P. Hayes. A NAND model for fault diagnosis in combinational logic networks. IEEE Transactions on Computers, C-20(12):1496–1506, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671755.

Hayes:1971:RBF

[Hay71b] J. P. Hayes. On realizations of Boolean functions requiring a minimal or near-minimal number of tests. IEEE Transactions on Computers, C-20(12):1506–1513, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671756.

Hayes:1974:MLN

[Hay74] J. P. Hayes. On modifying logic networks to improve their diagnosability. IEEE Transactions on Computers, C-23(1):56– 62, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672370.

Hayes:1975:DPS

[Hay75] J. P. Hayes. Detection of patternsensitive faults in random-access memories. IEEE Transactions on Computers, C-24(2):150–157, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672775.

Hayes:1976:GMF

[Hay76a] J. P. Hayes. A graph model for fault-tolerant computing systems. IEEE Transactions on Computers, C-25(9):875–884, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674712.

Hayes:1976:PIL

[Hay76b] J. P. Hayes. On the properties of irredundant logic networks. IEEE Transactions on Computers, C-25(9):884–892, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674713.

Hayes:1976:TCT

[Hay76c] J. P. Hayes. Transition count testing of combinational logic circuits. IEEE Transactions on Computers, C-25(6):613–620, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674661.

Hayes:1978:GOT

[Hay78a] J. P. Hayes. Generation of optimal transition count tests. IEEE Transactions on Computers, C-27 (1):36–41, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674950.

Hayes:1978:PCL

[Hay78b] J. P. Hayes. Path complexity of logic networks. IEEE Transactions on Computers, C-27(5):459– 462, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675127.

Hulina:1970:TIS

[HB70] P. T. Hulina and J. G. Bredeson. On Tracey’s internal state assignment method. IEEE Transactions on Computers, C-19(5): 458, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671539.

Hsiao:1975:OLS

[HB75a] M. Y. Hsiao and D. C. Bossen. Orthogonal latin square configuration for LSI memory yield and reliability enhancement. IEEE Transactions on Computers, C-24(5): 512–516, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672847.

Hunt:1975:SDC

[HB75b] B. R. Hunt and J. R. Breedlove.

Scan and display considerations in processing images by digital computer. IEEE Transactions on Computers, C-24(8):848– 853, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672915.

Heath:1970:SPC

[HC70] J. R. Heath and C. C. Carroll.

Special-purpose computer organization for double-precision realization of digital filters. IEEE Transactions on Computers, C-19(12):1146–1152, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671446.

Harlow:1971:IES

[HC71a] C. Harlow and C. L. Coates, Jr. Inessential errors in sequential machines. IEEE Transactions on Computers, C-20(6):688– 690, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671921.

Hsiao:1971:BDF

[HC71b] M. Y. Hsiao and D. K. Chia. Boolean difference for fault detection in asynchronous sequential machines. IEEE Transactions on Computers, C-20(11):1356–1361, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671731.

Harlow:1972:FSM

[HC72] Charles A. Harlow and Clarence L. Coates. Feedback in sequential machine realizations. IEEE Transactions on Computers, C-21(4):371– 381, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008978.

Ho:1973:MAR

[HC73] Irving T. Ho and Tien Chi Chen. Multiple addition by residue threshold functions and their representation by array logic. IEEE Transactions on Computers, C-22(8):762– 767, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009155.

Hamacher:1974:AR

[HC74] V. C. Hamacher and M. Cappa.

Authors’ reply. IEEE Transactions on Computers, C-23(3): 327, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672525.

Hemmati:1978:ACA

[HC78] F. Hemmati and D. J. Costello, Jr. An algebraic construction for q-ary shift register sequences. IEEE Transactions on Computers, C-27(12):1192–1195, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675025.

Hackbart:1971:AEF

[HD71] R. R. Hackbart and D. L. Dietmeyer. The avoidance and elimination of function hazards in asynchronous sequential circuits. IEEE Transactions on Computers, C-20(2):184–189, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671802.

Harlow:1973:ARI

[HE73] Charles A. Harlow and Sharon A. Eisenbeis. The analysis of radiographic images. IEEE Transactions on Computers, C-22(7): 678–689, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009135.

Hikita:1979:NMS

[HE79] T. Hikita and H. Enomoto. On the number of multivalued switching functions realizable by cascades. IEEE Transactions on Computers, C-28(5): 371–374, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675369.

Hecht:1973:FMF

[Hec73] H. Hecht. Figure of merit for faulttolerant space computers. IEEE Transactions on Computers, C-22 (3):246–251, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672295.

Hedetniemi:1970:RMA

[Hed70] S. Hedetniemi. R70-36 maximin automata. IEEE Transactions on Computers, C-19(9):859, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671657.

Heines:1979:BBC

[Hei79] T. S. Heines. Buffer behavior in computer communication systems. IEEE Transactions on Computers, C-28(8):573–576, August 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675413.

Hellerman:1972:MCW

[Hel72] L. Hellerman. A measure of computational work. IEEE Transactions on Computers, C-21(5):439– 446, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672132. See comments [Wel74] and discussion [Hel74].

Hellerman:1974:DMC

[Hel74] L. Hellerman. A discussion of “A Measure of Computational Work”. IEEE Transactions on Computers, C-23(2):208– 211, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672487. See [Hel72, Wel74, Mas78].

Henderson:1970:CCF

[Hen70] K. W. Henderson. Comment on “Computation of the Fast Walsh– Fourier Transform”. IEEE Transactions on Computers, C-19(9):850– 851, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671647. See [Lec70a].

Henderson:1971:PIR

[Hen71] K. W. Henderson. Previous investigation of realization of an arbitrary switching function with a network of threshold and parity elements. IEEE Transactions on Computers, C-20(8):941– 942, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671970.

Henschen:1976:ITR

[Hen76a] L. J. Henschen. Introduction tutorial on resolution. IEEE Transactions on Computers, C-25(8):769– 772, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674695.

Henschen:1976:SRH

[Hen76b] L. J. Henschen. Semantic resolution for Horn sets. IEEE Transactions on Computers, C-25(8):816– 822, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674700.

Henderson:1978:SSF

[Hen78] K. W. Henderson. Suggested standard forms for certain real Hadamard matrices. IEEE Transactions on Computers, C-27(5): 469–471, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675130.

Hall:1972:OMR

[HEWB72] D. J. Hall, R. M. Endlich, D. E.

Wolf, and A. E. Brain. Objective methods for registering landmarks and determining cloud motions from satellite data. IEEE Transactions on Computers, C-21 (7):768–776, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672173.

Heydorn:1971:RFE

[Hey71] R. P. Heydorn. Redundancy in feature extraction. IEEE Transactions on Computers, C-20(9):1051– 1054, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671994.

Henrichon:1970:AR

[HF70] E. G. Henrichon, Jr. and K. S.

Fu. Authors’ reply. IEEE Transactions on Computers, C-19(4):363– 364, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671521.

Hoffman:1971:PRS

[HF71] R. L. Hoffman and K. Fukunaga. Pattern recognition signal processing for mechanical diagnostics signature analysis. IEEE Transactions on Computers, C-20(9):1095– 1100, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672002.

Hallin:1972:PAF

[HF72] Thomas G. Hallin and Michael J. Flynn. Pipelining of arithmetic functions. IEEE Transactions on Computers, C-21(8):880– 886, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009044.

Hayes:1974:TPP

[HF74] J. P. Hayes and A. D. Friedman. Test point placement to simplify fault detection. IEEE Transactions on Computers, C-23(7):727– 735, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672614.

Halatsis:1978:INF

[HG78a] C. Halatsis and N. Gaitanis. Irredundant normal forms and minimal dependence sets of a Boolean function. IEEE Transactions on Computers, C-27(11):1064–1068, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674997.

Halatsis:1978:MCS

[HG78b] C. Halatsis and N. Gaitanis. On the minimization of the control store in microprogrammed computers. IEEE Transactions on Computers, C-27(12):1189–1192, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675024.

Halatsis:1979:PFS

[HG79] C. Halatsis and N. Gaitanis.

Positive fail-safe realization of synchronous sequential machines. IEEE Transactions on Computers, C-28(2):167–172, February 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675309.

Huskey:1976:CCD

[HH76] H. D. Huskey and V. R. Huskey.

Chronology of computing devices. IEEE Transactions on Computers, C-25(12):1190–1199, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674587.

Hill:1977:SSS

[HH77] F. J. Hill and B. Huey. SCIRTSS: a search system for sequential circuit test sequences. IEEE Transactions on Computers, C-26(5): 490–502, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674866.

Henle:1976:ATT

[HHJ^{+}76] R. A. Henle, I. T. Ho, W. S. Johnson, W. D. Pricer, and J. L. Walsh. The application of transistor technology to computers. IEEE Transactions on Computers, C-25(12):1289–1303, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674597.

Hu:1972:ATB

[HI72] Ming-Kuei Hu and A. Iosupovicz. Analysis of the terminal behavior of some classes of iterative arrays of linear machines. IEEE Transactions on Computers, C-21(12):1394–1399, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672105.

Hight:1971:MIS

[Hig71] S. L. Hight. Minimal input solutions. IEEE Transactions on Computers, C-20(8):923–925, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671965.

Hight:1973:CDD

[Hig73] S. L. Hight. Complex disjunctive decomposition of incompletely specified Boolean functions. IEEE Transactions on Computers, C-22(1):103–110, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672197.

Higbie:1976:VFP

[Hig76] L. C. Higbie. Vector floatingpoint data format. IEEE Transactions on Computers, C-25(1):25– 32, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009201.

Higbie:1978:OOM

[Hig78] L. C. Higbie. Overlapped operation with microprogramming. IEEE Transactions on Computers, C-27 (3):270–275, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675085.

Hill:1979:RBL

[Hil79] J. S. Hill. Revision of the buffer length derivation for a modified ek/D/1 systems by maritsas and Hartley. IEEE Transactions on Computers, C-28(3):269– 273, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675336.

Hutton:1974:SSA

[HK74] D. W. Hutton and W. H. Kautz. A simplified summation array for cellular logic modules. IEEE Transactions on Computers, C-23(2):203– 206, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672483.

Higuchi:1977:SHF

[HK77] T. Higuchi and M. Kameyama. Static-hazard-free T -gate for ternary memory element and its application to ternary counters. IEEE Transactions on Computers, C-26(12):1212–1221, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674782. See correction [HK78].

Higuchi:1978:CSH

[HK78] T. Higuchi and M. Kameyama.

Correction to “Static-Hazard-Free T -Gate for Ternary Memory Element and Its Application to Ternary Counters”. IEEE Transactions on Computers, C-27(12): 1227, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675035. See [HK77].

Hall:1971:SPF

[HKD^{+}71] E. L. Hall, R. P. Kruger, S. J.

Dwyer, D. L. Hall, R. W. Mclaren, and G. S. Lodwick. A survey of preprocessing and feature extraction techniques for radiographic images. IEEE Transactions on Computers, C-20(9):1032– 1044, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671992.

Hongyuan:1986:CSL

[HL86] Wang Hongyuan and S. C.

Lee. Comments on “Sign/ Logarithm Arithmetic for FFT Implementation”. IEEE Transactions on Computers, C-35(5):482– 484, May 1986. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1676792. See [SCNS83].

Hlavicka:1970:EHC

[Hla70] J. Hlavicka. Essential hazard correction without the use of delay elements. IEEE Transactions on Computers, C-19(3):232– 238, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671495.

Hlawiczka:1978:CPE

[Hla78] A. Hlawiczka. Comments on “Procedures for Eliminating Static and Dynamic Hazards in Test Generation”. IEEE Transactions on Computers, C-27(2):191, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675064. See [BH74b].

Hall:1970:GPQ

[HLD70] E. L. Hall, D. D. Lynch, and S. J. Dwyer. Generation of products and quotients using approximate binary logarithms for digital filtering applications. IEEE Transactions on Computers, C-19(2):97– 105, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671467.

Hashimoto:1971:PG

[HN71] A. Hashimoto and K. Noshita. A property of N-graphs. IEEE Transactions on Computers, C-20(1):95– 97, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671678.

Hao:1973:STM

[HN73] H. T. Hao and M. M. Newborn. A study of trigger machines. IEEE Transactions on Computers, C-22(12):1123–1131, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672251.

Harao:1978:SDP

[HN78] M. Harao and S. Noguchi. On some dynamical properties of finite cellular automaton. IEEE Transactions on Computers, C-27(1):42– 52, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674951.

Hagler:1975:POD

[HND75] M. O. Hagler, T. G. Newman, and H. Dirilten. A programmable optical-digital scanner-processor for automated two-dimensional data analysis. IEEE Transactions on Computers, C-24(10):1036–1038, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672716.

Hong:1972:CBF

[HO72] S. J. Hong and D. L. Ostapko. On complementation of Boolean functions. IEEE Transactions on Computers, C-21(9):1022, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009083.

Hakkala:1974:DEA

[HO74] L. Hakkala and L. Ojala. Discretization error analysis in linear DDA connections. IEEE Transactions on Computers, C-23(9):932– 941, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672650.

Hartimo:1979:CST

[HO79] I. Hartimo and L. Ojala. Comments on “A Stable, Two-Multiplies-per-Cycle Algorithm for Digital Generation of Sinusoids in Real Time”. IEEE Transactions on Computers, C-28(4):300, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675349. See [Sil77].

Hoagland:1976:MRS

[Hoa76] A. S. Hoagland. Magnetic recording storage. IEEE Transactions on Computers, C-25(12):1283–1288, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674596.

Hobbs:1973:BCT

[Hob73] L. C. Hobbs. B73-6 computing terminals: International computer state of the art report. IEEE Transactions on Computers, C-22 (4):429–430, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672331.

Hoevel:1974:IDE

[Hoe74] L. W. Hoevel. “Ideal” directly executed languages: An analytical argument for emulation. IEEE Transactions on Computers, C-23 (8):759–767, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672625.

Hoel:1976:SVL

[Hoe76] Jeffrey H. Hoel. Some variations of Lee’s algorithm. IEEE Transactions on Computers, C-25(1):19– 24, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009200.

Holborow:1972:IBL

[Hol72a] C. E. Holborow. An improved bound on the length of checking experiments for sequential machines with counter cycles. IEEE Transactions on Computers, C-21(6): 597–598, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009016.

Holst:1972:STC

[Hol72b] P. A. Holst. Symbolic treatment of certain selector functions. IEEE Transactions on Computers, C-21 (5):486–488, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672138.

Holden:1976:TAI

[Hol76] A. D. C. Holden. Trends in artificial intelligence. IEEE Transactions on Computers, C-25(4):313– 316, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674610.

Hollaar:1979:DLM

[Hol79] L. A. Hollaar. A design for a list merging network. IEEE Transactions on Computers, C-28(6):406– 413, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675380.

Hoogendoorn:1977:GMM

[Hoo77] C. H. Hoogendoorn. A general model for memory interference in multiprocessors. IEEE Transactions on Computers, C-26(10):998– 1005, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674734.

Hopcroft:1970:RNS

[Hop70] J. E. Hopcroft. R70-2 nested stack automata. IEEE Transactions on Computers, C-19(5): 463, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671546.

Hopkins:1971:FTI

[Hop71] A. L. Hopkins, Jr. A fault-tolerant information processing concept for space vehicles. IEEE Transactions on Computers, C-20(11):1394–1403, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671738.

Horelick:1971:CSB

[Hor71a] D. Horelick. Comment on “Serial Binary-to-Decimal and Decimal-to-Binary Conversion”. IEEE Transactions on Computers, C-20(4): 476, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671869. See [Rhy70].

Horna:1971:PRT

[Hor71b] O. A. Horna. Pulse reflection in transmission lines. IEEE Transactions on Computers, C-20(12): 1558–1563, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671762.

Horna:1972:NTT

[Hor72] Otakar A. Horna. Nonlinear termination of transmission lines. IEEE Transactions on Computers, C-21(9):1011–1015, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009080.

Hou:1972:AUL

[Hou72] H. S. Hou. Application of uniform loading theory to circuit packaging and memory arrays in highspeed computers. IEEE Transactions on Computers, C-21(5): 454–463, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672134.

Howden:1975:MGP

[How75] W. E. Howden. Methodology for the generation of program test data. IEEE Transactions on Computers, C-24(5):554– 560, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672852.

Hong:1972:GCM

[HP72] Se June Hong and A. M. Patel. A general class of maximal codes for computer applications. IEEE Transactions on Computers, C-21(12):1322–1331, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672096.

Hight:1973:DMV

[HP73] S. L. Hight and D. P. Petersen. Dissent in a majority voting system. IEEE Transactions on Computers, C-22(2):168– 171, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672272.

Hoehne:1975:DVR

[HP75] H. Hoehne and R. Piloty. Design verification at the register transfer language level. IEEE Transactions on Computers, C-24(9):861– 867, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672924.

Hsiao:1977:SAG

[HPP77] M. Y. Hsiao, A. M. Patel, and D. K. Pradhan. Store address generator with on-line fault-detection capability. IEEE Transactions on Computers, C-26(11):1144–1147, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674762.

Hammer:1979:ADR

[HPP79] P. L. Hammer, U. N. Peled, and M. A. Pollatschek. An algorithm to dualize a regular switching function. IEEE Transactions on Computers, C-28(3):238– 243, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675324.

House:1970:NRR

[HS70] R. W. House and D. W. Stevens. A new rule for reducing CC tables. IEEE Transactions on Computers, C-19(11):1108–1111, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671435.

Hopkins:1975:AES

[HS75] A. L. Hopkins, Jr. and T. B. Smith III. The architectural elements of a symmetric fault-tolerant multiprocessor. IEEE Transactions on Computers, C-24(5):498– 505, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672845.

Hsiao:1971:GRO

[Hsi71a] D. K. Hsiao. A generalized record organization. IEEE Transactions on Computers, C-20(12):1490–1495, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671754.

Hsieh:1971:CES

[Hsi71b] E. P. Hsieh. Checking experiments for sequential machines. IEEE Transactions on Computers, C-20(10):1152–1166, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671693. See comments [WY75].

Halatsis:1978:PDS

[HSP78] C. Halatsis, M. Sigala, and G. Philokyprou. Polylinear decomposition of synchronous sequential machines. IEEE Transactions on Computers, C-27(12):1144–1152, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675017.

Hartke:1978:DRD

[HSS78] D. H. Hartke, W. M. Sterling, and J. E. Shemer. Design of a raster display processor for office applications. IEEE Transactions on Computers, C-27(4):337– 348, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675106.

Hussain:1972:ROU

[HTD72] A. B. Shahidul Hussain, Godfried T. Toussaint, and Robert W. Donaldson. Results obtained using a simple character recognition procedure on Munson’s handprinted data. IEEE Transactions on Computers, C-21(2):201– 205, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008927.

Hu:1972:CSS

[Hu72] S. C. Hu. Cellular synthesis of synchronous sequential machines. IEEE Transactions on Computers, C-21(12):1399–1405, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672106.

Hu:1973:BCI

[Hu73] Sung C. Hu. Boundary conditions for iteratively realized sequential circuits. IEEE Transactions on Computers, C-22(9):867– 868, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009184.

Hu:1975:PAD

[Hu75] S. C. Hu. A probabilistic approach of designing more reliable logic gates with asymmetric input faults. IEEE Transactions on Computers, C-24(10): 1012–1014, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672706.

Huang:1971:UCA

[Hua71] Jung-Chang Huang. A universal cellular array. IEEE Transactions on Computers, C-20(3):317– 320, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671830.

Huber:1976:GES

[Hub76] Franz Huber. On the generation of English sentences. IEEE Transactions on Computers, C-25(1):90– 91, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009209.

Huffman:1976:CCP

[Huf76] D. A. Huffman. Curvature and creases: a primer on paper. IEEE Transactions on Computers, C-25(10):1010–1019, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674542.

Hunt:1971:SEU

[Hun71] B. R. Hunt. Spectral effects in the use of Newton–Cotes approximations for computing Discrete Fourier Transforms. IEEE Transactions on Computers, C-20(8):942– 943, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671971.

Hunt:1972:MCT

[Hun72] B. R. Hunt. Minimizing the computation time for using the technique of sectioning for digital filtering of pictures. IEEE Transactions on Computers, C-21(11):1219–1222, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672073.

Hunt:1973:ACL

[Hun73] B. R. Hunt. The application of constrained least squares estimation to image restoration by digital computer. IEEE Transactions on Computers, C-22(9):805– 812, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009169.

Hunt:1977:BMN

[Hun77] B. R. Hunt. Bayesian methods in nonlinear digital image restoration. IEEE Transactions on Computers, C-26(3):219– 229, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674810.

Hurd:1974:EGS

[Hur74] W. J. Hurd. Efficient generation of statistically good pseudonoise by linearly interconnected shift registers. IEEE Transactions on Computers, C-23(2):146–152, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672470.

Hussain:1972:CSS

[Hus72] A. B. Shahidul Hussain. On the correctness of some sequential classification schemes in pattern recognition. IEEE Transactions on Computers, C-21(3):318– 320, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008959.

Huskey:1973:PA

[Hus73] H. D. Huskey. The Polish assembler. IEEE Transactions on Computers, C-22(11):1035–1040, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672234.

Hussain:1974:CSP

[Hus74] A. B. S. Hussain. Compound sequential probability ratio test for the classification of statistically dependent patterns. IEEE Transactions on Computers, C-23(4):398– 410, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672548.

Habibi:1970:FM

[HW70] A. Habibi and P. A. Wintz. Fast multipliers. IEEE Transactions on Computers, C-19(2):153–157, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671474.

Hulme:1975:PIA

[HW75] B. L. Hulme and R. B. Worrell. A prime implicant algorithm with factoring. IEEE Transactions on Computers, C-24(11):1129–1131, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672739.

Hannington:1976:FPM

[HW76] G. Hannington and D. G. Whitehead. A floating-point multiplexed DDA system. IEEE Transactions on Computers, C-25(11):1074–1077, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674557. See comments [MMB78].

Hannington:1978:AR

[HW78] G. Hannington and D. G. Whitehead. Authors reply. IEEE Transactions on Computers, C-27 (12):1226–1227, December 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675034.

Hopper:1979:BRN

[HW79] A. Hopper and J. Wheeler. Binary routing networks. IEEE Transactions on Computers, C-28(10):699– 703, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675237.

Hwa:1972:CEM

[Hwa72] H. R. Hwa. Contradiction equations in a B matrix of vertex weight method and their correspondence with the k-summability property of vertices. IEEE Transactions on Computers, C-21(6):606– 610, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009019.

Hwang:1973:PRS

[Hwa73] Kai Hwang. Periodic realization of synchronous sequential machines. IEEE Transactions on Computers, C-22(10):923–927, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672212.

Hwa:1974:MGP

[Hwa74] H. R. Hwa. A method for generating prime implicants of a Boolean expression. IEEE Transactions on Computers, C-23(6): 637–641, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672596.

Hwang:1979:GMT

[Hwa79] Kai Hwang. Global and modular two’s complement cellular array multipliers. IEEE Transactions on Computers, C-28(4):300– 306, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1675350.

Irao:1976:RIT

[IA76] V. V. Bapeswara Irao and V. K.

Aatre. Reliability index of teleprocessing tree networks. IEEE Transactions on Computers, C-25(1):97– 98, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009213.

Ibaraki:1971:GIM

[Iba71] T. Ibaraki. Gate-interconnection minimization of switching networks using negative gates. IEEE Transactions on Computers, C-20(6): 698–706, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671924.

Ibaraki:1974:CMF

[Iba74] T. Ibaraki. Comments on “Monotone Functions in Sequential Circuits”. IEEE Transactions on Computers, C-23(9):989– 990, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672665. See [Mag73].

Ito:1978:CBC

[IC78] M. R. Ito and R. D. Cameron. Combined binary code translation and parallel-to-serial conversion using stored logic arrays. IEEE Transactions on Computers, C-27(9):833– 841, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675203.

Iosupovicz:1972:CAO

[IH72] A. Iosupovicz and Ming-Kuei Hu. A class of autonomous onedimensional iterative arrays of linear machines. IEEE Transactions on Computers, C-21(10): 1073–1086, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672047.

Ibaraki:1971:SNM

[IM71] T. Ibaraki and S. Muroga. Synthesis of networks with a minimum number of negative gates. IEEE Transactions on Computers, C-20 (1):49–58, January 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671674.

Iosupovicz:1975:AUF

[Ios75] A. Iosupovicz. Adaptive universal fault-tolerant logic networks. IEEE Transactions on Computers, C-24(10):1043–1048, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672720.

Iosupovicz:1978:ODB

[Ios78] A. Iosupovicz. Optimal detection of bridge faults and stuck-at faults in two-level logic. IEEE Transactions on Computers, C-27(5): 452–455, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675125.

Ikram:1976:STI

[IR76] M. Ikram and D. A. Roy. A simple technique to improve the pialgorithm for prime implicant determination. IEEE Transactions on Computers, C-25(11):1184–1187, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674579.

Ibarra:1975:PCF

[IS75] O. H. Ibarra and S. K. Sahni.

Polynomially complete fault detection problems. IEEE Transactions on Computers, C-24(3):242– 249, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672798.

Ingle:1976:RMV

[IS76] Ashok D. Ingle and Daniel P.

Siewiorek. A reliability model for various switch designs in hybrid redundancy. IEEE Transactions on Computers, C-25(2):115– 133, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009226.

Ishizuka:1977:MMN

[Ish77a] O. Ishizuka. On multivalued multithreshold networks composed of conventional threshold elements. IEEE Transactions on Computers, C-26(12):1251–1257, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674786.

Ishizuka:1977:SMT

[Ish77b] O. Ishizuka. Synthesis of multithreshold tree networks. IEEE Transactions on Computers, C-26 (12):1294–1297, December 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674796.

Irving:1976:FFM

[ISN76] Thurman A. Irving, Sajjan G. Shiva, and H. Troy Nagle. Flip-flops for multiple-valued logic. IEEE Transactions on Computers, C-25 (3):237–246, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
5009250.

Jain:1974:IRM

[JA74] A. K. Jain and E. Angel. Image restoration, modelling, and reduction of dimensionality. IEEE Transactions on Computers, C-23 (5):470–476, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672562.

Jump:1978:EPD

[JA78] J. R. Jump and S. R. Ahuja.

Effective pipelining of digital systems. IEEE Transactions on Computers, C-27(9):855–865, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675205.

Jacobsohn:1973:CDA

[Jac73] David H. Jacobsohn. A combinatoric division algorithm for fixedinteger divisors. IEEE Transactions on Computers, C-22(6):608– 610, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Jacobsohn.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009114.

Jain:1977:FKL

[Jai77a] A. K. Jain. A fast Karhunen– Loeve transform for digital restoration of images degraded by white and colored noise. IEEE Transactions on Computers, C-26(6): 560–571, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674881.

Jain:1977:OFM

[Jai77b] A. K. Jain. An operator factorization method for restoration of blurred images. IEEE Transactions on Computers, C-26(11):1061–1071, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674752.

Jain:1977:SMR

[Jai77c] A. K. Jain. A semicausal model for recursive filtering of twodimensional images. IEEE Transactions on Computers, C-26(4):343– 350, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674844.

Jayashree:1976:BMU

[JB76a] T. Jayashree and D. Basu. On binary multiplication using the quarter square algorithm. IEEE Transactions on Computers, C-25(9): 957–960, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1674723.

Jayasri:1976:AOM

[JB76b] T. Jayasri and D. Basu. An approach to organizing microinstructions which minimizes the width of control store words. IEEE Transactions on Computers, C-25(5): 514–521, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674639.

Jump:1972:MA

[JF72] J. Robert Jump and Dennis R.

Fritsche. Microprogrammed arrays. IEEE Transactions on Computers, C-21(9):974–984, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009074.

Jakubowski:1977:SDR

[JK77] R. Jakubowski and A. Kasprzak.

A syntactic description and recognition of rotary machine elements. IEEE Transactions on Computers, C-26(10):1039–1043, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674743.

Jordan:1973:IAG

[JLH73] B. W. Jordan, Jr., W. J. Lennon, and B. D. Holm. An improved algorithm for the generation of nonparametric curves. IEEE Transactions on Computers, C-22(12):1052–1060, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672243. See comment [Bel76].

Janisz:1970:IHS

[JM70] T. Janisz and R. C. Martin. Interconnection of high-speed logic circuits. IEEE Transactions on Computers, C-19(9):831–837, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671642.

Jones:1971:OST

[JM71] S. K. Jones and J. L. Melsa. Optimization and simulation of two classes of nonresetting data reconstructors. IEEE Transactions on Computers, C-20(1):12– 18, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671668.

Juliussen:1973:MMC

[JM73] J. E. Juliussen and F. J. Mowle. Multiple microprocessors with common main and control memories. IEEE Transactions on Computers, C-22(11):999–1007, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672229.

Jones:1974:TMS

[JM74] L. H. Jones and R. E. Merwin.

Trends in microprogramming: a second reading. IEEE Transactions on Computers, C-23(8):754– 759, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672624.

Johnson:1971:MIS

[Joh71] A. M. Johnson. The microdiagnostics for the IBM System/ 360 Model 30. IEEE Transactions on Computers, C-20(7):798– 803, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671942.

Joliat:1976:STP

[Jol76] M. L. Joliat. A simple technique for partial elimination of unit productions from LR(k) parsers. IEEE Transactions on Computers, C-25 (7):763–764, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674686.

Jorgensen:1974:BBP

[Jor74a] P. Jorgensen. B74-7 basic programming. IEEE Transactions on Computers, C-23(2):212, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672489.

Jorgensen:1974:BPP

[Jor74b] P. C. Jorgensen. B74-6 PL/I programming for engineering and science. IEEE Transactions on Computers, C-23(2):211–212, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672488.

Joseph:1972:BCT

[Jos72a] E. C. Joseph. B72-18 computers and their uses, 2nd ed. IEEE Transactions on Computers, C-21(10): 1135, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672058.

Joseph:1972:BEU

[Jos72b] E. C. Joseph. B72-4 expanding use of computers in the 70’s technology — markets, needs, technology. IEEE Transactions on Computers, C-21(1):110, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672038.

Joseph:1974:BDB

[Jos74] E. C. Joseph. B74-14 data base systems. IEEE Transactions on Computers, C-23(2):217– 218, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672496.

Joseph:1978:EDS

[Jos78] J. Joseph. On easily diagnosable sequential machines. IEEE Transactions on Computers, C-27(2):159– 162, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675049.

Jarvis:1973:CUS

[JP73] R. A. Jarvis and E. A. Patrick.

Clustering using a similarity measure based on shared near neighbors. IEEE Transactions on Computers, C-22(11):1025–1034, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672233.

Jagoe:1976:MPC

[JP76] J. Roger Jagoe and Keith A. Paton. Measurement of pneumoconiosis by computer. IEEE Transactions on Computers, C-25(1):95– 97, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009212.

Jarvis:1976:NTD

[JR76] J. F. Jarvis and C. S. Roberts. A new technique for displaying continuous tone images on a bilevel display. IEEE Transactions on Communications, 24(8):891–898, August 1976. CODEN IECMBT. ISSN 0090-6778.

Jullien:1978:RNS

[Jul78] G. A. Jullien. Residue number scaling and other operations using ROM arrays. IEEE Transactions on Computers, C-27(4):325– 336, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675105;   https://ieeexplore.
                                        ieee.org/xpl/RecentIssue.jsp?
                                        punumber=12;   https://ieeexplore.
                                        ieee.org/xpl/tocresult.jsp?isnumber=
                                        35168.

Jump:1970:RUS

[Jum70] J. R. Jump. R70-29 uniform synthesis of sequential circuits. IEEE Transactions on Computers, C-19 (8):765, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671627.

Jump:1974:ACA

[Jum74] J. R. Jump. Asynchronous control arrays. IEEE Transactions on Computers, C-23(10):1020–1029, October 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672396.

Jordan:1973:DSS

[JW73] H. F. Jordan and D. C. M.

Wood. On the distribution of sums of successive bits of shiftregister sequences. IEEE Transactions on Computers, C-22(4):400– 408, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672321.

Joshi:1976:SFM

[JW76] A. K. Joshi and R. M. Weischedel.

Some frills for modal tic-tac-toe: Semantics of predicate complement constructions. IEEE Transactions on Computers, C-25(4):374– 389, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674620.

Karlgaard:1977:AR

[KAA77] D. C. Karlgaard and A. M. Abd-Alla. Author’s reply. IEEE Transactions on Computers, C-26(9): 932, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674942.

Kain:1970:NSC

[Kai70] R. Y. Kain. Nonlinear sequential circuits. IEEE Transactions on Computers, C-19(3):249– 254, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671497. See comments [Du71].

Kameda:1970:RMT

[Kam70a] T. Kameda. R70-23 multitape and multi-head pushdown automata. IEEE Transactions on Computers, C-19(7):659–660, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671603.

Kameda:1970:RRH

[Kam70b] T. Kameda. R70-39 on the relational homomorphisms of automata. IEEE Transactions on Computers, C-19(10):987, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671405.

Kamal:1975:ADI

[Kam75] S. Kamal. An approach to the diagnosis of intermittent faults. IEEE Transactions on Computers, C-24 (5):461–467, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672840.

Kambayashi:1979:LDP

[Kam79] Y. Kambayashi. Logic design of programmable logic arrays. IEEE Transactions on Computers, C-28(9):609–617, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675428.

Kandel:1973:CMF

[Kan73a] A. Kandel. Comments on “Minimization of Fuzzy Functions”. IEEE Transactions on Computers, C-22 (2):217, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672281. See [SC72a].

Kandel:1973:NHE

[Kan73b] A. Kandel. Note on hazard elimination. IEEE Transactions on Computers, C-22(10):955–956, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672216.

Kandel:1973:MFF

[Kan73c] Abraham Kandel. On minimization of fuzzy functions. IEEE Transactions on Computers, C-22(9):826– 832, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009172.

Kang:1975:EUM

[Kan75] A. N. C. Kang. On the efficiency of universal machines. IEEE Transactions on Computers, C-24(10): 1010–1012, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672705.

Kandel:1976:DFF

[Kan76] A. Kandel. On the decomposition of fuzzy functions. IEEE Transactions on Computers, C-25(11):1124–1130, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674563.

Kobylarz:1979:ECF

[KAN79] T. Kobylarz and A. Al-Najjar. An examination of the cost function for programmable logic arrays. IEEE Transactions on Computers, C-28 (8):586–590, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675418.

Karplus:1970:RUF

[Kar70] W. J. Karplus. R70-6 use of functional approximation methods in the computer solution of initial value partial differential equation problems. IEEE Transactions on Computers, C-19(5): 465, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671550.

Kartashev:1974:TIM

[Kar74] S. P. Kartashev. Theory and implementation of p-multiple sequential machines. IEEE Transactions on Computers, C-23(5): 500–523, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672566.

Kartashev:1976:MSR

[Kar76] S. I. Kartashev. A microcomputer with a shift-register memory. IEEE Transactions on Computers, C-25 (5):470–484, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674635.

Karpovsky:1977:EDD

[Kar77a] M. G. Karpovsky. Error detection in digital devices and computer programs with the aid of linear recurrent equations over finite commutative groups. IEEE Transactions on Computers, C-26(3):208– 218, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674809.

Karpovsky:1977:FFT

[Kar77b] M. G. Karpovsky. Fast Fourier Transforms on finite non-Abelian groups. IEEE Transactions on Computers, C-26(10):1028–1030, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674739.

Kautz:1970:BSC

[Kau70a] W. H. Kautz. Bypass switching for cellular cascades. IEEE Transactions on Computers, C-19(9):837– 839, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671643.

Kautz:1970:NCC

[Kau70b] W. H. Kautz. The necessity of closed circuit loops in minimal combinational circuits. IEEE Transactions on Computers, C-19(2):162– 164, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671477.

Kautz:1970:RIS

[Kau70c] W. H. Kautz. A readily implemented single-error-correcting unitdistance counting code. IEEE Transactions on Computers, C-19 (10):972–975, October 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=

1671397. See comments [Dev79].

Kaufman:1973:EFR

[Kau73] A. B. Kaufman. An expandable ferroelectric random access memory. IEEE Transactions on Computers, C-22(2):154–158, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672270.

Kaufman:1974:AOA

[Kau74a] M. T. Kaufman. An almostoptimal algorithm for the assembly line scheduling problem. IEEE Transactions on Computers, C-23(11):1169–1174, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672418.

Kautz:1974:TFW

[Kau74b] W. H. Kautz. Testing for faults in wiring networks. IEEE Transactions on Computers, C-23(4):358– 363, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672543.

Kohavi:1975:FDC

[KB75] Z. Kohavi and I. Berger. Fault diagnosis in combinational tree networks. IEEE Transactions on Computers, C-24(12):1161–1167, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672752.

Kennedy:1971:HIU

[KC71] J. D. Kennedy and S. J. Clark. Halftone images using computer graphics. IEEE Transactions on Computers, C-20(5):584–585, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671890.

Kukreja:1973:CSC

[KC73] S. N. Kukreja and I-Ngo Chen. Combinational and sequential cellular structures. IEEE Transactions on Computers, C-22(9):813– 823, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009170.

Kelly:1971:FEV

[KCLB71] K. K. Kelly, T. W. Calvert, R. L. Longini, and J. P. Brown. Feature enhancement of vectorcardiograms by linear normalization. IEEE Transactions on Computers, C-20(9):1109–1111, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672006.

Khabra:1973:MPS

[KD73] N. S. Khabra and S. R. Das.

Multiform partial symmetry and linearity. IEEE Transactions on Computers, C-22(8):804, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009162.

Kam:1979:SDS

[KD79] John B. Kam and George I.

Davida. Structured design of substitution-permutation encryption networks. IEEE Transactions on Computers, C-28(10):747– 753, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675242.

Kella:1970:SMI

[Kel70] J. Kella. State minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-19(4):342– 348, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671513.

Kella:1971:SMI

[Kel71] J. Kella. Sequential machine identification. IEEE Transactions on Computers, C-20(3):332– 338, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671834. See comments [WZ72].

Keller:1974:TTU

[Kel74] R. M. Keller. Towards a theory of universal speed-independent modules. IEEE Transactions on Computers, C-23(1):21–33, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672366.

Kendall:1974:NAC

[Ken74] W. B. Kendall. A new algorithm for computing correlations. IEEE Transactions on Computers, C-23 (1):88–90, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672376.

Kent:1977:HSF

[Ken77a] J. G. Kent. Highlights of a study of floating-point instructions. IEEE Transactions on Computers, C-26 (7):660–666, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674899.

Kent:1977:HST

[Ken77b] S. A. Kent. A high-speed threshold gate multiplier. IEEE Transactions on Computers, C-26(12):1279–1283, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674790.

Kernighan:1973:BSP

[Ker73] B. W. Kernighan. B73-8 systems programming. IEEE Transactions on Computers, C-22(4):430– 431, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672333.

Kamal:1972:ACH

[KES72] A. A. Kamal and A. K. El-Sherif. Analysis and compensation of high-speed electronic analogcomputer errors. IEEE Transactions on Computers, C-21(8):886– 891, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009045.

Keyes:1977:PUI

[Key77] R. W. Keyes. Physical uncertainty and information. IEEE Transactions on Computers, C-26(10): 1017–1025, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674737.

Koontz:1972:NFE

[KF72a] W. L. G. Koontz and K. Fukunaga. A nonlinear feature extraction algorithm using distance transformation. IEEE Transactions on Computers, C-21(1):56– 63, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672024.

Koontz:1972:AAN

[KF72b] Warren L. G. Koontz and Keinosuke Fukunaga. Asymptotic analysis of a nonparametric clustering technique. IEEE Transactions on Computers, C-21(9):967– 974, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009073.

Koontz:1972:NVS

[KF72c] Warren L. G. Koontz and Keinosuke Fukunaga. A nonparametric valley-seeking technique for cluster analysis. IEEE Transactions on Computers, C-21(2):171–178, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008922.

Kreitzer:1973:VDS

[KF73] N. H. Kreitzer and W. J. Fitzgerald. A video display system for image processing by computer. IEEE Transactions on Computers, C-22(2):128–134, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672266.

Karunanithi:1978:SNT

[KF78] S. Karunanithi and A. D. Friedman. Some new types of logical completeness. IEEE Transactions on Computers, C-27(11):998– 1005, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674989.

Karunanithi:1979:ADS

[KF79] S. Karunanithi and A. D. Friedman. Analysis of digital systems using a new measure of system diagnosis. IEEE Transactions on Computers, C-28(2):121– 133, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675301.

Kamal:1972:HSM

[KG72] A. A. Kamal and M. A. N.

Ghannam. High-speed multiplication systems. IEEE Transactions on Computers, C-21(9):1017– 1021, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009082.

Kahveci:1974:SDD

[KH74] A. E. Kahveci and E. L. Hall.

Sequency domain design of frequency filters. IEEE Transactions on Computers, C-23(9):976– 981, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672658.

Kameyama:1977:SMV

[KH77a] M. Kameyama and T. Higuchi. Synthesis of multiple-valued logic networks based on tree-type universal logic module. IEEE Transactions on Computers, C-26(12):1297–1302, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674797.

Knaizuk:1977:ATR

[KH77b] J. Knaizuk, Jr. and C. R. P. Hartmann. An algorithm for testing random access memories. IEEE Transactions on Computers, C-26 (4):414–416, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674851.

Knaizuk:1977:OAT

[KH77c] J. Knaizuk, Jr. and C. R. P. Hartmann. An optimal algorithm for testing stuck-at faults in random access memories. IEEE Transactions on Computers, C-26(11):1141–1144, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=

1674761. See comments [Nai79].

Kime:1970:AMD

[Kim70] C. R. Kime. An analysis model for digital system diagnosis. IEEE Transactions on Computers, C-19(11):1063–1073, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671426.

Kimbleton:1974:BEM

[Kim74] S. R. Kimbleton. B74-43 evaluation and measurement techniques for digital computer systems. IEEE Transactions on Computers, C-23(12):1328–1329, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672457.

Kime:1975:FTC

[Kim75] C. R. Kime. Fault-tolerant computing: An introduction and a perspective. IEEE Transactions on Computers, C-24(5):457– 460, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672839.

Kime:1979:AMD

[Kim79] C. R. Kime. An abstract model for digital system fault diagnosis. IEEE Transactions on Computers, C-28(10):754–767, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675243.

Kinney:1970:DAS

[Kin70] L. L. Kinney. Decomposition of asynchronous sequential switching circuits. IEEE Transactions on Computers, C-19(6):515– 529, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671562.

King:1971:PPC

[Kin71a] J. C. King. Proving programs to be correct. IEEE Transactions on Computers, C-20(11):1331–1336, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671728.

King:1971:DAM

[Kin71b] W. K. King. Design of an associative memory. IEEE Transactions on Computers, C-20(6):671– 674, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671915.

Kinney:1971:CSA

[Kin71c] L. L. Kinney. A characterization of some asynchronous sequential networks and state assignments. IEEE Transactions on Computers, C-20 (4):426–436, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671852.

Kinney:1972:BIS

[Kin72] L. L. Kinney. B72-11 an introduction to switching system design. IEEE Transactions on Computers, C-21(5):518, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672154.

Kittler:1977:DVM

[Kit77] J. Kittler. On the discriminant vector method of feature selection. IEEE Transactions on Computers, C-26(6):604– 606, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674885.

Kittler:1978:CIP

[Kit78] J. Kittler. Classification of incomplete pattern vectors using modified discriminant functions. IEEE Transactions on Computers, C-27 (4):367–375, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675109.

Kulkarni:1979:OCA

[KJ79] K. G. Kulkarni and V. Jayakumar. Ordering of connections for automated routing. IEEE Transactions on Computers, C-28(10): 791–794, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675248.

Kjelkerud:1972:CPS

[Kje72] Eskil Kjelkerud. A computer program for the synthesis of switching circuits by decomposition. IEEE Transactions on Computers, C-21 (6):568–573, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009009.

Korn:1970:PHC

[KK70] G. A. Korn and H. Kosako. A proposed hybrid-computer method for functional optimization. IEEE Transactions on Computers, C-19 (2):149–153, February 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671473.

Kohavi:1972:DMF

[KK72] Igal Kohavi and Zvi Kohavi. Detection of multiple faults in combinational logic networks. IEEE Transactions on Computers, C-21 (6):556–568, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009008.

Koren:1977:DIF

[KK77a] I. Koren and Z. Kohavi. Diagnosis of intermittent faults in combinational networks. IEEE Transactions on Computers, C-26(11):1154–1158, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674766.

Koren:1977:SFD

[KK77b] I. Koren and Z. Kohavi. Sequential fault diagnosis in combinational networks. IEEE Transactions on Computers, C-26(4):334– 342, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674843.

Kartashev:1978:MNS

[KK78] S. P. Kartashev and S. I. Kartashev. On modular networks satisfying the shift-register rule. IEEE Transactions on Computers, C-27(12):1153–1176, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675018.

Kartashev:1979:MSD

[KK79a] S. I. Kartashev and S. P. Kartashev. A multicomputer system with dynamic architecture. IEEE Transactions on Computers, C-28(10): 704–721, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675238.

Koren:1979:PSP

[KK79b] I. Koren and Z. Kohavi. On the properties of sensitized paths. IEEE Transactions on Computers, C-28 (3):268–269, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675335.

Klinger:1971:CAC

[KKA71] A. Klinger, A. Kochman, and N. Alexandridis. Computer analysis of chromosome patterns: Feature encoding for flexible decision making. IEEE Transactions on Computers, C-20(9):1014–1022, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671990.

Kinoshita:1973:GDS

[KKK73] E. Kinoshita, H. Kosako, and Y. Kojima. General division in the symmetric residue number system. IEEE Transactions on Computers, C-22(2):134–142, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672267;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35073.

Kinoshita:1974:FPA

[KKK74] Eisuke Kinoshita, Hideo Kosako, and Yoshiaki Kojima. Floatingpoint arithmetic algorithms in the symmetric residue number system. IEEE Transactions on Computers, C-23(1):9–20, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672365;
 https://ieeexplore.ieee.org/
xpl/tocresult.jsp?isnumber=35077.

Klaschka:1971:MRS

[Kla71] T. F. Klaschka. A method for redundancy scheme performance assessment. IEEE Transactions on Computers, C-20(11):1371–1376, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671734.

Kang:1977:SRT

[KLCC77] A. N. C. Kang, R. C. T. Lee, Chin-Liang Chang, and Shi-Kuo Chang. Storage reduction through minimal spanning trees and spanning forests. IEEE Transactions on Computers, C-26(5):425– 434, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674859.

Kleinrock:1970:STC

[Kle70] L. Kleinrock. Swap-time considerations in time-shared systems. IEEE Transactions on Computers, C-19 (6):534–540, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671564.

Kleinrock:1976:CN

[Kle76] L. Kleinrock. On communications and networks. IEEE Transactions on Computers, C-25(12):1326–1335, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674601.

Klir:1971:ULP

[Kli71] G. J. Klir. On universal logic primitives. IEEE Transactions on Computers, C-20(4):467– 469, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671862.

Klir:1974:SBP

[Kli74] G. J. Klir. On the solution of Boolean and pseudo-Boolean relations. IEEE Transactions on Computers, C-23(10):1098–1100, October 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672403.

Kluge:1977:TBT

[Klu77] W. E. Kluge. Traversing binary tree structures with shift-register memories. IEEE Transactions on Computers, C-26(11):1112–1122, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674757.

Kashef:1973:APC

[KM73] R. S. Kashef and R. B. McGhee.

Augmented parity check codes for encoding of asynchronous sequential machines. IEEE Transactions on Computers, C-22(10):891– 896, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672207.

Kashyap:1975:PRP

[KM75a] R. L. Kashyap and M. C. Mittal. Picture reconstruction from projections. IEEE Transactions on Computers, C-24(9):915–923, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672930.

Ku:1975:BDM

[KM75b] Chia-Tai Ku and G. M. Masson. The Boolean difference and multiple fault analysis. IEEE Transactions on Computers, C-24(1):62– 71, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672676.

Kashyap:1978:RSW

[KM78] R. L. Kashyap and M. C. Mittal. Recognition of spoken words and phrases in multitalker environment using syntactic methods. IEEE Transactions on Computers, C-27 (5):442–452, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675124.

Kuck:1972:NOS

[KMC72] D. J. Kuck, Y. Muraoka, and Shyh-Ching Chen. On the number of operations simultaneously executable in FORTRANlike programs and their resulting speedup. IEEE Transactions on Computers, C-21(12):1293–1310, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672094.

Korn:1970:RAF

[KN70] I. Korn and A. Nathan. On the representation of arbitrary functions and their generation. IEEE Transactions on Computers, C-19 (6):483–486, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671557.

Kim:1972:SSM

[KN72] Joonki Kim and M. M. Newborn. The simplification of sequential machines with input restrictions. IEEE Transactions on Computers, C-21(12):1440–1443, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672114.

Koontz:1975:BBC

[KNF75] W. L. G. Koontz, P. M. Narendra, and K. Fukunaga. A branch and bound clustering algorithm. IEEE Transactions on Computers, C-24(9):908–915, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672929. See comments [Wet78].

Koontz:1976:GTA

[KNF76] W. L. G. Koontz, P. M. Narendra, and K. Fukunaga. A graphtheoretic approach to nonparametric cluster analysis. IEEE Transactions on Computers, C-25(9):936– 944, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674719.

Kanani:1973:NCS

[KO73] Dhirubhai V. Kanani and Kenneth H. O’Keefe. A note on conditional-sum addition for base −2 systems. IEEE Transactions on Computers, C-22(6):626, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        www.acsel-lab.com/arithmetic/
                                        arith2/papers/ARITH2_Kanani.
                                        pdf;   https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        5009117.

Kobayashi:1978:SML

[KO78] H. Kobayashi and H. Ohara. A synthesizing method for large parallel counters with a network of smaller ones. IEEE Transactions on Computers, C-27(8):753– 757, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675184.

Koczela:1971:TFT

[Koc71] L. J. Koczela. A three-failure-tolerant computer system. IEEE Transactions on Computers, C-20 (11):1389–1393, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671737.

Kock:1975:RTP

[Koc75] W. E. Kock. A real-time parallel optical processing technique. IEEE Transactions on Computers, C-24 (4):407–411, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672828.

Kodandapani:1974:NET

[Kod74] K. L. Kodandapani. A note on easily testable realizations for logic functions. IEEE Transactions on Computers, C-23(3):332– 333, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672528.

Koh:1971:MTT

[Koh71] Kyung Shik Koh. A minimization technique for TANT networks. IEEE Transactions on Computers, C-20(1):105–107, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671683. See comments [LD72a].

Koh:1972:AR

[Koh72a] Kyung Shik Koh. Author’s reply.

IEEE Transactions on Computers, C-21(4):408, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
5008988.

Kohonen:1972:CMM

[Koh72b] Teuvo Kohonen. Correlation matrix memories. IEEE Transactions on Computers, C-21(4):353– 359, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008975.

Kohonen:1974:AAM

[Koh74] T. Kohonen. An adaptive associative memory principle. IEEE Transactions on Computers, C-23 (4):444–445, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672553.

Kohler:1975:PEC

[Koh75] W. H. Kohler. A preliminary evaluation of the critical path method for scheduling tasks on multiprocessor systems. IEEE Transactions on Computers, C-24(12):1235–1238, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672764.

Kolp:1972:SMC

[Kol72] O. Kolp. The synthesis of multivalued cellular cascades and the decomposability of group functions. IEEE Transactions on Computers, C-21(5):489–492, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672139.

Kontos:1971:IAC

[Kon71] J. Kontos. An iterative analog computer using pulse-ratio modulation. IEEE Transactions on Computers, C-20(10):1188–1192, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671696.

Koontz:1978:AR

[Koo78] W. L. G. Koontz. Author’s reply. IEEE Transactions on Computers, C-27(2):189, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675060.

Koren:1976:ICN

[Kor76] Yoram Koren. Interpolator for a computer numerical control system. IEEE Transactions on Computers, C-25(1):32– 37, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009202.

Koren:1979:ASR

[Kor79] I. Koren. Analysis of the signal reliability measure and an evaluation procedure. IEEE Transactions on Computers, C-28(3):244– 249, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675326.

Kozlay:1971:FEO

[Koz71] D. Kozlay. Feature extraction in an optical character recognition machine. IEEE Transactions on Computers, C-20(9):1063–1067, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671997.

Kontos:1971:SMC

[KP71] J. Kontos and G. Papakonstantinou. A simulation method for computer control systems. IEEE Transactions on Computers, C-20(1):98– 100, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671680.

Kamal:1974:IFM

[KP74] S. Kamal and C. V. Page. Intermittent faults: a model and a detection procedure. IEEE Transactions on Computers, C-23(7): 713–719, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672612.

Kioustelidis:1975:PLA

[KP75] J. B. Kioustelidis and J. K. Petrou.

A piecewise linear approximation of log_{2} x with equal maximum errors in all intervals. IEEE Transactions on Computers, C-24(9):858– 861, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672923.

Kuck:1977:ARM

[KPS77] David J. Kuck, Douglass S. Parker Jr., and Ahmed H. Sameh. Analysis of rounding methods in floatingpoint arithmetic. IEEE Transactions on Computers, C-26(7): 643–650, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674897.

Karplus:1971:IDC

[KR71a] W. J. Karplus and R. A. Russell. Increasing digital computer efficiency with the aid of error-correcting analog subroutines. IEEE Transactions on Computers, C-20(8):831– 837, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671950.

Kleir:1971:OSM

[KR71b] R. L. Kleir and C. V. Ramamoorthy. Optimization strategies for microprograms. IEEE Transactions on Computers, C-20(7):783– 794, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671940.

Kohonen:1973:RAD

[KR73] Teuvo Kohonen and Matti Ruohonen. Representation of associated data by matrix operators. IEEE Transactions on Computers, C-22 (7):701–702, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009138.

Knuth:1975:AIM

[KR75] D. E. Knuth and G. S. Rao. Activity in an interleaved memory. IEEE Transactions on Computers, C-24(9):943–944, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672937.

Kuhl:1978:MST

[KR78a] J. G. Kuhl and S. M. Reddy. A multicode single transition-time state assignment for asynchronous sequential machines. IEEE Transactions on Computers, C-27(10): 927–934, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674972.

Kuhl:1978:DTS

[KR78b] J. G. Kuhl and S. M. Reddy.

On the detection of terminal stuck-faults. IEEE Transactions on Computers, C-27(5):467– 469, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675129.

Kraft:1974:BMI

[Kra74] G. D. Kraft. B74-35 minicomputersinternational computer state of the art report. IEEE Transactions on Computers, C-23(6):647– 648, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672600.

Krishnamurthy:1970:OIS

[Kri70a] E. V. Krishnamurthy. On optimal iterative schemes for highspeed division. IEEE Transactions on Computers, C-19(3):227– 231, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671494.

Krishnamurthy:1970:RTT

[Kri70b] E. V. Krishnamurthy. On range-transformation techniques for division. IEEE Transactions on Computers, C-19(2):157–160, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671475.

Krishnamurthy:1971:CTW

[Kri71a] E. V. Krishnamurthy. Complementary two-way algorithms for negative radix conversions. IEEE Transactions on Computers, C-20 (5):543–550, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671881.

Krishnamurthy:1971:EIR

[Kri71b] E. V. Krishnamurthy. Economical iterative and range-transformation schemes for division. IEEE Transactions on Computers, C-20(4): 470–472, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671863.

Kriz:1973:SBO

[Kri73] Thomas A. Kriz. Some binary output sequence properties of deterministic autonomous finite-state machines with probabilistic initialization. IEEE Transactions on Computers, C-22(9):868–869, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009185.

Krishnamurthy:1977:MPU

[Kri77] E. V. Krishnamurthy. Matrix processors using p-adic arithmetic for exact linear computations. IEEE Transactions on Computers, C-26 (7):633–639, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674895.

Kroft:1974:CTC

[Kro74] D. Kroft. Comments on “A Two’s Complement Parallel Array Multiplication Algorithm”. IEEE Transactions on Computers, C-23(12):1327–1328, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672456. See [BW73, Bla74].

Kruskal:1971:CNM

[Kru71] J. B. Kruskal. Comments on “A Nonlinear Mapping for Data Structure Analysis”. IEEE Transactions on Computers, C-20(12): 1614, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671777. See [Whi72].

Kruse:1973:PPP

[Kru73] B. Kruse. A parallel picture processing machine. IEEE Transactions on Computers, C-22(12):1075–1087, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672246.

Kohavi:1971:DSF

[KS71] Z. Kohavi and D. A. Spires. Designing sets of fault-detection tests for combinational logic circuits. IEEE Transactions on Computers, C-20(12):1463–1469, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671751.

Kitai:1972:CSD

[KS72] R. Kitai and K.-H. Siemens. Comments on “A Simplified Definition of Walsh Functions”. IEEE Transactions on Computers, C-21 (5):512, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672148. See [Car71b, LM71b].

Kogge:1973:PAE

[KS73] Peter M. Kogge and Harold S. Stone. A parallel algorithm for the efficient solution of a general class of recurrence equations. IEEE Transactions on Computers, C-22 (8):786–793, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009159.

Kodandapani:1975:RMC

[KS75] K. L. Kodandapani and R. V.

Setlur. Reed–Muller canonical forms in multivalued logic. IEEE Transactions on Computers, C-24 (6):628–636, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672868.

Keshavan:1977:SET

[KS77a] H. R. Keshavan and M. D.

Srinath. Sequential estimation technique for enhancement of noisy images. IEEE Transactions on Computers, C-26(10):971– 987, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674732.

Kodandapani:1977:NMR

[KS77b] K. L. Kodandapani and R. V.

Setlur. A note on minimal Reed– Muller canonical forms of switching functions. IEEE Transactions on Computers, C-26(3):310– 313, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674830.

Kornerup:1977:UNR

[KS77c] P. Kornerup and B. D. Shriver.

A unified numeric representation arithmetic unit and its language support. IEEE Transactions on Computers, C-26(7):651– 659, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674898.

Kanefsky:1978:DTA

[KS78a] M. Kanefsky and M. G. Strintzis. A decision theory approach to picture smoothing. IEEE Transactions on Computers, C-27(1):32– 36, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674949.

Kodandapani:1978:CNR

[KS78b] K. L. Kodandapani and S. C. Seth. On combinational networks with restricted fan-out. IEEE Transactions on Computers, C-27(4):309– 318, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675103.

Kodandapani:1978:CAM

[KS78c] K. L. Kodandapani and R. V. Setlur. A cellular array for multivalued logic functions. IEEE Transactions on Computers, C-27 (11):1055–1059, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674995.

Koren:1979:RAM

[KS79] I. Koren and S. Y. H. Su. Reliability analysis of N-modular redundancy systems with intermittent and permanent faults. IEEE Transactions on Computers, C-28 (7):514–520, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675397.

Kamal:1974:GPA

[KSA74] A. K. Kamal, H. Singh, and D. P. Agrawal. A generalized pipeline array. IEEE Transactions on Computers, C-23(5):533– 536, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672570.

Kinoshita:1976:MBL

[KSM76] Kozo Kinoshita, Tsutomu Sasao, and Jun Matsuda. On magnetic bubble logic circuits. IEEE Transactions on Computers, C-25(3): 247–253, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009251.

Kratz:1974:MAS

[KSW74] G. L. Kratz, W. W. Sproul, and E. T. Walendziewicz. A microprogrammed approach to signal processing. IEEE Transactions on Computers, C-23(8):808– 817, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672630.

Kitahashi:1972:OEM

[KT72] Tadahiro Kitahashi and Kohkichi Tanaka. Orthogonal expansion of many-valued logical functions and its application to their realization with a single-threshold element. IEEE Transactions on Computers, C-21(2):211–218, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008930.

Kuck:1970:PHS

[Kuc70] D. J. Kuck. A preprocessing highspeed memory system. IEEE Transactions on Computers, C-19(9): 793–802, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671635.

Kulisch:1977:MFC

[Kul77] Ulrich Kulisch. Mathematical foundation of computer arithmetic. IEEE Transactions on Computers, C-26(7):610–621, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674893.

Kulkarni:1978:MAH

[Kul78] A. V. Kulkarni. On the mean accuracy of hierarchical classifiers. IEEE Transactions on Computers, C-27 (8):771–776, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675190.

Kunt:1975:CHT

[Kun75] M. Kunt. On computation of the Hadamard transform and the R transform in ordered form. IEEE Transactions on Computers, C-24(11):1120–1121, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672736.

Kunz:1979:EBO

[Kun79] H. O. Kunz. On the equivalence between one-dimensional discrete Walsh–Hadamard and multidimensional discrete Fourier transforms. IEEE Transactions on Computers, C-28(3):267– 268, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675334.

Kuo:1974:BSS

[Kuo74] B. C. Kuo. B 74-31 signals, systems, and the computer. IEEE Transactions on Computers, C-23 (5):557–558, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672580.

Kurtzberg:1974:MCP

[Kur74] J. M. Kurtzberg. On the memory conflict problem in multiprocessor systems. IEEE Transactions on Computers, C-23(3):286– 293, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672517.

Kurtzberg:1973:BPA

[KV73] J. M. Kurtzberg and R. D. Villani. A balanced pipelining approach to multiprocessing on an instruction stream level. IEEE Transactions on Computers, C-22(2):143– 148, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672268.

Korn:1976:AHC

[KV76] G. A. Korn and R. Vichnevetsky. Analog/hybrid computation and digital simulation. IEEE Transactions on Computers, C-25(12):1312–1320, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674599.

Kameda:1970:SMN

[KW70] T. Kameda and P. Weiner. On the state minimization of nondeterministic finite automata. IEEE Transactions on Computers, C-19 (7):617–627, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671587.

Kambayashi:1970:FMS

[KYO70] Y. Kambayashi, S. Yajima, and I. Ohbayashi. On finite-memory sequential machines. IEEE Transactions on Computers, C-19(3):254– 258, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671498.

Lamagna:1979:CMN

[Lam79a] E. A. Lamagna. The complexity of monotone networks for certain bilinear forms, routing problems, sorting, and merging. IEEE Transactions on Computers, C-28(10):773– 782, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675245.

Lamport:1979:HMM

[Lam79b] L. Lamport. How to make a multiprocessor computer that correctly executes multiprocess programs. IEEE Transactions on Computers, C-28(9):690–691, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675439.

Lang:1976:IBP

[Lan76] T. Lang. Interconnections between processors and memory modules using the shuffle-exchange network. IEEE Transactions on Computers, C-25(5):496–503, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674637.

Landis:1977:MRR

[Lan77] D. Landis. Multiple-response resolution in associative systems. IEEE Transactions on Computers, C-26 (3):230–235, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674811.

Langdon:1978:DCT

[Lan78] G. G. Langdon, Jr. A decomposition chart technique to aid in realizations with multiplexers. IEEE Transactions on Computers, C-27(2):157–159, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675048.

Langdon:1979:DMI

[Lan79] G. G. Langdon, Jr. Database machines: An introduction. IEEE Transactions on Computers, C-28 (6):381–383, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675377.

Larsen:1971:CAP

[Lar71] R. P. Larsen. Computer-aided preliminary layout design of customized MOS arrays. IEEE Transactions on Computers, C-20(5): 512–523, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671878.

Lawler:1970:RGT

[Law70] E. L. Lawler. R70-9 A graphtheoretic model for periodic discrete structures. IEEE Transactions on Computers, C-19(6): 563, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671573.

Lawrie:1975:AAD

[Law75] D. H. Lawrie. Access and alignment of data in an array processor. IEEE Transactions on Computers, C-24(12):1145–1155, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672750.

Latino:1974:CCE

[LB74] C. D. Latino and J. G. Bredeson. Comments on “Cause-Effect Analysis for Multiple Fault Detection in Combinational Networks”. IEEE Transactions on Computers, C-23 (12):1326, December 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672452. See [BH71a].

Little:1972:DMA

[LC72] W. D. Little and A. C. Capel. Digital multiplexing analog signals. IEEE Transactions on Computers, C-21(8):920, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009054.

Lee:1979:ASB

[LC79] Share Young Lee and Hsu Chang. Associative-search bubble devices for content-addressable memory and array logic. IEEE Transactions on Computers, C-28(9):627– 636, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675430.

Lee:1972:CMT

[LD72a] Hsiao P. Lee and Edward S. Davidson. Comments on “A Minimization Technique for TANT Networks”. IEEE Transactions on Computers, C-21(4):407, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008987. See [Koh71].

Lee:1972:TNN

[LD72b] Hsiao-Peng Lee and E. S. Davidson. A transform for NAND network design. IEEE Transactions on Computers, C-21(1):12– 20, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672019.

Lee:1974:RTC

[LD74] H.-P. S. Lee and E. S. Davidson. Redundancy testing in combinational networks. IEEE Transactions on Computers, C-23(10): 1029–1047, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672397. See comments [Smi79b].

Lakshminarasimhan:1975:UAF

[LD75] A. L. Lakshminarasimhan and B. V. Dasarathy. A unified approach to feature selection and learning in unsupervised environments. IEEE Transactions on Computers, C-24(9):948–952, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672939.

Lempel:1971:HSG

[LE71] A. Lempel and W. L. Eastman. High speed generation of maximal length sequences. IEEE Transactions on Computers, C-20(2):227– 229, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671813.

Lee:1977:FNS

[LE77] S. C. Lee and A. D. Edgar. The Focus number system. IEEE Transactions on Computers, C-26 (11):1167–1170, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=

1674770. See comments [LE79, Swa79a].

Lee:1979:AFN

[LE79] S. C. Lee and A. D. Edgar. Addendum to “The Focus Number System”. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675442. See [LE77, Swa79a].

Lechner:1970:CCF

[Lec70a] R. J. Lechner. Comment on “Computation of the Fast Walsh– Fourier Transform”. IEEE Transactions on Computers, C-19(2): 174, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671483. See [Hen70].

Lechner:1970:TAL

[Lec70b] R. J. Lechner. A transform approach to logic design. IEEE Transactions on Computers, C-19(7): 627–640, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671588.

Ledley:1972:AC

[Led72] R. S. Ledley. Analysis of cells. IEEE Transactions on Computers, C-21 (7):740–753, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672170.

Leeman:1975:SPC

[Lee75] G. B. Leeman, Jr. Some problems in certifying microprograms. IEEE Transactions on Computers, C-24 (5):545–553, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672851.

Lee:1976:VBA

[Lee76] S. C. Lee. Vector Boolean algebra and calculus. IEEE Transactions on Computers, C-25(9):865– 874, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674711.

Lee:1978:MFP

[Lee78a] Deok Kyo Lee. Method of folding a piecewise polynomial function in the delta function integral representation. IEEE Transactions on Computers, C-27(12):1177–1178, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675019.

Lee:1978:AMT

[Lee78b] H.-P. S. Lee. An algorithm for minimal TANT network generation. IEEE Transactions on Computers, C-27(12):1202–1206, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675028.

Leinius:1974:BCT

[Lei74] R. P. Leinius. B 74-29 currents in the theory of computing. IEEE Transactions on Computers, C-23 (5):556–557, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672578.

Lempel:1970:HBG

[Lem70] A. Lempel. On a homomorphism of the de Bruijn graph and its applications to the design of feedback shift registers. IEEE Transactions on Computers, C-19(12):1204–1209, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671452.

Lenfant:1977:FRS

[Len77] J. Lenfant. Fast random and sequential access to dynamic memories of any size. IEEE Transactions on Computers, C-26(9):847– 855, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674933.

Lenfant:1978:PPD

[Len78] J. Lenfant. Parallel permutations of data: a Benes network control algorithm for frequently used permutations. IEEE Transactions on Computers, C-27(7): 637–647, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675164.

Lesser:1971:IDE

[Les71] V. R. Lesser. An introduction to the direct emulation of control structures by a parallel microcomputer. IEEE Transactions on Computers, C-20(7):751– 764, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671937.

Levitt:1970:RTR

[Lev70] K. N. Levitt. R70-38 the time required for group multiplication. IEEE Transactions on Computers, C-19(9):859–860, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671659.

Lewis:1974:BTP

[Lew74] P. M. Lewis. B74-17 the theory of parsing, translation, and compiling — Volume II: Compiling. IEEE Transactions on Computers, C-23(2):219–220, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672499.

Lew:1977:ACG

[Lew77] A. Lew. On analog computer generation of continuous functions. IEEE Transactions on Computers, C-26 (3):193–195, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674807.

Lee:1972:SSA

[LF72] H. C. Lee and King-Sun Fu. A stochastic syntax analysis procedure and its application to pattern classification. IEEE Transactions on Computers, C-21(7): 660–666, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672164.

Levy:1974:EFS

[LF74] L. S. Levy and M. Freeman. Every finite-state machine can be simulated (realized) by a synchronous (asynchronous) binary feedback shift-register machine. IEEE Transactions on Computers, C-23(2): 174–178, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672474.

Lu:1977:SEC

[LF77] Shin-Yee Lu and King-Sun Fu.

Stochastic error-correcting syntax analysis for recognition of noisy patterns. IEEE Transactions on Computers, C-26(12):1268–1276, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674788.

Lu:1978:ECT

[LF78] Shin-Yee Lu and King-Sun Fu.

Error-correcting tree automata for syntactic pattern recognition. IEEE Transactions on Computers, C-27(11):1040–1053, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674993.

Locs:1974:FED

[LG74] G. Locs and J. M. Gary. A Fortran extension for data display. IEEE Transactions on Computers, C-23(12):1257–1263, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672437.

Lanzl:1975:DCO

[LH75] F. L. Lanzl and H. Heitmann. Direct coherent Optical Fourier Transform of curves. IEEE Transactions on Computers, C-24(4):434– 442, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672832. See comments [FHTS78].

Lee:1976:ATA

[LH76] Der Tsai Lee and Se June Hong. An algorithm for transformation of an arbitrary switching function to a completely symmetric function. IEEE Transactions on Computers, C-25(11):1117–1123, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674562.

Little:1979:ASG

[LH79] W. D. Little and R. Heuft.

An area shading graphics display system. IEEE Transactions on Computers, C-28(7):528– 531, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675399.

Liu:1974:OOB

[LHSM74] Tso-Kai Liu, K. R. Hohulin, Lih-Er Shiau, and S. Muroga. Optimal one-bit full adders with different types of gates. IEEE Transactions on Computers, C-23(1):63– 70, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672371.

Li:1977:STP

[Li77] H. F. Li. Scheduling trees in parallel/pipelined processing environments. IEEE Transactions on Computers, C-26(11):1101–1112, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674756.

Ligomenides:1977:SSF

[Lig77] P. A. Ligomenides. The skipand-set fast-division algorithm.

IEEE Transactions on Computers, C-26(10):1030–1032, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674740.

Lillestrand:1972:TCD

[Lil72] R. L. Lillestrand. Techniques for change detection. IEEE Transactions on Computers, C-21(7): 654–659, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672163.

Ling:1970:HSC

[Lin70] H. Ling. High-speed computer multiplication using a multiplebit decoding algorithm. IEEE Transactions on Computers, C-19 (8):706–709, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith1/papers/ARITH1_
Ling.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=1671613.

Liou:1976:SFM

[Lio76] Ming-Lei Liou. Spline fit made easy. IEEE Transactions on Computers, C-25(5):522–527, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674640.

Lipovski:1977:VAM

[Lip77] G. Jack Lipovski. On a varistructured array of microprocessors. IEEE Transactions on Computers, C-26(2):125–138, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009291.

Little:1973:SHC

[Lit73] W. D. Little. Serial hybrid computation and errors in hybrid loops. IEEE Transactions on Computers, C-22(4):367– 370, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672316.

Little:1974:AHS

[Lit74] W. D. Little. An algorithm for high-speed digital filters. IEEE Transactions on Computers, C-23 (5):466–469, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672561.

Liu:1975:SAL

[Liu75] Tso-Kai Liu. Synthesis algorithms for 2-level MOS networks. IEEE Transactions on Computers, C-24 (1):72–79, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672677.

Liu:1977:SFF

[Liu77a] T. K. Liu. Synthesis of feed-forward MOS networks with cells of similar complexities. IEEE Transactions on Computers, C-26(8):826– 831, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674923.

Liu:1977:SMF

[Liu77b] Tso-Kai Liu. Synthesis of multilevel feed-forward MOS networks. IEEE Transactions on Computers, C-26 (6):581–588, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674883.

Liu:1979:IMB

[LJ79] J. W. S. Liu and M. Jino. Intelligent magnetic bubble memories and their applications in data base management systems. IEEE Transactions on Computers, C-28(12):888– 906, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675279.

Leonard:1975:MFI

[LK75] M. S. Leonard and K. E. Kilpatrick. A modified fixed-increment training algorithm. IEEE Transactions on Computers, C-24(6):

673–675, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672874.

Lee:1972:MSF

[LL72] Samuel C. Lee and Edward T.

Lee. On multivalued symmetric functions. IEEE Transactions on Computers, C-21(3):312– 317, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008957.

Landauro:1975:ODC

[LL75] A. Landauro and J. Lienard. On overflow detection and correction in digital filters. IEEE Transactions on Computers, C-24(12):1226–1228, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672761.

Leung:1979:MCR

[LL79] J. Y.-T. Leung and E. K. Lai.

On minimum cost recovery from system deadlock. IEEE Transactions on Computers, C-28(9):671– 677, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675435.

Loomis:1970:SSHa

[LM70] Herschel H. Loomis and Michael R. McCoy. A scheme for synchronizing high-speed logic: Part I. IEEE Transactions on Computers, C-19 (1):39–47, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008898.

Lackey:1971:AR

[LM71a] R. B. Lackey and D. Meltzer. Authors’ reply. IEEE Transactions on Computers, C-20(12):1617, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671783.

Lackey:1971:SDW

[LM71b] R. B. Lackey and D. Meltzer. A simplified definition of Walsh functions. IEEE Transactions on Computers, C-20(2):211–213, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671807. See comments [Car71b, KS72].

Liu:1978:TPE

[LM78] P. S. Liu and F. J. Mowle. Techniques of program execution with a writable control memory. IEEE Transactions on Computers, C-27(9):816–827, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675201.

Lai:1979:MPB

[LM79] Hung Chi Lai and S. Muroga. Minimum parallel binary adders with NOR (NAND) gates. IEEE Transactions on Computers, C-28(9): 648–659, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675433.

Lange-Nielsen:1972:PSA

[LNL72] T. Lange-Nielsen and G. M. Lance.

A pattern search algorithm for feedback-control system parameter optimization. IEEE Transactions on Computers, C-21(11):1222–1227, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672074.

Lomet:1974:BSE

[Lom74] D. B. Lomet. B74-22 software engineering: International computer state of the art report. IEEE Transactions on Computers, C-23(2): 223–224, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672504.

Loomis:1970:SSHb

[Loo70] H. H. Loomis, Jr. A scheme for synchronizing high-speed logic. Part II. IEEE Transactions on Computers, C-19(2):116–124, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671469.

Losq:1976:HER

[Los76] J. Losq. A highly efficient redundancy scheme: Self-purging redundancy. IEEE Transactions on Computers, C-25(6):569– 578, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674656.

Losq:1978:ERC

[Los78] J. Losq. Efficiency of random compact testing. IEEE Transactions on Computers, C-27(6):516– 525, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675142.

Loutrel:1970:SHL

[Lou70] P. P. Loutrel. A solution to the hidden-line problem for computerdrawn polyhedra. IEEE Transactions on Computers, C-19(3):205– 213, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671491.

Lowe:1973:AIS

[Low73] T. C. Lowe. Analysis of an information system model with transfer penalties. IEEE Transactions on Computers, C-22(5):469– 480, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672343.

Luccio:1976:HHB

[LP76] Fabrizio Luccio and Linda Pagli.

On the height of height-balanced trees. IEEE Transactions on Computers, C-25(1):87–90, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009208.

Luccio:1978:RHB

[LP78] F. Luccio and L. Pagli. Rebalancing height balanced trees. IEEE Transactions on Computers, C-27 (5):386–396, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675117.

Landman:1971:PVB

[LR71a] B. S. Landman and R. L. Russo.

On a pin versus block relationship for partitions of logic graphs. IEEE Transactions on Computers, C-20(12):1469–1479, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671752.

Laws:1971:CAM

[LR71b] B. A. Laws, Jr. and C. K. Rushforth. A cellular-array multiplier for GF(2^{m}). IEEE Transactions on Computers, C-20(12):1573–1578, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671766.

Larsen:1972:RCV

[LR72] Ronald W. Larsen and Irving S. Reed. Redundancy by coding versus redundancy by replication for failure-tolerant sequential circuits. IEEE Transactions on Computers, C-21(2):130–137, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008917.

Liu:1975:DMC

[LR75] J. J. Liu and L. D. Rudolph. A direct method of computing the GNAF of an integer. IEEE Transactions on Computers, C-24(10): 1042–1043, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672719.

Luczak:1976:DHG

[LR76] E. Luczak and A. Rosenfeld. Distance on a hexagonal grid. IEEE Transactions on Computers, C-25 (5):532–533, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674642.

Lawson:1971:FCM

[LS71] H. W. Lawson, Jr. and B. K.

Smith. Functional characteristics of a multilingual processor. IEEE Transactions on Computers, C-20 (7):732–742, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671935.

Lawrence:1972:RPS

[LS72] J. P. Lawrence and Kenneth Steiglitz. Randomized pattern search. IEEE Transactions on Computers, C-21(4):382– 385, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008979.

Lang:1976:SEN

[LS76a] Tomas Lang and Harold S. Stone.

A shuffle-exchange network with simplified control. IEEE Transactions on Computers, C-25(1):55– 65, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009205.

Loveland:1976:HGT

[LS76b] D. W. Loveland and M. E. Stickel. A hole in goal trees: Some guidance from resolution theory. IEEE Transactions on Computers, C-25 (4):335–341, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674614.

Lee:1977:TMS

[LSB77] R. C. T. Lee, J. R. Slagle, and H. Blum. A triangulation method for the sequential mapping of points from N-space to two-space. IEEE Transactions on Computers, C-26 (3):288–292, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674822.

Lakshmivarahan:1970:PCU

[LT70] S. Lakshmivarahan and M. A. L. Thathachar. Pattern classification using stochastic approximation techniques. IEEE Transactions on Computers, C-19(7):649– 651, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671591.

Larson:1973:PPA

[LT73] Robert E. Larson and Edison Tse.

Parallel processing algorithms for the optimal control of nonlinear dynamic systems. IEEE Transactions on Computers, C-22(8):777– 786, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009158. See comments [Tab74].

Lavington:1977:MMC

[LTE77] Simon H. Lavington, Gareth Thomas, and David B. G. Edwards. The MU5 multicomputer communication system. IEEE Transactions on Computers, C-26(1):19– 28, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009269.

Lucas:1972:BCB

[Luc72] H. C. Lucas, Jr. B72-20 computers in business — an introduction, 2nd ed. IEEE Transactions on Computers, C-21(10): 1136, October 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672060.

Luderer:1974:BDL

[Lud74] G. W. R. Luderer. B74-16 design of on-line computer systems. IEEE Transactions on Computers, C-23(2):218–219, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672498.

LeVan:1975:DUL

[LV75] T. Le Van and N. Van Houtte. Delayed universal logic modules and sequential machine synthesis. IEEE Transactions on Computers, C-24 (8):853–855, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672916. See comments [Sur78].

Liu:1978:ECC

[LW78] Chao-Kai Liu and Tse Lin Wang. Error-correcting codes in binarycoded-decimal arithmetic. IEEE Transactions on Computers, C-27(11):977–984, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674987.

Landgraff:1971:DDI

[LY71] R. W. Landgraff and S. S. Yau. Design of diagnosable iterative arrays. IEEE Transactions on Computers, C-20(8):867– 877, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671956. See comments [FM72b].

Lynn:1972:NRR

[Lyn72] Donald S. Lynn. New results for Rado’s sigma function for binary Turing Machines. IEEE Transactions on Computers, C-21(8):894– 896, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009047.

Lyons:1970:SRT

[Lyo70] R. E. Lyons. The synthesis of redundant threshold-logic elements. IEEE Transactions on Computers, C-19(5):429–443, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671534.

Marihugh:1971:DGA

[MA71] G. E. Marihugh and R. E. Anderson. The H diagram: a graphical approach to logic design. IEEE Transactions on Computers, C-20(10):1192–1196, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671697.

McKee:1977:CRP

[MA77] J. W. McKee and J. K. Aggarwal. Computer recognition of partial views of curved objects. IEEE Transactions on Computers, C-26 (8):790–800, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674917.

Maritsas:1978:PSA

[MAB78] D. G. Maritsas, A. C. Arvillias, and A. C. Bounas. Phase-shift analysis of linear feedback shift register structures generating pseudorandom sequences. IEEE Transactions on Computers, C-27(7): 660–669, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675166.

Macleod:1970:POL

[Mac70] I. D. G. Macleod. Pictorial output with a line printer. IEEE Transactions on Computers, C-19(2):160– 162, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671476. See correction [Ano70y].

Mage:1970:AIC

[Mag70] J. M. Mage. Application of iterative consensus to multipleoutput functions. IEEE Transactions on Computers, C-19(4): 359, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671517.

Mago:1971:RMA

[Mag71] G. Mago. Realization methods for asynchronous sequential circuits. IEEE Transactions on Computers, C-20(3):290– 297, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671826.

Mago:1973:MFS

[Mag73] G. Mago. Monotone functions in sequential circuits. IEEE Transactions on Computers, C-22(10): 928–933, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672213. See comments [Iba74].

Mahl:1972:VSA

[Mah72] Robert Mahl. Visible surface algorithms for quadric patches. IEEE Transactions on Computers, C-21 (1):1–4, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672017.

Maheshwari:1974:BCT

[Mah74] S. N. Maheshwari. B74-36 communication, transmission, and transportation networks. IEEE Transactions on Computers, C-23(8): 878, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672637.

Maison:1971:MSR

[Mai71] F. P. Maison. The MECRA: a selfreconfigurable computer for highly reliable process. IEEE Transactions on Computers, C-20(11):1382–1388, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671736.

Majithia:1972:CAE

[Maj72a] J. C. Majithia. Cellular array for extraction of squares and square roots of binary numbers. IEEE Transactions on Computers, C-21(9):1023– 1024, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009084.

Majithia:1972:CFD

[Maj72b] J. C. Majithia. Comments on “A Fast Digital Computer Method for Recursive Estimation of the Mean”. IEEE Transactions on Computers, C-21(5):511– 512, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672147. See [Whi70].

Majithia:1972:STD

[Maj72c] J. C. Majithia. A simple technique for determination of essential multiple output prime implicants. IEEE Transactions on Computers, C-21(9):1024–1026, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009085.

Majithia:1973:CAB

[Maj73] J. C. Majithia. A cellular array for binary to BCD conversion. IEEE Transactions on Computers, C-22 (4):423–424, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672325.

Majithia:1976:SCC

[Maj76] J. C. Majithia. Some comments concerning design of pipeline arithmetic arrays. IEEE Transactions on Computers, C-25(11):1132–1134, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674565.

Makhoul:1971:SAL

[Mak71] J. Makhoul. Speaker adaptation in a limited speech recognition system. IEEE Transactions on Computers, C-20(9):1057–1063, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671996.

Mallach:1972:JMM

[Mal72] E. G. Mallach. Job-mix modeling and system analysis of an aerospace multiprocessor. IEEE Transactions on Computers, C-21 (5):446–454, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672133.

Mallach:1973:UIW

[Mal73] E. G. Mallach. On the use of “Importance weights” in assessing subsystem reliability. IEEE Transactions on Computers, C-22(4):424– 425, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672326.

Mamdani:1977:AFL

[Mam77] E. H. Mamdani. Application of fuzzy logic to approximate reasoning using linguistic synthesis. IEEE Transactions on Computers, C-26(12):1182–1191, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674779.

Mandelbaum:1971:PAI

[Man71] D. Mandelbaum. On parallelacting index registers. IEEE Transactions on Computers, C-20(3): 361, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671839.

Mandelbaum:1972:ECS

[Man72a] D. Mandelbaum. On error control in sequential machines. IEEE Transactions on Computers, C-21 (5):492–495, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672140.

Mandelbaum:1972:ECR

[Man72b] David Mandelbaum. Error correction in residue arithmetic. IEEE Transactions on Computers, C-21 (6):538–545, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009006.

Manning:1977:AHI

[Man77] F. B. Manning. An approach to highly integrated, computermaintained cellular arrays. IEEE Transactions on Computers, C-26 (6):536–552, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674879. See comments [Agr79c].

Marino:1970:LDS

[Mar70] P. J. Marino. A linear decomposition for sequential machines. IEEE Transactions on Computers, C-19(10):956–963, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671395.

Marinos:1971:DMC

[Mar71] P. N. Marinos. Derivation of minimal complete sets of testinput sequences using Boolean differences. IEEE Transactions on Computers, C-20(1):25–32, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671670.

Marino:1972:NAA

[Mar72] D. Marino. New algorithms for the approximate evaluation in hardware of binary logarithms and elementary functions. IEEE Transactions on Computers, C-21(12):1416–1421, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672109.

Maritsaa:1973:SPC

[Mar73a] D. G. Maritsaa. On the statistical properties of a class of linear product feedback shift-register sequences. IEEE Transactions on Computers, C-22(10):961–962, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672218.

Maritsas:1973:AFT

[Mar73b] D. G. Maritsas. The autocorrelation function of the two feedback shift-register pseudorandom source. IEEE Transactions on Computers, C-22(10):962– 964, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672219.

Maritsas:1973:HSA

[Mar73c] Dimitris G. Maritsas. A high speed and accuracy digital Gaussian generator of pseudorandom numbers. IEEE Transactions on Computers, C-22(7):629– 634, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009127.

Maruyama:1973:PEP

[Mar73d] K. Maruyama. On the parallel evaluation of polynomials. IEEE Transactions on Computers, C-22 (1):2–5, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672186.

Marjanovic:1974:MDC

[Mar74] S. Marjanovic. Multiplying D/A converter. IEEE Transactions on Computers, C-23(9):986–988, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672661.

Marinos:1975:AR

[Mar75] P. N. Marinos. Author’s reply. IEEE Transactions on Computers, C-24(1):109, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672685.

Marinos:1976:AR

[Mar76] P. N. Marinos. Author’s reply. IEEE Transactions on Computers, C-25(10):1056, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674549.

Marino:1977:EAI

[Mar77] L. R. Marino. The effect of asynchronous inputs on sequential network reliability. IEEE Transactions on Computers, C-26(11):1082–1090, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674754.

Marovac:1978:SMR

[Mar78] N. Marovac. Sieve method for realtime computer-aided extraction of multivalued analog signals in the presence of noise. IEEE Transactions on Computers, C-27(10):935– 944, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674973.

Markowsky:1979:DSF

[Mar79] G. Markowsky. Diagnosing single faults in fanout-free combinational circuits. IEEE Transactions on Computers, C-28(11):863– 864, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675266.

Mase:1978:CMC

[Mas78] K. Mase. Comments on “A Measure of Computational Work” and “Logical Network Cost and Entropy”. IEEE Transactions on Computers, C-27(1):94–95, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674960. See [Wel74, Hel74, CF73].

Matula:1970:FFP

[Mat70] D. W. Matula. A formalization of floating-point numeric base conversion. IEEE Transactions on Computers, C-19(8):681–692, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Matula.

                                        pdf;   https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671610.

Matheson:1971:PEC

[Mat71a] W. S. Matheson. PCN equivalence class invariants and information quantities. IEEE Transactions on Computers, C-20(6):691– 694, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671922.

Matheson:1971:RMU

[Mat71b] W. S. Matheson. Recognition of monotonic and unate cascade realizable functions using an informational model of switching circuits. IEEE Transactions on Computers, C-20(10):1214–1219, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671702.

Mathur:1971:RMA

[Mat71c] F. P. Mathur. On reliability modeling and analysis of ultrareliable fault-tolerant digital systems. IEEE Transactions on Computers, C-20(11):1376–1382, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671735.

Maxwell:1977:CDR

[Max77] P. C. Maxwell. Correct DDA register transfers for trapezoidal integration when solving nonlinear equations of the form ẏ = Ay^{n}. IEEE Transactions on Computers, C-26(11):1151–1153, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674764.

Mazer:1971:TSB

[Maz71] L. Mazer. Topological solution of bilateral switching networks. IEEE Transactions on Computers, C-20(2):234–238, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671816.

McCrea:1975:DDA

[MB75] P. G. McCrea and P. W. Baker.

On Digital Differential Analyzer (DDA) circle generation for computer graphics. IEEE Transactions on Computers, C-24(11):1109–1110, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672733.

Maryanski:1977:IFS

[MB77] F. J. Maryanski and T. L. Booth.

Inference of finite-state probabilistic grammars. IEEE Transactions on Computers, C-26(6):521– 536, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674878.

Maryanski:1979:AR

[MB79] F. Maryanski and T. L. Booth. Authors’ reply. IEEE Transactions on Computers, C-28(1): 64, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675225.

Miller:1977:STD

[MBM77] A. J. Miller, A. W. Brown, and P. Mars. A simple technique for the determination of delayed maximal length linear binary sequences. IEEE Transactions on Computers, C-26(8):808– 811, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674919.

Minnick:1975:CRM

[MBSS75] R. C. Minnick, P. T. Bailey, R. M. Sandfort, and W. L. Semon. Cascade realizations of magnetic bubble logic using a small set of primitives. IEEE Transactions on Computers, C-24(2):215– 217, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672782.

Moore:1970:SCA

[MC70] J. B. Moore and K. T. Clark.

A simple convergent algorithm for rapid solution of polynomial equations. IEEE Transactions on Computers, C-19(1):79– 80, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008904.

McCluskey:1971:FEC

[MC71] E. J. McCluskey and F. W. Clegg.

Fault equivalence in combinational logic networks. IEEE Transactions on Computers, C-20(11):1286–1293, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671722.

Menon:1978:DFS

[MC78] P. R. Menon and S. G. Chappell.

Deductive fault simulation with functional blocks. IEEE Transactions on Computers, C-27(8):689– 695, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675175.

McCluskey:1979:LDM

[McC79] E. J. McCluskey. Logic design of multivalued I^{2}L logic circuits. IEEE Transactions on Computers, C-28 (8):546–559, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675410.

Mcdougall:1972:NIE

[Mcd72] M. H. Mcdougall. A note on the interruption of extended core storage transfers. IEEE Transactions on Computers, C-21(1):87– 90, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672028.

McLean:1972:BMD

[McL72] E. R. McLean. B72-23 modem data processors and systems. IEEE Transactions on Computers, C-21(11):1251–1252, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672083.

McCaughern:1972:VSF

[MCRD72] R. W. McCaughern, D. M. Caughey, H. Rombeek, and W. S. Davidson. A video simulation facility. IEEE Transactions on Computers, C-21(7):738– 740, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672169.

Mei:1974:BSF

[Mei74] K. C. Y. Mei. Bridging and stuck-at faults. IEEE Transactions on Computers, C-23(7):720– 727, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672613.

Menger:1970:AR

[Men70a] K. S. Menger. Author’s reply. IEEE Transactions on Computers, C-19 (6):560, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671571.

Mengert:1970:SLI

[Men70b] P. H. Mengert. Solution of linear inequalities. IEEE Transactions on Computers, C-19(2):124– 131, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671470.

Meo:1975:ANT

[Meo75] A. R. Meo. Arithmetic networks and their minimization using a new line of elementary units. IEEE Transactions on Computers, C-24 (3):258–280, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672800.

Metropolis:1973:ABC

[Met73] Nicholas C. Metropolis. Analyzed binary computing. IEEE Transactions on Computers, C-22(6): 573–576, June 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009109.

Meyer:1971:FTS

[Mey71] J. F. Meyer. Fault tolerant sequential machines. IEEE Transactions on Computers, C-20(10): 1167–1177, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671694.

Meyer:1976:CBR

[Mey76] J. F. Meyer. Computation-based reliability analysis. IEEE Transactions on Computers, C-25(6): 578–584, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674657.

Menon:1971:FDI

[MF71] P. R. Menon and A. D. Friedman. Fault detection in iterative logic arrays. IEEE Transactions on Computers, C-20(5): 524–535, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671879.

Merlin:1975:PMD

[MF75] P. M. Merlin and D. J. Farber.

A parallel mechanism for detecting curves in pictures. IEEE Transactions on Computers, C-24(1):96– 98, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672680.

Malhotra:1976:DEC

[MF76a] Vinod K. Malhotra and Robert D.

Fisher. A double error-correction scheme for peripheral systems. IEEE Transactions on Computers, C-25(2):105–115, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009225.

Manning:1976:SCC

[MF76b] Frank B. Manning and Robert R.

Fenichel. Synchronous counters constructed entirely of J– K flip-flops. IEEE Transactions on Computers, C-25(3):300– 306, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009257.

Moayer:1976:TSA

[MF76c] Bijan Moayer and King-Sun Fu.

A tree system approach for fingerprint pattern recognition. IEEE Transactions on Computers, C-25 (3):262–274, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009253.

Montalvo:1976:AIC

[MF76d] Fanya S. Montalvo and Caxton C. Foster. An algorithm for intercell communication in a tesselated automaton. IEEE Transactions on Computers, C-25(1):99– 102, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009214.

Marouf:1978:EDS

[MF78] M. A. Marouf and A. D. Friedman. Efficient design of selfchecking checker for any mout-of-n code. IEEE Transactions on Computers, C-27(6):482– 490, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675138.

Michales:1971:HCS

[MG71a] G. C. Michales and V. Gourishankar. Hybrid computer solution of optimal control problems. IEEE Transactions on Computers, C-20(2):209–211, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671806.

Mucciardi:1971:CST

[MG71b] A. N. Mucciardi and E. E. Gose. A comparison of seven techniques for choosing subsets of pattern recognition properties. IEEE Transactions on Computers, C-20(9):1023– 1031, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671991. See comments [Tou72a].

McDonald:1975:TCQ

[MG75] T. G. McDonald and R. K. Guha.

The two’s complement quasi-serial multiplier. IEEE Transactions on Computers, C-24(12):1233–1235, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672763.

Maritsas:1970:BLE

[MH70a] D. G. Maritsas and M. G. Hartley.

Buffer length for Erlang input and constant removal rate. IEEE Transactions on Computers, C-19(9): 839–843, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671644.

Maritsas:1970:CSV

[MH70b] D. G. Maritsas and M. G. Hartley. A case study of a versatile generator of repeatable non-Poisson sequences of pseudorandom pulses. IEEE Transactions on Computers, C-19(10):924– 938, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671393.

Maritsas:1970:DCG

[MH70c] D. G. Maritsas and M. G. Hartley. Design criteria for a generator of repeatable non-Poisson sequences of pseudorandom pulses. IEEE Transactions on Computers, C-19(9): 812–817, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671638.

Maruoka:1973:LNF

[MH73] A. Maruoka and N. Honda. Logical networks of flexible cells. IEEE Transactions on Computers, C-22 (4):347–358, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672314.

Maruoka:1975:RLF

[MH75] A. Maruoka and N. Honda. The range of logical flexibility of tree networks. IEEE Transactions on Computers, C-24(1):9– 28, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672673.

Maheshwari:1976:MDS

[MH76] Shachindra N. Maheshwari and S. Louis Hakimi. On models for diagnosable systems and probabilistic fault diagnosis. IEEE Transactions on Computers, C-25(3):228– 236, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009249. See corrections and comments [MH78].

Maheshwari:1978:CCM

[MH78] S. N. Maheshwari and S. L.

Hakimi. Corrections and comments on “On Models for Diagnosable Systems and Probabilistic Fault Diagnosis”. IEEE Transactions on Computers, C-27(3):287– 288, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675092. See [MH76].

Muroga:1972:DOS

[MI72] Saburo Muroga and Toshihide Ibaraki. Design of optimal switching networks by integer programming. IEEE Transactions on Computers, C-21(6):573– 582, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009010.

Michalopouos:1974:BPR

[Mic74] D. A. Michalopouos. B74-46 pattern recognition, learning and thought. IEEE Transactions on Computers, C-23(12):1330, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672460.

Miller:1973:CST

[Mil73] Raymond E. Miller. A comparison of some theoretical models of parallel computation. IEEE Transactions on Computers, C-22(8):710– 717, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009149.

Miller:1974:BOS

[Mil74a] E. F. Miller, Jr. B74-25 operating systems: a pragmatic approach. IEEE Transactions on Computers, C-23(3):336, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672534.

Miller:1974:MSR

[Mil74b] E. F. Miller, Jr. A multiplestream registerless shared-resource processor. IEEE Transactions on Computers, C-23(3):277–285, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672516.

Milgram:1975:CMC

[Mil75] D. L. Milgram. Computer methods for creating photomosaics. IEEE Transactions on Computers, C-24(11):1113–1119, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672735.

Milgram:1977:ATP

[Mil77] D. L. Milgram. Adaptive techniques for photomosaicking. IEEE Transactions on Computers, C-26 (11):1175–1180, November 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674772.

Minnick:1975:SMB

[Min75] R. C. Minnick. A system of magnetic bubble logic. IEEE Transactions on Computers, C-24(2):217– 218, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672783.

Mouftah:1977:DTC

[MJ77] H. T. Mouftah and I. B. Jordan.

Design of ternary COS/MOS memory and sequential circuits. IEEE Transactions on Computers, C-26 (3):281–288, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674821. See correction [MJ78].

Mouftah:1978:CDT

[MJ78] H. T. Mouftah and I. B. Jordan. Correction to “Design of Ternary COS/MOS Memory and Sequential Circuits”. IEEE Transactions on Computers, C-27(2): 192, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675066. See [MJ77].

McMahon:1975:HOC

[MJTW75] D. H. McMahon, G. L. Johnson, S. L. Teeter, and C. G. Whitney. A hybrid optical computer processing technique for fingerprint identification. IEEE Transactions on Computers, C-24(4):358– 369, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672822.

Majithia:1971:CAN

[MK71a] J. C. Majithia and R. Kitai. A cellular array for the nonrestoring extraction of square roots. IEEE Transactions on Computers, C-20(12):1617–1618, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671784.

Majithia:1971:IAM

[MK71b] J. C. Majithia and R. Kitai. An iterative array for multiplication of signed binary numbers. IEEE Transactions on Computers, C-20(2):214–216, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671809.

McPherson:1979:TLD

[MK79] J. A. McPherson and C. R. Kime.

A two-level diagnostic model for digital systems. IEEE Transactions on Computers, C-28(1):16– 27, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675218.

Murakami:1970:SMC

[MKO70] S. Murakami, K. Kinoshita, and H. Ozaki. Sequential machines capable of fault diagnosis. IEEE Transactions on Computers, C-19(11):1079–1085, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671428.

Muroga:1976:MLN

[ML76] S. Muroga and Hung Chi Lai.

Minimization of logic networks under a generalized cost function. IEEE Transactions on Computers, C-25(9):893–907, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674714.

Morreale:1970:CEP

[MM70] E. Morreale and M. Mennucci. Computer experience on partitioned list algorithms. IEEE Transactions on Computers, C-19 (11):1099–1105, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671433.

Marasa:1973:SSC

[MM73a] John D. Marasa and David W. Matula. A simulative study of correlated error propagation in various finite-precision arithmetics. IEEE Transactions on Computers, C-22(6):587–597, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Marasa.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009111.

Meisel:1973:PAA

[MM73b] W. S. Meisel and D. A. Michalopoulos. A partitioning algorithm with application in pattern classification and the optimization of decision trees. IEEE Transactions on Computers, C-22(1):93– 103, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672196.

Mallela:1978:DSI

[MM78a] S. Mallela and G. M. Masson.

Diagnosable systems for intermittent faults. IEEE Transactions on Computers, C-27(6):560– 566, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675147.

Meyer:1978:EFD

[MM78b] G. G. L. Meyer and G. M.

Masson. An efficient fault diagnosis algorithm for symmetric multiple processor architectures. IEEE Transactions on Computers, C-27(11):1059–1063, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674996.

Mitchell:1977:MMM

[MMB77] O. R. Mitchell, C. R. Myers, and W. Boyne. A max-min measure for image texture analysis. IEEE Transactions on Computers, C-26 (4):408–414, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674850.

McCrea:1978:CFP

[MMB78] P. G. McCrea, P. C. Maxwell, and P. W. Baker. Comments on “A Floating-Point Multiplexed DDA System”. IEEE Transactions on Computers, C-27(12):1226, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675033. See [HW76].

Matsushita:1974:MCC

[MMO74] S. Matsushita and T. Moto-Oka. Magnitude of cross-coupling noise in digital multiwire transmission lines. IEEE Transactions on Computers, C-23(11):1122–1132, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672412.

McGhee:1970:ERD

[MN70] Robert B. McGhee and Ragnar N. Nilsen. The extended resolution digital differential analyzer: a new computing structure for solving differential equations. IEEE Transactions on Computers, C-19(1):1– 9, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008893.

Mark:1978:SPS

[MN78] J. W. Mark and S. F. W. Ng. Satellite packet switching with global assignments and batch Poisson arrivals. IEEE Transactions on Computers, C-27(12):1216–1221, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675031.

Mohn:1971:TSF

[Moh71] W. S. Mohn, Jr. Two statistical feature evaluation techniques applied to speaker identification. IEEE Transactions on Computers, C-20(9):979–987, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671985.

Montangero:1974:AOS

[Mon74] C. Montangero. An approach to the optimal specification of readonly memories in microprogrammed digital computers. IEEE Transactions on Computers, C-23(4):375– 389, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672546.

Moore:1971:BSS

[Moo71] F. R. Moore. On the bounds for state-set size in the proofs of equivalence between deterministic, nondeterministic, and twoway finite automata. IEEE Transactions on Computers, C-20(10): 1211–1214, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671701.

Moore:1976:DSC

[Moo76] R. C. Moore. D-Script: a computational theory of descriptions. IEEE Transactions on Computers, C-25 (4):366–373, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674619.

Morishita:1970:AAT

[Mor70a] I. Morishita. Analysis of an adaptive threshold logic unit. IEEE Transactions on Computers, C-19(12):1181–1192, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671450.

Morreale:1970:CCP

[Mor70b] E. Morreale. Computational complexity of partitioned list algorithms. IEEE Transactions on Computers, C-19(5):421– 428, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671533.

Morreale:1970:ROP

[Mor70c] E. Morreale. Recursive operators for prime implicant and irredundant normal form determination. IEEE Transactions on Computers, C-19 (6):504–509, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671560.

Morris:1971:TFP

[Mor71] R. Morris. Tapered floating point: a new floating-point representation. IEEE Transactions on Computers, C-20(12):1578–1579, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671767.

Morrin:1974:BWR

[Mor74] T. H. Morrin. A black-white representation of a gray-scale picture. IEEE Transactions on Computers, C-23(2):184–186, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672476.

Morgan:1975:DCU

[Mor75] D. R. Morgan. A/D conversion using geometric feedback AGC. IEEE Transactions on Computers, C-24(11):1074–1078, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672728.

Mor:1976:TVS

[Mor76a] M. Mor. On the three-valued simulation of digital systems. IEEE Transactions on Computers, C-25(11):1152–1156, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674571.

Morgan:1976:MAT

[Mor76b] C. G. Morgan. Methods for automated theorem proving in nonclassical logics. IEEE Transactions on Computers, C-25(8):852– 862, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674704.

Moravec:1979:FIM

[Mor79] H. P. Moravec. Fully interconnecting multiple computers with pipelined sorting nets. IEEE Transactions on Computers, C-28(10): 795–798, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675249.

McCharen:1976:PEA

[MOW76] J. D. McCharen, R. A. Overbeek, and L. A. Wos. Problems and experiments for and with automated theorem-proving programs. IEEE Transactions on Computers, C-25 (8):773–782, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674696.

McCluskey:1978:BNP

[MPS78] E. J. McCluskey, K. P. Parker, and J. J. Shedletsky. Boolean network probabilities and network design. IEEE Transactions on Computers, C-27(2):187– 189, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675058.

Miczo:1972:RBO

[MR72] A. Miczo and L. D. Rudolph. A relationship between output symbol occurrence rate and observability of autonomous machines. IEEE Transactions on Computers, C-21 (8):911–913, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009052.

Mennone:1974:ECL

[MR74] A. Mennone and R. L. Russo.

An example computer logic graph and its partitions and mappings. IEEE Transactions on Computers, C-23(11):1198–1204, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672423.

Molberg:1975:NMI

[MR75] J. R. Molberg and D. K. Reynolds. A new method for initial estimates in the numerical solution of Laplace’s equation. IEEE Transactions on Computers, C-24(6): 678–680, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672876.

Meyer:1976:FTC

[MR76] J. F. Meyer and J.-C. Rault. Fault-tolerant computing: a introduction. IEEE Transactions on Computers, C-25(6):553– 556, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674654.

Mukhopadhyay:1970:MEL

[MS70] A. Mukhopadhyay and G. Schmitz. Minimization of exclusive or and logical equivalence switching circuits. IEEE Transactions on Computers, C-19(2):132–140, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671471.

Maki:1973:CNS

[MS73a] G. K. Maki and D. H. Sawin III. Comments on “Next-State Equations of Asynchronous Sequential Machines”. IEEE Transactions on Computers, C-22(2): 218, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672282. See [Sau72a].

Mar:1973:FHT

[MS73b] H. Y. L. Mar and C. L. Sheng.

Fast Hadamard Transform using the H diagram. IEEE Transactions on Computers, C-22(10):957– 960, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672217.

Maki:1974:FTA

[MS74] G. K. Maki and D. H. Sawin III. Fault-tolerant asynchronous sequential machines. IEEE Transactions on Computers, C-23(7): 651–657, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672606.

Meyer:1975:LDU

[MS75a] J. F. Meyer and R. J. Sundstrom. On-line diagnosis of unrestricted faults. IEEE Transactions on Computers, C-24(5):468– 475, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672841.

Mizoguchi:1975:AUL

[MS75b] R. Mizoguchi and M. Shimura. An approach to unsupervised learning classification. IEEE Transactions on Computers, C-24(10):979– 983, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672697.

Mizoguchi:1976:NLT

[MS76] R. Mizoguchi and M. Shimura. Nonparametric learning without a teacher based on mode estimation. IEEE Transactions on Computers, C-25(11):1109–1117, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674561.

Mehta:1977:SOA

[MS77a] M. A. Mehta and W. B. Smith. Synthesis of optimal ambiguity resolver functions. IEEE Transactions on Computers, C-26(8):782– 789, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674916.

Mitchell:1977:TLM

[MS77b] R. L. Mitchell and C. R. Stone. Table-lookup methods for generating arbitrary random numbers. IEEE Transactions on Computers, C-26(10):1006–1008, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674735.

Mitra:1978:ITD

[MS78] S. K. Mitra and G. K. Sorknes.

On the implementation of a twodimensional FIR filter using a single multiplier. IEEE Transactions on Computers, C-27(8):762– 764, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675186.

Maki:1972:ISA

[MSJ72] G. K. Maki, D. H. Sawin III, and B.-R. A. Jeng. Improved state assignment selection tests. IEEE Transactions on Computers, C-21(12):1443–1449, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672115.

Metze:1975:CDM

[MST^{+}75] G. Metze, D. R. Schertz, Kilin To, G. Whitney, C. R. Kime, and J. D. Russell. Comments on “Derivation of Minimal Complete Sets of Test-Input Sequences Using Boolean Differences”. IEEE Transactions on Computers, C-24 (1):108, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672684. See [DDSM76].

Maki:1970:SAS

[MT70] G. K. Maki and J. H. Tracey. State assignment selection in asynchronous sequential circuits. IEEE Transactions on Computers, C-19 (7):641–644, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671589.

Maki:1971:SAP

[MT71] G. K. Maki and J. H. Tracey. A state assignment procedure for asynchronous sequential circuits. IEEE Transactions on Computers, C-20(6):666–668, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671913.

Marinkovic:1974:AMP

[MT74a] S. B. Marinkovic and Z. Tosic. Algorithm for minimal polarized polynomial form determination. IEEE Transactions on Computers, C-23(12):1313–1315, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672447.

Mukai:1974:MRF

[MT74b] Y. Mukai and Y. Tohma. A method for the realization of fail-safe asynchronous sequential circuits. IEEE Transactions on Computers, C-23 (7):736–739, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672615.

Muroga:1970:ETF

[MTB70] S. Muroga, T. Tsuboi, and C. R. Baugh. Enumeration of threshold functions of eight variables. IEEE Transactions on Computers, C-19(9):818–825, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671639.

Muir:1974:BDB

[Mui74] D. Muir. B74-37 data base management-international computer state of the art report 15. IEEE Transactions on Computers, C-23(8):878–879, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672638.

Mukhopadhyay:1979:HAN

[Muk79] A. Mukhopadhyay. Hardware algorithms for nonnumeric computation. IEEE Transactions on Computers, C-28(6):384–394, June 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675378.

Murry:1970:GAG

[Mur70] Herschell F. Murry. A general approach for generating natural random variables. IEEE Transactions on Computers, C-19(12):1210–1213, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671453.

Muth:1976:NVC

[Mut76] P. Muth. A nine-valued circuit model for test generation. IEEE Transactions on Computers, C-25 (6):630–636, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674663.

Morris:1979:PRS

[MVW79] S. B. Morris, A. Valliere, and R. A. Wisniewski. Processes for random and sequential accessing in dynamic memories. IEEE Transactions on Computers, C-28(3):225– 237, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675323.

Mykkeltveit:1975:GCD

[Myk75] J. Mykkeltveit. Generating and counting the double adjacencies in a pure circulating shift register. IEEE Transactions on Computers, C-24 (3):299–304, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672803.

Newborn:1972:UMB

[NA72] M. M. Newborn and T. F. Arnold.

Universal modules for bounded signal fan-out synchronous sequential circuits. IEEE Transactions on Computers, C-21(1):63– 79, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672025.

Ni:1974:TDD

[NA74] Ming-Duenn Ni and J. K. Aggarwal.

Two-dimensional digital filtering and its error analysis. IEEE Transactions on Computers, C-23(9): 942–954, September 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672651.

Ng:1976:CFF

[NA76a] Y. W. Ng and A. Avizienis. Comments on “Fault Folding for Irredundant and Redundant Combinational Circuits”. IEEE Transactions on Computers, C-25(2): 207, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009237. See [To73].

Ni:1976:EAT

[NA76b] Ming Duenn Ni and J. K. Aggarwal. Error analysis of two-dimensional recursive digital filters employing floating-point arithmetic. IEEE Transactions on Computers, C-25 (7):755–759, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674684.

Nadler:1971:ERR

[Nad71] M. Nadler. Error and reject rates in a hierarchical pattern recognizer. IEEE Transactions on Computers, C-20(12):1598–1601, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671773.

Nair:1979:COA

[Nai79] R. Nair. Comments on “An Optimal Algorithm for Testing Stuck-at Faults in Random Access Memories”. IEEE Transactions on Computers, C-28(3):258– 261, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675331. See [KH77c].

Nakamura:1979:SGM

[Nak79] K. Nakamura. Synthesis of gate-minimum multi-output twolevel negative gate networks.

IEEE Transactions on Computers, C-28(10):768–772, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675244.

Nambiar:1970:SHP

[Nam70] K. K. Nambiar. Solution to Harrison’s problem. IEEE Transactions on Computers, C-19(2):166– 167, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671479.

Naylor:1971:SSI

[Nay71] W. C. Naylor. Some studies in the interactive design of character recognition systems. IEEE Transactions on Computers, C-20(9):1075– 1086, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671999.

Naccarato:1978:RMM

[NC78] D. Naccarato and Y. T. Chien.

Recursive methods for matrix inversion in pattern recognition environments. IEEE Transactions on Computers, C-27(9):866–869, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675206.

Newman:1973:NTD

[ND73] Thomas G. Newman and Hudai Dirilten. A nonlinear transformation for digital picture processing. IEEE Transactions on Computers, C-22(9):869–873, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009186.

Necula:1979:AFE

[Nec79] M.-A. Necula. An algorithm for fast evaluation of time functions. IEEE Transactions on Computers, C-28 (7):506–513, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675396.

Nelson:1974:SEC

[Nel74] T. J. Nelson. A study of the electrochemical cell as a storage element for the memory access gap. IEEE Transactions on Computers, C-23(12):1277–1290, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672440.

Nevatia:1976:LOB

[Nev76] R. Nevatia. Locating object boundaries in textured environments. IEEE Transactions on Computers, C-25(11):1170–1175, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674576.

Newborn:1970:RFD

[New70] M. Newborn. R70-19 frequency of decomposability among machines with a large number of states. IEEE Transactions on Computers, C-19 (7):658, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671599.

Newman:1975:CRP

[New75a] W. M. Newman. Comments on the review of “Principles of Interactive Computer Graphics”. IEEE Transactions on Computers, C-24 (3):336, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672812. See [Bib74].

Newton:1975:SPT

[New75b] T. A. Newton. Some parametric techniques in the analog solution of ordinary differential equations. IEEE Transactions on Computers, C-24(1):1– 8, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672672.

Newman:1976:TGD

[New76] W. M. Newman. Trends in graphic display design. IEEE Transactions on Computers, C-25(12):1321–1325, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674600.

Narendra:1977:BBA

[NF77] P. M. Narendra and K. Fukunaga. A branch and bound algorithm for feature subset selection. IEEE Transactions on Computers, C-26(9): 917–922, September 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674939.

Nakamura:1977:OOE

[NFS77] Y. Nakamura, M. Furuya, and S. Sunohara. An optimal orthogonal expansion for classification of patterns. IEEE Transactions on Computers, C-26(12):1288–1290, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674793.

Nicoud:1971:IAR

[Nic71] Jean-Daniel Nicoud. Iterative arrays for radix conversion. IEEE Transactions on Computers, C-20 (12):1479–1489, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671753.

Nieh:1970:UMS

[Nie70a] T. T. Nieh. On the uniqueness of minimal-state stochastic sequential machines. IEEE Transactions on Computers, C-19(2):164– 166, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671478.

Nievergelt:1970:TRT

[Nie70b] J. Nievergelt. On the time required for timing — the halting problem rephrased. IEEE Transactions on Computers, C-19(5): 458–459, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671540.

Nieminen:1975:NMF

[Nie75] J. Nieminen. A new method of formulating a minimum edge set. IEEE Transactions on Computers, C-24(10):1015–1018, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672708.

Nahi:1977:IBE

[NJ77] N. E. Nahi and M. H. Jahanshahi.

Image boundary estimation. IEEE Transactions on Computers, C-26 (8):772–781, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674915.

Nagaraja:1974:ASL

[NK74] G. Nagaraja and G. Krishna. An algorithm for the solution of linear inequalities. IEEE Transactions on Computers, C-23(4):421– 427, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672550.

Noural:1975:UFS

[NK75] F. Noural and R. S. Kashef. A universal four-state cellular computer. IEEE Transactions on Computers, C-24(8):766–776, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672901.

Neuman:1976:AST

[NM76a] F. Neuman and C. F. Martin. The autocorrelation structure of Tausworthe pseudorandom number generators. IEEE Transactions on Computers, C-25(5):460– 464, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674633.

Neuman:1976:APC

[NM76b] F. Neuman and R. B. Merrick. Autocorrelation peaks in congruential pseudorandom number generators. IEEE Transactions on Computers, C-25(5):457– 460, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674632.

Nordmann:1977:MAC

[NM77] B. J. Nordmann, Jr. and B. H. Mc-Cormick. Modular asynchronous control design. IEEE Transactions on Computers, C-26(3):196– 207, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674808.

Noe:1973:MNR

[NN73] Jerre D. Noe and Gary J. Nutt.

Macro E-nets for representation of parallel systems. IEEE Transactions on Computers, C-22(8):718– 727, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009150.

Noe:1973:RSO

[Noe73] Philip S. Noe. Remarks on the SHR-optimal state assignment procedure. IEEE Transactions on Computers, C-22(9):873–875, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009187.

Noe:1974:MSO

[NR74] P. S. Noe and V. T. Rhyne. A modification to the SHR-optimal state assignment procedure. IEEE Transactions on Computers, C-23 (3):327–329, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672526.

Neumann:1975:ECC

[NR75] P. G. Neumann and T. R. N. Rao. Error-correcting codes for byte-organized arithmetic processors. IEEE Transactions on Computers, C-24(3):226–232, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672796. See comments [Gri78b].

Nitzan:1976:PIA

[NR76a] D. Nitzan and C. A. Rosen. Programmable industrial automation. IEEE Transactions on Computers, C-25(12):1259–1270, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674593.

Noe:1976:OSA

[NR76b] P. S. Noe and V. T. Rhyne.

Optimum state assignment for the D flip-flop. IEEE Transactions on Computers, C-25(3):306– 311, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009258.

Nagel:1977:CDF

[NR77a] R. N. Nagel and A. Rosenfeld.

Computer detection of freehand forgeries. IEEE Transactions on Computers, C-26(9):895–905, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674937.

Ninke:1977:SRB

[NR77b] W. H. Ninke and G. R. Ritchie.

Shift register binary rate multipliers. IEEE Transactions on Computers, C-26(3):276–278, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674819.

Noe:1974:CWM

[NRS74] P. S. Noe, V. T. Rhyne, and S. Suraratrungsi. Comments on “Weighting Method for the Determination of the Irredundant Set of Prime Implicants”. IEEE Transactions on Computers, C-23(6):

646, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672599. See [Bub72].

Niederjohn:1975:CIE

[NS75] R. J. Niederjohn and P. P. Stick. A computer interface for efficient zero-crossing interval measurement. IEEE Transactions on Computers, C-24(3):329– 331, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672810.

Nassimi:1979:BSM

[NS79] D. Nassimi and S. Sahni. Bitonic sort on a mesh-connected parallel computer. IEEE Transactions on Computers, C-28(1):2– 7, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675216.

Nutter:1974:ETP

[NSR74] R. S. Nutter, Jr., R. E. Swartwout, and D. C. Rine. Equivalence and transformations for post multivalued algebras. IEEE Transactions on Computers, C-23(3):294– 300, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672518.

Nordquist:1970:NRS

[NT70] W. R. Nordquist and W. N. Toy. A novel rotate and shift circuit using bidirectional gates. IEEE Transactions on Computers, C-19(9):802– 808, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671636.

Nanya:1978:UST

[NT78] T. Nanya and Y. Tohma. On universal single transition time asynchronous state assignments. IEEE Transactions on Computers, C-27 (8):781–782, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675192.

Nanya:1979:UMS

[NT79] T. Nanya and Y. Tohma. Universal multicode STT state assignments for asynchronous sequential machines. IEEE Transactions on Computers, C-28(11):811– 818, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675261.

Nair:1978:EAT

[NTA78] R. Nair, S. M. Thatte, and J. A.

Abraham. Efficient algorithms for testing semiconductor randomaccess memories. IEEE Transactions on Computers, C-27(6):572– 576, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675150.

Nakamura:1972:MNG

[NTK72] K. Nakamura, N. Tokura, and T. Kasami. Minimal negative gate networks. IEEE Transactions on Computers, C-21(1):5– 11, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672018.

Nussabaumer:1978:FMN

[Nus78] H. J. Nussabaumer. Fast multipliers for number theoretic transforms. IEEE Transactions on Computers, C-27(8):764–765, August 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675187.

Nutt:1977:MBC

[Nut77] G. J. Nutt. Memory and bus conflict in an array processor. IEEE Transactions on Computers, C-26 (6):514–521, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674877.

Niemann:1979:FCA

[NW79] H. Niemann and J. Weiss. A fastconverging algorithm for nonlinear mapping of high-dimensional data to a plane. IEEE Transactions on Computers, C-28(2):142– 147, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675303.

Oberman:1972:FRM

[Obe72] R. M. M. Oberman. A flexible rate multiplier circuit with uniform pulse distribution outputs. IEEE Transactions on Computers, C-21 (8):896–899, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009048.

Oberman:1974:NER

[Obe74] R. M. M. Oberman. A new explanation of the reflected binary code. IEEE Transactions on Computers, C-23(6):641– 642, June 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672597.

Oberman:1976:JKG

[Obe76] R. M. M. Oberman. The J– K gate. IEEE Transactions on Computers, C-25(11):1156–1159, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674572.

OCallaghan:1975:ADN

[O’C75] J. F. O’Callaghan. An alternative definition for “Neighborhood of a point”. IEEE Transactions on Computers, C-24(11):1121–1125, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672737.

OGorman:1976:FPE

[OC76] Frank O’Gorman and M. B. Clowes. Finding picture edges through collinearity of feature points. IEEE Transactions on Computers, C-25 (4):449–456, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674627.

Olsen:1973:RND

[OF73] D. R. Olsen and K. Fukunaga. Representation of nonlinear data surfaces. IEEE Transactions on Computers, C-22(10):915–922, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672211.

Ogus:1974:FTI

[Ogu74] R. C. Ogus. Fault-tolerance of the iterative cell array switch for hybrid redundancy. IEEE Transactions on Computers, C-23(7): 667–681, July 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672608.

Ogus:1975:PCO

[Ogu75] R. C. Ogus. The probability of a correct output from a combinational circuit. IEEE Transactions on Computers, C-24(5):534– 544, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672850.

Ostapko:1979:FAT

[OH79] D. L. Ostapko and Se June Hong.

Fault analysis and test generation for programmable logic arrays (PLA’s). IEEE Transactions on Computers, C-28(9):617– 627, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675429.

Oja:1979:CPU

[Oja79] E. Oja. On the construction of projectors using products of elementary matrices. IEEE Transactions on Computers, C-28(1):65– 66, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675226.

OKeefe:1973:MDS

[O’K73] K. H. O’Keefe. Modularity in design: Skit registers and counters used as system building blocks. IEEE Transactions on Computers, C-22(2):194–197, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672277.

OKeefe:1975:NFB

[O’K75] K. H. O’Keefe. A note on fast base extension for residue number systems with three moduli. IEEE Transactions on Computers, C-24(11):1132–1133, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672740;   https://ieeexplore.
                                        ieee.org/xpl/tocresult.jsp?isnumber=
                                        35091.

Olson:1970:NFI

[Ols70] R. R. Olson. Note on feedforward inverses for linear sequential circuits. IEEE Transactions on Computers, C-19(12):1216–1221, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671455.

Opsahl:1972:OLM

[Ops72] G. I. Opsahl. Optimum logic modules. IEEE Transactions on Computers, C-21(1):90–96, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672029.

Orcutt:1976:IPF

[Orc76] S. E. Orcutt. Implementation of permutation functions in Illiac IVtype computers. IEEE Transactions on Computers, C-25(9):929– 936, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674718.

Onoe:1976:ATS

[OS76] M. Onoe and M. Saito. Automatic threshold setting for the sequential similarity detection algorithm. IEEE Transactions on Computers, C-25(10):1052–1053, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674547.

Osborne:1977:SLL

[Osb77] M. Osborne. Seniority logic:

a logic for a committee machine. IEEE Transactions on Computers, C-26(12):1302–1306, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674798.

Okuda:1976:MCG

[OTK76] Teruo Okuda, Eiichi Tanaka, and Tamotsu Kasai. A method for the correction of garbled words based on the Levenshtein metric. IEEE Transactions on Computers, C-25(2):172–178, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009232.

Osman:1972:UBF

[OW72] Mohamed Y. Osman and C. Dennis Weiss. Universal base functions and modules for realizing arbitrary switching functions. IEEE Transactions on Computers, C-21(9):985– 995, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009075.

OKeefe:1973:RBE

[OW73a] Kenneth H. O’Keefe and John L. Wright. Remarks on base extension for modular arithmetic. IEEE Transactions on Computers, C-22(9):833–835, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009173.

Osman:1973:SLR

[OW73b] M. Y. Osman and C. D. Weiss. Shared logic realizations of dynamically self-checked and faulttolerant logic. IEEE Transactions on Computers, C-22(3):298– 306, March 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672303.

Ostapko:1970:RAS

[OY70] D. L. Ostapko and S. S. Yau. Realization of an arbitrary switching function with a two-level network of threshold and parity elements. IEEE Transactions on Computers, C-19(3):262– 269, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671500.

Ozawa:1978:CKS

[Oza78] K. Ozawa. Classification of the keyhole shaped tombs by template matching method. IEEE Transactions on Computers, C-27(5): 462–467, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675128.

Parhami:1978:DSE

[PA78] B. Parhami and A. Avizienis. Detection of storage errors in mass memories using low-cost arithmetic error codes. IEEE Transactions on Computers, C-27(4):302– 308, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675102.

Paaske:1979:CNR

[Paa79] E. Paaske. Comments on “A New Random-Error-Correction Code”. IEEE Transactions on Computers, C-28(3):255–257, March 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=

1675329. See [En77].

Page:1970:RSL

[Pag70a] C. V. Page. R70-3 on stochastic languages. IEEE Transactions on Computers, C-19(5):463– 464, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671547.

Page:1970:SDP

[Pag70b] C. V. Page. The search for a definition of partition pair for stochastic automata. IEEE Transactions on Computers, C-19(12):1222–1223, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671457.

Pager:1972:SNS

[Pag72] D. Pager. Some notes on speeding up certain loops by software, firmware, and hardware means. IEEE Transactions on Computers, C-21(1):97– 100, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672031.

Paige:1973:SDF

[Pai73] M. R. Paige. Synthesis of diagnosable FET networks. IEEE Transactions on Computers, C-22 (5):513–516, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672347.

Paker:1971:BFP

[Pak71] Y. Paker. A binary floatingpoint resistor. IEEE Transactions on Computers, C-20(1):7– 11, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671667.

Papakonstantinou:1972:SMC

[Pap72] G. K. Papakonstantinou. A synthesis method for cutpoint cellular arrays. IEEE Transactions on Computers, C-21(12):1286–1292, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672093.

Papakonstantinou:1976:CT

[Pap76] G. Papakonstantinou. Cascade transformation. IEEE Transactions on Computers, C-25(1):93– 95, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009211.

Papaioannou:1977:OTG

[Pap77a] S. G. Papaioannou. Optimal test generation in combinational networks by pseudo-Boolean programming. IEEE Transactions on Computers, C-26(6):553– 560, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674880.

Papakonstantinou:1977:MGP

[Pap77b] G. Papakonstantinou. A method to generate the prime cascades of an arbitrary switching function. IEEE Transactions on Computers, C-26 (5):503–506, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674867.

Papachristou:1978:AON

[Pap78] C. A. Papachristou. An algorithm for optimal NAND cascade logic synthesis. IEEE Transactions on Computers, C-27(12):1099–1111, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675012.

Papakonstantinou:1979:MMS

[Pap79] G. Papakonstantinou. Minimization of modulo-2 sum of products. IEEE Transactions on Computers, C-28(2):163–167, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675308.

Parks:1971:BFT

[Par71] H. L. Parks. Batch-fabricated three-dimensional planar coaxial interconnections for microelectronic systems. IEEE Transactions on Computers, C-20(5):504– 511, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671877. See correction [Ano72-39].

Parchmann:1972:NSA

[Par72a] Rainer Parchmann. The number of state assignments for sequential machines. IEEE Transactions on Computers, C-21(6): 613–614, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009022.

Parhami:1972:SAP

[Par72b] Behrooz Parhami. Stochastic automata and the problems of reliability in sequential machines. IEEE Transactions on Computers, C-21 (4):388–391, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008981.

Parker:1979:CMH

[Par79] D. S. Parker, Jr. Combinatorial merging and Huffman’s algorithm. IEEE Transactions on Computers, C-28(5):365–367, May 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675367.

Patel:1971:GRL

[Pat71] A. R. Patel. Generation of right-linear grammars from regular expressions. IEEE Transactions on Computers, C-20(4):472– 473, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671864.

Patt:1972:MST

[Pat72] Yale N. Patt. Minimum search tree structures for data partitioned into pages. IEEE Transactions on Computers, C-21(9):961–967, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009072.

Patt:1973:ONO

[Pat73] Y. N. Patt. Optimal and near-optimal universal logic modules with interconnected external terminals. IEEE Transactions on Computers, C-22(10):903– 907, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672209.

Patterson:1976:SSM

[Pat76] D. A. Patterson. Strum: Structured microprogram development system for correct firmware. IEEE Transactions on Computers, C-25(10): 974–985, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674538.

Pavlidis:1973:WST

[Pav73] Theodosios Pavlidis. Waveform segmentation through functional approximation. IEEE Transactions on Computers, C-22(7):689– 697, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009136.

Pavlidis:1974:BPC

[Pav74] T. Pavlidis. B74-21 patern classification and scene analysis. IEEE Transactions on Computers, C-23 (2):223, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672503.

Pavlidis:1975:BNM

[Pav75a] T. Pavlidis. B75-3 numerical methods. IEEE Transactions on Computers, C-24(2):222, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672787.

Pavlidis:1975:OPP

[Pav75b] T. Pavlidis. Optimal piecewise polynomial L^{2} approximation of functions of one and two variables. IEEE Transactions on Computers, C-24 (1):98–102, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672681.

Pavlidis:1977:PAN

[Pav77] T. Pavlidis. Polygonal approximations by Newton’s method. IEEE Transactions on Computers, C-26 (8):800–807, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674918.

Pawlak:1971:ACN

[Paw71] Z. Pawlak. Another comment on “Negative Radix Conversion”. IEEE Transactions on Computers, C-20(5):587, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671894. See comments [Paw71, Wad71, Zoh70].

Paz:1970:RFE

[Paz70a] A. Paz. R70-33 fuzzy events realized by finite probabilistic automata. IEEE Transactions on Computers, C-19(9):858, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671654.

Paz:1970:RDR

[Paz70b] A. Paz. R70-45 on decompositions of regular events. IEEE Transactions on Computers, C-19(10): 990, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671411.

Paz:1970:RES

[Paz70c] A. Paz. Regular events in stochastic sequential machines. IEEE Transactions on Computers, C-19 (5):456–457, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671538.

Paz:1971:WDS

[Paz71] A. Paz. Whirl decomposition of stochastic systems. IEEE Transactions on Computers, C-20(10): 1208–1211, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671700.

Plantz:1971:AON

[PB71] A. R. Plantz and M. Berman. Adoption of the octal number system. IEEE Transactions on Computers, C-20(5):593– 598, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671900.

Polge:1976:EFF

[PB76a] R. J. Polge and B. K. Bhagavan. Efficient Fast Fourier Transform programs for arbitrary factors with one step loop unscrambling. IEEE Transactions on Computers, C-25 (5):534–539, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674643.

Prusinkiewicz:1976:DTE

[PB76b] P. Prusinkiewicz and S. Budkowski. A double track error-correction code for magnetic tape. IEEE Transactions on Computers, C-25 (6):642–645, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674665.

Polge:1974:FCA

[PBC74] R. J. Polge, B. K. Bhagavan, and J. M. Carswell. Fast computational algorithms for bit reversal. IEEE Transactions on Computers, C-23 (1):1–9, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672364.

Preston:1972:DOS

[PC72] K. Preston, Jr. and J. R. Carvalko. On determining optimum simple Golay marking transformations for binary image processing. IEEE Transactions on Computers, C-21(12):1430–1433, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672112.

Prather:1978:RBE

[PC78] R. E. Prather and H. T. Casstevens.

Realization of Boolean expressions by atomic digraphs. IEEE Transactions on Computers, C-27(8):681– 688, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675174.

Patrick:1970:DDE

[PCM70] E. A. Patrick, J. P. Costello, and F. C. Monds. Decision-directed estimation of a two-class decision boundary. IEEE Transactions on Computers, C-19(3):197– 205, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671490.

Pratt:1977:FCT

[PD77] W. K. Pratt and F. Davarian. Fast computational techniques for pseudoinverse and Wiener image restoration. IEEE Transactions on Computers, C-26(6):571– 580, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674882.

Plavsic:1979:SEB

[PD79] V. M. Plavsic and P. E. Danielsson. Sequential evaluation of Boolean functions. IEEE Transactions on Computers, C-28(12): 879–887, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675278.

Pearl:1975:ODM

[Pea75] J. Pearl. Optimal dyadic models of time-invariant systems. IEEE Transactions on Computers, C-24 (6):598–603, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672865.

Pearl:1976:MVE

[Pea76] J. Pearl. Memory versus error characteristics for inexact representations of linear orders. IEEE Transactions on Computers, C-25(9): 922–928, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674717.

Pease:1977:IBC

[Pea77] M. C. Pease III. The indirect binary n-cube microprocessor array. IEEE Transactions on Computers, C-26 (5):458–473, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674863.

Pechoucek:1976:ART

[Pec76] Miroslav Pechoucek. Anomalous response times of input synchronizers. IEEE Transactions on Computers, C-25(2):133–139, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009227.

Perlman:1970:DSL

[Per70] M. Perlman. The decomposition of the states of a linear feedback shift register into cycles of equal length. IEEE Transactions on Computers, C-19(11):1029–1035, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671421.

Perkins:1978:MBV

[Per78] W. A. Perkins. A modelbased vision system for industrial parts. IEEE Transactions on Computers, C-27(2):126–143, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675046.

Pezaris:1971:NBB

[Pez71a] S. D. Pezaris. A 40-ns 17-bit by 17-bit array multiplier. IEEE Transactions on Computers, C-20 (4):442–447, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671854.

Pezaris:1971:BBA

[Pez71b] S. D. Pezaris. A 40ns 17-bit by 17bit array multiplier. IEEE Transactions on Computers, C-20(4):442– 447 (??), April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pavlidis:1972:STW

[PF72] Theodosios Pavlidis and Geng Seng Fang. A segmentation technique for waveform classification. IEEE Transactions on Computers, C-21 (8):901–904, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009050.

Pfleeger:1973:SRI

[Pfl73] C. P. Pfleeger. State reduction in incompletely specified finitestate machines. IEEE Transactions on Computers, C-22(12):1099–1102, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672248.

Pearce:1975:AAM

[PFL75] R. C. Pearce, J. A. Field, and W. D. Little. Asynchronous arbiter module. IEEE Transactions on Computers, C-24(9):931– 932, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672932.

Prasad:1975:MFD

[PG75] B. A. Prasad and F. G. Gray. Multiple fault detection in arrays of combinational cells. IEEE Transactions on Computers, C-24(8):794– 802, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672904.

Podraza:1970:EMP

[PGS70] G. V. Podraza, R. S. Gregg, Jr., and J. R. Slager. Efficient MSI partitioning for a digital computer. IEEE Transactions on Computers, C-19(11):1020–1028, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671420.

Polge:1973:NTC

[PH73] R. J. Polge and R. D. Hays. Numerical technique for the convolution of piecewise polynomial functions. IEEE Transactions on Computers, C-22(11):970–975, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672226.

Pavlidis:1974:SPC

[PH74] T. Pavlidis and S. L. Horowitz. Segmentation of plane curves. IEEE Transactions on Computers, C-23 (8):860–870, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672634.

Philokyprou:1977:PFT

[PHG77] G. Philokyprou, C. Halatsis, and N. Gaitanis. On the problem of finding d, t, r, and j type partition pairs of a sequential machine. IEEE Transactions on Computers, C-26 (3):294–297, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674824.

Picchia:1974:NAR

[Pic74] W. Del Picchia. A numerical algorithm for the resolution of Boolean equations. IEEE Transactions on Computers, C-23(9):983– 986, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672660.

Pinter:1971:STF

[Pin71] C. C. Pinter. On simplifying truth functions: a preliminary reduction of coreless formulas. IEEE Transactions on Computers, C-20(8):938– 941, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671969.

Plauger:1973:BHL

[Pla73] P. J. Plauger. B73-4 high level languages — international computer state of the art report. IEEE Transactions on Computers, C-22(2): 221–222, February 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672285.

Pless:1977:ESC

[Ple77] Vera S. Pless. Encryption schemes for computer confidentiality. IEEE Transactions on Computers, C-26(11):1133–1136, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674759.

Plummer:1972:AA

[Plu72] W. W. Plummer. Asynchronous arbiters. IEEE Transactions on Computers, C-21(1):37–42, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672022.

Preparata:1971:DRR

[PM71] F. P. Preparata and D. E. Muller. On the delay required to realize Boolean functions. IEEE Transactions on Computers, C-20(4):459– 461, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671859.

Patterson:1974:FSA

[PM74] W. W. Patterson and G. Metze. A fail-safe asynchronous sequential machine. IEEE Transactions on Computers, C-23(4):369– 374, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672545.

Parker:1975:ALC

[PM75a] K. P. Parker and E. J. Mc-Cluskey. Analysis of logic circuits with faults using input signal probabilities. IEEE Transactions on Computers, C-24(5): 573–578, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672857.

Parker:1975:PTG

[PM75b] K. P. Parker and E. J. McCluskey.

Probabilistic treatment of general combinational networks. IEEE Transactions on Computers, C-24 (6):668–670, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672872.

Parrish:1976:ICC

[PM76a] E. A. Parrish, Jr. and E. S.

McVey. Implications of chargecoupled devices for pattern recognition. IEEE Transactions on Computers, C-25(11):1146–1152, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674570.

Preparata:1976:EPE

[PM76b] F. P. Preparata and D. E. Muller.

Efficient parallel evaluation of Boolean expressions. IEEE Transactions on Computers, C-25(5): 548–549, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674647.

Page:1977:PAR

[PM77a] E. W. Page and P. N. Marinos. Programmable array realizations of synchronous sequential machines. IEEE Transactions on Computers, C-26(8):811– 818, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674920.

Payne:1977:ACO

[PM77b] H. J. Payne and W. S. Meisel. An algorithm for constructing optimal binary decision trees. IEEE Transactions on Computers, C-26(9):905– 916, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674938.

Parker:1978:SCO

[PM78a] K. P. Parker and E. J. McCluskey. Sequential circuit output probabilities from regular expressions. IEEE Transactions on Computers, C-27 (3):222–231, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675075.

Pearce:1978:ASR

[PM78b] R. C. Pearce and J. C. Majithia. Analysis of a shared resource MIMD computer organization. IEEE Transactions on Computers, C-27 (1):64–67, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674953.

Preparata:1977:RDB

[PMB77] F. P. Preparata, D. E. Mulller, and A. B. Barak. Reduction of depth of Boolean networks with a fan-in constraint. IEEE Transactions on Computers, C-26(5): 474–479, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674864.

Poe:1974:BAA

[Poe74] A. T. Poe. B74-11 abstract automata. IEEE Transactions on Computers, C-23(2):214–216, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672493.

Pols:1971:RTR

[Pol71] L. C. W. Pols. Real-time recognition of spoken words. IEEE Transactions on Computers, C-20(9): 972–978, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671984.

Pradhan:1975:RML

[PP75] D. K. Pradhan and A. M. Patel. Reed–Muller like canonic forms for multivalued functions. IEEE Transactions on Computers, C-24(2): 206–210, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672779.

Putzolu:1971:HAT

[PR71] G. R. Putzolu and J. P. Roth. A heuristic algorithm for the testing of asynchronous circuits. IEEE Transactions on Computers, C-20 (6):639–647, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671908.

Pradhan:1972:ECT

[PR72] D. K. Pradhan and S. M. Reddy. Error-control techniques for logic processors. IEEE Transactions on Computers, C-21(12):1331–1336, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672097.

Pradhan:1973:FTA

[PR73] Dhiraj K. Pradhan and Sudhakar M. Reddy. Fault-tolerant asynchronous networks. IEEE Transactions on Computers, C-22 (7):662–669, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009132. See comments [SM75a].

Paz:1974:LAA

[PR74a] A. Paz and M. Rabinovitz.

Linear automata approximation problem. IEEE Transactions on Computers, C-23(3):249–255, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672512.

Pradhan:1974:DTL

[PR74b] D. K. Pradhan and S. M.

Reddy. Design of two-level faulttolerant networks. IEEE Transactions on Computers, C-23(1):41– 48, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672368.

Pradhan:1975:AR

[PR75] D. K. Pradhan and S. M. Reddy.

Authors’ reply. IEEE Transactions on Computers, C-24(7):758– 759, July 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672893.

Pradhan:1976:TC

[PR76] D. K. Pradhan and S. M.

Reddy. Techniques to construct.

IEEE Transactions on Computers, C-25(9):945–949, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674720.

Panda:1978:ISP

[PR78] D. P. Panda and A. Rosenfeld. Image segmentation by pixel classification in (gray level, edge value) space. IEEE Transactions on Computers, C-27(9):875–879, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675208.

Prather:1971:APP

[Pra71a] R. E. Prather. An algebraic proof of the Paull–Unger theorem. IEEE Transactions on Computers, C-20 (5):578–580, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671887.

Pratt:1971:AR

[Pra71b] W. K. Pratt. Author’s reply. IEEE Transactions on Computers, C-20 (5):588, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671897.

Pratt:1972:GWF

[Pra72] W. K. Pratt. Generalized Wiener filtering computation techniques. IEEE Transactions on Computers, C-21(7):636–641, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672160.

Pradhan:1974:FTC

[Pra74] D. K. Pradhan. Fault-tolerant carry-save adders. IEEE Transactions on Computers, C-23(12): 1320–1322, December 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672450.

Pradhan:1978:ASA

[Pra78a] D. K. Pradhan. Asynchronous state assignments with unateness properties and fault-secure design. IEEE Transactions on Computers, C-27 (5):396–404, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675118.

Pradhan:1978:FTA

[Pra78b] D. K. Pradhan. Fault-tolerant asynchronous networks using readonly memories. IEEE Transactions on Computers, C-27(7):674– 679, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675168.

Pradhan:1978:TGS

[Pra78c] D. K. Pradhan. A theory of Galois switching functions. IEEE Transactions on Computers, C-27(3):239– 248, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675077.

Pradhan:1978:UTS

[Pra78d] D. K. Pradhan. Universal test sets for multiple fault detection in AND–EXOR arrays. IEEE Transactions on Computers, C-27(2): 181–187, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675057.

Pardee:1971:GRT

[PRD71] S. Pardee, P. E. Rosenfeld, and P. G. Dowd. G101 — a remote time share terminal with graphic output capabilities. IEEE Transactions on Computers, C-20(8):878– 881, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671957.

Preparata:1970:RND

[Pre70] F. P. Preparata. R70-28 A note on definite stochastic sequential machines. IEEE Transactions on Computers, C-19(8):764– 765, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671626.

Preparata:1971:DUB

[Pre71a] F. P. Preparata. On the design of universal Boolean functions. IEEE Transactions on Computers, C-20 (4):418–423, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671850.

Preston:1971:FEG

[Pre71b] K. Preston, Jr. Feature extraction by Golay hexagonal pattern transforms. IEEE Transactions on Computers, C-20(9):1007–1014, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671989.

Preparata:1972:ULM

[Pre72] Franco P. Preparata. Universal logic modules of a new type. IEEE Transactions on Computers, C-21 (6):585–588, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009012.

Preparata:1978:NPS

[Pre78a] F. P. Preparata. New parallelsorting schemes. IEEE Transactions on Computers, C-27(7):669– 673, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675167.

Preston:1978:NAT

[Pre78b] F. S. Preston. A new algorithm for the tangent. IEEE Transactions on Computers, C-27(2): 167, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675052.

Patrick:1971:IUP

[PS71] E. A. Patrick and L. Y. L. Shen. Interactive use of problem knowledge for clustering and decision making. IEEE Transactions on Computers, C-20(2):216– 222, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671810.

Pavlidis:1978:ACA

[PS78] T. Pavlidis and K. Steiglitz. The automatic counting of asbestos fibers in air samples. IEEE Transactions on Computers, C-27(3):258– 261, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675082.

Pu:1972:CSF

[Pu72] A. T. Pu. On complete systems and finite automata. IEEE Transactions on Computers, C-21(10): 1109–1113, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672050.

Pangratz:1979:PRN

[PW79a] H. Pangratz and H. Weinrichter.

Pseudo-random number generator based on binary and quinary maximal-length sequences. IEEE Transactions on Computers, C-28(9):637–642, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675431.

Patil:1979:PLA

[PW79b] S. S. Patil and T. A. Welch.

A programmable logic approach for VLSI. IEEE Transactions on Computers, C-28(9):594–601, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675426.

Quan:1974:NHS

[Qua74] A. Y. C. Quan. A note on high-speed generation of maximal length sequences. IEEE Transactions on Computers, C-23(2):201– 203, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672482.

Rauscher:1978:DPO

[RA78] T. G. Rauscher and A. K. Agrawala. Dynamic problemoriented redefinition of computer architecture via microprogramming. IEEE Transactions on Computers, C-27(11):1006–1014, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674990.

Rader:1972:DCM

[Rad72] C. M. Rader. Discrete convolutions via Mersenne transforms. IEEE Transactions on Computers, C-21(12):1269–1273, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672090. See correction [Rad73].

Rader:1973:CDC

[Rad73] Charles M. Rader. Correction to “Discrete Convolutions via Mersenne Transforms”. IEEE Transactions on Computers, C-22 (7):703, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009140. See [Rad72].

Ramamoorthy:1971:FTC

[Ram71] C. V. Ramamoorthy. Fault-tolerant computing: An introduction and an overview. IEEE Transactions on Computers, C-20(11):1241–1244, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671715.

Ramamoorthy:1973:BIC

[Ram73a] C. V. Ramamoorthy. B73-9 introduction to computer organization and data structures. IEEE Transactions on Computers, C-22 (5):545–546, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672356.

Ramanujam:1973:DPN

[Ram73b] H. R. Ramanujam. Decomposition of permutation networks. IEEE Transactions on Computers, C-22 (7):639–643, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009129.

Ramapriyan:1975:GEA

[Ram75] H. K. Ramapriyan. A generalization of Eklundh’s algorithm for transposing large matrices. IEEE Transactions on Computers, C-24(12):1221–1226, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672760. See [Ekl72].

Ramamoorthy:1976:CSE

[Ram76a] C. V. Ramamoorthy. Computer science and engineering education. IEEE Transactions on Computers, C-25(12):1200–1206, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674588.

Ramapriyan:1976:MAS

[Ram76b] H. K. Ramapriyan. A multilevel approach to sequential detection of pictorial features. IEEE Transactions on Computers, C-25(1):66– 78, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009206.

Ramot:1976:NC

[Ram76c] J. Ramot. Nonparametric curves. IEEE Transactions on Computers, C-25(1):103–104, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009218.

Ramapriyan:1977:ACM

[Ram77a] H. K. Ramapriyan. An algorithm for constrained maximization of the trace of a matrix. IEEE Transactions on Computers, C-26(12):1306–1308, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674799.

Ramapriyan:1977:DHG

[Ram77b] H. K. Ramapriyan. Data handling for the geometric correction of large images. IEEE Transactions on Computers, C-26(11):1163–1167, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674769.

Rao:1970:BEC

[Rao70] T. R. N. Rao. Biresidue error-correcting codes for computer arithmetic. IEEE Transactions on Computers, C-19(5): 398–402, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671530.

Rao:1972:ECA

[Rao72] Thammavarapu R. N. Rao. Error correction in adders using systematic subcodes. IEEE Transactions on Computers, C-21(3):254– 259, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008947.

Rao:1978:ARC

[Rao78] T. R. N. Rao. Author’s reply and comments on the arithmetic codes generated by A = (r − 1)p^{2}. IEEE Transactions on Computers, C-27 (2):190, February 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675062.

Ratz:1976:MTT

[Rat76] H. C. Ratz. Matrix transformations for N-tuple analysis of binary patterns. IEEE Transactions on Computers, C-25(5):542– 546, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674645.

Rau:1979:IMB

[Rau79a] B. R. Rau. Interleaved memory bandwidth in a model of a multiprocessor computer system. IEEE Transactions on Computers, C-28(9):678–681, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675436.

Rau:1979:PBP

[Rau79b] B. R. Rau. Program behavior and the performance of interleaved memories. IEEE Transactions on Computers, C-28(3):191– 199, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675319.

Ravi:1972:BII

[Rav72] C. V. Ravi. On the bandwidth and interference in interleaved memory systems. IEEE Transactions on Computers, C-21(8):899– 901, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009049.

Ricardi:1972:RTE

[RB72] L. J. Ricardi and M. L. Burrows. A recurrence technique for expanding a function in spherical harmonics. IEEE Transactions on Computers, C-21(6):583– 585, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009011.

Rao:1975:MIS

[RB75] C. V. S. Rao and N. N. Biswas.

Minimization of incompletely specified sequential machines. IEEE Transactions on Computers, C-24 (11):1089–1100, November 1975.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672730.

Rao:1976:FCC

[RB76] C. V. S. Rao and N. N. Biswas. Further comments on “Closure Partition Method for Minimizing Incomplete Sequential Machines”. IEEE Transactions on Computers, C-25 (7):767–768, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674690. See [YB75].

Rao:1978:FCP

[RB78] C. V. Kameswara Rao and K. Balck. Finding the core point in a fingerprint. IEEE Transactions on Computers, C-27(1):77– 81, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674955.

Reed:1970:CTF

[RC70] I. S. Reed and A. C. L. Chiang. Coding techniques for failuretolerant counters. IEEE Transactions on Computers, C-19(11): 1035–1038, November 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671422.

Ramamoorthy:1971:SSP

[RC71a] C. V. Ramamoorthy and Lih-Chung Chang. System segmentation for the parallel diagnosis of computers. IEEE Transactions on Computers, C-20(3):261– 270, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671823.

Ramamoorthy:1971:EOL

[RC71b] C. V. Ramamoorthy and Yeh-Hao Chin. An efficient organization or large frequency-dependent files for binary searching. IEEE Transactions on Computers, C-20(10): 1178–1187, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671695.

Rigas:1971:PAC

[RC71c] H. B. Rigas and D. J. Coombs.

Patch: Analog computer patching from a digital simulation language. IEEE Transactions on Computers, C-20(10):1140–1146, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671691.

Ramamoorthy:1972:SMT

[RC72] C. V. Ramamoorthy and L. C.

Chang. System modeling and testing procedures for microdiagnostics. IEEE Transactions on Computers, C-21(11):1169–1183, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672067.

Ryan:1974:PMS

[RC74] T. A. Ryan, Jr. and E. G. Coffman, Jr. A problem in multiprogrammed storage allocation. IEEE Transactions on Computers, C-23(11):1116–1122, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672411.

Ramamoorthy:1972:OSS

[RCG72] C. V. Ramamoorthy, K. M. Chandy, and Mario J. Gonzalez. Optimal scheduling strategies in a multiprocessor system. IEEE Transactions on Computers, C-21(2):137– 146, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008918.

Riseman:1971:CWR

[RE71] E. M. Riseman and R. W. Ehrich. Contextual word recognition using binary digrams. IEEE Transactions on Computers, C-20(4):397– 403, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671848.

Redfield:1971:SMP

[Red71a] S. R. Redfield. A study in microprogrammed processors: a medium sized microprogrammed processor. IEEE Transactions on Computers, C-20(7):743–750, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671936.

Redinpo:1971:ITW

[Red71b] G. R. Redinpo. An implementation technique for Walsh functions. IEEE Transactions on Computers, C-20(6):706–707, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671925.

Reddy:1972:DPF

[Red72a] S. M. Reddy. A design procedure for fault-locatable switching circuits. IEEE Transactions on Computers, C-21(12):1421–1426, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672110.

Reddy:1972:ETR

[Red72b] S. M. Reddy. Easily testable realizations for logic functions. IEEE Transactions on Computers, C-21(11):1183–1188, November 1972. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672068.

Reddy:1973:CTS

[Red73a] S. M. Reddy. Complete test sets for logic functions. IEEE Transactions on Computers, C-22(11):1016–1020, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672231.

Redinbo:1973:BSC

[Red73b] G. R. Redinbo. Bounds on the sequency content of the circular functions. IEEE Transactions on Computers, C-22(4):426– 427, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672327.

Reddy:1974:NSC

[Red74] S. M. Reddy. Note on selfchecking checkers. IEEE Transactions on Computers, C-23(10): 1100–1102, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672404.

Reddy:1977:CMF

[Red77a] S. M. Reddy. Comments on “Minimal Fault Tests for Combinational Networks”. IEEE Transactions on Computers, C-26(3):318– 319, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674835. See [FD74b].

Reddy:1977:NTL

[Red77b] S. M. Reddy. A note on testing logic circuits by transition counting. IEEE Transactions on Computers, C-26(3):313– 314, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674831.

Reddy:1978:CLC

[Red78a] S. M. Reddy. A class of linear codes for error control in byte-per-card organized digital systems. IEEE Transactions on Computers, C-27 (5):455–459, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675126.

Reddy:1978:FTC

[Red78b] S. M. Reddy. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-27 (6):481–482, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675137.

Redinbo:1979:FFA

[Red79] G. R. Redinbo. Finite field arithmetic on an array processor. IEEE Transactions on Computers, C-28 (7):461–471, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675390.

Reeves:1970:CTL

[Ree70] R. Reeves. Comment on “A Transform for Logic Networks”. IEEE Transactions on Computers, C-19 (6):559–560, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671570.

Reddy:1976:HSU

[REFN76] D. R. Reddy, L. D. Erman, R. D. Fennell, and R. B. Neely. The Hearsay-I speech understanding system: An example of the recognition process. IEEE Transactions on Computers, C-25(4):422– 431, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674624.

Regis:1973:MQM

[Reg73] Renaud C. Regis. Multiserver queueing models of multiprocessing systems. IEEE Transactions on Computers, C-22(8):736– 745, August 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009152.

Reiliy:1974:BID

[Rei74a] K. D. Reiliy. B74-42 introduction to discrete structures. IEEE Transactions on Computers, C-23(11):1215–1216, November 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672427.

Reilly:1974:BMB

[Rei74b] K. D. Reilly. B74-41 the metaphorical brain, an introduction to cybernetics as artificial intelligence and brain theory. IEEE Transactions on Computers, C-23(10): 1103–1104, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672405.

Reiter:1976:SGD

[Rei76] R. Reiter. A semantically guided deductive system for automatic theorem proving. IEEE Transactions on Computers, C-25(4):328– 334, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674613.

Rejchrt:1972:SFG

[Rej72] Vladimir J. Rejchrt. Signal flow graph and a Fortran program for Haar–Fourier transform. IEEE Transactions on Computers, C-21(9):1026–1027, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009086.

Remy:1972:BPP

[Rem72] E. H. Remy. B72-10 PL/I programming in technological applications. IEEE Transactions on Computers, C-21(5):517, May 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672153.

Remler:1974:PRC

[Rem74] M. P. Remler. Pattern recognition by convolution polynomial. IEEE Transactions on Computers, C-23 (5):528–530, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672568.

Reusch:1972:NMC

[Reu72] B. Reusch. Note on minimal congruences on transition graphs. IEEE Transactions on Computers, C-21 (1):96–97, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672030.

Reusch:1975:GPI

[Reu75] B. Reusch. Generation of prime implicants from subfunctions and a unifying approach to the covering problem. IEEE Transactions on Computers, C-24(9):924– 930, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672931.

Reynolds:1970:GRE

[Rey70] B. G. Reynolds. General repetitive events and machines. IEEE Transactions on Computers, C-19(2): 167–169, February 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671480.

Riseman:1972:IPP

[RF72] E. M. Riseman and C. C. Foster. The inhibition of potential parallelism by conditional jumps. IEEE Transactions on Computers, C-21(12):1405–1411, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672107.

Reid:1975:CIF

[RF75] R. J. Reid and J. S. Frame. Convergence in iteratively formed correlation matrix memories. IEEE Transactions on Computers, C-24 (8):827–830, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672907.

Reddi:1978:RCS

[RF78] S. S. Reddi and E. A. Feustel. A restructurable computer system. IEEE Transactions on Computers, C-27(1):1–20, January 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674947.

Ramamoorthy:1976:SPP

[RFL76] C. V. Ramamoorthy, T. F. Fox, and H. F. Li. Scheduling parallel processable tasks for a uniprocessor. IEEE Transactions on Computers, C-25(5):485– 495, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674636. See corrections [RL78].

Ramamoorthy:1972:SPI

[RGK72] C. V. Ramamoorthy, James R. Goodman, and K. H. Kim. Some properties of iterative squarerooting methods using high-speed multiplication. IEEE Transactions on Computers, C-21(8):837– 847, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009039.

Rowland:1971:VAD

[RH71] J. R. Rowland and W. M. Holmes.

A variational approach to digital integration. IEEE Transactions on Computers, C-20(8):894– 900, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671960.

Rowland:1973:DPP

[RH73] J. R. Rowland and W. M.

Holmes. Dynamic programming for the partitioning of numerical simulations with hybrid computer applications. IEEE Transactions on Computers, C-22(5): 516–523, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672348.

Riseman:1974:CPS

[RH74] E. M. Riseman and A. R. Hanson. A contextual postprocessing system for error correction using binary n-grams. IEEE Transactions on Computers, C-23(5): 480–493, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672564.

Ramamoorthy:1975:RAS

[RH75a] C. V. Ramamoorthy and Yih-Wu Han. Reliability analysis of systems with concurrent error detection. IEEE Transactions on Computers, C-24(9):868– 878, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672925.

Robinson:1975:ETT

[RH75b] J. P. Robinson and C. W. Hoffner II. Easily tested three-level gate networks for T or more of N symmetric functions. IEEE Transactions on Computers, C-24(3):331– 335, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672811.

Rabbat:1976:CIS

[RHR76a] N. B. Rabbat, S. Q. A. M. A. Hossain, and W. D. Ryan. Corrections to “Improved Simulation of Pulse Propagation Between Digital Circuits”. IEEE Transactions on Computers, C-25(10): 1056, October 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674550. See [RHR76b].

Rabbat:1976:ISP

[RHR76b] N. B. Rabbat, S. Q. A. M. A. Hossain, and W. D. Ryan. Improved simulation of pulse propagation between digital circuits. IEEE Transactions on Computers, C-25 (3):221–228, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009248. See corrections [RHR76a].

Rhyne:1970:SBD

[Rhy70] V. T. Rhyne. Serial binaryto-decimal and decimal-to-binary conversion. IEEE Transactions on Computers, C-19(9):808–812, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671637. See comment [Hor71a].

Rhyne:1971:SPN

[Rhy71] V. T. Rhyne. A simple postcorrection for nonrestoring division. IEEE Transactions on Computers, C-20(2):213–214, February 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671808.

Ricco:1972:CSC

[Ric72] R. Ricco. Comments on “Single-Channel and Multichannel Finite-State Machines”. IEEE Transactions on Computers, C-21(11): 1247, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672078. See [Gil70a].

Rice:1974:BCC

[Ric74a] J. R. Rice. B74-1 complexity of computer computations.

IEEE Transactions on Computers, C-23(1):109–110, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672384.

Richardson:1974:RWS

[Ric74b] R. L. Richardson. Reciprocal Walsh series. IEEE Transactions on Computers, C-23(9):964–966, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672653.

Rips:1972:BCA

[Rip72] E. M. Rips. B72-17 computer analysis of circuits. IEEE Transactions on Computers, C-21(10): 1133–1135, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672057.

Ristenbatt:1974:BCC

[Ris74] M. P. Ristenbatt. B74-45 computer-communication networks. IEEE Transactions on Computers, C-23(12):1329–1330, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672459.

Rivest:1977:NFM

[Riv77] R. L. Rivest. The necessity of feedback in minimal monotone combinational circuits. IEEE Transactions on Computers, C-26(6): 606–607, June 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674886.

Read:1972:AGT

[RJ72] J. S. Read and S. N. Jayaramamurthy. Automatic generation of texture feature detectors. IEEE Transactions on Computers, C-21 (7):803–812, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672178.

Rosenfeld:1973:ADD

[RJ73] Azriel Rosenfeld and Emily Johnston. Angle detection on digital curves. IEEE Transactions on Computers, C-22(9):875–878, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009188.

Russell:1971:SFF

[RK71] J. D. Russell and C. R. Kime.

Structural factors in the fault diagnosis of combinational networks. IEEE Transactions on Computers, C-20(11):1276–1285, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671721.

Russell:1975:SFDa

[RK75a] J. D. Russell and C. R. Kime. System fault diagnosis: Closure and diagnosability with repair. IEEE Transactions on Computers, C-24(11):1078–1089, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672729.

Russell:1975:SFDb

[RK75b] J. D. Russell and C. R. Kime. System fault diagnosis: Masking, exposure, and diagnosability without repair. IEEE Transactions on Computers, C-24(12):1155–1161, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672751.

Rosenfeld:1972:CHL

[RL72] Azriel Rosenfeld and Yung H. Lee. A clustering heuristic for linedrawing analysis. IEEE Transactions on Computers, C-21(8):904– 911, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009051.

Ramamoorthy:1978:CSP

[RL78] C. V. Ramamoorthy and H. F.

Li. Corrections to “Scheduling Parallel Processable Tasks for a Uniprocessor”. IEEE Transactions on Computers, C-27(3): 288, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675094. See [RFL76].

Ramamoorthy:1971:CDU

[RM71] C. V. Ramamoorthy and W. Mayeda.

Computer diagnosis using the blocking gate approach. IEEE Transactions on Computers, C-20 (11):1294–1299, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671723.

Reynolds:1978:FDC

[RM78] D. A. Reynolds and G. Metze.

Fault detection capabilities of alternating logic. IEEE Transactions on Computers, C-27(12):1093–1098, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675011.

Rhyne:1977:NDS

[RN77] V. T. Rhyne and P. S. Noe. On the number of distinct state assignments for a sequential machine. IEEE Transactions on Computers, C-26(1):73–75, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009276.

Rhyne:1977:NTF

[RNMP77] V. T. Rhyne, P. S. Noe, M. H. Mckinney, and U. W. Pooch. A new technique for the fast minimization of switching functions. IEEE Transactions on Computers, C-26 (8):757–764, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674913.

Rao:1975:IDP

[RNR75] K. R. Rao, M. A. Narasimhan, and K. Revuluri. Image data processing by Hadamard–Haar transform. IEEE Transactions on Computers, C-24(9):888–896, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672927.

Robertson:1970:CBM

[Rob70a] J. E. Robertson. The correspondence between methods of digital division and multiplier recoding procedures. IEEE Transactions on Computers, C-19(8):692–701, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Robertson.

pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671611.

Robinson:1970:RPS

[Rob70b] J. P. Robinson. R70-24 A procedure for selecting diagnostic tests. IEEE Transactions on Computers, C-19(7):660, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671604.

Robertson:1979:MBO

[Rob79a] E. L. Robertson. Microcode bit optimization is NP-complete. IEEE Transactions on Computers, C-28 (4):316–319, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675353.

Robinson:1979:OGR

[Rob79b] J. P. Robinson. Optimum Golomb rulers. IEEE Transactions on Computers, C-28(12):943–944, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675287.

Rochkind:1970:TFS

[Roc70] M. M. Rochkind. Transferable FORTRAN subroutine for rapid extended sorting. IEEE Transactions on Computers, C-19(3):270– 272, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671501.

Roesser:1978:TDM

[Roe78] R. P. Roesser. Two-dimensional microprocessor pipelines for image processing. IEEE Transactions on Computers, C-27(2):144– 156, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675047.

Rom:1978:DRB

[Rom78] R. Rom. Distribution of runs in binary words. IEEE Transactions on Computers, C-27(11):1087–1089, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675003.

Roochvarg:1973:BTM

[Roo73] A. C. Roochvarg. B73-7 theory of machines and computations. IEEE Transactions on Computers, C-22 (4):430, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672332.

Rossol:1971:GHC

[Ros71] L. Rossol. Generation of halftones by computer-controlled microfilm recorder. IEEE Transactions on Computers, C-20(6):662– 664, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671911.

Rosen:1974:BCS

[Ros74a] S. Rosen. B74-10 computer semantics, studies of algorithms, processors and languages. IEEE Transactions on Computers, C-23(2): 214, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672492.

Rosenfeld:1974:DSL

[Ros74b] A. Rosenfeld. Digital straight line segments. IEEE Transactions on Computers, C-23(12):1264–1269, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672438.

Rosenfeld:1975:NAD

[Ros75] A. Rosenfeld. A note on automatic detection of texture gradients. IEEE Transactions on Computers, C-24(10):988–991, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672699.

Roth:1970:RDS

[Rot70] J. P. Roth. R70-12 diagnosis of single-gate failures in combinational circuits. IEEE Transactions on Computers, C-19(6):564– 565, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671576.

Roth:1977:HV

[Rot77] J. P. Roth. Hardware verification. IEEE Transactions on Computers, C-26(12):1292–1294, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674795.

Roth:1978:PLA

[Rot78] J. P. Roth. Programmed logic array optimization. IEEE Transactions on Computers, C-27(2):174– 176, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675055.

Russo:1971:HPP

[ROW71] R. L. Russo, P. H. Oden, and P. K. Wolff, Sr. A heuristic procedure for the partitioning and mapping of computer logic graphs. IEEE Transactions on Computers, C-20(12):1455–1462, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671750.

Roy:1974:DFE

[Roy74] B. K. Roy. Diagnosis and fault equivalence in combinational circuits. IEEE Transactions on Computers, C-23(9):955– 963, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672652.

Ramamoorthy:1973:CTR

[RPL73] C. V. Ramamoorthy, J. H. Park, and H. F. Li. Compilation techniques for recognition of parallel processable tasks in arithmetic expressions. IEEE Transactions on Computers, C-22(11):986–998, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672228.

Robbins:1974:TLE

[RR74] W. E. Robbins and L. D. Rudolph.

On two-level exclusive-or majority networks. IEEE Transactions on Computers, C-23(1):34– 41, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672367.

Revuluri:1974:CDS

[RRNA74] K. Revuluri, K. R. Rao, M. A. Narasimhan, and N. Ahmed. Cyclic and dyadic shifts, Walsh– Hadamard transform, and the H diagram. IEEE Transactions on Computers, C-23(12):1303–1306, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672444.

Reischer:1971:AAS

[RS71a] C. Reischer and D. A. Simovici. Associative algebraic structures in the set of Boolean functions and some applications in automata theory. IEEE Transactions on Computers, C-20(3):298– 303, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671827.

Rink:1971:ADC

[RS71b] R. A. Rink and W. Streifer. Application of digital computers to solve analytically a class of secondorder non-linear ordinary differential equations. IEEE Transactions on Computers, C-20(8):901– 910, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671961.

Reischer:1972:EPA

[RS72a] C. Reischer and D. A. Simovici.

On the existence of a periodic analog of a finite connected automaton. IEEE Transactions on Computers, C-21(2):208–211, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008929.

Roy:1972:DMD

[RS72b] P. K. Sinha Roy and C. L. Sheng. A decomposition method of determining maximum compatibles. IEEE Transactions on Computers, C-21 (3):309–312, March 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008956. See comments [RS73b].

Roy:1973:AR

[RS73a] P. K. Sinha Roy and C. L. Sheng.

Authors’ reply. IEEE Transactions on Computers, C-22(6): 627, June 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009119.

Roy:1973:ARC

[RS73b] P. K. Sinha Roy and C. L.

Sheng. Author’s reply to comments on “Decomposition Method of Determining Maximum Compatibles”. IEEE Transactions on Computers, C-22(6):627, June 1973.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Sureshchander.pdf. See [RS72b,

Sur73].

Ramamoorthy:1974:ATC

[RS74] C. V. Ramamoorthy and K. S. Shankar. Automatic testing for the correctness and equivalence of loopfree microprograms. IEEE Transactions on Computers, C-23 (8):768–782, August 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672626.

Robach:1975:DTM

[RS75] C. Robach and G. Saucier. Diversified test methods for local control units. IEEE Transactions on Computers, C-24(5):562– 567, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672854.

Robach:1978:DTC

[RS78] C. Robach and G. Saucier. Dynamic testing of control units. IEEE Transactions on Computers, C-27 (7):617–623, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675161.

Rao:1979:ATD

[RSH79] G. S. Rao, H. S. Stone, and T. C.

Hu. Assignment of tasks in a distributed processor system with limited memory. IEEE Transactions on Computers, C-28(4):291– 299, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675348.

Robach:1976:PTD

[RSL76] C. Robach, G. Saucier, and J. Lebrun. Processor testability and design consequences. IEEE Transactions on Computers, C-25(6): 645–652, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674666.

Rao:1970:BLR

[RT70] T. R. N. Rao and A. K. Trehan. Binary logic for residue arithmetic using magnitude index. IEEE Transactions on Computers, C-19(8):752–757, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Rao.pdf;
https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671619.

Rosenfeld:1971:ECD

[RT71] A. Rosenfeld and M. Thurston.

Edge and curve detection for visual scene analysis. IEEE Transactions on Computers, C-20(5): 562–569, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671883.

Robertson:1973:SIC

[RT73a] James E. Robertson and Kishor S. Trivedi. The status of investigations into computer hardware design based on the use of continued fractions. IEEE Transactions on Computers, C-22(6):555– 560, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith2/papers/ARITH2_Robertson.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009106.

Roome:1973:AMS

[RT73b] W. D. Roome and H. C. Toorng. Algorithms for multiple shift register realizations of sequential machines. IEEE Transactions on Computers, C-22(10):933–943, October 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672214.

Ramamoorthy:1974:HLL

[RT74] C. V. Ramamoorthy and M. Tsuchiya. A high-level language for horizontal microprogramming. IEEE Transactions on Computers, C-23(8):791– 801, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672628.

Reed:1979:NHA

[RT79] I. S. Reed and T. K. Truong.

A new hybrid algorithm for computing a fast discrete Fourier transform. IEEE Transactions on Computers, C-28(7):487–492, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675393.

Reed:1977:IPT

[RTKH77] I. S. Reed, T. K. Truong, Y. S.

Kwoh, and E. L. Hall. Image processing by transforms over a finite field. IEEE Transactions on Computers, C-26(9):874–881, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674935.

Rosenfeld:1972:ECD

[RTL72] A. Rosenfeld, M. Thurston, and Yung-Han Lee. Edge and curve detection: Further experiments. IEEE Transactions on Computers, C-21 (7):677–715, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672166.

Ramamoorthy:1978:DFC

[RTW78] C. V. Ramamoorthy, J. L. Turner, and B. W. Wah. A design of a fast cellular associative memory for ordered retrieval. IEEE Transactions on Computers, C-27(9):800– 815, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675200.

Rubin:1974:LPC

[Rub74] F. Rubin. The Lee path connection algorithm. IEEE Transactions on Computers, C-23(9):907– 914, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672647. See correction [Rub76a].

Rubin:1975:IAT

[Rub75a] F. Rubin. An improved algorithm for testing the planarity of a graph. IEEE Transactions on Computers, C-24(2):113–121, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672772.

Rubin:1975:WCB

[Rub75b] F. Rubin. Worst case bounds for maximal compatible subsets. IEEE Transactions on Computers, C-24 (8):830–831, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672908.

Rubinfield:1975:PMB

[Rub75c] L. P. Rubinfield. A proof of the modified Booth’s algorithm for multiplication. IEEE Transactions on Computers, C-24(10): 1014–1015, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672707.

Rubin:1976:CLP

[Rub76a] Frank Rubin. Correction to “The Lee Path Connection Algorithm”. IEEE Transactions on Computers, C-25(2):208, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009239.

See [Rub74].

Rubin:1976:GNC

[Rub76b] Frank Rubin. Generation of nonparametric curves. IEEE Transactions on Computers, C-25(1): 103, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009216.

Rubin:1979:DSC

[Rub79] F. Rubin. Decrypting a stream cipher based on J–K flip-flops. IEEE Transactions on Computers, C-28 (7):483–487, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675392.

Rudeanu:1974:AAB

[Rud74] S. Rudeanu. An algebraic approach to Boolean equations. IEEE Transactions on Computers, C-23(2): 206–207, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672484.

Rudeanu:1976:SRF

[Rud76] S. Rudeanu. Square roots and functional decompositions of Boolean functions. IEEE Transactions on Computers, C-25(5): 528–532, May 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674641.

Rumbaugh:1977:DFM

[Rum77] James Rumbaugh. A data flow multiprocessor. IEEE Transactions on Computers, C-26(2):138– 146, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009292.

Russo:1972:TBL

[Rus72] Roy L. Russo. On the tradeoff between logic performance and circuit-to-pin ratio for LSI.

IEEE Transactions on Computers, C-21(2):147–153, February 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5008919.

Rux:1971:BMP

[Rux71] P. T. Rux. B71-3 microprogramming: Principles and practices. IEEE Transactions on Computers, C-20(10):1221–1222, October 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671705.

Rosenfeld:1973:MAM

[RV73] J. L. Rosenfeld and R. D. Villani.

Micromultiprocessing: An approach to multiprocessing at the level of very small tasks. IEEE Transactions on Computers, C-22(2):149– 153, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672269.

Rey:1974:SSA

[RV74] C. A. Rey and J. Vaucher. Selfsynchronized asynchronous sequential machines. IEEE Transactions on Computers, C-23(12):1306–1311, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672445.

Reddy:1974:ETC

[RW74] S. M. Reddy and J. R. Wilson. Easily testable cellular realizations for the (exactly P)- out-of-n and (p or more)-out-of-n logic functions. IEEE Transactions on Computers, C-23(1):98– 100, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672380.

Rosenfeld:1975:IMA

[RW75] A. Rosenfeld and J. S. Weszka. An improved method of angle detection on digital curves. IEEE Transactions on Computers, C-24(9):940– 941, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672935.

Santos:1970:GMS

[SA70] J. Santos and H. Arango. A graphic method for the synthesis of threshold ternary functions. IEEE Transactions on Computers, C-19(10): 975–976, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671398.

Satake:1974:FTN

[SA74] S. Satake and T. Aibara. A formula for ternary network cost. IEEE Transactions on Computers, C-23 (3):329–332, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672527.

Swartzlander:1975:SLN

[SA75] E. E. Swartzlander, Jr. and A. G.

Alexopoulos. The sign/logarithm number system. IEEE Transactions on Computers, C-24(12):1238–1242, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672765. See [HL86].

Schiller:1971:MIG

[SAFV71] W. L. Schiller, R. L. Abraham, R. M. Fox, and A. Van Dam. A microprogrammed intelligent graphics terminal. IEEE Transactions on Computers, C-20(7):775– 782, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671939.

Sagalovich:1977:CNC

[Sag77] Y. L. Sagalovich. Comments on “On the Number of Classes of Binary Matrices”. IEEE Transactions on Computers, C-26(5): 512, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674871. See [Har73].

Sahin:1974:MBR

[Sah74] K. E. Sahin. Message-based response routing with Selcuk networks. IEEE Transactions on Computers, C-23(12):1250–1257, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672436.

Sahin:1977:SSC

[Sah77] K. E. Sahin. Segmented Selcuk communication networks and response routing. IEEE Transactions on Computers, C-26(8):831– 833, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674924.

Sammon:1970:IPA

[Sam70a] J. W. Sammon, Jr. Interactive pattern analysis and classification. IEEE Transactions on Computers, C-19(7):594–616, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671586.

Sammon:1970:ODP

[Sam70b] J. W. Sammon, Jr. An optimal discriminant plane. IEEE Transactions on Computers, C-19(9):826– 829, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671640.

Sandewall:1976:CPC

[San76] E. Sandewall. Conversion of predicate-calculus axioms to corresponding deterministic programs. IEEE Transactions on Computers, C-25(4):342–346, April 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674615.

Santos:1970:CAS

[SAPR70] J. Santos, H. Arango, M. Pascual, and G. Roing. A cyclic algebra for the synthesis of ternary digital systems. IEEE Transactions on Computers, C-19(7):651– 653, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671592.

Sarma:1971:ADC

[Sar71] M. S. Sarma. Application of digital computers to three-dimensional field determination. IEEE Transactions on Computers, C-20(8):914– 917, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671963.

Sarwate:1978:SFF

[Sar78] D. V. Sarwate. Semi-fast Fourier transforms over GF(2^{m}). IEEE Transactions on Computers, C-27 (3):283–285, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675088.

Saucier:1972:NSE

[Sau72a] Gabriele Saucier. Next-state equations of asynchronous sequential machines. IEEE Transactions on Computers, C-21(4):397– 399, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008984. See comments [MS73a].

Saucier:1972:SAA

[Sau72b] Gabriele Saucier. State assignment of asynchronous sequential machines using graph techniques. IEEE Transactions on Computers, C-21(3):282–288, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008950.

Sawin:1974:OAS

[Saw74] D. H. Sawin III. Optimization of asynchronous sequential circuit realizations. IEEE Transactions on Computers, C-23(2):186– 188, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672477.

Sawin:1975:DRS

[Saw75] D. H. Sawin III. Design of reliable synchronous sequential circuits. IEEE Transactions on Computers, C-24(5):567– 570, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672855.

Sawchuk:1977:RTC

[Saw77] Alexander A. Sawchuk. Real-time correction of intensity nonlinearities in imaging systems. IEEE Transactions on Computers, C-26(1):34– 39, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009271.

Schoen:1971:BEI

[SB71] T. A. Schoen and M. R. Belsole, Jr. A Burroughs 220 emulator for the IBM 360/25. IEEE Transactions on Computers, C-20(7): 795–798, July 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671941.

Sidhu:1972:PEA

[SB72] G. S. Sidhu and R. T. Boute. Property encoding: Application in binary picture encoding and boundary following. IEEE Transactions on Computers, C-21(11):1206–1216, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672071.

Sureshchander:1974:NRS

[SB74a] Sureshchander and P. C. P. Bhatt. A new representation of sequential machines for determining all incompatible pairs. IEEE Transactions on Computers, C-23(3):325– 326, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672523.

Sureshchander:1974:NUL

[SB74b] Sureshchander and P. C. P. Bhatt. A note on Q-universal logic modules. IEEE Transactions on Computers, C-23(3):319– 320, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672521.

Sarje:1977:ATA

[SB77a] A. K. Sarje and N. N. Biswas. An algorithm for testing 2-asummability of Boolean functions. IEEE Transactions on Computers, C-26(10): 1049–1053, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674745.

Snir:1977:DAP

[SB77b] M. Snir and A. B. Barak. A direct approach to the parallel evaluation of rational expressions with a small number of processors. IEEE Transactions on Computers, C-26(10): 933–937, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674728.

Sarje:1979:NAA

[SB79] A. K. Sarje and N. N. Biswas. A new approach to 2-asummability testing. IEEE Transactions on Computers, C-28(10):798–801, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675250.

Siy:1972:MFF

[SC72a] P. Siy and C. S. Chen. Minimization of fuzzy functions. IEEE Transactions on Computers, C-21(1):100– 102, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672032. See comments [Kan73a].

Su:1972:NAF

[SC72b] S. Y. H. Su and Yun-Chung Cho.

A new approach to the fault location of combinational circuits. IEEE Transactions on Computers, C-21 (1):21–30, January 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672020.

Su:1972:CMM

[SC72c] Stephen Y. H. Su and Peter T. Cheung. Computer minimization of multivalued switching functions. IEEE Transactions on Computers, C-21(9):995–1003, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009076.

Sung:1974:TAC

[SC74] Chia-Hsiaing Sung and C. L. Coates, Jr. Tessellation aspect of combinational cellular array testing. IEEE Transactions on Computers, C-23(4):363– 369, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672544.

Shliferstein:1977:SPI

[SC77] A. R. Shliferstein and Y. T. Chien. Some properties of imageprocessing operations on projection sets obtained from digital pictures. IEEE Transactions on Computers, C-26(10):958– 970, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674731.

Sklansky:1972:MPP

[SCH72] Jack Sklansky, Robert L. Chazin, and Bruce J. Hansen. Minimumperimeter polygons of digitized silhouettes. IEEE Transactions on Computers, C-21(3):260– 268, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008948.

Schumann:1973:CFC

[Sch73] U. Schumann. Comments on “A Fast Computer Method for Matrix Transposing” and application to the solution of Poisson’s equation. IEEE Transactions on Computers, C-22(5):542– 543, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672353. See [Ekl72].

Schertz:1974:FTC

[Sch74a] D. R. Schertz. Fault-tolerant computing: An introduction. IEEE Transactions on Computers, C-23 (7):649–650, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672605.

Schoeffler:1974:BMD

[Sch74b] J. D. Schoeffler. B74-4 minicomputers in data processing and simulation. IEEE Transactions on Computers, C-23(1):111–112, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672387.

Schkolnick:1975:CPP

[Sch75] M. Schkolnick. On a covering problem for partially specified switching functions. IEEE Transactions on Computers, C-24(11):1139–1141, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672742.

Schachter:1978:DPC

[Sch78a] B. Schachter. Decomposition of polygons into convex sets. IEEE Transactions on Computers, C-27(11):1078–1082, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675001.

Schaffner:1978:PDP

[Sch78b] M. R. Schaffner. Processing by data and program blocks. IEEE Transactions on Computers, C-27(11):1015–1028, November 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674991.

Schurmann:1978:MWR

[Sch78c] J. Schurmann. A multifont word recognition system for postal address reading. IEEE Transactions on Computers, C-27(8):721– 732, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675179.

Schneck:1979:CWU

[Sch79a] P. B. Schneck. Comment on “When to Use Random Testing”. IEEE Transactions on Computers, C-28 (8):580–581, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675415. See [Agr78b].

Schwartz:1979:TTD

[Sch79b] M. Schwartz. Throughput and time delay analysis for a common queue configuration in a multiprocessor enviromnent. IEEE Transactions on Computers, C-28(12): 939–941, December 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675285.

Sankar:1973:AAN

[SCK73a] P. V. Sankar, S. Chakrabarti, and E. V. Krishnamurthy. Arithmetic algorithms in a negative base. IEEE Transactions on Computers, C-22(2):120–125, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672264.

Sankar:1973:DDA

[SCK73b] P. V. Sankar, S. Chakrabarti, and E. V. Krishnamurthy. Deterministic division algorithm in a negative base. IEEE Transactions on Computers, C-22(2):125– 128, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672265.

Slagle:1970:NAG

[SCL70] J. R. Slagle, Chin-Liang Chang, and R. C. T. Lee. A new algorithm for generating prime implicants. IEEE Transactions on Computers, C-19 (4):304–310, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671510. See comments [Das71c].

Sklansky:1976:PDC

[SCL76] Jack Sklansky, Luigi P. Cordella, and Stefano Levialdi. Parallel detection of concavities in cellular blobs. IEEE Transactions on Computers, C-25(2):187–196, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009234.

Swartzlander:1983:SLA

[SCNS83] E. E. Swartzlander, Jr., D. V. Satish Chandra, H. T. Nagle, Jr., and S. A.

Starks. Sign/logarithm arithmetic for FFT implementation. IEEE Transactions on Computers, C-32 (6):526–534, June 1983. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1676274. See comments [HL86].

Schmidt:1974:ECT

[SD74] D. C. Schmidt and L. E. Druffel.

An extension of the clause table approach to multi-output combinational switching networks. IEEE Transactions on Computers, C-23 (4):338–346, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672540.

Sethi:1979:IMS

[SD79] A. S. Sethi and N. Deo. Interference in multiprocessor systems with localized memory access probabilities. IEEE Transactions on Computers, C-28(2):157– 163, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675307.

Selkow:1974:DKR

[Sel74] S. M. Selkow. Diagnostic keys as a representation for context in pattern recognition. IEEE Transactions on Computers, C-23(9):970– 971, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672656.

Servit:1973:HCA

[Ser73] M. Servit. Hazard correction in asynchronous sequential circuits using inertial delay elements. IEEE Transactions on Computers, C-22(11):1041–1042, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672235.

Servit:1975:HCS

[Ser75] M. Servit. Hazard correction in synchronous sequential circuits. IEEE Transactions on Computers, C-24 (3):305–310, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672804.

Sethares:1970:CPM

[Set70a] G. C. Sethares. Completely periodic multithreshold functions. IEEE Transactions on Computers, C-19 (4):355–358, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671516.

Setlur:1970:MDE

[Set70b] R. V. Setlur. A method to determine the expressive power of a set of connectives. IEEE Transactions on Computers, C-19(12):1223–1225, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671458.

Sanathanan:1971:HCT

[SF71] C. K. Sanathanan and J. D.

Ferguson. Hybrid computation techniques inferred from functional analysis. IEEE Transactions on Computers, C-20(1):19– 24, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671669.

Sengbush:1972:DAO

[SF72] R. L. Sengbush and M. R. Foster. Design and application of optimal velocity filters in seismic exploration. IEEE Transactions on Computers, C-21(7):648– 654, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672162.

Sadeh:1974:MCS

[SF74] E. Sadeh and M. A. Franklin. Monte Carlo solution of partial differential equations by special purpose digital computer. IEEE Transactions on Computers, C-23(4):389– 397, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672547.

Shapira:1977:COP

[SF77] R. Shapira and H. Freeman. A cyclic-order property of bodies with three-face vertices. IEEE Transactions on Computers, C-26(10): 1035–1039, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674742.

Shapira:1978:CDB

[SF78] R. Shapira and H. Freeman. Computer description of bodies bounded by quadric surfaces from a set of imperfect projections. IEEE Transactions on Computers, C-27(9):841– 854, September 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675204.

Short:1971:SPD

[SG71] R. A. Short and J. Goldberg. Soviet progress in the design of fault-tolerant digital machines. IEEE Transactions on Computers, C-20(11):1337–1352, November 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671729.

Salama:1974:CAS

[SG74] A. I. A. Salama and V. Gourishankar. A computational algorithm for solving a system of coupled algebraic matrix Riccati equations. IEEE Transactions on Computers, C-23(1):100– 102, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672381.

Swartzlander:1978:IPC

[SGR78] E. E. Swartzlander, Jr., B. K.

Gilbert, and I. S. Reed. Inner product computers. IEEE Transactions on Computers, C-27(1):21– 31, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674948. See comments [Bla79].

Sutton:1972:TMA

[SH72] R. N. Sutton and E. L. Hall. Texture measures for automatic classification of pulmonary disease. IEEE Transactions on Computers, C-21 (7):667–676, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672165.

Saxton:1977:OFS

[SH77] Terrence L. Saxton and Cheng-Chi Huang. An optimal frame synchronization technique using an associative processor. IEEE Transactions on Computers, C-26(2):170– 174, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009296.

Swartzlander:1979:RAS

[SH79] E. E. Swartzlander, Jr. and D. J. Heath. A routing algorithm for signal processing networks. IEEE Transactions on Computers, C-28 (8):567–572, August 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675412.

Shanmugam:1975:CDC

[Sha75] K. S. Shanmugam. Comments on “Discrete Cosine Transform”. IEEE Transactions on Computers, C-24 (7):759, July 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672894. See [ANR74].

Shapiro:1978:TLE

[Sha78a] H. D. Shapiro. Theoretical limitations on the efficient use of parallel memories. IEEE Transactions on Computers, C-27(5):421– 428, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675122.

Shapiro:1978:TMC

[Sha78b] S. D. Shapiro. Transform method of curve detection for textured image data. IEEE Transactions on Computers, C-27(3):254– 255, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675080.

Shelly:1970:REA

[She70a] J. H. Shelly. R70-37 extensions of asynchronous circuits and the delay problem I. good extensions and the delay problem of the first kind. IEEE Transactions on Computers, C-19(9):859, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671658.

Sheng:1970:RRS

[She70b] C. L. Sheng. R70-15 realization of sequential machines with threshold elements. IEEE Transactions on Computers, C-19(6): 565–566, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671579.

Shepherdson:1970:RSB

[She70c] J. C. Shepherdson. R70-20 sequential Boolean equations. IEEE Transactions on Computers, C-19 (7):658, July 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671600.

Sherlock:1970:SMS

[She70d] J. F. Sherlock. The simulation of a multicomputer system. IEEE Transactions on Computers, C-19(11):1114–1118, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671437.

Shepherd:1971:RSL

[She71] B. J. Shepherd. Right shift for low-cost multiply and divide. IEEE Transactions on Computers, C-20(12):1586–1589, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671770.

Shedletsky:1977:CSI

[She77] J. J. Shedletsky. Comment on the sequential and indeterminate behavior of an end-around-carry adder. IEEE Transactions on Computers, C-26(3):271– 272, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674817.

Shedletsky:1978:ECA

[She78a] J. J. Shedletsky. Error correction by alternate-data retry. IEEE Transactions on Computers, C-27(2):106– 112, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675044.

Shedletsky:1978:RIN

[She78b] J. J. Shedletsky. A rollback interval for networks with an imperfect self-checking property. IEEE Transactions on Computers, C-27 (6):500–508, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675140.

Shively:1970:PFI

[Shi70] Richard R. Shively. Preface:

First IEEE Conference on Computer Arithmetic, Minneapolis, Minnesota, June 16, 1969. IEEE Transactions on Computers, C-19 (8):679–680, August 1970. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith1/papers/ARITH1_
contents.pdf;   http://www.acsel-
lab.com/arithmetic/arith1/papers/
ARITH1_Shively.pdf;   https:
//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671609.

Shively:1975:BRE

[Shi75] R. R. Shively. Book reviews Editor’s comments. IEEE Transactions on Computers, C-24(3): 336, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672814.

Sholl:1974:DTM

[Sho74] H. A. Sholl. Direct transition memory and its application in computer design. IEEE Transactions on Computers, C-23(10): 1048–1061, October 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672398.

Shriver:1971:MNA

[Shr71] B. D. Shriver, Sr. Microprogramming and numerical analysis. IEEE Transactions on Computers, C-20 (7):808–812, July 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671944.

Story:1972:OSA

[SHR72] J. R. Story, H. J. Harrison, and E. A. Reinhard. Optimum state assignment for synchronous sequential circuits. IEEE Transactions on Computers, C-21(12):1365–1373, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672101.

Si:1978:DTR

[Si78]  Siu-Chong Si.  Dynamic test-

ing of redundant logic networks. IEEE Transactions on Computers, C-27(9):828–832, September 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675202.

Sickel:1976:STC

[Sic76] S. Sickel. A search technique for clause interconnectivity graphs. IEEE Transactions on Computers, C-25(8):823– 835, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674701.

Siewiorek:1974:BDS

[Sie74] D. Siewiorek. B74-40 digital systems: Hardware organization and design. IEEE Transactions on Computers, C-23(9):991–992, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672667.

Siewiorek:1975:RMC

[Sie75] D. P. Siewiorek. Reliability modeling of compensating module failures in majority voted redundancy. IEEE Transactions on Computers, C-24(5):525– 533, May 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672849.

Siegel:1977:ATS

[Sie77] Howard Jay Siegel. Analysis techniques for SIMD machine interconnection networks and the effects of processor address masks. IEEE Transactions on Computers, C-26(2):153– 161, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009294.

Siegel:1979:MSM

[Sie79] H. J. Siegel. A model of SIMD machines and a comparison of various interconnection networks. IEEE Transactions on Computers, C-28(12):907–917, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675280.

Sugino:1976:NLS

[SIF76] K. Sugino, Y. Inagaki, and T. Fukumura. A note on the linear space automata stability problem. IEEE Transactions on Computers, C-25 (7):678–683, July 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674677.

Sikdar:1970:DMM

[Sik70] K. Sikdar. Determination of multipliers mapping an arbitrary integer into a range of certain type. IEEE Transactions on Computers, C-19(12):1221–1222, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671456.

Siler:1973:BIC

[Sil73] K. F. Siler. B73-1 an introduction to Cobol programming. IEEE Transactions on Computers, C-22 (1):112, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672200.

Silverston:1977:STM

[Sil77] S. M. Silverston. A stable, twomultiplies-per-cycle algorithm for digital generation of sinusoids in real time. IEEE Transactions on Computers, C-26(11):1153–1154, November 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674765. See comments [HO79].

Silio:1979:ESC

[Sil79] C. B. Silio, Jr. An efficient simplex coverability algorithm in E2 with application to stochastic sequential machines. IEEE Transactions on Computers, C-28(2):109– 120, February 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675300.

Singh:1971:DIA

[Sin71] S. Singh. On delayed-input asynchronous sequential circuits. IEEE Transactions on Computers, C-20 (5):500–503, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671876.

Sintonen:1972:RFV

[Sin72] Leo Sintonen. On the realization of functions in N-valued logic. IEEE Transactions on Computers, C-21 (6):610–612, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009020.

Sintonen:1977:CMF

[Sin77] L. Sintonen. A clocked multivalued flip-flop. IEEE Transactions on Computers, C-26(3):292– 294, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674823.

Sites:1974:SBD

[Sit74] R. L. Sites. Serial binary division by ten. IEEE Transactions on Computers, C-23(12):1299–1301, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672442.

Sarkar:1971:EPP

[SK71] B. P. Sarkar and E. V. Krishnamurthy. Economic pseudodivision processes for obtaining square root, logarithm, and arctan. IEEE Transactions on Computers, C-20(12):1589–1593, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671771.

Sen:1974:RAA

[SK74] S. K. Sen and E. V. Krishnamurthy. Rank-augmented LUalgorithm for computing generalized matrix inverses. IEEE Transactions on Computers, C-23(2):199– 201, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672481.

Sastry:1975:PCM

[SK75] K. V. Sastry and R. Y. Kain.

On the performance of certain multiprocessor computer organizations. IEEE Transactions on Computers, C-24(11):1066–1074, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672727.

Sameh:1977:CPA

[SK77a] A. H. Sameh and D. J. Kuck.

Correction to “A Parallel QR Algorithm for Symmetric Tridiagonal Matrices”. IEEE Transactions on Computers, C-26(8): 836, August 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674926. See [SK77b].

Sameh:1977:PAS

[SK77b] Ahmed H. Sameh and David J. Kuck. A parallel QR algorithm for symmetric tridiagonal matrices. IEEE Transactions on Computers, C-26(2):147–153, February 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009293. See correction [SK77a].

Seth:1977:DFL

[SK77c] Sharad C. Seth and K. L. Kodandapani. Diagnosis of faults in linear tree networks. IEEE Transactions on Computers, C-26(1):29– 33, January 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009270.

Sasao:1978:CRI

[SK78a] T. Sasao and K. Kinoshita. Cascade realization of 3-input 3-output conservative logic circuits. IEEE Transactions on Computers, C-27 (3):214–221, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675074.

Sasao:1978:RMC

[SK78b] T. Sasao and K. Kinoshita. Realization of minimum circuits with two-input conservative logic elements. IEEE Transactions on Computers, C-27(8):749–752, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675183.

Sasao:1979:CLE

[SK79a] T. Sasao and K. Kinoshita. Conservative logic elements and their universality. IEEE Transactions on Computers, C-28(9):682–685, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675437.

Sasao:1979:NFF

[SK79b] T. Sasao and K. Kinoshita. On the number of fanout-free functions and unate cascade functions. IEEE Transactions on Computers, C-28 (1):66–72, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675227.

Stenzel:1977:CHS

[SKG77] W. J. Stenzel, W. J. Kubitz, and G. H. Garcia. A compact high-speed parallel multiplication scheme. IEEE Transactions on Computers, C-26(10):948– 957, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674730.

Suenaga:1979:HSA

[SKK79] Y. Suenaga, T. Kamae, and T. Kobayashi. A high-speed algorithm for the generation of straight lines and circular arcs. IEEE Transactions on Computers, C-28(10): 728–736, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675240.

Sklansky:1970:RUC

[Skl70] J. Sklansky. R70-16 unate cellular logic. IEEE Transactions on Computers, C-19(6): 566, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671580.

Sklansky:1972:MCR

[Skl72] J. Sklansky. Measuring concavity on a rectangular mosaic. IEEE Transactions on Computers, C-21(12):1355–1364, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672100.

Sklansky:1978:HTC

[Skl78] J. Sklansky. On the Hough technique for curve detection. IEEE Transactions on Computers, C-27(10):923–926, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674971.

Su:1978:CPM

[SKM78] S. Y. H. Su, I. Koren, and Y. K.

Malaiya. A continuous-parameter Markov model and detection procedures for intermittent faults. IEEE Transactions on Computers, C-27 (6):567–570, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675148.

Schnurmann:1975:WRT

[SLC75] H. D. Schnurmann, E. Lindbloom, and R. G. Carpenter. The weighted random test-pattern generator. IEEE Transactions on Computers, C-24(7):695–700, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672883.

Slivinski:1970:ETL

[Sli70] T. A. Slivinski. An extension of threshold logic. IEEE Transactions on Computers, C-19(4):319– 341, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671512.

Shen:1970:ADD

[SM70a] V. Y.-S. Shen and A. C. McKellar. An algorithm for the disjunctive decomposition of switching functions. IEEE Transactions on Computers, C-19(3):239– 248, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671496.

Stein:1970:SIO

[SM70b] M. L. Stein and E. J. Mundstock. Sorting implicit outputs in digital simulation. IEEE Transactions on Computers, C-19(9):844– 847, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671645.

Schertz:1971:DMF

[SM71a] D. R. Schertz and G. Metze. On the design of multiple fault diagnosable networks. IEEE Transactions on Computers, C-20(11):1361–1364, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671732.

Schmitt:1971:TAI

[SM71b] N. M. Schmitt and J. L. Melsa.

Two approaches for increasing storage density in modern digital computing systems. IEEE Transactions on Computers, C-20(2):167– 175, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671800.

Stein:1971:SMA

[SM71c] M. L. Stein and W. D. Munro.

Scaling machine arithmetic. IEEE Transactions on Computers, C-20 (6):675–678, June 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671917.

Schaefer:1972:EDC

[SM72a] L. F. Schaefer and A. Macovski.

Encoding and decoding of color information using two-dimensional spatial filtering. IEEE Transactions on Computers, C-21(7):642– 647, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672161.

Schertz:1972:NRF

[SM72b] Donald R. Schertz and Gernot Metze. A new representation for faults in combinational digital circuits. IEEE Transactions on Computers, C-21(8):858– 866, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009041.

Siewiorek:1973:ICS

[SM73a] D. P. Siewiorek and E. J. Mc-Cluskey. An iterative cell switch design for hybrid redundancy. IEEE Transactions on Computers, C-22 (3):290–297, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672302.

Siewiorek:1973:SCS

[SM73b] D. P. Siewiorek and E. J. Mc-Cluskey. Switch complexity in systems with hybrid redundancy. IEEE Transactions on Computers, C-22 (3):276–282, March 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672300.

Sawin:1974:ASM

[SM74] D. H. Sawin III and G. K. Maki. Asynchronous sequential machines designed for fault detection. IEEE Transactions on Computers, C-23 (3):239–249, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672511. See comments [Wak75a].

Sawin:1975:CFT

[SM75a] D. H. Sawin III and G. K. Maki.

Comments on “Fault-Tolerant Asynchronous Networks”. IEEE Transactions on Computers, C-24 (7):756–758, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672892. See [PR73].

Sawin:1975:FSA

[SM75b] D. H. Sawin III and G. K.

Maki. Fail-safe asynchronous sequential machines. IEEE Transactions on Computers, C-24(6): 675–677, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672875.

Schmidt:1975:MRC

[SM75c] D. C. Schmidt and G. Metze. Modular replacement of combinational switching networls. IEEE Transactions on Computers, C-24(1):29– 48, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672674.

Shedletsky:1976:ELF

[SM76] J. J. Shedletsky and E. J. Mc-Cluskey. The error latency of a fault in a sequential digital circuit. IEEE Transactions on Computers, C-25 (6):655–659, June 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1674668. See comments [DTG79].

Smith:1978:SFS

[SM78] J. E. Smith and G. Metze. Strongly fault secure logic networks. IEEE Transactions on Computers, C-27 (6):491–499, June 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675139.

Shedletsky:1979:AR

[SM79] J. J. Shedletsky and E. J. Mc-Cluskey. Authors’ reply. IEEE Transactions on Computers, C-28 (1):86, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675231.

Smith:1971:GSR

[Smi71a] A. R. Smith. General shiftregister sequences of arbitrary cycle length. IEEE Transactions on Computers, C-20(4):456– 459, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671858.

Smith:1971:CPD

[Smi71b] D. R. Smith. Complexity of partially defined combinational switching functions. IEEE Transactions on Computers, C-20(2):204– 208, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671805.

Smith:1974:GIS

[Smi74] R. J. Smith II. Generation of internal state assignments for large asynchronous sequential machines. IEEE Transactions on Computers, C-23(9):924– 932, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672649.

Smith:1975:DFT

[Smi75] T. B. Smith III. A damageand fault-tolerant input/output network. IEEE Transactions on Computers, C-24(5):505–512, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672846.

Smith:1976:MWS

[Smi76] A. J. Smith. A modified working set paging algorithm. IEEE Transactions on Computers, C-25(9):907– 914, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674715.

Smith:1978:ECL

[Smi78] J. E. Smith. On the existence of combinational logic circuits exhibiting multiple redundancy. IEEE Transactions on Computers, C-27(12):1221–1225, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675032.

Smith:1979:AES

[Smi79a] A. J. Smith. An analytic and experimental study of multiple channel controllers. IEEE Transactions on Computers, C-28(1):38– 49, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675220.

Smith:1979:CRT

[Smi79b] J. E. Smith. Comments on “Redundancy Testing in Combinational Networks”. IEEE Transactions on Computers, C-28(3):261– 262, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675332. See [LD74].

Smith:1979:DFP

[Smi79c] J. E. Smith. Detection of faults in programmable logic arrays. IEEE Transactions on Computers, C-28(11):845–853, November 1979. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675264.

Smith:1979:NSC

[Smi79d] J. E. Smith. On necessary and sufficient conditions for multiple fault undetectability. IEEE Transactions on Computers, C-28(10): 801–802, October 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675251.

Smith:1979:USD

[Smi79e] J. E. Smith. Universal system diagnosis algorithms. IEEE Transactions on Computers, C-28(5): 374–378, May 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675370.

Shen:1971:FAD

[SMW71] V. Y.-S. Shen, A. C. McKellar, and P. Weiner. A fast algorithm for the disjunctive decomposition of switching functions. IEEE Transactions on Computers, C-20(3):304– 309, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671828.

Su:1971:CAS

[SN71] S. Y. H. Su and Chong-Woo Nam. Computer-aided synthesis or multiple-output multilevel NAND networks with fan-in and fan-out constraints. IEEE Transactions on Computers, C-20(12):1445–1455, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671749.

Sklansky:1972:PMD

[SN72a] J. Sklansky and P. J. Nahin. A parallel mechanism for describing silhouettes. IEEE Transactions on Computers, C-21(11):1233–1239, November 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672076.

Som:1972:MSP

[SN72b] A. Som and A. K. Nath. On a method of sequential pattern recognition. IEEE Transactions on Computers, C-21(12):1433–1440, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672113.

Speelpenning:1979:SMP

[SN79] B. Speelpenning and J. Nievergelt. A simple model of processor — resource utilization in networks of communicating modules. IEEE Transactions on Computers, C-28(12):927–929, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675282.

Su:1979:AFI

[SNEL79] S. Y. W. Su, Le Huu Nguyen, A. Emam, and G. J. Lipovski. The architectural features and implementation techniques of the multicell CASSM. IEEE Transactions on Computers, C-28(6):430– 445, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675382.

Sakaguchi:1970:NAM

[SNN70] M. Sakaguchi, N. Nishida, and T. Nemoto. A new associative memory system utilizing holography. IEEE Transactions on Computers, C-19(12):1174–1181, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671449.

Schuster:1979:RAP

[SNOS79] S. A. Schuster, H. B. Nguyen, E. A. Ozkarahan, and K. C. Smith. RAP.2 — an associative processor for databases and its applications. IEEE Transactions on Computers, C-28(6):446– 458, June 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675383. Also published in/as: ASCA, Apr.1978, pp.52–59.

Sutherland:1973:HBS

[SO73] I. E. Sutherland and D. Oestreicher. How big should a printed circuit board be? IEEE Transactions on Computers, C-22(5): 537–542, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672352.

Saluja:1979:MRM

[SO79a] K. K. Saluja and E. H. Ong. Minimization of Reed–Muller canonic expansion. IEEE Transactions on Computers, C-28(7):535–537, July 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675401.

Srihari:1979:EAD

[SO79b] S. N. Srihari and M. K. Ohanesian. An efficient algorithm for determining Hadamard sequency vectors. IEEE Transactions on Computers, C-28(3):243– 244, March 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675325.

Sokal:1972:OCN

[Sok72] Nathan O. Sokal. Optimum choice of noise frequency band and sampling rate for generating random binary digits from clipped white noise. IEEE Transactions on Computers, C-21(6): 614–615, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009023.

Smith:1972:ACT

[SP72a] E. A. Smith and D. R. Phillips.

Automated cloud tracking using precisely aligned digital ATS pictures. IEEE Transactions on Computers, C-21(7):715–729, July 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672167.

Sobolewski:1972:PAAa

[SP72b] John S. Sobolewski and W. H.

Payne. Pseudonoise with arbitrary amplitude distribution — Part I: Theory. IEEE Transactions on Computers, C-21(4):337– 345, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008973.

Sobolewski:1972:PAAb

[SP72c] John S. Sobolewski and W. H.

Payne. Pseudonoise with arbitrary amplitude distribution — Part II: Hardware implementation. IEEE Transactions on Computers, C-21 (4):346–352, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008974.

Steer:1977:DHS

[SP77] D. G. Steer and S. R. Penstone. Digital hardware for sine–cosine functions. IEEE Transactions on Computers, C-26(12):1283–1286, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674791.

Spira:1971:TNC

[Spi71] P. M. Spira. On the time necessary to compute switching functions. IEEE Transactions on Computers, C-20(1):104–105, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671682.

Spira:1973:CTA

[Spi73] Philip M. Spira. Computation times of arithmetic and Boolean functions in (d, r) circuits. IEEE Transactions on Computers, C-22(6):552– 555, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        www.acsel-lab.com/arithmetic/
                                        arith2/papers/ARITH2_Spira.pdf;
                                         https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009105.

Smith:1971:ASA

[SR71] J. R. Smith, Jr. and C. H. Roth, Jr. Analysis and synthesis of asynchronous sequential networks using edge-sensitive flip-flops. IEEE Transactions on Computers, C-20 (8):847–855, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671953.

Shaham:1972:NDA

[SR72a] Z. Shaham and Z. Riesel. A note on division algorithms based on multiplication. IEEE Transactions on Computers, C-21(5): 513–514, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672150.

Sheng:1972:ASM

[SR72b] C. L. Sheng and P. K. Sinha Roy. An approach for the synthesis of multithreshold threshold elements. IEEE Transactions on Computers, C-21(8):913– 920, August 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009053.

Scherba:1973:CTM

[SR73] M. B. Scherba and R. P. Roesser.

Computation of the transition matrix of a linear sequential circuit. IEEE Transactions on Computers, C-22(4):427–428, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672328.

Saluja:1974:ETT

[SR74a] K. K. Saluja and S. M. Reddy. Easily testable two-dimensional cellular logic arrays. IEEE Transactions on Computers, C-23(11):1204–1207, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672424.

Saluja:1974:MTL

[SR74b] K. K. Saluja and S. M. Reddy. On minimally testable logic networks. IEEE Transactions on Computers, C-23(5):552–554, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672574.

Saluja:1975:FDT

[SR75] K. K. Saluja and S. M. Reddy. Fault detecting test sets for Reed–Muller canonic networks. IEEE Transactions on Computers, C-24(10):995– 998, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672701.

Sundaramurthy:1977:SRF

[SR77] M. Sundaramurthy and V. U.

Reddy. Some results in fixedpoint Fast Fourier Transform error analysis. IEEE Transactions on Computers, C-26(3):305– 308, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674827.

Srinivasan:1971:CECa

[Sri71a] C. V. Srinivasan. Codes for error correction in high-speed memory systems — Part I: Correction of cell defects in integrated memories. IEEE Transactions on Computers, C-20(8):882– 888, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671958.

Srinivasan:1971:CECb

[Sri71b] C. V. Srinivasan. Codes for error correction in high-speed memory systems. Part II: Correction of temporary and catastrophic errors. IEEE Transactions on Computers, C-20(12):1514–1520, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671757.

Srini:1975:RFF

[Sri75] V. P. Srini. Realization of fuzzy forms. IEEE Transactions on Computers, C-24(9):941–943, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672936.

Srinivasan:1976:ACI

[Sri76] C. V. Srinivasan. The architecture of coherent information system: a general problem solving system. IEEE Transactions on Computers, C-25(4):390– 402, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674621.

Srini:1978:FLS

[Sri78] V. P. Srini. Fault location in a semiconductor random-access memory unit. IEEE Transactions on Computers, C-27(4):349– 358, April 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675107.

Srini:1979:IRM

[Sri79] V. P. Srini. Iterative realization of multivalued logic systems. IEEE Transactions on Computers, C-28 (4):306–310, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675351.

Singh:1971:AUD

[SS71] S. Singh and S. G. S. Shiva. Asynchronous unit delays. IEEE Transactions on Computers, C-20(5): 493–499, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671875.

Su:1972:RBM

[SS72] S. Y. H. Su and A. A. Sarris.

The relationship between multivalued switching algebra and Boolean algebra under different definitions of complement. IEEE Transactions on Computers, C-21(5):479– 485, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672137. See comments [Eps73b, DM76].

Si:1975:MOS

[SS75a] Siu-Chong Si and A. K. Susskind.

A method for obtaining SPOOF’s. IEEE Transactions on Computers, C-24(5):560–562, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672853.

Su:1975:AR

[SS75b] S. Y. H. Su and A. A. Sarris.

Authors’ reply. IEEE Transactions on Computers, C-24(2):220– 221, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672786.

Spillman:1977:DSS

[SS77] R. J. Spillman and S. Y. H. Su. Detection of single, stuck-type failures in multivalued combinational networks. IEEE Transactions on Computers, C-26(12):1242–1251, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674785.

Sankar:1972:SIM

[SSK72] P. V. Sankar, S. K. Sen, and E. V. Krishnamurthy. Simply invertible matrices. IEEE Transactions on Computers, C-21(5): 512–513, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672149.

Suryanarayanan:1972:LCM

[SSR72] K. L. Suryanarayanan, A. C. Soudack, and S. G. Rao. An on-line computer method for parameter tracking of sampleddata systems. IEEE Transactions on Computers, C-21(3):292– 299, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008952.

Szygenda:1976:MDS

[ST76] S. A. Szygenda and E. W. Thompson. Modeling and digital simulation for design verification and diagnosis. IEEE Transactions on Computers, C-25(12):1242–1253, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674591.

Sloan:1979:PRR

[ST79] K. R. Sloan, Jr. and S. L. Tanimoto. Progressive refinement of raster images. IEEE Transactions on Computers, C-28(11):871– 874, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675269.

Stabler:1970:MT

[Sta70a] E. P. Stabler. Microprogram transformations. IEEE Transactions on Computers, C-19(10):908– 916, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671391.

Stabler:1970:SDL

[Sta70b] E. P. Stabler. System description languages. IEEE Transactions on Computers, C-19(12):1160–1173, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671448.

Stamopoulos:1974:PAJ

[Sta74] C. D. Stamopoulos. Parallel algorithms for joining two points by a straight-line segment. IEEE Transactions on Computers, C-23 (6):642–646, June 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672598.

Stamopoulos:1975:PIP

[Sta75a] C. D. Stamopoulos. Parallel image processing. IEEE Transactions on Computers, C-24(4):424– 433, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672831.

Stark:1975:ODC

[Sta75b] H. Stark. An optical-digital computer for parallel processing of images. IEEE Transactions on Computers, C-24(4):340– 347, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672820.

Stefanelli:1972:SHS

[Ste72] R. Stefanelli. A suggestion for a high-speed parallel binary divider. IEEE Transactions on Computers, C-21(1):42– 55, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672023.

Stellhorn:1977:IFP

[Ste77] W. H. Stellhorn. An inverted file processor for information retrieval. IEEE Transactions on Computers, C-26(12):1258–1267, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674787.

Shoenfelt:1971:TEE

[STG71] J. E. Shoenfelt, C. C. Tappert, and A. J. Goetze. Techniques for efficient encoding of features in pattern recognition. IEEE Transactions on Computers, C-20(9):1104– 1106, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672004.

Stiffler:1978:CRA

[Sti78a] J. J. Stiffler. Coding for randomaccess memories. IEEE Transactions on Computers, C-27(6):526– 531, June 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675143.

Stiffler:1978:RFT

[Sti78b] J. J. Stiffler. The reliability of a fault-tolerant configuration having variable coverage. IEEE Transactions on Computers, C-27(12):1195–1197, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675026.

Stone:1970:BTS

[Sto70a] H. S. Stone. B70-4 theory of scheduling. IEEE Transactions on Computers, C-19(9):854–855, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671651.

Stone:1970:LMC

[Sto70b] Harold S. Stone. A logic-in-memory computer. IEEE Transactions on Computers, C-19(1):73– 78, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008902.

Stone:1970:OHS

[Sto70c] Harold S. Stone. The organization of high-speed memory for parallel block transfer of data. IEEE Transactions on Computers, C-19 (1):47–53, January 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008899.

Storwick:1970:ICT

[Sto70d] R. M. Storwick. Improved construction techniques for (d, k) graphs. IEEE Transactions on Computers, C-19(12):1214–1216, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671454.

Stone:1971:PPP

[Sto71] H. S. Stone. Parallel processing with the perfect shuffle. IEEE Transactions on Computers, C-20(2):153– 161, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671798.

Stone:1972:DME

[Sto72] Harold S. Stone. Dynamic memories with enhanced data access. IEEE Transactions on Computers, C-21 (4):359–366, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008976.

Stoffel:1974:CDT

[Sto74a] J. C. Stoffel. A classifier design technique for discrete variable pattern recognition problems. IEEE Transactions on Computers, C-23 (4):428–441, April 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672551.

Stoffers:1974:SAD

[Sto74b] K. E. Stoffers. Sequential algorithm for the determination of maximum compatibles. IEEE Transactions on Computers, C-23(1):95– 98, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672379.

Stone:1975:DMF

[Sto75] H. S. Stone. Dynamic memories with fast random and sequential access. IEEE Transactions on Computers, C-24(12):1167–1174, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672753.

Stoffers:1977:PSE

[Sto77] K. E. Stoffers. Partitioning of separating edges: a new approach to combinational logic design. IEEE Transactions on Computers, C-26 (8):833–836, August 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674925.

Stucki:1975:SSA

[Stu75] P. Stucki. Solid-state area scanning arrays as interface devices between optical and digital computing systems — a simulation study. IEEE Transactions on Computers, C-24 (4):370–380, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672823.

Su:1970:RFL

[Su70] Yueh-Hsung Su. R70-30 fuzzy logic and its application to switching systems. IEEE Transactions on Computers, C-19(8): 766, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671628.

Su:1973:BFD

[Su73] S. Y. H. Su. B73-12 fault detection in digital circuits. IEEE Transactions on Computers, C-22 (5):547, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672359.

Su:1975:BLL

[Su75] S. Y. H. Su. B75-1 logic and logic design. IEEE Transactions on Computers, C-24(1):110, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672686.

Sung:1976:TSC

[Sun76] Chia-Hsiaing Sung. Testable sequential cellular arrays. IEEE Transactions on Computers, C-25 (1):11–18, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009199.

Sundberg:1978:EED

[Sun78] C.-E. W. Sundberg. Erasure and error decoding for semiconductor memories. IEEE Transactions on Computers, C-27(8):696– 705, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675176.

Sundberg:1979:PTS

[Sun79] C.-E. Sundberg. Properties of transparent shortened codes for memories with stuck-at faults. IEEE Transactions on Computers, C-28(9):686–690, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1675438.

Sureshchander:1970:CSB

[Sur70a] Sureshchander. Comments on “The Synthesis of Binary Sequence Detectors”. IEEE Transactions on Computers, C-19(2):173–174, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671482.

Sureshchander:1970:RFF

[Sur70b] Sureshchander. RST flip-flop input equations. IEEE Transactions on Computers, C-19(11):1118–1120, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671438.

Sureshchander:1973:CDM

[Sur73] Sureshchander. Comments on “Decomposition Method of Determining Maximum Compatibles”. IEEE Transactions on Computers, C-22(6):627, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Sureshchander.pdf;   https://
ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5009118. See

[RS72b, RS73b].

Sureshchander:1974:ATA

[Sur74] Sureshchander. An algorithm for testing asummability of Boolean functions. IEEE Transactions on Computers, C-23(2):188–191, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672478.

Sureshchander:1975:MSF

[Sur75] Sureshchander. Minimization of switching functions — a fast technique. IEEE Transactions on Computers, C-24(7):753–756, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672891.

Sureshchander:1978:CDU

[Sur78] Sureshchander. Comments on “Delayed Universal Logic Modules and Sequential Machine Synthesis”. IEEE Transactions on Computers, C-27(2):191, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675065. See [LV75].

Suter:1974:MAA

[Sut74] B. W. Suter. The modular arithmetic of arbitrarily long sequences of digits. IEEE Transactions on Computers, C-23(12):1301–1303, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672443.

Sheppard:1974:FDB

[SV74] D. A. Sheppard and Z. G. Vranesic. Fault detection of binary sequential machines using R-valued test machines. IEEE Transactions on Computers, C-23(4):352– 358, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672542.

Schwartzmann:1975:ADT

[SV75] D. H. Schwartzmann and J. J. Vidal. An algorithm for determining the topological dimensionality of point clusters. IEEE Transactions on Computers, C-24(12):1175–1182, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672754.

Svoboda:1970:ADC

[Svo70] A. Svoboda. Adder with distributed control. IEEE Transactions on Computers, C-19(8):749–751, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Svoboda.
pdf;   https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671618.

Svoboda:1973:PPB

[Svo73] Antonin Svoboda. Parallel processing in Boolean algebra. IEEE Transactions on Computers, C-22(9):848–851, September 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
5009177.

Sasaki:1970:CCS

[SW70a] A. Sasaki and S. Watanabe. Comment on “Computer Simulation of Pulse Propagation Through a Periodic Loaded Transmission Line”. IEEE Transactions on Computers, C-19(10): 983, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671402. See [SW70b].

Sasaki:1970:CSP

[SW70b] Akio Sasaki and S. Watanabe. Computer simulation of pulse propagation through a periodic loaded transmission line. IEEE Transactions on Computers, C-19(1):25– 33, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

5008896. See comment [SW70a].

Schmookler:1971:HSD

[SW71a] Martin S. Schmookler and Arnold Weinberger. High speed decimal addition. IEEE Transactions on Computers, C-20(8):862– 867, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671955.

Smith:1971:ASP

[SW71b] F. W. Smith and M. H. Wright. Automatic ship photo interpretation by the method of moments. IEEE Transactions on Computers, C-20(9):1089–1095, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672001.

Singh:1973:MOA

[SW73] S. Singh and R. Waxman. Multiple operand addition and multiplication. IEEE Transactions on Computers, C-22(2):113–120, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672263.

Smith:1976:MHH

[SW76] Carl H. Smith and Larry D. Wittie. Memory hardware for high speed job selection. IEEE Transactions on Computers, C-25(2):148– 156, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009229.

Swamy:1972:GRM

[Swa72a] Sowmitri Swamy. On generalized Reed–Muller expansions. IEEE Transactions on Computers, C-21(9):1008–1009, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009078.

Swartz:1972:BFI

[Swa72b] D. Swartz. B72-22 FORTRAN IV: a modem approach. IEEE Transactions on Computers, C-21(11): 1251, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672082.

Swartzlander:1973:PC

[Swa73a] E. E. Swartzlander, Jr. Parallel counters. IEEE Transactions on Computers, C-22(11):1021–1024, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672232.

Swartzlander:1973:QSM

[Swa73b] E. E. Swartzlander, Jr. The quasiserial multiplier. IEEE Transactions on Computers, C-22(4):317– 321, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672310.

Swanson:1974:IPM

[Swa74a] R. C. Swanson. Interconnections for parallel memories to unscramble pordered vectors. IEEE Transactions on Computers, C-23(11):1105–1115, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1672410.

Swartwout:1974:BDC

[Swa74b] R. E. Swartwout. B74-9 digital computer fundamentals. IEEE Transactions on Computers, C-23(2):213– 214, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672491.

Swartzlander:1979:CFN

[Swa79a] E. E. Swartzlander, Jr. Comment on “The Focus Number System”. IEEE Transactions on Computers, C-28(9):693, September 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675441. See [LE77, LE79].

Swartzlander:1979:MCS

[Swa79b] E. E. Swartzlander, Jr. Microprogrammed control for specialized processors. IEEE Transactions on Computers, C-28(12):930– 934, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675283.

Swartzlander:1990:CAa

[Swa90] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 1 of IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8931-5 (hardcover), 0-8186-5931-9 (microfiche). xiii + 378 pp. LCCN QA76.6 .C633 1990.

Sholl:1975:DAS

[SY75] H. A. Sholl and Shou-Chung Yang. Design of asynchronous sequential networks using readonly memories. IEEE Transactions on Computers, C-24(2):195– 206, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672778.

Tohma:1971:FTS

[TA71] Y. Tohma and S. Aoyagi. Failuretolerant sequential machines with past information. IEEE Transactions on Computers, C-20(4):392– 396, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671847.

Tabak:1970:ANP

[Tab70a] D. Tabak. An algorithm for nonlinear process stabilization and control. IEEE Transactions on Computers, C-19(6):487– 492, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671558.

Tabak:1970:BCM

[Tab70b] Daniel Tabak. B70-6: Computing Methods in Optimization Problems — 2, L. A. Zadeh, L. W. Neustadt, and A. V. Balakrishnam, Eds. (Academic Press, Inc., N.Y., 1969, 393 pp.). IEEE Transactions on Computers, C-19(10): 986, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671404.

Tabak:1974:CPP

[Tab74] D. Tabak. Comments on “Parallel Processing Algorithms for the Optimal Control of Nonlinear Dynamic Systems”. IEEE Transactions on Computers, C-23(12): 1327, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672454. See [LT73].

Takiyama:1976:RBT

[Tak76] R. Takiyama. A relationship between two types of twolayer threshold-element patternrecognition networks. IEEE Transactions on Computers, C-25(11): 1136–1139, November 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674567.

Tan:1971:SAA

[Tan71] Chung-Jen Tan. State assignments for asynchronous sequential machines. IEEE Transactions on Computers, C-20(4):382– 391, April 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671846.

Tanimoto:1978:OAC

[Tan78] S. L. Tanimoto. An optimal algorithm for computing Fourier texture descriptors. IEEE Transactions on Computers, C-27(1):81– 84, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674956.

Tumbush:1974:SAT

[TB74] G. L. Tumbush and J. E. Brandeberry. A state assignment technique for sequential machines using J–K flip-flops. IEEE Transactions on Computers, C-23(1):85– 86, January 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672374.

Tylaska:1975:IBC

[TB75] T. T. Tylaska and J. D. Bargainer. An improved bound for checking experiments that use simple input-output and characterizing sequences. IEEE Transactions on Computers, C-24(6):670– 673, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672873.

Tzidon:1978:PAF

[TBY78] A. Tzidon, I. Berger, and M. Yoeli.

A practical approach to fault detection in combinational networks. IEEE Transactions on Computers, C-27(10):968–971, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674978.

Toussaint:1970:ARC

[TD70] G. T. Toussaint and R. W. Donaldson. Algorithms for recognizing contour-traced handprinted characters. IEEE Transactions on Computers, C-19(6): 541–546, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671565.

Thayse:1973:BDC

[TD73] A. Thayse and M. Davio. Boolean differential calculus and its application to switching theory. IEEE Transactions on Computers, C-22 (4):409–420, April 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672322.

Thayse:1977:LPU

[TD77] A. Thayse and J.-P. Deschamps. Logic properties of unate discrete and switching functions. IEEE Transactions on Computers, C-26(12):1202–1212, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674781.

Twogood:1976:EEM

[TE76] R. E. Twogood and M. P. Ekstrom. An extension of Eklundh’s matrix transposition algorithm and its application in digital image processing. IEEE Transactions on Computers, C-25(9):950–952, September 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1674721. See [Ekl72].

Trivedi:1977:LAD

[TE77] Kishor S. Trivedi and Miloš D. Ercegovac. On-line algorithms for division and multiplication. IEEE Transactions on Computers, C-26 (7):681–687, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674901.

Teitelman:1976:CCL

[Tei76] W. Teitelman. Clisp: Conversational Lisp. IEEE Transactions on Computers, C-25(4):354– 357, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674617.

Tjaden:1970:DPE

[TF70] G. S. Tjaden and M. J. Flynn.

Detection and parallel execution of independent instructions. IEEE Transactions on Computers, C-19(10):889–895, October 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671388.

Tjaden:1973:RCO

[TF73] Garold S. Tjaden and Michael J.

Flynn. Representation of concurrency with ordering matrices. IEEE Transactions on Computers, C-22 (8):752–761, August 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009154.

Tanaka:1978:ECP

[TF78] E. Tanaka and King-Sun Fu.

Error-correcting parsers for formal languages. IEEE Transactions on Computers, C-27(7):605– 616, July 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675160.

Tam:1972:ASW

[TG72a] Le Dinh Chon Tam and R. Y. Goulet. On arithmetical shift for Walsh functions. IEEE Transactions on Computers, C-21(12): 1451–1452, December 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1672117.

Tou:1972:RHC

[TG72b] J. T. Tou and R. C. Gonzalez. Recognition of handwritten characters by topological feature extraction and multilevel categorization. IEEE Transactions on Computers, C-21(7):776– 785, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672174.

Thomason:1975:SRI

[TG75] M. G. Thomason and R. C. Gonzalez. Syntactic recognition of imperfectly specified patterns. IEEE Transactions on Computers, C-24 (1):93–95, January 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672679.

Tsuchiya:1976:TOH

[TG76] M. Tsuchiya and M. J. Gonzalez. Toward optimization of horizontal microprograms. IEEE Transactions on Computers, C-25(10):

992–999, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674540.

Thompson:1978:UMT

[TG78] R. A. Thompson and F. G. Gray.

Universal modular trees: a design procedure. IEEE Transactions on Computers, C-27(1):53– 63, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674952.

Tront:1979:DMV

[TG79] J. G. Tront and D. D. Givone. A design for multiple-valued logic gates based on MESFET’s. IEEE Transactions on Computers, C-28(11): 854–862, November 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675265.

Trussell:1979:IMM

[TH79] H. J. Trussell and B. R. Hunt. Improved methods of maximum a posteriori restoration. IEEE Transactions on Computers, C-28(1):57– 62, January 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675223.

Thayse:1978:MJD

[Tha78] A. Thayse. Meet and join derivatives and their use in switching theory. IEEE Transactions on Computers, C-27(8):713– 720, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675178.

Therrien:1975:EPP

[The75] C. W. Therrien. Eigenvalue properties of projection operators and their application to the subspace method of feature extraction. IEEE Transactions on Computers, C-24(9):944–948, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672938.

Thomas:1974:OEU

[Tho74] R. T. Thomas. Organization for execution of user microprograms from main memory: Synthesis and analysis. IEEE Transactions on Computers, C-23(8):783– 791, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672627.

Thomason:1975:SSD

[Tho75] M. G. Thomason. Stochastic syntax-directed translation schemata for correction of errors in context-free languages. IEEE Transactions on Computers, C-24(12):1211–1216, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672758.

Thompson:1976:LCU

[Tho76] Richard A. Thompson. Language correction using probabilistic grammars. IEEE Transactions on Computers, C-25(3):275– 286, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009254.

Thompson:1977:TBA

[Tho77] W. B. Thompson. Textural boundary analysis. IEEE Transactions on Computers, C-26(3):272– 276, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674818.

Thompson:1978:GCN

[Tho78] C. D. Thompson. Generalized connection networks for parallel processor intercommunication. IEEE Transactions on Computers, C-27(12):1119–1125, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675014.

Takaoka:1972:FSS

[TI72a] T. Takaoka and T. Ibaraki. N-fail-safe sequential machines. IEEE Transactions on Computers, C-21(11):1189–1196, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672069.

Thorington:1972:ARA

[TI72b] J. M. Thorington, Jr. and J. D. Irwin. An adaptive replacement algorithm for paged-memory computer systems. IEEE Transactions on Computers, C-21(10): 1053–1061, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672044.

Tison:1971:ALS

[Tis71] P. L. Tison. An algebra for logic systems-switching circuits application. IEEE Transactions on Computers, C-20(3):339– 351, March 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671835.

Tseng:1979:IFS

[TJM79] Ben-Dau Tseng, G. A. Jullien, and W. C. Miller. Implementation of FFT structures using the residue number system. IEEE Transactions on Computers, C-28(11):831– 845, November 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675263;   https://ieeexplore.
ieee.org/xpl/RecentIssue.jsp?
punumber=12;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35176.

Tjostheim:1977:RWU

[Tjo77] D. Tjostheim. Recognition of waveforms using autoregressive feature extraction. IEEE Transactions on Computers, C-26(3):268– 270, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674816.

Torode:1974:LMM

[TK74] J. Q. Torode and T. H. Kehl. The logic machine: a modular computer design system. IEEE Transactions on Computers, C-23(11):1164–1169, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672417.

Tokura:1971:FLN

[TKH71] N. Tokura, T. Kasami, and A. Hashimoto. Failsafe logic nets. IEEE Transactions on Computers, C-20(3):323–330, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL https://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&arnumber=
                                        1671832.

Tacker:1973:DHC

[TL73a] E. C. Tacker and T. D. Linton. Digital and hybrid computational aspects of the discrete representation theorem of nonlinear estimation. IEEE Transactions on Computers, C-22(1):61– 67, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672194.

Tang:1973:DCC

[TL73b] D. T. Tang and C. N. Liu. Distance-2 cyclic chaining of constantweight codes. IEEE Transactions on Computers, C-22(2):176– 180, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672274.

Thurber:1970:SDC

[TM70] K. J. Thurber and J. W. Myrna. System design of a cellular APL computer. IEEE Transactions on Computers, C-19(4):291–303, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671509.

Takaoka:1971:FSL

[TM71] T. Takaoka and H. Mine. N-fail-safe logical systems. IEEE Transactions on Computers, C-20(5): 536–542, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671880.

Tabloski:1976:NET

[TM76] T. F. Tabloski and F. J. Mowle. A numerical expansion technique and its application to minimal multiplexer logic circuits. IEEE Transactions on Computers, C-25(7): 684–702, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674678.

Tseng:1978:CDC

[TM78a] B. D. Tseng and W. C. Miller.

On computing the Discrete Cosine Transform. IEEE Transactions on Computers, C-27(10):966– 968, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674977.

Tsuji:1978:DEM

[TM78b] S. Tsuji and F. Matsumoto. Detection of ellipses by a modified Hough transformation. IEEE Transactions on Computers, C-27(8):777– 781, August 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675191.

To:1973:FFI

[To73] Kilin To. Fault folding for irredundant and redundant combinational circuits. IEEE Transactions on Computers, C-22(11):1008–1015, November 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672230. See comments [NA76a].

To:1976:AR

[To76] Kilin To. Author’s reply. IEEE Transactions on Computers, C-25 (2):207, February 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009238.

Tohma:1974:DTF

[Toh74] Y. Tohma. Design technique of fail-safe sequential circuits using flip-flops for internal memory. IEEE Transactions on Computers, C-23(11):1149–1154, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672415.

Toida:1976:SDR

[Toi76] S. Toida. System diagnosis and redundant tests. IEEE Transactions on Computers, C-25(11):1167–1170, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674575.

Tomescu:1970:MMA

[Tom70] Ioan Tomescu. A modified matrix algorithm for determining the complete connection matrix of a switching network. IEEE Transactions on Computers, C-19(1):78– 79, January 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008903.

Tomescu:1972:MMD

[Tom72] I. Tomescu. A matrix method for determining all pairs of compatible states of a sequential machine. IEEE Transactions on Computers, C-21(5):502– 503, May 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672143.

Tomescu:1973:CAS

[Tom73] I. Tomescu. A combinatorial algorithm for solving covering problems. IEEE Transactions on Computers, C-22(2):218–220, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672283.

Tomek:1974:TAP

[Tom74] Ivan Tomek. Two algorithms for piecewise-linear continuous approximation of functions of one variable. IEEE Transactions on Computers, C-23(4):445– 448, April 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672554.

Toma:1975:CLA

[Tom75] C. I. Toma. Cellular logic array for high-speed signed binary number multiplication. IEEE Transactions on Computers, C-24(9):932– 935, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672933.

Tohma:1971:RFS

[TOS71] Y. Tohma, Y. Ohyama, and R. Sakai. Realization of fail-safe sequential machines by using a kout-of-n code. IEEE Transactions on Computers, C-20(11):1270–1275, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671720.

Toussaint:1971:SUB

[Tou71] G. T. Toussaint. Some upper bounds on error probability for multiclass pattern recognition. IEEE Transactions on Computers, C-20 (8):943–944, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671973.

Toussaint:1972:CCS

[Tou72a] Godfried T. Toussaint. Comments on “A Comparison of Seven Techniques for Choosing Subsets of Pattern Recognition Properties”. IEEE Transactions on Computers, C-21(9):1028–1029, September 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009089. See [MG71b].

Toussaint:1972:CEB

[Tou72b] Godfried T. Toussaint. Comments on “Error Bounds for a Contextual Recognition Procedure”. IEEE Transactions on Computers, C-21 (9):1027, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009087. See [Chu71].

Toussaint:1972:CTC

[Tou72c] Godfried T. Toussaint. Comments on “Theoretical Comparison of a Class of Feature Selection Criteria in Pattern Recognition”. IEEE Transactions on Computers, C-21 (6):615–616, June 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/

                                        stamp/stamp.jsp?tp=&arnumber=

5009024. See [Che71a].

Toussaint:1972:PRC

[Tou72d] Godfried T. Toussaint. Polynomial representation of classifiers with independent discretevalued features. IEEE Transactions on Computers, C-21(2):205– 208, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008928.

Toussaint:1972:SIB

[Tou72e] Godfried T. Toussaint. Some inequalities between distance measures for feature evaluation. IEEE Transactions on Computers, C-21 (4):409–410, April 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008991.

Toussaint:1974:CEP

[Tou74] G. T. Toussaint. Comments on “The Extraction of Pattern Features from Imperfectly Identified Samples”. IEEE Transactions on Computers, C-23(9):988–989, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672662. See [Bab72].

Toussaint:1975:CNC

[Tou75] G. T. Toussaint. Comments on “On a New Class of Bounds on Bayes’ Risk in Multihypothesis Pattern Recognition”. IEEE Transactions on Computers, C-24(8):855– 856, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672917. See [Dev74].

Toy:1971:MLC

[Toy71a] W. N. Toy. Modular LSI control logic design with error detection. IEEE Transactions on Computers, C-20(2):161–166, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671799.

Toy:1971:NPB

[Toy71b] W. N. Toy. A novel parallel binary counter design with parity prediction and error detection scheme. IEEE Transactions on Computers, C-20(1):44– 48, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671673.

Thedchanamoorthy:1972:VMD

[TP72] N. Thedchanamoorthy and J. B.

Plant. A versatile multiplying digital-to-analog converter. IEEE Transactions on Computers, C-21(10):1113–1116, October 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672051.

Thurber:1973:FPP

[TP73] K. J. Thurber and P. C. Patton. The future of parallel processing. IEEE Transactions on Computers, C-22(12):1140–1143, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672255.

Trivedi:1976:PAA

[Tri76] K. S. Trivedi. Prepaging and applications to array algorithms. IEEE Transactions on Computers, C-25(9):915–921, September 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1674716.

Trivedi:1977:PPA

[Tri77a] K. S. Trivedi. On the paging performance of array algorithms. IEEE Transactions on Computers, C-26(10):938–947, October 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674729.

Trivedi:1977:UCF

[Tri77b] K. S. Trivedi. On the use of continued fractions for digital computer arithmetic. IEEE Transactions on Computers, C-26(7):

700–704, July 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674903. See corrections [Tri78].

Trivedi:1978:CUC

[Tri78] K. S. Trivedi. Corrections to “On the Use of Continued Fractions for Digital Computer Arithmetic”. IEEE Transactions on Computers, C-27(3):288, March 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675093. See [Tri77b].

Troxel:1975:SIM

[Tro75] D. E. Troxel. Serial interfaces for minicomputers. IEEE Transactions on Computers, C-24(10): 1027–1028, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672711.

Trunk:1971:CAF

[Tru71] G. V. Trunk. Comments on “An Algorithm for Finding Intrinsic Dimensionality of Data”. IEEE Transactions on Computers, C-20(12): 1615, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671779. See [FO71a].

Trunk:1976:SEI

[Tru76] Gerard V. Trunk. Stastical estimation of the intrinsic dimensionality of a noisy signal collection. IEEE Transactions on Computers, C-25(2):165–171, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009231.

Taniguchi:1975:NVF

[TS75] K. Taniguchi and T. Sakai. A new voltage to frequency converter. IEEE Transactions on Computers, C-24(10):1035–1036, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672715.

Toueg:1979:DSD

[TS79] S. Toueg and K. Steiglitz. The design of small-diameter networks by local search. IEEE Transactions on Computers, C-28(7):537– 542, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675402.

Tsichritzis:1970:RCA

[Tsi70] D. Tsichritzis. R70-35 classes of automata and transitive closure. IEEE Transactions on Computers, C-19(9):858–859, September 1970. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671656.

Tung:1970:ACP

[Tun70a] Chin Tung. On the apparent continuity of processing in a paging environment. IEEE Transactions on Computers, C-19(11):1047–1054, November 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671424.

Tung:1970:SDD

[Tun70b] Chin Tung. Signed-digit division using combinational arithmetic nets. IEEE Transactions on Computers, C-19(8):746–748, August 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

www.acsel-lab.com/arithmetic/
arith1/papers/ARITH1_Tung.pdf;
https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671617.

Turner:1970:RSD

[Tur70] J. Turner. R70-22 subdirect decompositions of transformation graphs. IEEE Transactions on Computers, C-19(7):659, July 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671602.

Toussaint:1972:CFS

[TV72] Godfried T. Toussaint and Toomas R. Vilmansen. Comments on “Feature Selection with a Linear Dependence Measure”. IEEE Transactions on Computers, C-21(4): 408, April 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008989. See [Das71a].

Turcat:1976:RTC

[TV76] C. Turcat and A. Verdillon. Recursion and testing of combinational circuits. IEEE Transactions on Computers, C-25(6):652– 654, June 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674667.

Turcat:1979:SAT

[TV79] C. Turcat and A. Verdillon. Symmetry, automorphism, and test. IEEE Transactions on Computers, C-28 (4):319–325, April 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675354.

Turn:1976:PSI

[TW76] R. Turn and W. H. Ware. Privacy and security issues in information systems. IEEE Transactions on Computers, C-25(12):1353–1361, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674604.

Torng:1977:OIC

[TW77] H. C. Torng and N. C. Wilhelm.

The optimal interconnection of circuit modules in microprocessor and digital system design. IEEE Transactions on Computers, C-26(5): 450–457, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674862.

Tang:1978:OTN

[TWB78] D. T. Tang, L. S. Woo, and L. R. Bahl. Optimization of teleprocessing networks with concentrators and multiconnected terminals. IEEE Transactions on Computers, C-27(7):594–604, July 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675159.

Ulug:1974:UTA

[UB74] M. E. Ulug and B. A. Bowen. A unified theory of the algebraic topological methods for the synthesis of switching systems. IEEE Transactions on Computers, C-23(3):255– 267, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672513.

Underwood:1975:VLM

[UC75] S. A. Underwood and C. L. Coates, Jr. Visual learning from multiple views. IEEE Transactions on Computers, C-24(6):651– 661, June 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672870.

Uhr:1972:LRC

[Uhr72] L. Uhr. Layered “Recognition Cone” networks that preprocess, classify, and describe. IEEE Transactions on Computers, C-21(7): 758–768, July 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672172.

Uhrig:1973:BDP

[Uhr73] J. Uhrig. B73-3 dynamic programming and partial differential equations. IEEE Transactions on Computers, C-22(2):221, February 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672284.

Ullmann:1972:TLB

[Ull72] J. R. Ullmann. Transference of learning between recognition classes. IEEE Transactions on Computers, C-21(2):219–220, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008932.

Ulman:1970:CHT

[Ulm70] L. J. Ulman. Computation of the Hadamard transform and the Rtransform in ordered form. IEEE Transactions on Computers, C-19 (4):359–360, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671518.

Ulrich:1971:NSS

[Ulr71] J. W. Ulrich. A note on the solution of sequential Boolean equations. IEEE Transactions on Computers, C-20(2):231–234, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671815.

Udupa:1977:NCT

[UM77] K. J. Udupa and I. S. N. Murthy.

New concepts for three-dimensional shape analysis. IEEE Transactions on Computers, C-26(10): 1043–1049, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674744.

Unger:1971:ASS

[Ung71] S. H. Unger. Asynchronous sequential switching circuits with unrestricted input changes. IEEE Transactions on Computers, C-20 (12):1437–1444, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671748.

Unger:1977:CRA

[Ung77a] B. W. Unger. A computer resource allocation model with some measured and simulation results. IEEE Transactions on Computers, C-26 (3):243–259, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674813.

Unger:1977:SSC

[Ung77b] S. H. Unger. Self-synchronizing circuits and nonfundamental mode operation. IEEE Transactions on Computers, C-26(3):278–281, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674820.

Unger:1977:TRI

[Ung77c] S. H. Unger. Tree realizations of iterative circuits. IEEE Transactions on Computers, C-26(4):365– 383, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674846.

Unger:1977:GCS

[Ung77d] Stephen H. Unger. The generation of completion signals in iterative combinational circuits. IEEE Transactions on Computers, C-26 (1):13–18, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009268.

Usas:1975:TSC

[Usa75] A. M. Usas. A totally self-checking checker design for the detection of errors in periodic signals. IEEE Transactions on Computers, C-24 (5):483–489, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672843.

Vairavan:1971:RCI

[Vai71] K. Vairavan. A remark on the concepts of input-memory and output-memory of sequential machines. IEEE Transactions on Computers, C-20(2):239, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671817.

Vairavan:1972:MIM

[Vai72] K. Vairavan. Minimal inputmemory and output-memory finitestate machines. IEEE Transactions on Computers, C-21(6):598– 602, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009017.

VanVoorhis:1972:ILB

[Van72] David C. Van Voorhis. An improved lower bound for sorting networks. IEEE Transactions on Computers, C-21(6):612– 613, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009021.

VanLantschoot:1973:DAB

[Van73a] E. J. Van Lantschoot. Double adjacencies between cycles of a circulating shift register. IEEE Transactions on Computers, C-22(10): 944–955, October 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672215.

VanTill:1973:CDG

[Van73b] J. W. J. Van Till. Comments on “On the Definition and Generation of Walsh Functions”. IEEE Transactions on Computers, C-22 (7):702–703, July 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

5009139. See [Dav72a].

Vandewalle:1975:CPL

[Van75] J. Vandewalle. On the calculation of the piecewise linear approximation to a discrete function. IEEE Transactions on Computers, C-24 (8):843–846, August 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672913.

VanVoorhis:1977:CCA

[Van77] D. C. Van Voorhis. Comments on “A Computer Algorithm for Transposing Nonsquare Matrices”. IEEE Transactions on Computers, C-26 (6):607–608, June 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674887. See [All75a].

Varshney:1979:AMI

[Var79] P. K. Varshney. On analytical modeling of intermittent faults in digital systems. IEEE Transactions on Computers, C-28(10):786– 791, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675247.

Vairavan:1976:CCG

[VD76a] K. Vairavan and R. A. Demillo.

On the computational complexity of a generalized scheduling problem. IEEE Transactions on Computers, C-25(11):1067–1073, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674556.

VanHaersmaBuma:1976:CCP

[VD76b] Chris E. Van Haersma Buma and Robert P. W. Duin. Computation of concave piecewise linear discriminant functions using Chebyshev polynomials. IEEE Transactions on Computers, C-25(2):178– 186, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009233.

Veelenturf:1978:ISM

[Vee78] L. P. J. Veelenturf. Inference of sequential machines from sample computations. IEEE Transactions on Computers, C-27(2):167– 170, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675053.

Vichnevetsky:1971:MDU

[Vic71] R. Vichnevetsky. The method of decomposition for unstable two-point boundary value problems. IEEE Transactions on Computers, C-20 (8):910–914, August 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671962.

Vichnevetsky:1974:BCS

[Vic74] R. Vichnevetsky. B74-24 computer simulation of dynamic systems. IEEE Transactions on Computers, C-23(3):335–336, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672533.

Vilfan:1970:IUB

[Vil70] B. Vilfan. An improved upper bound for the finite delay of graphs. IEEE Transactions on Computers, C-19(5):453– 456, May 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671537.

Vilmansen:1972:DDP

[Vil72] Toomas R. Vilmansen. On dependence and discrimination in pattern recognition. IEEE Transactions on Computers, C-21(9):1029– 1031, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009091.

Vilfan:1973:APT

[Vil73a] B. Vilfan. Another proof of the two-dimensional Cayley–Hamilton theorem. IEEE Transactions on Computers, C-22(12):1140, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672253.

Vilmansen:1973:FEM

[Vil73b] T. R. Vilmansen. Feature evaluation with measures of probabilistic dependence. IEEE Transactions on Computers, C-22(4):381– 388, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672318.

Vilmansen:1974:AMI

[Vil74] T. Vilmansen. On an approximation of mutual information. IEEE Transactions on Computers, C-23(12):1311–1313, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672446.

Vink:1978:MTN

[Vin78] H. A. Vink. Minimal TANT networks of functions with DON’T CARE’S and some complemented input variables. IEEE Transactions on Computers, C-27(11):1073–1078, November 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675000.

Vranesic:1970:MVA

[VLS70] Z. G. Vranesic, E. S. Lee, and K. C.

Smith. A many-valued algebra for switching systems. IEEE Transactions on Computers, C-19(10):964– 971, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671396.

VanVoorhis:1978:MSI

[VM78] D. C. Van Voorhis and T. H.

Morrin. Memory systems for image processing. IEEE Transactions on Computers, C-27(2):113– 125, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675045.

VanCampenhout:1978:SMC

[VN78] J. M. Van Campenhout and P. Notredame. A stochastic model for closed-loop preemptive microprocessor I/O organizations. IEEE Transactions on Computers, C-27(12):1126–1136, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675015.

Voith:1974:BDD

[Voi74] R. P. Voith. B 74-32 digital design. IEEE Transactions on Computers, C-23(5):558– 559, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672581.

Voith:1975:BCO

[Voi75] R. P. Voith. B75-4 computer organization and programming. IEEE Transactions on Computers, C-24(2):222–223, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672788.

Voith:1977:MUL

[Voi77a] R. P. Voith. Minimum universal logic module sequential circuits with decoders. IEEE Transactions on Computers, C-26(10): 1032–1035, October 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674741.

Voith:1977:UIM

[Voi77b] R. P. Voith. ULM implicants for minimization of universal logic module circuits. IEEE Transactions on Computers, C-26(5):417– 424, May 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674858.

VonPuttkamer:1975:SHB

[Von75] E. Von Puttkamer. A simple hardware buddy system memory allocator. IEEE Transactions on Computers, C-24(10):953– 957, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672693.

Vanderbrug:1977:TST

[VR77] G. J. Vanderbrug and A. Rosenfeld. Two-stage template matching. IEEE Transactions on Computers, C-26(4):384–393, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674847.

Vlasenko:1979:UMT

[VR79] V. Vlasenko and K. R. Rao. Unified matrix treatment of discrete transforms. IEEE Transactions on Computers, C-28(12):934–938, December 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675284.

Vranesic:1977:MVL

[Vra77] Z. G. Vranesic. Multiple-valued logic: An introduction and overview. IEEE Transactions on Computers, C-26(12):1181–1182, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674778.

Vinea:1971:DCF

[VV71] A. Vinea and V. Vinea. A distance criterion for figural pattern recognition. IEEE Transactions on Computers, C-20(6):680– 685, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671919.

VanLantschoot:1974:DWC

[VV74] E. J. Van Lantschoot and J. P. Vandewalle. Design of weighted counters with rational scale using continued fraction expansion. IEEE Transactions on Computers, C-23 (3):232–238, March 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672510.

Vink:1978:RCT

[VVA78] H. A. Vink, B. Van Den Dolder, and J. Al. Reduction of CC-tables using multiple implication. IEEE Transactions on Computers, C-27(10): 961–966, October 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674976.

Vranesic:1972:FTS

[VW72] Z. G. Vranesic and K. M. Waliuzzaman. Functional transformation in simplification of multivalued switching functions. IEEE Transactions on Computers, C-21(1):102– 105, January 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672033.

Williams:1973:ETL

[WA73] M. J. Y. Williams and J. B. Angell. Enhancing testability of largescale integrated circuits via test points and additional logic. IEEE Transactions on Computers, C-22 (1):46–60, January 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672193.

Wadel:1971:CNR

[Wad71] L. B. Wadel. Comment on “Negative Radix Conversion”. IEEE Transactions on Computers, C-20 (5):587, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671893. See [Paw71, Zoh70].

Wagner:1970:RTA

[Wag70] E. G. Wagner. R70-32 tessellation automata. IEEE Transactions on Computers, C-19(9):857– 858, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671653.

Wait:1970:SVT

[Wai70] J. V. Wait. State-variable techniques for digital simulation of bandpass systems. IEEE Transactions on Computers, C-19(9):830– 831, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671641.

Waksman:1970:CI

[Wak70a] A. Waksman. On the complexity of inversions. IEEE Transactions on Computers, C-19(12):1225–1226, December 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671459.

Waksman:1970:WAI

[Wak70b] A. Waksman. On Winograd’s algorithm for inner products. IEEE Transactions on Computers, C-19 (4):360–361, April 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671519.

Waksman:1970:RGF

[Wak70c] A. Waksman. R-70-31 A generalized firing squad problem. IEEE Transactions on Computers, C-19 (8):766, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=
1671629.

Wakerly:1974:PSC

[Wak74] J. F. Wakerly. Partially selfchecking circuits and their use in performing logical operations. IEEE Transactions on Computers, C-23(7):658–666, July 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1672607.

Wakerly:1975:CAS

[Wak75a] J. F. Wakerly. Comments on “Asynchronous Sequential Machines Designed for Fault Detection”. IEEE Transactions on Computers, C-24 (7):759–760, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672895. See [SM74].

Wakerly:1975:DUM

[Wak75b] J. F. Wakerly. Detection of unidirectional multiple errors using lowcost arithmetic codes. IEEE Transactions on Computers, C-24(2): 210–212, February 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672780.

Wakerly:1975:TFT

[Wak75c] J. F. Wakerly. Transient failures in triple modular redundancy systems with sequential modules. IEEE Transactions on Computers, C-24 (5):570–573, May 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672856.

Wallach:1974:BCA

[Wal74] S. J. Wallach. B 74-39 A course in APL/360 with applications. IEEE Transactions on Computers, C-23(9):990–991, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672666.

Walker:1976:SUT

[Wal76] Donald E. Walker. Speech understanding through syntactic and semantic analysis. IEEE Transactions on Computers, C-25(4):432– 439, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674625.

Wang:1971:ANS

[Wan71] K. C. Wang. On autonomous NOR sequential machines. IEEE Transactions on Computers, C-20 (5):580–582, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671888.

Wang:1972:BDS

[Wan72] C. P. Wang. B72-21 data structures and management. IEEE Transactions on Computers, C-21(11): 1250, November 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672081.

Wang:1975:AGT

[Wan75a] D. T. Wang. An algorithm for the generation of test sets for combinational logic networks. IEEE Transactions on Computers, C-24 (7):742–746, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672888.

Wang:1975:PFC

[Wan75b] D. T. Wang. Properties of faults and criticalities of values under tests for combinational networks. IEEE Transactions on Computers, C-24 (7):746–750, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672889.

Wang:1975:CBP

[Wan75c] Y. R. Wang. Characterization of binary patterns and their projections. IEEE Transactions on Computers, C-24(10):1032–1035, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672714.

Wang:1977:CSR

[Wan77] L. Wang. Cross-section reconstruction with a fan-beam scanning geometry. IEEE Transactions on Computers, C-26(3):264– 268, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674815.

Warner:1976:CCT

[WBRG76] Harold Warner, Charles L. Bell, Duane M. Rumbaugh, and Timothy V. Gill. Computercontrolled teaching instrumentation for linguistic studies with the Great Apes. IEEE Transactions on Computers, C-25(1):38– 43, January 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009203.

Wong:1974:PGB

[WC74] C. K. Wong and Shi-Kuo Chang. Parallel generation of binary search trees. IEEE Transactions on Computers, C-23(3):268– 271, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672514.

Wong:1976:GPM

[WC76] C. K. Wong and Don Coppersmith.

The generation of permutations in magnetic bubble memories. IEEE Transactions on Computers, C-25 (3):254–262, March 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009252.

Wernecke:1977:MEI

[WD77] S. J. Wernecke and L. R. D’Addario.

Maximum entropy image reconstruction. IEEE Transactions on Computers, C-26(4):351– 364, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674845.

Weaver:1975:SPT

[Wea75] C. S. Weaver. Some properties of threshold logic unit pattern recognition networks. IEEE Transactions on Computers, C-24(3):290– 298, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672802.

Wegbreit:1972:SEL

[Weg72] Ben Wegbreit. A space-efficient list structure tracing algorithm. IEEE Transactions on Computers, C-21(9):1009–1010, September 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=5009079.

Wegner:1976:PLF

[Weg76] P. Wegner. Programming languages — the first 25 years. IEEE Transactions on Computers, C-25(12):1207–1225, December 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674589.

Wehrfritz:1974:TTL

[Weh74] H. C. Wehrfritz. Techniques for the transformation of logic equations. IEEE Transactions on Computers, C-23(5):477–480, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1672563.

Weiss:1972:BLT

[Wei72] C. Dennis Weiss. Bounds on the length of terminal stuckfault test. IEEE Transactions on Computers, C-21(3):305– 309, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5008955.

Weiss:1975:NPD

[Wei75] H. Weiss. Nonredundant point distribution for coded aperture imaging with application to threedimensional on-line X-ray information retrieving. IEEE Transactions on Computers, C-24(4):391– 394, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672825.

Welch:1974:CMC

[Wel74] T. A. Welch. Comments on “A Measure of Computational Work”. IEEE Transactions on Computers, C-23(2):208, February 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672486.

See [Hel72, Hel74, Mas78].

Welch:1978:MHC

[Wel78] T. A. Welch. Memory hierarchy configuration analysis. IEEE Transactions on Computers, C-27 (5):408–413, May 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675120.

Wesselkamper:1978:DDM

[Wes78] T. C. Wesselkamper. Divided difference methods for Galois switching functions. IEEE Transactions on Computers, C-27(3):232– 238, March 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675076.

Wetjen:1978:CBB

[Wet78] A. Wetjen. Comments on “A Branch and Bound Clustering Algorithm”. IEEE Transactions on Computers, C-27(2):189, February 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1675059. See [KNF75].

Wexelblatt:1971:BPS

[Wex71] R. L. Wexelblatt. B71-1 problem solving by digital computers with PL/I programming. IEEE Transactions on Computers, C-20 (8):948, August 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671975.

Warmack:1973:AOS

[WG73] R. E. Warmack and R. C. Gonzalez. An algorithm for the optimal solution of linear inequalities and its application to pattern recognition. IEEE Transactions on Computers, C-22(12):1065–1075, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672245.

Waksman:1974:CRP

[WG74] A. Waksman and M. W. Green. On the consecutive retrieval property in file organization. IEEE Transactions on Computers, C-23(2):173– 174, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672473.

Will:1975:ESC

[WG75] P. M. Will and D. D. Grossman. An experimental system for computer controlled mechanical assembly. IEEE Transactions on Computers, C-24(9):879– 888, September 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672926.

Wendling:1978:SIW

[WGS78] S. Wendling, G. Gagneux, and G. Stamon. A set of invariants within the power spectrum of unitary transformations. IEEE Transactions on Computers, C-27(12):1213–1216, December 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675030.

Wong:1978:SHS

[WH78] R. Y. Wong and E. L. Hall. Sequential hierarchical scene matching. IEEE Transactions on Computers, C-27(4):359–366, April 1978. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1675108.

White:1970:FDC

[Whi70] R. C. White, Jr. A fast digital computer method for recursive estimation of the mean. IEEE Transactions on Computers, C-19(9):847– 850, September 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671646. See comments [Maj72b].

White:1971:FES

[Whi71a] H. S. White. Finding events in a sea of bubbles. IEEE Transactions on Computers, C-20(9):988– 995, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671986.

Whitney:1971:DMN

[Whi71b] A. W. Whitney. A direct method of nonparametric measurement selection. IEEE Transactions on Computers, C-20(9):1100–1103, September 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672003.

Whitney:1971:AFA

[Whi71c] G. E. Whitney. Algebraic fault analysis for constrained combinational networks. IEEE Transactions on Computers, C-20(2):141– 148, February 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671796.

White:1972:CNM

[Whi72] I. White. Comment on “A Nonlinear Mapping for Data Structure Analysis”. IEEE Transactions on Computers, C-21(2):220– 221, February 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5008933. See [Kru71].

Wilkes:1970:SEL

[Wil70] M. A. Wilkes. Scroll editing:

An on-line algorithm for manipulating long character strings. IEEE Transactions on Computers, C-19(11):1009–1015, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1671418.

Wilkes:1971:SMS

[Wil71] M. V. Wilkes. Slave memories and segmentation. IEEE Transactions on Computers, C-20(6): 674–675, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671916.

Williams:1975:UDI

[Wil75] G. H. Williams. Uniform decomposition of incompletely specified sequential machines. IEEE Transactions on Computers, C-24(8):840– 843, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672912.

Williams:1977:SIA

[Wil77a] R. P. Williams. Serial integer arithmetic with magnetic bubbles. IEEE Transactions on Computers, C-26 (3):260–264, March 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674814.

Williams:1977:CLM

[Wil77b] T. P. T. Williams. Comments on “A Least Mean Squares Cubic Algorithm for On-Line Differential of Sampled Analog Signals”. IEEE Transactions on Computers, C-26 (4):416, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674852. See [All75b].

Wilkens:1978:RSM

[Wil78] E. J. Wilkens. Realizations of sequential machines using random access memory. IEEE Transactions on Computers, C-27(5):429– 441, May 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675123.

Winder:1970:RMS

[Win70a] R. O. Winder. R70-13 multigate synthesis of general Boolean functions by threshold logic elements. IEEE Transactions on Computers, C-19(6):565, June 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671577.

Winder:1970:TLA

[Win70b] R. O. Winder. Threshold logic asymptotes. IEEE Transactions on Computers, C-19(4):349– 353, April 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671514.

Winzer:1975:CRC

[Win75] G. Winzer. Character recognition with a coherent optical multichannel correlator. IEEE Transactions on Computers, C-24(4):419– 423, April 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672830.

Winker:1976:EIQ

[Win76] S. K. Winker. An evaluation of an implementation of qualified hyperresolution. IEEE Transactions on Computers, C-25(8):835– 843, August 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674702.

Witsenhausen:1970:HSP

[Wit70] H. S. Witsenhausen. Hybrid solution of partial differential equations. IEEE Transactions on Computers, C-19(9):851, September 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671648.

Webb:1971:SCF

[WK71] S. R. Webb and J. D. Kennedy. Some comments on the Fast Hadamard Transform of order twelve. IEEE Transactions on Computers, C-20(5):588–590, May 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671898.

Wasserstorm:1973:ASN

[WK73] E. Wasserstorm and A. Kirszenblat. Analog solutions of nonlinear boundary-value problems by the continuation method. IEEE Transactions on Computers, C-22(11): 966–970, November 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672225.

Weiler:1970:RTO

[WKD70] P. W. Weiler, R. S. Kopp, and R. G. Dorman. A real-time operating system for manned spaceflight. IEEE Transactions on Computers, C-19 (5):388–398, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671529.

Wade:1975:NEA

[WKP75] G. Wade, H. Keyani, and Soo-Chang Pei. New experiments and analysis in Bragg-diffraction imaging. IEEE Transactions on Computers, C-24(4):395–401, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672826.

Wang:1974:WPS

[WL74] Tse Lin Wang and Chao-Kai Liu. Weight-preserved singleerror-correcting scheme for binary adders. IEEE Transactions on Computers, C-23(10):1002–1007, October 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672394.

Wong:1975:TDF

[WL75] A. K. C. Wong and T. S. Liu. Typicality, diversity, and feature pattern of an ensemble. IEEE Transactions on Computers, C-24(2):158– 181, February 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672776.

Wong:1977:DDC

[WL77] Andrew K. C. Wong and T. S. Liu. A decision-directed clustering algorithm for discrete data. IEEE Transactions on Computers, C-26 (1):75–82, January 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009277.

Wojcik:1971:CBA

[WM71] A. S. Wojcik and G. Metze. On the cost of base N adders. IEEE Transactions on Computers, C-20(10): 1196–1203, October 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671698.

Wickersham:1974:SAS

[WM74] R. S. Wickersham and G. K. Maki. Safe asynchronous sequential circuits. IEEE Transactions on Computers, C-23(5):494– 500, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672565.

Wilson:1976:RRS

[WM76] G. A. Wilson and J. Minker. Resolution, refinements, and search strategies: a comparative study. IEEE Transactions on Computers, C-25(8):782–801, August 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1674697.

Waterman:1976:IIT

[WN76] D. A. Waterman and A. Newell.

PAS-II: An interactive task-free version of an automatic protocol analysis system. IEEE Transactions on Computers, C-25(4):402– 413, April 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674622.

Weszka:1974:TST

[WNR74] J. S. Weszka, R. N. Nagel, and A. Rosenfeld. A threshold selection technique. IEEE Transactions on Computers, C-23(12):1322–1326, December 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672451.

Wood:1979:HDP

[Woo79] R. A. Wood. A high density programmable logic array chip. IEEE Transactions on Computers, C-28(9):602–608, September 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675427.

Worland:1976:PMN

[Wor76] P. B. Worland. Parallel methods for the numerical solution of ordinary differential equations. IEEE Transactions on Computers, C-25(10): 1045–1048, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674545.

Wormald:1977:NSI

[Wor77] E. G. Wormald. A note on synchronizer or interlock maloperation. IEEE Transactions on Computers, C-26(3):317–318, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674833. See comments [Cha79, Wor79].

Wormald:1979:SJC

[Wor79] E. G. Wormald. Support for T. J. Chaney’s comments on “A Note on Synchronizer or Interlock Maloperation”. IEEE Transactions on Computers, C-28(10): 804, October 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1675253. See [Wor77, Cha79].

Waxman:1977:LCC

[WR77] J. Waxman and J. Rootenberg. Logic circuit for cycle detection in a state diagram. IEEE Transactions on Computers, C-26(3):303– 305, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674826.

Wright:1973:TSS

[Wri73] Thomas J. Wright. A twospace solution to the hidden line problem for plotting functions of two variables. IEEE Transactions on Computers, C-22(1):28– 33, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672190.

Williams:1976:SBC

[WS76] Stephen Williams and Ivor R.

Smith. SCR bridge convertor computation using tensor methods. IEEE Transactions on Computers, C-25(1):1–6, January 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009197.

Walker:1979:RSM

[WSB79] W. K. S. Walker, C.-E. W. Sundberg, and C. J. Black. A reliable spaceborne memory with a single error and erasure correction scheme. IEEE Transactions on Computers, C-28(7):493– 500, July 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675394.

Wulf:1974:BPT

[Wul74] W. A. Wulf. B 74-27 program test methods. IEEE Transactions on Computers, C-23(5):555– 556, May 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672576.

Wong:1973:ACC

[WY73] C. K. Wong and P. C. Yue. The anticipatory control of a cyclically permutable memory. IEEE Transactions on Computers, C-22(5): 481–488, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672344.

Wakimura:1975:CCE

[WY75] Y. Wakimura and N. Yoshida. Comments on “Checking Experiments for Sequential Machines”. IEEE Transactions on Computers, C-24(11):1141–1142, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672743. See [Hsi71b].

Wyman:1973:LDD

[Wym73] Robert H. Wyman. Loading of decoder drivers. IEEE Transactions on Computers, C-22(9):878– 879, September 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009189.

Wyner:1971:NCC

[Wyn71] A. D. Wyner. Note on circuits and chains of spread k in the n-cube. IEEE Transactions on Computers, C-20(4):474, April 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671865.

Williams:1972:CSM

[WZ72] George H. Williams and E. G.

Zavisca. Comments on “Sequential Machine Identification”. IEEE Transactions on Computers, C-21 (6):616, June 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

5009026. See [Kel71].

Yamada:1970:RRE

[Yam70] H. Yamada. R70-11 representation of events in the von Neumann cellular model. IEEE Transactions on Computers, C-19(6): 563–564, June 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671575.

Yang:1973:CPM

[Yan73] Chao-Chih Yang. Closure partition method for minimizing incomplete sequential machines.

IEEE Transactions on Computers, C-22(12):1109–1122, December 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672250.

Yang:1974:MDP

[Yan74a] C. C. Yang. On the modeling of demand paging algorithms by finite automata. IEEE Transactions on Computers, C-23(8):870– 874, August 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672635.

Yang:1974:GAC

[Yan74b] Chao-Chih Yang. Generation of all closed partitions on the state set of a sequential machine. IEEE Transactions on Computers, C-23 (5):530–533, May 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672569.

Yang:1975:ETA

[Yan75a] Chao-Chih Yang. On the equivalence of two algorithms for finding all maximal compatibles. IEEE Transactions on Computers, C-24(10):977–979, October 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:

//ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672696.

Yang:1975:SPM

[Yan75b] Chao-Chih Yang. Structural preserving morphisms of finite automata and an application to graph isomorphism. IEEE Transactions on Computers, C-24(11):1133–1139, November 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672741.

Yang:1975:CCP

[YB75] Chao-Chih Yang and M. K. Babinski. Comments on “Closure Partition Method for Minimizing Incomplete Sequential Machines”. IEEE Transactions on Computers, C-24(1):106–108, January 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672683. See [RB76].

Yau:1976:DMA

[YC76] S. S. Yau and J. Chung. On the design of modulo arithmetic units based on cyclic groups. IEEE Transactions on Computers, C-25(11):1057–1067, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674555.

Yeh:1970:RLF

[Yeh70] R. T. Yeh. R70-43 lattice functions, pair algebras, and finitestate machines. IEEE Transactions on Computers, C-19(10): 989, October 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671409.

Yenersoy:1979:SAM

[Yen79] O. Yenersoy. Synthesis of asynchronous machines using mixedoperation mode. IEEE Transactions on Computers, C-28(4):325– 329, April 1979. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1675355.

Yajima:1974:PMP

[YI74] S. Yajima and K. Inagaki. Power minimization problems of logic networks. IEEE Transactions on Computers, C-23(2):153–165, February 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672471.

Yip:1976:SAZ

[Yip76] P. C. Y. Yip. Some aspects of the zoom transform. IEEE Transactions on Computers, C-25(3):287– 296, March 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009255.

Yelowitz:1976:NRT

[YK76] L. Yelowitz and A. Kandel. New results and techniques in resolution theory. IEEE Transactions on Computers, C-25(7):673– 677, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674676.

Yau:1973:ECR

[YL73] S. S.-S. Yau and Yu-Cheng Liu. Error correction in redundant residue number systems. IEEE Transactions on Computers, C-22(1):5– 11, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672187;   https://ieeexplore.
ieee.org/xpl/tocresult.jsp?isnumber=
35069.

Yang:1978:CSC

[YM78] Chao-Chin Yang and C. P. May.

A correction and some comments concerning graph isomorphism by finite automata. IEEE Transactions on Computers, C-27(1):95– 96, January 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674961.

Yau:1970:FDR

[YO70] S. S. Yau and M. Orsic. Fault diagnosis and repair of cutpoint cellular arrays. IEEE Transactions on Computers, C-19(3):259– 262, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671499.

Yoeli:1970:RPA

[Yoe70a] M. Yoeli. R70-26 probabilistic aspects of machine decomposition theory. IEEE Transactions on Computers, C-19(8): 764, August 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671624.

Yoeli:1970:SMC

[Yoe70b] M. Yoeli. The synthesis of multivalued cellular cascades. IEEE Transactions on Computers, C-19 (11):1089–1090, November 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=
1671430.

Yohe:1973:RFP

[Yoh73] J. Michael Yohe. Roundings in floating-point arithmetic. IEEE Transactions on Computers, C-22(6):577–586, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://www.acsel-lab.com/
arithmetic/arith2/papers/ARITH2_
Yohe.pdf;   https://ieeexplore.
ieee.org/stamp/stamp.jsp?tp=&
arnumber=5009110.

Young:1970:RAM

[You70] P. C. Young. R70-8 analog methods for on-line system identification using noisy measurements. IEEE Transactions on Computers, C-19 (5):465–466, May 1970. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671552.

Young:1971:RLF

[You71] T. Y. Young. The reliability of linear feature extractors. IEEE Transactions on Computers, C-20(9): 967–971, September 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671983.

Yue:1975:TRM

[YRT75] On-Ching Yue, E. L. Rope, and G. Tricoles. Two reconstruction methods for microwave imaging of buried dielectric anomalies. IEEE Transactions on Computers, C-24 (4):381–390, April 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672824.

Yau:1970:ULM

[YT70] S. S. Yau and C. K. Tang. Universal logic modules and their applications. IEEE Transactions on Computers, C-19(2):141–149, February 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1671472. See comments [Ale71].

Yau:1971:IRS

[YT71a] S. S. Yau and Y. S. Tang. On identification of redundancy and symmetry of switching functions. IEEE Transactions on Computers, C-20(12):1609–1613, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671776.

Yau:1971:TAS

[YT71b] S. S. Yau and Y. S. Tang. Transformation of an arbitrary switching function to a totally symmetric function. IEEE Transactions on Computers, C-20(12):1606–1609, December 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671775.

Yau:1971:EAG

[YT71c] S. S. Yau and Yu-Shan Tang. An efficient algorithm for generating complete test sets for combinational logic circuits. IEEE Transactions on Computers, C-20(11):1245–1251, November 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671716.

Yang:1976:ADA

[YT76] Chao-Chih Yang and Marilyn A.

Tarpy. An algorithm for deriving all pairs of compatible states by closure classes. IEEE Transactions on Computers, C-25(2):202– 207, February 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009236.

Yachida:1977:VMV

[YT77] M. Yachida and S. Tsuji. A versatile machine vision system for complex industrial parts. IEEE Transactions on Computers, C-26(9):882– 894, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674936.

Yuan:1974:MMI

[Yua74] Fu-Min Yuan. Minimal memory inverses of linear sequential circuits. IEEE Transactions on Computers, C-23(11):1155–1163, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672416.

Yuen:1972:UBW

[Yue72] Chung-Kwong Yuen. Upper bounds on Walsh transforms. IEEE Transactions on Computers, C-21(12):1273–1280, December 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672091.

Yuen:1973:TF

[Yue73a] C. K. Yuen. On the twiddling factors. IEEE Transactions on Computers, C-22(5):544– 545, May 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672355.

Yuen:1973:PON

[Yue73b] Chung-Kwong Yuen. The printing of octal numerals. IEEE Transactions on Computers, C-22(1): 111, January 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672199.

Yuen:1974:CSN

[Yue74a] C. K. Yuen. Comment on “Some New Results on Average Worst Case Carry”. IEEE Transactions on Computers, C-23(3): 333, March 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672529. See [Bri73].

Yuen:1974:EWP

[Yue74b] Chung Kwong Yuen. Evaluation of Walsh power spectrum of nearly white signals. IEEE Transactions on Computers, C-23(11):1213–1214, November 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672426.

Yuen:1975:CRW

[Yue75a] C. K. Yuen. Computing robust Walsh–Fourier transform by error product minimization. IEEE Transactions on Computers, C-24(3): 313–317, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672806.

Yuen:1975:NBA

[Yue75b] C. K. Yuen. A note on base −2 arithmetic logic. IEEE Transactions on Computers, C-24(3):325– 329, March 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672809. See comments [Agr77].

Yuen:1975:FPR

[Yue75c] C. K. Yuen. On the floating point representation of complex numbers. IEEE Transactions on Computers, C-24(8):846– 848, August 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1672914. See comments [Bau77].

Yuen:1975:TCC

[Yue75d] C. K. Yuen. A theorem on the cyclic correlations of Walsh functions. IEEE Transactions on Computers, C-24(10):1028–1029, October 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672712.

Yuen:1975:FAW

[Yue75e] Chung-Kwong Yuen. Function approximation by Walsh series. IEEE Transactions on Computers, C-24 (6):590–598, June 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672864.

Yuen:1976:CAR

[Yue76] C.-K. Yuen. Comments on “The Application of the Rademacher– Walsh Transform to Boolean Function Classification and Threshold Logic Synthesis”. IEEE Transactions on Computers, C-25(7): 766–767, July 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1674688. See [Edw75].

Yuen:1977:ARa

[Yue77a] C. K. Yuen. Author’s reply. IEEE Transactions on Computers, C-26 (2):192, February 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5009302.

Yuen:1977:ARb

[Yue77b] C. K. Yuen. Author’s reply. IEEE Transactions on Computers, C-26 (5):511–512, May 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674870.

Yuen:1977:NRD

[Yue77c] C. K. Yuen. A new representation for decimal numbers. IEEE Transactions on Computers, C-26(12):1286–1288, December 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674792.

Yuen:1977:LDF

[Yue77d] C. K. Yuen. On Little’s digital filtering algorithm. IEEE Transactions on Computers, C-26(3): 309, March 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1674829.

Yuen:1977:TRN

[Yue77e] Chung-Kwong Yuen. Testing random number generators by Walsh transform. IEEE Transactions on Computers, C-26(4):329– 333, April 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674842.

Yuen:1978:AGC

[Yue78] C. K. Yuen. Analog-to-gray code conversion. IEEE Transactions on Computers, C-27(10):971– 973, October 1978. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674979.

Yau:1975:MFD

[YY75] S. S. Yau and Shih-Chien Yang. Multiple fault detection for combinational logic circuits. IEEE Transactions on Computers, C-24 (3):233–242, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1672797.

Zahn:1971:GTM

[Zah71] C. T. Zahn. Graph-theoretical methods for detecting and describing Gestalt clusters. IEEE Transactions on Computers, C-20(1):68– 86, January 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1671676.

Zalcstein:1971:NFC

[Zal71] Y. Zalcstein. A note on fast cyclic convolution. IEEE Transactions on Computers, C-20(6):

665–666, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671912.

Zeigler:1971:FHR

[Zei71] B. P. Zeigler. Feedback in homomorphic realizations. IEEE Transactions on Computers, C-20(6): 685–688, June 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671920.

Zucker:1977:ARL

[ZHR77a] S. W. Zucker, R. A. Hummel, and A. Rosenfeld. An application of relaxation labeling to line and curve enhancement. IEEE Transactions on Computers, C-26 (4):394–403, April 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674848. See correction [ZHR77b].

Zucker:1977:CAR

[ZHR77b] S. W. Zucker, R. A. Hummel, and A. Rosenfeld. Correction to “An Application of Relaxation Labeling to Line and Curve Enhancement”. IEEE Transactions on Computers, C-26(9):922–929, September 1977. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=

1674940. See [ZHR77a].

Zisapel:1976:SGM

[Zis76] Y. Zisapel. A simplified general method for static hazard detection. IEEE Transactions on Computers, C-25(11):1160–1164, November 1976. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674573.

Zisapel:1979:DHC

[ZKK79] Y. Zisapel, M. Krieger, and J. Kella. Detection of hazards in combinational switching circuits. IEEE Transactions on Computers, C-28 (1):52–56, January 1979. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1675222.

Zohar:1970:NRC

[Zoh70] S. Zohar. Negative radix conversion. IEEE Transactions on Computers, C-19(3):222–226, March 1970. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1671493. See comments [Paw71, Wad71].

Zohar:1971:AR

[Zoh71] S. Zohar. Author’s reply. IEEE Transactions on Computers, C-20 (5):587, May 1971. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1671895.

Zohar:1973:CRD

[Zoh73a] S. Zohar. The counting recursive digital filter. IEEE Transactions on Computers, C-22(4):338– 347, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672313.

Zohar:1973:FHF

[Zoh73b] S. Zohar. Fast hardware Fourier transformation through counting. IEEE Transactions on Computers, C-22(5):433–441, May 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL https://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=

1672338. See comments [Zoh74].

Zohar:1973:NHR

[Zoh73c] S. Zohar. New hardware realizations of nonrecursive digital filters. IEEE Transactions on Computers, C-22(4):328– 338, April 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672312.

Zohar:1973:DCR

[Zoh73d] Shalhav Zohar. A/D conversion for radix (−2). IEEE Transactions on Computers, C-22(7):698– 701, July 1973. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        5009137.

Zohar:1974:CFH

[Zoh74] S. Zohar. Comments on “Fast Hardware Fourier Transformation Through Counting”. IEEE Transactions on Computers, C-23(9): 989, September 1974. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=

1672664. See [Zoh73b].

Zohar:1976:RRD

[Zoh76a] S. Zohar. A realization of the RAM digital filter. IEEE Transactions on Computers, C-25(10): 1048–1052, October 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674546.

Zohar:1976:RTR

[Zoh76b] S. Zohar. Rounding and truncation in radix (-2) systems. IEEE Transactions on Computers, C-25 (5):464–469, May 1976. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?tp=&arnumber=
                                        1674634.

Zahn:1972:FDP

[ZR72] Charles T. Zahn and Ralph Z. Roskies. Fourier descriptors for plane closed curves. IEEE Transactions on Computers, C-21(3):269– 281, March 1972. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
5008949.

Zucker:1975:PST

[ZRD75] S. W. Zucker, A. Rosenfeld, and L. S. Davis. Picture segmentation by texture discrimination. IEEE Transactions on Computers, C-24(12):1228–1233, December 1975. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672762.

Zobrist:1975:BDF

[ZT75] A. L. Zobrist and W. B. Thompson. Building a distance function for Gestalt grouping. IEEE Transactions on Computers, C-24 (7):718–728, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=
1672885.